### v1

# Statistical Inference for Temporal Difference Learning with Linear Function Approximation

Abstract
Statistical inference with finite-sample validity for the value function of a given policy in Markov decision processes (MDPs) is crucial for ensuring the reliability of reinforcement learning. Temporal Difference (TD) learning, arguably the most widely used algorithm for policy evaluation, serves as a natural framework for this purpose. In this paper, we study the consistency properties of TD learning with Polyak-Ruppert averaging and linear function approximation, and obtain three significant improvements over existing results. First, we derive a novel sharp high-dimensional probability convergence guarantee that depends explicitly on the asymptotic variance and holds under weak conditions. We further establish refined high-dimensional Berry-Esseen bounds over the class of convex sets that guarantee faster rates than those in the literature. Finally, we propose a plug-in estimator for the asymptotic covariance matrix, designed for efficient online computation. These results enable the construction of confidence regions and simultaneous confidence intervals for the linear parameters of the value function, with guaranteed finite-sample coverage. We demonstrate the applicability of our theoretical findings through numerical experiments.
1 Introduction
Statistical inference tasks, such as constructing confidence regions or simultaneous confidence intervals, are often addressed by deriving distributional theory such as central limit theorems (CLTs) for the estimator in use. Due to the high dimensionality of modern science and engineering applications, there has been a surge of interests in deriving convergence results that are valid in a finite-sample manner. In Reinforcement Learning (RL), a discipline that underpins many recent machine learning breakthroughs (Sutton and Barto, (2018); Agarwal et al., (2019)) one central question is to evaluate with confidence the quality of a given policy, measured by its value function. As RL is often modeled as decision making in Markov decision processes (MDPs), the task of statistical inference needs to accommodate the online nature of the sampling mechanism.
Temporal Difference (TD) learning (Sutton, (1988)) is arguably the most widely used algorithm designed for this purpose. TD learning, which is an instance of stochastic approximation (SA) method (Robbins and Monro, (1951)), approximates the value function of a given policy in an iterative manner. Towards understanding the non-asymptotic convergence rate of TD to the target value function, extensive recent efforts have been made (see, e.g. Nemirovski et al., (2009); Rakhlin et al., (2012); Li et al., 2024a ; Bhandari et al., (2021); Dalal et al., (2018); Khamaru et al., (2021); Srikant and Ying, (2019); Lakshminarayanan and Szepesvari, (2018); Mou et al., (2020); Durmus et al., (2022); Patil et al., (2023); Li et al., 2024b ; Samsonov et al., (2023) and references therein) and it is known that for the tabular case, where both state and action spaces are assumed to be finite, the TD algorithm is minimax optimal (Li et al., 2024a ). Another thread of work considers the asymptotic performance of TD where it has been shown that with Polyak-Ruppert averaging (Polyak and Juditsky, (1992); Ruppert, (1998)), TD converges to a Gaussian distribution in the large sample limit (Fort, (2015); Mou et al., (2020); Li et al., (2023)). However, these two types of results remain insufficient for conducting statistical inference that is valid in the finite-sample fashion.
Figure 1: Emergence of the first entry of TD estimation error on an MDP with a 3-dimensional feature map.
To close this gap, in this paper, based on the TD learning algorithm, we design confidence regions and simultaneous confidence intervals for the linear coefficients of the value function when the it is approximated by a linear combination of features. To achieve this goal, we first provide a novel sharp high-dimensional probability convergence guarantee for the TD iterates, and follow up with refined high-dimensional Berry-Esseen bounds over the class of convex sets. Finally, we propose a plug-in estimator for the asymptotic covariance matrix, improving upon other alternatives proposed by recent literature (Chen et al., (2020), White and White, (2010), Hanna et al., (2017), Hao et al., (2021), Ramprasad et al., (2022)) in both computational and statistical efficiency. We summarize our main contributions in details in the following section. While we are writing this paper, we become aware of a great recent paper by Samsonov et al., (2024), which considers similar problems to ours and derives a Berry-Esseen bound for linear SA with Polyak-Ruppert averaging. It remains unclear whether their general framework gives tight results when specialized to TD learning. We offer more detailed comparisons between our results and theirs in the following sections.
1.1 Contributions
Focusing on TD learning with linear function approximation, polynomial-decaying stepsize, Polyak-Ruppert averaging and $i.i.d.$ samples, this paper makes the following theoretical and methodological contributions.
  1. 1.
    Improved high-probability guarantees. We show that, for an appropriate choice of the initial stepsize, the $L_{2}$ norm of the estimation error is of order
    $O\left(\sqrt{\frac{\mathsf{Tr}(\bm{\Lambda}^{\star})}{T}\log\frac{d}{\delta}}\right),$
    with probability at least $1-\delta$, where $T$ denotes the number of iterations, $\bm{\Lambda}^{\star}$ represents the asymptotic variance matrix of the estimation error, and $\mathsf{Tr}(\bm{\Lambda}^{\star})$ its trace. To the best of our knowledge, this is the first high-probability convergence guarantee for this problem that matches the asymptotic variance, up to logarithm factors, without requiring a projection step to control the iterative trajectory. Table 1 compares our high-probability convergence guarantees with the latest results in the literature.
  2. 2.
    Sharper Berry-Esseen bounds. We derive a new high-dimensional Berry-Esseen bound for the estimation error over the class of convex sets, achieving a $O(T^{-\frac{1}{3}})$ rate as the function of the number of iterations. Our results improve upon the sharpest rate to date, or order $O(T^{-\frac{1}{4}})$, due to Samsonov et al., (2024).
  3. 3.
    Novel variance estimator. We develop a computationally efficient online algorithm to construct an estimator $\widehat{\bm{\Lambda}}_{T}$ of the asymptotic variance $\bm{\Lambda}^{\star}$, and establish a converge rate for $\widehat{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star}$ of order $O(T^{-\frac{1}{3}})$ in Frobenius norm.
We also verify our theoretical results by numerical simulations. As an intuitive illustration, Figure 1 shows how the empirical distribution of TD estimation error in $1,000$ independent trials (as shown in blue histograms), scaled by the square-root of the number of iterations $T$, converges to its asymptotic Gaussian distribution (as shown in red curves).
1.2 Notation
For any complex number $z\in\mathbb{C}$, we use $\mathsf{Re}(z)$ to denote its real part. Throughout the paper, we use boldface small letters to denote vectors and capital letters to denote matrices. For any vector $\bm{x}$, we use $\|\bm{x}\|_{2}$ to denote its $L_{2}$ norm; for any matrix $\bm{M}$, we use $\mathsf{Tr}(\bm{M})$ to denote its trace, $\mathsf{det}(\bm{M})$ to denote its determinant, $\mathsf{cond}(\bm{M})$ to denote its conditional number, $\|\bm{M}\|$ to denote its spectral norm (i.e. the largest singular value), and $\|\bm{M}\|_{\mathsf{F}}$ to denote its Frobenius norm, i.e. $\|\bm{M}\|_{\mathsf{F}}=\sqrt{\mathsf{Tr}(\bm{M}^{\top}\bm{M})}$. For any symmetric matrix $\bm{M}\in\mathbb{R}^{d\times d}$, we use $\lambda_{\max}(\bm{M})=\lambda_{1}(\bm{M})\geq\lambda_{2}(\bm{M})\geq...\geq\lambda_{d}(\bm{M})=\lambda_{\min}(\bm{M})$ to denote its eigenvalues. If a matrix $\bm{M}$ is positive (semi-)definite, i.e. $\lambda_{\min}(\bm{M})>(\geq)0$, we write $\bm{M}\succ(\succeq)\bm{0}$; when two matrices $\bm{X}$ and $\bm{Y}$ satisfy $\bm{X}-\bm{Y}\succ(\succeq)\bm{0}$, we write $\bm{X}\succ(\succeq)\bm{Y}$ or equivalently $\bm{Y}\prec(\preceq)\bm{X}$. For a vector $\bm{x}\in\mathbb{R}^{d}$, we use $\{x_{i}\}_{1\leq i\leq d}$ to denote its entries; for a matrix $\bm{X}\in\mathbb{R}^{m\times n}$, we use $\{X_{ij}\}_{1\leq i\leq m,1\leq j\leq n}$ to denote its entries. The vectorization of matrix $\bm{X}\in\mathbb{R}^{m\times n}$, denoted as $\textbf{vec}(\bm{X})$, is defined as a vector in $\mathbb{R}^{mn}$ with entries $[\textbf{vec}(\bm{X})]_{(i-1)m+j}=\bm{X}_{ij}$ for every $1\leq i\leq m$ and $1\leq j\leq n$. For any matrices $\bm{X}\in\mathbb{R}^{m\times n}$, and $\bm{Y}\in\mathbb{R}^{p\times q}$, their Kronecker product, denoted as $\bm{X}\otimes\bm{Y}$, is a matrix of the size of $mp\times nq$, and can be written in blocked form as
$\bm{X}\otimes\bm{Y}=\begin{bmatrix}X_{11}\bm{Y}&X_{12}\bm{Y}&\ldots&X_{1n}\bm{Y}\\ X_{21}\bm{Y}&X_{22}\bm{Y}&\ldots&X_{2n}\bm{Y}\\ \ldots&\ldots&\ldots&\ldots\\ X_{m1}\bm{Y}&X_{m2}\bm{Y}&\ldots&X_{mn}\bm{Y}\end{bmatrix}.$
With subscripts, it can be represented as $(\bm{X}\otimes\bm{Y})_{(i-1)m+j,(k-1)p+\ell}=X_{ik}Y_{j\ell}$.
For sequences $\{f_{t}\}_{t\in\mathbb{N}}$ and $\{g_{t}\}_{t\in\mathbb{N}}$ of numbers, we write $f_{t}\lesssim g_{t}$, or $f_{t}=O(g_{t})$, to rsignify that there exists a quantity $C>0$ independent of $t$ such that $f_{t}\leq Cg_{t}$ for all $t$. We will use the notatin $\widetilde{C}$ to express a quantity independent of $T$, the number of iterations/sample size, but possibly dependent on other problem-related parameters. We will measure the difference between two probability distributions on $\mathbb{R}^{d}$ using the convex distance, the maximal distance convex sets. Specifically, for any two probability distributions $P$ and $Q$, we set
$d_{\mathsf{C}}(P,Q)=\sup_{\mathcal{A}\in\mathscr{C}_{d}}|P(\mathcal{A})-Q(\mathcal{A})|.$
where $\mathscr{C}_{d}$ is the class of all convex set in $\mathbb{R}^{d}$. We use $d_{\mathsf{TV}(P,Q)}$ to denote the Total Variation (TV) distance between $P$ and $Q$.
paper stepsize convergence guarantee constraint on $t,T,\delta$
[…]
Li et al., 2024b $\eta_{t}=\eta$ $\|\overline{\bm{\Delta}}_{T}\|_{2}\leq O\left(T^{-\frac{1}{2}}\log^{\frac{1}{2}}\frac{T}{\delta}\right)$ $\eta\sqrt{\log\frac{T}{\delta}}\leq\widetilde{C}$
Samsonov et al., (2023) $\eta_{t}=\eta$ $\|\overline{\bm{\Delta}}_{T}\|_{2}\leq O\left(T^{-\frac{1}{2}}\log\frac{T}{\delta}\right)$ $\eta\log\frac{T}{\delta}\leq\widetilde{C}$
This work, Theorem 3.1 $\eta_{t}=\eta_{0}t^{-\alpha},\alpha\in\left(\frac{1}{2},1\right)$ $\|\overline{\bm{\Delta}}_{T}\|_{2}\leq O\left(T^{-\frac{1}{2}}\sqrt{\log\frac{d}{\delta}}\right)$ $\eta_{0}\sqrt{\log\frac{1}{\delta\eta_{0}}}\leq\widetilde{C}$
Table 1: Comparisons between our results on convergence rates for TD with prior work that derived with high probability control. The quantity $\widetilde{C}$, whose value may change at each occurrence, may depend on problem-related parameters other than $T$ and $\delta$.
2 Background
Markov decision process.
An infinite-horizon Markov Decision Process (MDP) (Bertsekas,, 2017; Sutton and Barto,, 2018) can be represented mathematically by the tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},P,r,\gamma)$. Here, $\mathcal{S}$ and $\mathcal{A}$ represent the state space of the environment and the action space available to an agent respectively, both assumed finite throughout this paper; the function $P:\mathcal{S}\times\mathcal{A}\mapsto\Delta(\mathcal{S})$ is a transition kernel that associates to any state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$ the probability distribution $P(\cdot\mid s,a)$ of the next state of the environment when the agent executes action $a$ while the environment is in state $s$; the reward function $r:\mathcal{S}\times\mathcal{A}\mapsto[0,1]$ measures the immediate reward to the agent after taking action $a$ at state $s$; finally, $\gamma\in(0,1]$ is the discount factor, used to assess the impact of future rewards.
In choosing their actions, the agent follows a randomized action selection rule, or policy, represented as a (time-invariant) conditional probability distribution $\pi:\mathcal{S}\mapsto\Delta(\mathcal{A})$ over the action space given the state of the environment. For any policy $\pi$, its associated reward function $r^{\pi}:\mathcal{S}\mapsto[0,1]$ is the expected reward of the next action selected according to the policy $\pi$, i.e.
$r^{\pi}(s):=\mathbb{E}_{a\sim\pi(\cdot\mid s)}[r(s,a)],\quad s\in\mathcal{S}.$
For readability, we will omit the superscript $\pi$ and use $r(\cdot)$ to denote $r^{\pi}(\cdot)$. Any policy $\pi$ induces a transition kernel $P^{\pi}:\mathcal{S}\mapsto\Delta(\mathcal{S})$, representing the probability distribution of the next state, given the current state $s$ and the policy $\pi$. Formally,
$P^{\pi}(s^{\prime}\mid s)=\mathbb{E}_{a\sim\pi(\cdot\mid s)}P(s^{\prime}\mid s,a),\quad s^{\prime}\in\mathcal{S}.$
We assume throughout that there exists a unique stationary distribution $\mu^{\pi}\in\Delta(\mathcal{S})$ for the transition matrix ${P}^{\pi}$, assigning positive probability to each state. That is,
$0<\mu^{\pi}(i)=\mathbb{P}_{s\sim\mu^{\pi},s^{\prime}\sim P^{\pi}(\cdot\mid s)}(s^{\prime}=i),\quad\forall i\in\mathcal{S}.$
The value function and the Bellman equation.
The worth of a policy $\pi$ is measured through its value function $V^{\pi}:\mathcal{S}\mapsto\mathbb{R}_{+}$ returning, for each initial state, the expected value of all cumulative future rewards, appropriately discounted, when policy $\pi$ is executed under the MDP $\mathcal{M}$. Formally,
$V^{\pi}(s):=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\,\big{|}\,s_{0}=s\right],\quad s\in\mathcal{S}.$ (1)
A fundamental, defining feature of the value function is that it satisfies a certain self-consistency property, known as the Bellman equation, a functional fixed-point equation of the form
$V^{\pi}(s)=r(s)+\mathbb{E}_{s^{\prime}\sim P^{\pi}(\cdot\mid s)}V^{\pi}(s^{\prime})=:\mathcal{T}V^{\pi}(s),$ (2)
where $\mathcal{T}$ is the Bellman’s operator. Indeed, the value function is the unique solution to the Bellman equation. With a slight abuse of notation, we shall omit the superscript $\pi$ in $V^{\pi}$ and abbreviate it as $V$ as the target policy is considered fixed.
Linear policy approximation and the projected Bellman equation.
In most practical scenarios, though finite, the state space $\mathcal{S}$ is prohibitively large. As a result, the value function is often very complex and impractical to compute. To overcome this issue, it is customary to consider a parametric approximation of $V$ through a linear function, which is computationally tractable and amenable to statistical estimation (see, e.g. Tsitsiklis and Van Roy, (1997); Dalal et al., (2018); Bhandari et al., (2021); Srikant and Ying, (2019); Lakshminarayanan and Szepesvari, (2018)) In detail, for a given feature map $\bm{\phi}:\mathcal{S}\mapsto\mathbb{R}^{d}$, mapping each state in $\mathcal{S}$ to a feature vector in $\mathbb{R}^{d}$, where $d$ is usually smaller than $|\mathcal{S}|$, the value function at $s$, $V(s)$, is expproximated by
$V_{\bm{\theta}}(s):={\bm{\phi}}(s)^{\top}{\bm{\theta}},\qquad\quad s\in\mathcal{S},$ (3)
where ${\bm{\theta}}\in\mathbb{R}^{d}$ is a vector of linear coefficients.
Computing the optimal linear approximation to $V$ can be cast as a least-squares minimization problem over $\mathbb{R}^{d}$ of finding a coefficient vector $\bm{\theta}$ minimizing the expected difference between $V_{\bm{\theta}}$ and its Bellman projection $\mathcal{T}V_{\bm{\theta}}$. That is,
$\bm{\theta}^{\star}:=\underset{\bm{\theta}\in\mathbb{R}^{d}}{\text{argmin }}\mathbb{E}_{s\sim\mu}(V_{\bm{\theta}}(s)-\mathcal{T}V_{\bm{\theta}}(s))^{2},$ (4)
where we recall that $\mu$ that denote the stationary distribution of $P^{\pi}$. It turns out that the problem (4) admits a unique solution (Tsitsiklis and Van Roy,, 1997), given by
$\bm{\theta}^{\star}:=\bm{A}^{-1}\bm{b},$ (5)
where
$\bm{A}$ $:=\mathop{\mathbb{E}}\limits_{s\sim\mu,s^{\prime}\sim P(\cdot\mid s)}\left[\bm{\phi}(s)\left(\bm{\phi}(s)-\gamma\bm{\phi}(s^{\prime})\right)^{\top}\right]\in\mathbb{R}^{d\times d},$ (6a)
[…]
For notional simplicity, we further introduce the notation $\bm{\Sigma}$ as feature covariance matrix
$\bm{\Sigma}:=\mathbb{E}_{s\sim\mu}[\bm{\phi}(s)\bm{\phi}^{\top}(s)],$ (7)
and use $\lambda_{\Sigma}$ and $\lambda_{0}$ to denote its largest and smallest eigenvalue, namely
$\lambda_{0}:=\lambda_{\min}(\bm{\Sigma}),\quad\text{and}\quad\lambda_{\Sigma}:=\lambda_{\max}(\bm{\Sigma}).$ (8)
The Temporal Difference (TD) learning algorithm.
To estimate the optimal vector $\bm{\theta}^{*}$ given by (4) we use the observe data, consisting of $T$ triplets of (state, action, reward),
$\{(s_{0},a_{0},r_{0}),(s_{1},a_{1},r_{1}),\ldots(s_{T},a_{T},r_{T})\},$
where, for each $t=0,1,\ldots,T$, $a_{t}\sim\pi(\cdot\mid s_{t})$, $r_{t}=r(s_{t},a_{t})$ and $s_{t+1}\sim P(\cdot\mid s_{t},a_{t})$.
The TD learning algorithm is a stochastic approximation algorithm that iteratively approximates the solution to the linear system $\bm{A}\bm{\theta}=\bm{b}$. Specifically, given a sequence of pre-selected step size $\{\eta_{t}\}_{t\geq 1},$ TD proceeds using the updating rule
$\bm{\theta}_{t}$ $=\bm{\theta}_{t-1}-\eta_{t}(\bm{A}_{t}\bm{\theta}_{t-1}-\bm{b}_{t}),$ (9a)
where
$\bm{A}_{t}$ $:=\bm{\phi}(s_{t-1})\left(\bm{\phi}(s_{t-1})-\gamma\bm{\phi}(s_{t})\right)^{\top}$ (9b)
and $\text{and }\qquad\bm{b}_{t}$ $:=\bm{\phi}(s_{t-1})r_{t-1}.$ (9c)
After a total number of $T$ iterations, the final iteration $\bm{\theta}_{T}$ is returned as the estimator for $\bm{\theta}^{\star}$.
Polyak-Ruppert averaging.
The Polyak-Ruppert averaging (Polyak and Juditsky,, 1992; Ruppert,, 1998) is a stochastic approximation methodology known to reduce the variability and accelerate the convergence of algorithms by averaging across the iterates. In addition, with a proper choice of stepsize, the averaged iterates converge to a Gaussian distribution asymptotically, which enables valid statistical inference that holds with large samples. When applied to the TD algorithm, Polyak-Ruppert averaging returns an estimator of $\bm{\theta}^{*}$ of the form
$\overline{\bm{\theta}}_{T}$ $=\frac{1}{T}\sum_{t=1}^{T}\bm{\theta}_{t},$ (10)
with associated estimation error
$\overline{\bm{\Delta}}_{T}=\overline{\bm{\theta}}_{T}-\bm{\theta}^{\star}.$ (11)
TD with Polyak-Ruppert averaging has been shown to deliver significant improvements, both of computational and statistical nature. While there is a recent strand of non-asymptotic results established for stochastic approximation (e.g. Mou et al., (2021); Moulines and Bach, (2011); Nemirovski et al., (2009); Lan, (2012)), the current results on Berry-Esseen bounds tailored to policy evaluation with linear function approximation still remain inadequate. In the work, we derive improved consistency and Gaussian approximation rates for $\overline{\bm{\Delta}}_{T}$. It is important to mention that the choice of learning rates $\{\eta_{t}\}_{t\geq 0}$ in the update (9) plays a critical role in determining the statistical properties of both the original and the averaged TD estimator, including their convergence rates and asymptotic distributions. This work focuses on the typical choice — the polynomial-decaying stepsizes — where $\eta_{t}=\eta_{0}t^{-\alpha}$ for $\alpha\in(1/2,1)$ and $\eta_{0}$ satisfying
$\eta_{0}<{1}/{(2\lambda_{\Sigma})}.$ (12)
3 Main results
While in most real-world RL scenarios the state-action-reward tuples are generated following a Markov decision process, there are applications where $i.i.d.$ samples become available, such as in game playing and autonomous driving. Furthermore, it is illuminating to first understand the statistical properties of this simpler setting before considering the more complicated setting of Markov samples. In this paper, like in most of the RL literature, we will make this simplifying assumption. Specifically, we will analyze the performance of the TD learning algorithm with Polyak-Ruppert averaging assuming a sample consisting of $T$ $i.i.d.$ state-action-reward tuples of the form
$\{(s_{t},a_{t},s_{t}^{\prime},r_{t})\}_{t=1}^{T},$ (13)
where $(s_{1},\ldots,s_{T})\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mu$ and, for eact $t$, $a_{t}\sim\pi(s_{t}),r_{t}=r(s_{t},a_{t})$ and $s_{t}^{\prime}\sim P(\cdot\mid s_{t},a_{t})$. Correspondingly, $\bm{A}_{t}$ and $\bm{b}_{t}$ are defined as
$\bm{A}_{t}$ $=\bm{\phi}(s_{t})\left(\bm{\phi}(s_{t})-\gamma\bm{\phi}(s_{t}^{\prime})\right)^{\top},$
$\text{and}\qquad\bm{b}_{t}$ $=r_{t}\bm{\phi}(s_{t}).$
In this setting, with polynomial-decaying stepsizes, it has been shown (Polyak and Juditsky, (1992); Borkar, (2008); Fort, (2015)) that the estimation error satisfies
[…]
where the covariance matrix $\bm{\Lambda}^{\star}$ is given as
$\bm{\Lambda}^{\star}=\bm{A}^{-1}\bm{\Gamma}\bm{A}^{-\top},\quad\text{with}$ (15)
$\bm{\Gamma}=\mathbb{E}[(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}].$ (16)
Thus, in the fixed-dimensional setting, the above result implies the rate
$\|\overline{\bm{\Delta}}_{T}\|_{2}=O_{P}\left(\sqrt{\frac{\mathsf{Tr}(\bm{\Lambda}^{\star})}{T}}\right),$ (17)
which may serve as a comparison term. Our main goal is to demonstrate how the TD algorithm can be used to construct confidence regions and simultaneous confidence intervals for $\bm{\theta}^{*}$, the linear parameters of the best value function linear approximation (see (5)). Towards this goal, we shall establish a high-probability convergence guarantee for $\overline{\bm{\Delta}}_{T}$, develop high-dimensional Berry-Esseen bounds over the class of convex sets, and finally build confidence regions/intervals for $\bm{\theta}^{\star}$ using the estimator $\overline{\bm{\theta}}_{T}$. These topics will be addressed in order.
3.1 High-probability convergence guarantees
We start by controlling the convergence rate of the TD iterates with Polyak-Ruppert averaging in high probability. Before stating our main results, let us first introduce additional quantities
$\overline{\bm{\Lambda}}_{T}=\frac{1}{T}\sum_{t=1}^{T}\bm{Q}_{t}\bm{\Gamma}\bm{Q}_{t}^{\top},$ (18)
where $\bm{\Gamma}$ is defined in (16) an, for each $t$, we set
$\bm{Q}_{t}=\eta_{t}\sum_{j=t}^{T}\prod_{k=t+1}^{j}(\bm{I}-\eta_{k}\bm{A}).$ (19)
The matrix $\overline{\bm{\Lambda}}_{T}$ plays the role of a non-asymptotic covariance matrix for $\overline{\bm{\theta}}_{T}$.
Theorem 3.1.
Consider TD with Polyak-Ruppert averaging (9) with independent samples and decaying stepsizes $\eta_{t}=\eta_{0}t^{-\alpha}$ for $\alpha\in(\frac{1}{2},1)$. For any $\delta\in(0,1)$, there exists $\eta_{0}=\eta_{0}(\delta)>0$, such that
$\|\overline{\bm{\Delta}}_{T}\|_{2}$ $\lesssim\sqrt{\frac{\mathsf{Tr}(\overline{\bm{\Lambda}}_{T})}{T}\log\frac{d}{\delta}}+\widetilde{C}T^{-\frac{\alpha+1}{2}}\log\frac{T}{\delta}$
with probability at least $1-\delta$, where $\widetilde{C}$ is a problem-related quantity independent of $T$ and $\delta$.
The proof can be found in Appendix B.2, where the value of $\widetilde{C}$ is given in equation (75). Throughout the paper, for the sake of readability, we have opted to express unwieldy problem-dependent quantities that do not directly depend on $T$ in implicit form; their values can be tracked in the proofs. Accordingly, when writing our finite-sample bounds, we will first present the leading terms in $T$.
In a nutshell, Theorem 3.1 states that $\overline{\bm{\theta}}_{T}$ converges to $\bm{\theta}^{*}$ in $\ell_{2}$ error at a rate $O(\sqrt{{\mathsf{Tr}(\overline{\bm{\Lambda}}_{T})}/{T}})$ up to log factor. It is worth comparing the high-probability sample complexity bound of Theorem 3.1 with those in the literature. To the best of our knowledge, the tightest results were attained by Dalal et al., (2018) for inverse-decaying stepsizes and Li et al., 2024b ; Samsonov et al., (2023) for constant stepsize as summarized in Table 1. With polynomial-decaying stepsizes, Theorem 3.1 is the first of its kind to achieve tight high-probability convergence guarantees up to log factors. Another subtle but consequential difference concerns the restriction on the step size. For our bound to hold, the choice of the initial stepsize $\eta_{0}$ depends on $\delta$, and is required to be sufficiently small to satisfy $\eta_{0}\sqrt{\log{1}/(\delta\eta_{0}})\leq\widetilde{C}$, where $\widetilde{C}$ is independent of $T$ and $\delta$ but may depend on the other parameters of interest. Though not ideal, this condition is a considerable weakening of the stepsize requirements imposed by Samsonov et al., (2023) and Li et al., 2024b , which depend on both $\delta$ and $T$; see Table 1.
The bound of Theorem 3.1 can be made dependent on the asymptotic variance $\bm{\Lambda}^{\star}$ without affecting the rate. Indeed, as we will see below, the difference between $\overline{\bm{\Lambda}}_{T}$ and $\bm{\Lambda}^{\star}$ converges at the rate of $O(T^{\alpha-1})$. Thus, the dominant terms in Theorem 3.1 match the convergence rate (17) implied by the asymptotic distribution up to a logarithm term.
3.2 High-dimensional Berry-Esseen bounds
Next we derive high-dimensional Berry-Esseen bounds for TD with Polyak-Ruppert averaging. We first characterize the convex distance between $\sqrt{T}\overline{\bm{\Delta}}_{T}$ and $\overline{\bm{\Lambda}}_{T}^{1/2}\bm{z}$, where $\bm{z}$ is a $d$-dimensional standard gaussian. The proof of this result is included in Appendix B.3.
Theorem 3.2.
Consider the TD with Polyak-Ruppert averaging with independent samples and decaying stepsizes $\eta_{t}=\eta_{0}t^{-\alpha}$ for $\alpha\in(\frac{1}{2},1)$. Then,
$d_{\mathsf{C}}(\sqrt{T}\overline{\bm{\Delta}}_{T},\overline{\bm{\Lambda}}_{T}^{1/2}\bm{z})\lesssim\sqrt{\frac{\eta_{0}}{(1-\alpha)(1-\gamma)\lambda_{0}}}\mathsf{Tr}(\bm{\Gamma})\|\bm{\Gamma}^{-1}\|T^{-\frac{\alpha}{2}}+\widetilde{C}\;T^{-\frac{1}{2}}\log T,$
where $\widetilde{C}$ is a problem related quantity indepedent of $T$.
To relate this result with the classical CLT (as in (14)), we will use the following comparison bound for the total variation distance between $\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T})$ and $\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})$; see Appendix B.4 for the proof.
Theorem 3.3.
Let $\bm{\Lambda}^{\star}$ and $\overline{\bm{\Lambda}}_{T}$ be defined as in (15) and (18), respectively, with the stepsizes $\eta_{t}=\eta_{0}t^{-\alpha}$, for all $t$ and any $\alpha\in(\frac{1}{2},1)$. Then,
$d_{\mathsf{C}}(\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T}),\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\leq d_{\mathsf{TV}}(\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T}),\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\lesssim\frac{\sqrt{d\mathsf{cond}(\bm{\Gamma})}}{(1-\gamma)\lambda_{0}\eta_{0}}T^{\alpha-1}+\widetilde{C}\;T^{2\alpha-2},$
where $\widetilde{C}$ is a problem related quantity indepedent of $T$.
Comparison with existing results.
The latest and sharpest Berry-Esseen bounds for stochastic approximation under condition (12) are du to Samsonov et al., (2024). Specifically, the authors show that
$d_{\mathsf{C}}(\sqrt{T}\overline{\bm{\Delta}}_{T},{\bm{\Lambda}}^{\star 1/2}\bm{z})\leq\widetilde{C}\;T^{\frac{\alpha-1}{2}},$
where $\widetilde{C}$ depend on other problem parameters. Upon choosing $\alpha=\frac{1}{2}$, they obtain a convergence rate of order $O(T^{-\frac{1}{4}})$. In contrast, the Berry-Esseen bounds of Theorem 3.2 and the comparison bound of Theorem 3.3 indicate that, selecting the optimal $\alpha=\frac{2}{3}$ leads to the improved rate
$d_{\mathsf{C}}(\sqrt{T}\overline{\bm{\Delta}}_{T},\bm{\Lambda}^{\star 1/2}\bm{z})\leq\widetilde{C}\;T^{-\frac{1}{3}}.$
The proof of Theorem 3.2 relies on Theorem 3.1 of Shao and Zhang, (2021), a high-dimensional Berry-Esseen bound for non-linear statistics, which is also used by Samsonov et al., (2024) to establish their main result about Gaussian approximations for the TD error, Theorem 3.4 therein. We are able to improve their the rate of convergence from $O(T^{\frac{1}{2}-\alpha})$ to $O(T^{-\frac{\alpha}{2}})$ by carefully handling some of the quantities appearing in the bound of Theorem 3.1 of Shao and Zhang, (2021) [^2] Specifically, the term $I_{3}$ of Equation (79) in Appendix B.3.. This sharpening is non-trivial and may be of independent interest, as it may be potentially generalized to analyze stochastic gradient algorithms applied to non-linear functions (as of Theorem 3.4 of Shao and Zhang, (2021)). We leave this extension to future work.
3.3 Online estimation of the covariance matrix
Thus far, we have provided high-dimensional Berry-Esseen bounds to quantify the non-asymptotic discrepancy between $\overline{\bm{\Delta}}_{T}$ and its Gaussian limit. In order to invoke these results for constructing confidence region or confidence intervals for $\bm{\theta}^{\star}$, it is necessary to estimate the covariance matrix $\bm{\Lambda}^{\star}$. Two types of approaches have been proposed recently, namely plug-in estimation (Chen et al., (2020); Zhu et al., (2023); Li et al., (2022)) and bootstraping (White and White, (2010); Hanna et al., (2017); Hao et al., (2021); Ramprasad et al., (2022); Fang et al., (2018); Samsonov et al., (2024)); nontheless, the statistical and computational efficiency of these estimations can both be further enhanced.
In this section, we introduce a computationally feasible online plug-in estimator for $\bm{\Lambda}^{\star}$, and establish its non-asymptotic rate of convergence. We will use
$\overline{\bm{A}}_{T}=\frac{1}{T}\sum_{t=1}^{T}\bm{A}_{t};$ (20)
$\widehat{\bm{\Gamma}}_{T}=\frac{1}{T}\sum_{t=1}^{T}(\bm{A}_{t}\overline{\bm{\theta}}_{T}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\theta}}_{T}-\bm{b}_{t})^{\top},$ (21)
as estimators for $\bm{A}$ and $\bm{\Gamma}$ respectively, and set
$\widehat{\bm{\Lambda}}_{T}=\overline{\bm{A}}_{T}^{-1}\widehat{\bm{\Gamma}}_{T}\overline{\bm{A}}_{T}^{-\top}$ (22)
as the final estimator of $\bm{\Lambda}^{\star}$.
Crucially, because of the linear nature of the problem, the estimator $\widehat{\bm{\Lambda}}_{T}$ can be evaluated in an online manner. Specifically, the basic properties of Kronecker products indicate
$\textbf{vec}\left(\bm{A}_{t}\overline{\bm{\theta}}_{T}\overline{\bm{\theta}}_{T}^{\top}\bm{A}_{t}\right)=(\bm{A}_{t}\otimes\bm{A}_{t})(\overline{\bm{\theta}}_{T}\otimes\overline{\bm{\theta}}_{T});$
$\textbf{vec}\left(\bm{A}_{t}\overline{\bm{\theta}}_{T}\overline{\bm{b}}_{t}^{\top}\right)=(\bm{A}_{t}\otimes\bm{b}_{t})\overline{\bm{\theta}}_{T};$
$\textbf{vec}\left(\bm{b}_{t}\bm{\overline{\theta}}_{T}^{\top}\bm{A}_{t}^{\top}\right)=(\bm{b}_{t}\otimes\bm{A}_{t})\overline{\bm{\theta}}_{T};$
$\textbf{vec}\left(\bm{b}_{t}\bm{b}_{t}^{\top}\right)=\bm{b}_{t}\otimes\bm{b}_{t}.$
Therefore, the vectorization of $\widehat{\bm{\Gamma}}_{T}$ can be represented as
$\textbf{vec}(\widehat{\bm{\Gamma}}_{T})$ $=\frac{1}{T}\sum_{t=1}^{T}\textbf{vec}\left(\bm{A}_{t}\overline{\bm{\theta}}_{T}\overline{\bm{\theta}}_{T}^{\top}\bm{A}_{t}-\bm{A}_{t}\overline{\bm{\theta}}_{T}\overline{\bm{b}}_{t}^{\top}-\bm{b}_{t}\bm{\overline{\theta}}_{T}^{\top}\bm{A}_{t}^{\top}+\bm{b}_{t}\bm{b}_{t}^{\top}\right)$
$=\frac{1}{T}\sum_{t=1}^{T}\left((\bm{A}_{t}\otimes\bm{A}_{t})(\overline{\bm{\theta}}_{T}\otimes\overline{\bm{\theta}}_{T})-(\bm{A}_{t}\otimes\bm{b}_{t})\overline{\bm{\theta}}_{T}-(\bm{b}_{t}\otimes\bm{A}_{t})\overline{\bm{\theta}}_{T}+\bm{b}_{t}\otimes\bm{b}_{t}\right)$
=($=\underset{I_{1}}{\underbrace{\left(\frac{1}{T}\sum_{t=1}^{T}\bm{A}_{t}\otimes\bm{A}_{t}\right)}}(\overline{\bm{\theta}}_{T}\otimes\overline{\bm{\theta}}_{T})-\underset{I_{2}}{\underbrace{\left(\frac{1}{T}\sum_{t=1}^{T}[\bm{A}_{t}\otimes\bm{b}_{t}+\bm{b}_{t}\otimes\bm{A}_{t}]\right)}}\overline{\bm{\theta}}_{T}$
$+\underset{I_{3}}{\underbrace{\frac{1}{T}\sum_{t=1}^{T}\bm{b}_{t}\otimes\bm{b}_{t}}},$
where $I_{1}$, $I_{2}$ and $I_{3}$ are sample means and can hence be calculated using online updates. Based on this, we propose Algorithm 1 for the calculation $\widehat{\bm{\Lambda}}_{T}$ in an online manner.
Since Algorithm 1 iterates from $t=0$ to $t=T-1$ in a single loop and does not store the samples for each iteration, it has a space complexity is $O(Td^{4})$ and a sapce complexity of $O(d^{4})$. Here, the $d^{4}$ term stems from calculating the Kronecker product $\bm{A}\otimes\bm{A}$. Not only is the estimator $\widehat{\bm{\Lambda}}_{T}$ computationally feasible, but it also enjoys good statistical properties. This is made precise by the following Gaussian comparison result, whose proof is in Appendix B.5, showing that the unknown asymptotic variance $\bm{\Lambda}^{\star}$ in the Berry-Esseen bound of Theorem 3.3 can be replaced by $\widehat{\bm{\Lambda}}_{T}$ without compromising the accuracy of the approximation.
Theorem 3.4.
Consider the TD algorithm (9) with independent samples and decaying stepsizes $\eta_{t}=\eta_{0}t^{-\alpha}$ for all $t$ and any $\alpha\in(1/2,1)$, and let $\widehat{\bm{\Lambda}}_{T}$, $\bm{\Lambda}^{\star}$ be defined as in (22), (15) respectively. Then,
$d_{\mathsf{TV}}(\mathcal{N}(\bm{0},\widehat{\bm{\Lambda}}_{T}),\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\lesssim\frac{\sqrt{\mathsf{Tr}(\bm{\Gamma})}}{\lambda_{0}(1-\gamma)}(\|\bm{\theta}^{\star}\|_{2}+1)\|\bm{\Gamma}^{-1}\|T^{-\frac{1}{3}}+o(T^{-\frac{1}{3}}).$
with probability at least $1-3T^{-\frac{1}{3}}-\frac{\sqrt{\mathsf{Tr}(\bm{\Gamma})}}{\lambda_{0}(1-\gamma)\sqrt{T}}$.
Above, the term $o(T^{-\frac{1}{3}})$ depends on all the parameters of the problem (and, in particular, to $\alpha$) but, as a function of $T$ alone, is of smaller order than $T^{-\frac{1}{3}}$. See Appendix B.5 for the proof and details.
Comparison with Chen et al., (2020).
In different settings, Chen et al., (2020) considered the problem of carrying out statistical inference for the true model parameters using stochastic gradient descent and assuming a strongly convex and smooth population loss function. In their work, a plug-in estimator of the form $\widehat{\bm{\Lambda}}_{T}^{\prime}=\overline{\bm{A}}^{-1}\widehat{\bm{\Gamma}}_{T}^{\prime}\overline{\bm{A}}^{\top}$, where
$\widehat{\bm{\Gamma}}_{T}^{\prime}:=\frac{1}{T}\sum_{t=1}^{T}(\bm{A}_{t}{\bm{\theta}}_{t}-\bm{b}_{t})(\bm{A}_{t}{\bm{\theta}}_{t}-\bm{b}_{t})^{\top},$ (23)
is proposed, which differs slightly from ours. The authors proved that the estimation error of $\widehat{\bm{\Lambda}}_{T}^{\prime}$ is bounded in expectatin by
$\mathbb{E}\|\widehat{\bm{\Lambda}}_{T}^{\prime}-\bm{\Lambda}^{\star}\|\leq\|\bm{\Gamma}\|\cdot(\widetilde{C}_{1}T^{-\frac{\alpha}{2}}+\widetilde{C}_{2}T^{-\alpha}),$ (24)
for problem-related constants $\widetilde{C}_{1}$ and $\widetilde{C}_{2}$. It is possible to modify the proof of Theorem 3.4 to obtain a guarantee in expectation for our problem as well, albeit in the TV metric. Specifically, we can show that
$\mathbb{E}\left[d_{\mathsf{TV}}(\mathcal{N}(\bm{0},\widehat{\bm{\Lambda}}_{T}),\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\right]$ $\lesssim\widetilde{C}\|\bm{\Gamma}\|T^{-\frac{1}{2}}$ (25)
where $\widetilde{C}$ depends on $\lambda_{0},\gamma,\eta_{0},\alpha,\bm{\theta}^{\star}$ and $d$. Comparing (24) with (25), we have tightened the bound from $O\left(T^{-\frac{\alpha}{2}}\right)$ to $O\left(T^{-\frac{1}{2}}\right)$, independent of $\alpha\in(1/2,1)$. This is a significant improvement, at least in our setting, not only because of the different metric but because a root-$T$ rate based on (24) would require $\alpha=1$, which is not permissible in our problem. It is noteworthy that the online algorithm 1 takes advantage of the linearity of the problem; in general, potentially non-linear settings, replacing $\bm{\theta}_{t}$ in (23) by $\overline{\bm{\theta}}_{T}$ in (21) would require storing all the samples throughout the iterative trajectory, resulting in a larger space complexity.
Comparison with Samsonov et al., (2024).
Using a multiplier-bootstrap technique, Samsonov et al., (2024) constructs a random estimator based on the samples $\mathcal{S}_{T}:=\{s_{t},s_{t}^{\prime}\}_{t=1}^{T}$, denoted as $\overline{\bm{\theta}}_{T}^{(b)}$.Theorem 4 and 5 in Samsonov et al., (2024)) imply that
$d_{\mathsf{C}}(\sqrt{T}(\overline{\bm{\theta}}_{T}^{(b)}-\bm{\theta}^{\star})|\mathcal{S}_{T},\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\lesssim\frac{\|\bm{\Gamma}^{-1}\|}{[\lambda_{0}(1-\gamma)]^{\frac{3}{2}}}(2\|\bm{\theta}^{\star}\|_{2}+1)^{2}\frac{\log T}{T^{\frac{1}{4}}}$ (26)
Compared with the bootstrap-based procedure of Samsonov et al., (2024), our plug-in estimator $\widehat{\bm{\Lambda}}_{T}$ differ in three fundamental ways. (i) The upper bound in Theorem 3.4 vanishes, as a function of $T$, at the rate of $O(T^{-\frac{1}{3}})$, which is faster than the $O(T^{-\frac{1}{4}}\log T)$ convergence rate obtained from the bound (26). (ii) The leading term in the upper bound in Theorem 3.4 entails a smaller problem-related coefficient than that of (26). (iii) Perhaps most importantly, in order to estimate the bootstrap distribution $\sqrt{T}(\overline{\bm{\theta}}_{T}^{(b)}-\bm{\theta}^{\star})|\mathcal{S}_{T}$ with an $O(T^{-\frac{1}{4}})$ error rate, it is necessary to store the sample trajectory $\mathcal{S}_{T}$ and generate $O(T^{\frac{1}{2}})$ perturbed trajectories. Consequently, the bootstrap procedure has a time complexity of order $O(T^{\frac{3}{2}}d^{2})$ and a space complexity of order $O(T)$. When $d\ll T$, this results in a significant computational border compared to our online procedure of Algorithm 1.
3.4 Statistical inference for $\bm{\theta}^{\star}$
By combining the results from Sections 3.2 and 3.3, we obtain a general methodology for constructing convex confidence sets for $\bm{\theta}^{\star}$ with guaranteed finite sample coverage, as shown in our final result, whose proof is in Appendix B.6.
Theorem 3.5.
Consider TD with Polyak-Ruppert averaging (9) with independent samples and decaying stepsizes $\eta_{t}=\eta_{0}t^{-\alpha}$, for all $t$ and with $\alpha=2/3$.For a given significance level $\delta\in(0,1)$, it holds that
$\mathbb{P}(\bm{\theta}^{\star}\in\mathcal{C}_{\delta})$ $\geq 1-\delta-\frac{\sqrt{\mathsf{Tr}(\bm{\Gamma})}}{\lambda_{0}(1-\gamma)}(2\|\bm{\theta}^{\star}\|_{2}+1)\|\bm{\Gamma}^{-1}\|T^{-\frac{1}{3}}-\frac{\sqrt{d\mathsf{cond}(\bm{\Gamma})}}{2(1-\gamma)\lambda_{0}\eta_{0}}T^{-\frac{1}{3}}-o(T^{-\frac{1}{3}}),$
uniformly over all convex sets $\mathcal{C}_{\delta}\subset\mathbb{R}^{d}$ satisfying
$\mathbb{P}_{\bm{z}\sim\mathcal{N}(\bm{0},\bm{I}_{d})}\left(\overline{\bm{\theta}}_{T}-\frac{\widehat{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}}{\sqrt{T}}\in\mathcal{C}_{\delta}\bigg{|}\overline{\bm{\theta}}_{T},\widehat{\bm{\Lambda}}_{T}\right)\geq 1-\delta,$ (27)
where the variance matrix estimator $\widehat{\bm{\Lambda}}_{T}$ is given in (22).
It is important to highlight that the above coverage guarantee holds uniformly over any choice of the convex set $\mathcal{C}_{\delta}$ meeting condition (27). In particular, this class includes all hyper-rectangles (cartesian products of univariate intervals), returning simultaneous confidence intervals that are especially useful due to their simplicity and interpretability. The construction of $\mathcal{C}_{\delta}$ requires computing appropriate quantiles of the $\mathcal{N}(\overline{\bm{\theta}}_{T},\widehat{\bm{\Lambda}}_{T})$ distribution, conditionally on $\overline{\bm{\theta}}_{T}$ and $\widehat{\bm{\Lambda}}_{T}$, depending on the choice of convex set. For the case of hyper-rectangles, this task is reduced to finding the quantile of the maximum of Gaussian variables.
Input: MDP transition kernel $\mathcal{P}$, stationary distribution $\mu$, discount factor $\gamma$, feature matrix $\bm{\Phi}$, initial stepsize $\eta_{0}$, stepsize decaying parameter $\alpha\in(0.5,1)$, sample size $T$
1 Initialize $\bm{\theta}=\bm{0}$, $\overline{\bm{\theta}}=\bm{0}$, $\overline{\bm{A}}=\bm{0}$, $\overline{\bm{AA}}=\bm{0}$, $\overline{\bm{Ab}}=\bm{0}$, $\overline{\bm{bb}}=\bm{0}$;
2 for $t=0:(T-1)$ do
3       sample $s\sim\mu$, $s^{\prime}\sim\mathcal{P}(\cdot\mid s)$;
4       compute
$\bm{A}\leftarrow\bm{\phi}(s)[\bm{\phi}(s)-\gamma\bm{\phi}(s^{\prime})];$
$\bm{b}\leftarrow r(s)\bm{\phi}(s);$
5       perform TD update
$\bm{\theta}\leftarrow\bm{\theta}-\eta_{0}t^{-\alpha}(\bm{A}\bm{\theta}-\bm{b});$
6       perform online updates
$\overline{\bm{\theta}}\leftarrow\overline{\bm{\theta}}+\frac{1}{t}(\bm{\theta}-\overline{\bm{\theta}});$
$\overline{\bm{A}}\leftarrow\overline{\bm{A}}+\frac{1}{t}(\bm{A}-\overline{\bm{A}});$
$\overline{\bm{AA}}\leftarrow\overline{\bm{AA}}+\frac{1}{t}(\bm{A}\otimes\bm{A}-\overline{\bm{AA}});$
$\overline{\bm{Ab}}\leftarrow\overline{\bm{Ab}}+\frac{1}{t}(\bm{A}\otimes\bm{b}+\bm{b}\otimes\bm{A}-\overline{\bm{Ab}});$
$\overline{\bm{bb}}\leftarrow\overline{\bm{bb}}+\frac{1}{t}(\bm{b}\otimes\bm{b}-\overline{\bm{bb}});$
7compute
$\widehat{\bm{\Gamma}}\leftarrow\overline{\bm{AA}}(\overline{\bm{\theta}}\otimes\overline{\bm{\theta}})-2\overline{\bm{Ab}}(\overline{\bm{\theta}})+\overline{\bm{bb}};$
8 reshape
$\widehat{\bm{\Gamma}}\leftarrow\mathsf{reshape}(\widehat{\bm{\Gamma}},(d\times d))$
compute
$\widehat{\bm{\Lambda}}\leftarrow\overline{\bm{A}}^{-1}\widehat{\bm{\Gamma}}\overline{\bm{A}}^{-\top}$
Output: averaged TD estimator $\overline{\bm{\theta}}$, covariance matrix estimator $\widehat{\bm{\Lambda}}$
Algorithm 1 Online algorithm for computing $\overline{\bm{\theta}}_{T}$ and $\widehat{\bm{\Lambda}}_{T}$
4 Numerical experiments
We illustrate our theoretical results with some numerical experiments. In order to test our proposed framework for statistical inference, we consider a member of a family of MDPs that are designed to be difficult to distinguish among each other. The specifics of the MDP follow the precedent of Li et al., 2024b and we refer readers to Appendix D for the construction details. We examine the high-probability convergence of the norm of TD errors, the rate by which they converge to their asymptotic Gaussian distributions, the precision of our estimator for the covariance matrix, and the validity of our statistical inference procedure. While the full details and results of the experiments are postponed to Appendix D, we demonstrate some important findings in the simulations that verify our theoretical conclusions.
4.1 Uncertainty quantification with different choices of $\alpha$
In order to explore the role played by the parameter $\alpha$ that decides the decaying speed of the stepsizes, we run TD iterates using different choices of $\alpha$ on the MDP with feature dimension $d=3$, record the $95\%$ quantile of $\|\overline{\bm{\Delta}}_{T}\|_{2}^{2}$ and measure the difference between the empirical and asymptotic distributions of $\overline{\bm{\Delta}}_{T}$  [^3] For computational simplicity, this difference between distributions is not measured by the distance on convex sets as we used in Theorems 3.2 and 3.3. Details are showin in Appendix D. .
Figure 2(a) shows the decay of the $95\%$ quantile of the $L_{2}$ norm of $\overline{\bm{\Delta}}_{T}$; the slope of this log-log plot verifies our theoretical claim in Theorem 3.1 that $\|\overline{\bm{\Delta}}_{T}\|_{2}$ converges by the rate of $O(T^{-\frac{1}{2}})$ with high probability. Notably, the convergence rate is similar for different choices of $\alpha>\frac{1}{2}$ when $T$ is large. Figure 2 (b) reveals that the empirical distribution of TD estimation error converges to the asymptotic Gaussian distribution slowly when $\alpha$ is set to ${1}/{2}$ as suggested by Samsonov et al., (2024). Both plots appear to indicate that choosing $\alpha=\frac{1}{2}$ yields a slower converging rate both in $\ell_{2}$ norm and in distribution.
Refer to caption Refer to caption
Figure 2: (a) $95\%$ quantile of $\|\overline{\bm{\Delta}}_{T}\|_{2}^{2}$, and (b) distance between empirical and asymptotic distributions of the TD estimation error $\overline{\bm{\Delta}}_{T}$, for an MDP with $d=3$ with different choices of $\alpha$.
4.2 Confidence intervals
Furthermore, we validate our statistical inference framework on the MDP with $d=3$ and TD learning with $\alpha=\frac{2}{3}$. We run 1000 independent trials of Algorithm 1, and at each iteration $T$, we construct both the individual and the simultaneous confidence intervals. Specifically, the individual confidence intervals for the three entries of $\bm{\theta}^{\star}$ are constructed by
$\mathcal{C}_{j}=\left[\overline{\bm{\theta}}_{T}-\Phi(1-\delta/2)\cdot\sqrt{\frac{(\widehat{\bm{\Lambda}}_{T})_{jj}}{T}},\overline{\bm{\theta}}_{T}+\Phi(1-\delta/2)\cdot\sqrt{\frac{(\widehat{\bm{\Lambda}}_{T})_{jj}}{T}}\right],\quad\text{for}\quad j=1,2,3.$
Here, $\Phi(1-\delta/2)$ is the $1-\frac{\delta}{2}$ quantile of the standard Gaussian distribution. In our experiment, we take $\delta=0.05$ and therefore $\Phi(1-\delta/2)=0.96$. Figure 3 illustrates the emergence of $\overline{\bm{\theta}}_{T}$ and $\{\mathcal{C}_{j}\}_{j=1,2,3}$ in one single trial. As $T$ grows larger, the three entries of the point estimate $\overline{\bm{\theta}}_{T}$ (blue curves) converges to the corresponding entries of the optimal coefficients $\bm{\theta}^{\star}$ (black dotted lines), while the individual confidence intervals $\{\mathcal{C}_{j}\}_{j=1,2,3}$ (blue shades) also concentrate around $\bm{\theta}^{\star}$.
Figure 3: The emergence of the point estimate $\overline{\bm{\theta}}_{T}$ (blue curves) and individual confidence intervals (blue shades) on the three entries of the optimal coefficients $\bm{\theta}^{\star}$ (black dotted lines) for the MDP with $d=3$, using TD with Polyak-Ruppert averaging, $\alpha=\frac{2}{3}$
Meanwhile, the simultaneous confidence intervals are constructed as
$\mathcal{C}=\{\bm{\theta}:\|\bm{\theta}-\overline{\bm{\theta}}_{T}\|_{\infty}\leq x^{\star}\},\quad\text{where}$
$x^{\star}:=\sup\left\{x:\mathbb{P}_{z\sim\mathcal{N}(\bm{0},\bm{I}_{3})}\left(\left\|\widehat{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}\right\|_{\infty}\leq x\right)\leq 1-\delta\right\}.$
Essentially, $x^{\star}$ is the $1-\delta$ quantile of the $\ell_{\infty}$ norm of the $\mathcal{N}(\bm{0},\widehat{\bm{\Lambda}}_{T})$ random variable. Calculating this quantity is a non-trivial task, so we use $1000$ independent simulations from $\mathcal{N}(\bm{0},\widehat{\bm{\Lambda}}_{T})$ and take the $1-\delta$ quantile of the $\ell_{\infty}$ norm of the samples as an estimate.
Figure 4 shows the coverage rates of the independent and simultaneous confidence intervals, defined as the percentage of times when the optimal coefficient $\bm{\theta}^{\star}$ (or its entries) are covered by the confidence intervals in the $1000$ independent trials. As is indicated by the plot, the coverage rates all approach the target confidence level of $1-\delta=95\%$ when $T$ grows large.
Figure 4: The emergence coverage rates of independent and simultaneous confidence intervals on the three entries of the optimal coefficients $\bm{\theta}^{\star}$ of the MDP with $d=3$, using TD with Polyak-Ruppert averaging, $\alpha=\frac{2}{3}$. Black dotted line represents the target confidence level of $95\%$; blue, orange and green curves represent the coverage rates of the individual confidence intervals $\mathcal{C}_{1}$, $\mathcal{C}_{2}$ and $\mathcal{C}_{3}$ correspondingly; the red curve represents the coverage rate of the confidence region (simultaneous confidence intervals) $\mathcal{C}$.
5 Discussion
We develop high-probability convergence gaurantees and Berry-Esseen bounds for TD learning with linear function approximation and $i.i.d.$ samples in a non-asymptotic manner. We also propose an efficient way to construct a covariance estimator, thus paving the way for statistical inference on policy evaluation. One may naturally be interested in whether and how our techniques can be generalized to the more realistic scenario of Markov samples, where the TD estimation error is also gauranteed to be asymptotically normal (Fort, (2015)). On this front, the answer is multifaceted. For the high-probability convergence guarantee, our induction-based technique used in the proof of Theorem 3.1 may be generalized to Markovian samples under appropriate assumptions on the mixing time of the Markov chain. As for the CLT, a cornerstone result invoked in the proof of Theorem 3.2 holds only with $i.i.d.$ samples, and a new framework is necessary for generalizing the Berry-Esseen bound to Markovian samples. Furthermore, it would be of interest to derive specialized Berry-Esseen bounds for the sub-class of hyper-rectangles. In this setting, we expect to find a better dependence on the dimension $d$ of the feature space, possibly of poly-logarithmic, as opposed to polynmial, order. Finally, the plug-in estimator $\widehat{\bm{\Lambda}}_{T}$ may be generalized to a batch-mean estimator. Its efficiency, however, needs to be further scrutinized, as the variance estimation problem for samples generated from a Markov chain remains still largely unsolved.
Acknowledgements
W. Wu and A.Rinaldo are supported in part by NIH under Grant R01 NS121913. Gen Li is supported in part by the Chinese University of Hong Kong Direct Grant for Research. Y. Wei is supported in part by the NSF grants DMS-2147546/2015447, CAREER award DMS-2143215, CCF-2106778, CCF-2418156 and the Google Research Scholar Award.
Appendix A Preliminary facts
In this section, we summarize some preliminary facts that would be useful in the proof of the theoretical results in this paper.
A.1 Useful algebriac facts
The following lemma captures some basic properties in linear algebra regarding the relaitionship between the Frobenius and operator norms of matrices.
Lemma A.1.
The following properties hold true:
  1. 1.
    For any sequence of matrices $\bm{X}_{1},\bm{X}_{2},\ldots,\bm{X}_{n}\in\mathbb{R}^{d}$:
    $\left\|\prod_{i=1}^{n}\bm{X}_{i}\right\|_{\mathsf{F}}\leq\|\bm{X}_{j}\|_{\mathsf{F}}\prod_{\begin{subarray}{c}1\leq i\leq n\\ i\neq j\end{subarray}}\|\bm{X}_{i}\|.$ (28)
  2. 2.
    For any matrix $\bm{X}\in\mathbb{R}^{d}$,
    $\|\bm{X}\|_{\mathsf{F}}\leq\sqrt{d}\|\bm{X}\|.$ (29)
The following lemma, introduced as Lemma 4.1 in Chen et al., (2020), bounds the difference between inverse matrices and would be useful for our analysis of the plug-in variance estimator.
Lemma A.2.
Let $\bm{X},\bm{Y}$ be two invertible matrices satifying $\|\bm{X}^{-1}(\bm{Y}-\bm{X})\|\leq\frac{1}{2}$. Then the difference between $\bm{Y}^{-1}$ and $\bm{X}^{-1}$ can be bounded by
$\left\|\bm{Y}^{-1}-\bm{X}^{-1}\right\|\leq 2\|\bm{Y}-\bm{X}\|\|\bm{X}^{-2}\|.$
The following lemma reveals several critical algebriac features of the matrix $\bm{A}$.
Lemma A.3.
Let $\bm{A}_{t},\bm{A}$ be defined as in (9b) and (6a) respectively, and $\bm{\Sigma}$, $\lambda_{0}$ and $\lambda_{\Sigma}$ be defined as in (7) and (8). Then the following features hold:
$2(1-\gamma)\bm{\Sigma}\preceq\bm{A}+\bm{A}^{\top}\preceq 2(1+\gamma)\bm{\Sigma};$ (30)
$\min_{1\leq i\leq d}\mathsf{Re}(\lambda_{i}(\bm{A}))\geq(1-\gamma)\lambda_{0};$ (31)
$\mathbb{E}[\bm{A}_{t}^{\top}\bm{A}_{t}]\leq\bm{A}+\bm{A}^{\top};$ (32)
$\left\|\bm{I}-\eta\bm{A}\right\|\leq 1-\frac{1-\gamma}{2}\lambda_{0}\eta,\quad\forall\eta\in\left(0,\frac{1+\gamma}{2\lambda_{\Sigma}}\right).$ (33)
Proof: We address these properties in order.
Proof of Equation (30).
By definition, $\bm{A}+\bm{A}^{\top}$ is featured by
$\bm{A}+\bm{A}^{\top}$ $=\mathbb{E}_{s\sim\mu,s^{\prime}\sim P(\cdot\mid s)}[\bm{\phi}(s)(\bm{\phi}(s)-\gamma\bm{\phi}(s^{\prime}))^{\top}+(\bm{\phi}(s)-\gamma\bm{\phi}(s^{\prime}))\bm{\phi}(s)^{\top}]$
$=2\mathbb{E}_{s\sim\mu}[\bm{\phi}(s)\bm{\phi}(s)^{\top}]-\gamma\mathbb{E}_{s\sim\mu,s^{\prime}\sim P(\cdot\mid s)}[\bm{\phi}(s)\bm{\phi}(s^{\prime})^{\top}+\bm{\phi}(s^{\prime})\bm{\phi}(s)^{\top}].$
By Cauchy-Schwartz inequality, the second term is bounded by
$-\left(\bm{\phi}(s)\bm{\phi}(s)^{\top}+\bm{\phi}(s^{\prime})\bm{\phi}(s^{\prime})^{\top}\right)$
$\preceq\bm{\phi}(s)\bm{\phi}(s^{\prime})^{\top}+\bm{\phi}(s^{\prime})\bm{\phi}(s)^{\top}$
$\preceq\bm{\phi}(s)\bm{\phi}(s)^{\top}+\bm{\phi}(s^{\prime})\bm{\phi}(s^{\prime})^{\top};$
Hence, $\bm{A}+\bm{A}^{\top}$ is bounded above by
$\bm{A}+\bm{A}^{\top}$ $\preceq 2\mathbb{E}_{s\sim\mu}[\bm{\phi}(s)\bm{\phi}(s)^{\top}]+\gamma\mathbb{E}_{s\sim\mu,s^{\prime}\sim P(\cdot\mid s)}\left[\bm{\phi}(s)\bm{\phi}(s)^{\top}+\bm{\phi}(s^{\prime})\bm{\phi}(s^{\prime})^{\top}\right]$
$=2\mathbb{E}_{s\sim\mu}[\bm{\phi}(s)\bm{\phi}(s)^{\top}]+\gamma\left(\mathbb{E}_{s\sim\mu}[\bm{\phi}(s)\bm{\phi}(s)^{\top}]+\mathbb{E}_{s^{\prime}\sim\mu}[\bm{\phi}(s)\bm{\phi}(s)^{\top}]\right)$
$=(2+2\gamma)\mathbb{E}_{s\sim\mu}[\bm{\phi}(s)\bm{\phi}(s)^{\top}]=2(1+\gamma)\bm{\Sigma},$
and similarly below by
$\bm{A}+\bm{A}^{\top}\succeq 2(1-\gamma)\bm{\Sigma}.$
Notice we applied the fact that when $s\sim\mu$ and $s^{\prime}\sim P(\cdot\mid s)$, the marginal distribution of $s$ is also $\mu$.
Proof of Equation (31).
This result is a direct corollary of (31). Specifically, let $z=\lambda+i\mu\in\mathbb{C}$ denote any eigenvalue of $\bm{A}$, where $\lambda,\mu\in\mathbb{R}$; and $\bm{x}+i\bm{y}$ denote the corresponding eigenvector in $\mathbb{C}^{d}$, where $\bm{x},\bm{y}\in\mathbb{R}^{d}$ satisfy $\|\bm{x}\|_{2}=\|\bm{y}\|_{2}=1$. It then follows that
$\bm{A}(\bm{x}+i\bm{y})=(\lambda+i\mu)(\bm{x}+i\bm{y})=(\lambda\bm{x}-\mu\bm{y})+i(\mu\bm{x}+\lambda\bm{y}).$
Comparing the real and imaginary parts of the left-most and right-most terms, we obtain
$\bm{A}\bm{x}=\lambda\bm{x}-\mu\bm{y},\quad\text{and}\quad\bm{A}\bm{y}=\mu\bm{x}+\lambda\bm{y}.$
Direct calculation yields
$\bm{x}^{\top}\bm{A}\bm{x}=\lambda\|\bm{x}\|_{2}^{2}-\mu\bm{x}^{\top}\bm{y}=\lambda-\mu\bm{x}^{\top}\bm{y},\quad\text{and}$
$\bm{y}^{\top}\bm{A}\bm{y}=\mu\bm{y}^{\top}\bm{x}+\lambda\|\bm{y}\|_{2}^{2}=\mu\bm{y}^{\top}\bm{x}+\lambda.$
Hence, Equation (30) implies
$\lambda$ $=\frac{1}{2}(\bm{x}^{\top}\bm{A}\bm{x}+\bm{y}^{\top}\bm{A}\bm{y})$
$=\frac{1}{4}\bm{x}^{\top}(\bm{A}+\bm{A}^{\top})\bm{x}+\frac{1}{4}\bm{y}^{\top}(\bm{A}+\bm{A}^{\top})\bm{y}$
$\geq\frac{1}{4}\bm{x}^{\top}(1-\gamma)\bm{\Sigma}\bm{x}+\frac{1}{4}\bm{y}^{\top}(1-\gamma)\bm{\Sigma}\bm{y}$
$\geq\frac{1}{4}(2(1-\gamma)\lambda_{0}+2(1-\gamma)\lambda_{0})=(1-\gamma)\lambda_{0}.$
This completes the proof of (31).
Proof of Equation (32).
The left-hand-side of (32) can be bounded by
$\mathbb{E}[\bm{A}_{t}^{\top}\bm{A}_{t}]$ $=\mathbb{E}[(\bm{\phi}(s_{t})-\gamma\bm{\phi}(s_{t}^{\prime}))\bm{\phi}(s_{t})^{\top}\bm{\phi}(s_{t})(\bm{\phi}(s_{t})-\gamma\bm{\phi}(s_{t}^{\prime}))^{\top}]$
$=\mathbb{E}\|\bm{\phi}(s_{t})\|_{2}^{2}[(\bm{\phi}(s_{t})-\gamma\bm{\phi}(s_{t}^{\prime}))(\bm{\phi}(s_{t})-\gamma\bm{\phi}(s_{t}^{\prime}))^{\top}]$
$\leq\max_{s\in\mathcal{S}}\|\bm{\phi}(s)\|_{2}^{2}\mathbb{E}[(\bm{\phi}(s_{t})-\gamma\bm{\phi}(s_{t}^{\prime}))(\bm{\phi}(s_{t})-\gamma\bm{\phi}(s_{t}^{\prime}))^{\top}]$
$\leq\mathbb{E}[\bm{\phi}(s_{t})\bm{\phi}(s_{t})^{\top}]-\gamma\mathbb{E}[\bm{\phi}(s_{t})\bm{\phi}(s_{t}^{\prime})^{\top}+\bm{\phi}(s_{t}^{\prime})\bm{\phi}(s_{t})^{\top}]+\gamma^{2}\mathbb{E}[\bm{\phi}(s_{t}^{\prime})\bm{\phi}(s_{t}^{\prime})^{\top}]$
$\leq 2\mathbb{E}[\bm{\phi}(s_{t})\bm{\phi}(s_{t})^{\top}]-\gamma\mathbb{E}[\bm{\phi}(s_{t})\bm{\phi}(s_{t}^{\prime})^{\top}+\bm{\phi}(s_{t}^{\prime})\bm{\phi}(s_{t})^{\top}]$
$=\bm{A}+\bm{A}^{\top},$
Notice that the inequality on the fourth line follows from the assumption that $\max_{s\in\mathcal{S}}\|\bm{\phi}(s)\|_{2}\leq 1$ and the inequality on the fifth line follows from the fact that since $s_{t}$ and $s_{t}^{\prime}$ and both drawn from the stationary distribution, $\mathbb{E}[\phi(s_{t}^{\prime})\phi(s_{t}^{\prime})^{\top}]=\mathbb{E}[\phi(s_{t})\phi(s_{t})^{\top}]$.
Proof of Equation (33).
We firstly notice that
$(\bm{I}-\eta\bm{A})^{\top}(\bm{I}-\eta\bm{A})=\bm{I}-\eta(\bm{A}+\bm{A}^{\top})+\eta^{2}\bm{A}^{\top}\bm{A},$
where $\bm{A}^{\top}\bm{A}$ can be bounded above by
$\bm{A}^{\top}\bm{A}$ $\preceq\bm{A}^{\top}\bm{A}+\bm{A}\bm{A}^{\top}\preceq\frac{1}{2}(\bm{A}+\bm{A}^{\top})^{2}\preceq(1+\gamma)\lambda_{\Sigma}(\bm{A}+\bm{A}^{\top}),$
where we applied Equation (30) in the last inequality. Hence,
$(\bm{I}-\eta\bm{A})^{\top}(\bm{I}-\eta\bm{A})\preceq\bm{I}-(\eta-(1+\gamma)\lambda_{\Sigma}\eta^{2})(\bm{A}+\bm{A}^{\top}).$
For any $\eta\in(0,\frac{1+\gamma}{2\lambda_{\Sigma}})$, this upper bound is controlled by
$(\bm{I}-\eta\bm{A})^{\top}(\bm{I}-\eta\bm{A})\preceq\bm{I}-\frac{\eta}{2}(\bm{A}+\bm{A}^{\top})\preceq(1-(1-\gamma)\lambda_{0})\bm{I}.$
Hence, the spectral norm of $\bm{I}-\eta\bm{A}$ is controlled by
$\|\bm{I}-\eta\bm{A}\|\leq\sqrt{1-(1-\gamma)\lambda_{0}}\leq 1-\frac{1-\gamma}{2}\lambda_{0}\eta.$
The following lemma further features the compressive power of the matrix $\prod_{k=1}^{t}(\bm{I}-\eta_{k}\bm{A})$.
Lemma A.4.
Assume that $\beta\in(0,1)$ and $\alpha\in(\frac{1}{2},1)$. Let $\eta_{t}=\eta_{0}t^{-\alpha}$ for all positive integer $t$, then the following inequalities hold:
$t^{\alpha}\prod_{k=1}^{t}\left(1-\beta k^{-\alpha}\right)\leq e\left(\frac{\alpha}{\beta}\right)^{\frac{\alpha}{1-\alpha}};$ (34)
$\sum_{i=1}^{t}i^{-\nu}\prod_{k=i+1}^{t}\left(1-\beta k^{-\alpha}\right)\leq\frac{1}{\nu-1}\left(\frac{2(\nu-\alpha)}{\beta}\right)^{\frac{\nu-\alpha}{1-\alpha}}t^{\alpha-\nu},\quad\forall\nu\in(1,\alpha+1];$ (35)
$\sum_{i=1}^{t}i^{-2\alpha}\prod_{k=i+1}^{t}\left(1-\beta k^{-\alpha}\right)=\frac{1}{\beta}t^{-\alpha}+5\left(\frac{2}{\beta}\right)^{\frac{2-\alpha}{1-\alpha}}t^{-1};$ (36)
$\max_{1\leq i\leq t}i^{-\alpha}\prod_{k=i+1}^{t}\left(1-\beta k^{-\alpha}\right)\leq e\left(\frac{\alpha}{\beta}\right)^{\frac{\alpha}{1-\alpha}}t^{-\alpha}.$ (37)
Proof: We prove these four inequalities in order.
Proof of Equation (34).
We firstly observe that
$t^{\alpha}\prod_{k=1}^{t}\left(1-\beta k^{-\alpha}\right)<t^{\alpha}\exp\left(-\beta\sum_{k=1}^{t}k^{-\alpha}\right),$
in which
$\sum_{k=1}^{t}k^{-\alpha}>\int_{1}^{t}x^{-\alpha}\mathrm{d}x=\frac{t^{1-\alpha}-1}{1-\alpha}.$
Hence, the left hand side of (34) is upper bounded by
$t^{\alpha}\prod_{k=1}^{t}\left(1-\beta k^{-\alpha}\right)$ $<t^{\alpha}\exp\left(-\beta\cdot\frac{t^{1-\alpha}-1}{1-\alpha}\right)$
$=e^{\frac{\beta}{1-\alpha}}\cdot t^{\alpha}\exp\left(-\frac{\beta}{1-\alpha}t^{1-\alpha}\right).$
By letting $u=t^{1-\alpha}$, we obtain
$t^{\alpha}\prod_{k=1}^{t}\left(1-\beta k^{-\alpha}\right)$ $<e^{\frac{\beta}{1-\alpha}}\cdot u^{\frac{\alpha}{1-\alpha}}\exp\left(-\frac{\beta}{1-\alpha}u\right),$
where the right hand side is a function of $u$. It is easy to verify that the maximum of this function is reached when $u=\alpha/\beta$, and hence
$t^{\alpha}\prod_{k=1}^{t}\left(1-\beta k^{-\alpha}\right)$ $<e^{\frac{\beta}{1-\alpha}}\cdot\left(\frac{\alpha}{\beta}\right)^{\frac{\alpha}{1-\alpha}}\exp\left(-\frac{\alpha}{1-\alpha}\right).$
The upper bound (34) follows naturally.
Proof of Equation (35).
In order to derive an iterative relationship with respect to the left-hand-side of (35), we denote
$X_{t}=\sum_{i=1}^{t}i^{-\nu}\prod_{k=i+1}^{t}\left(1-\beta k^{-\alpha}\right)$
for any positive integer $t$. Then by definition, $X_{1}=1$ and the sequence $\{X_{t}\}_{t\geq 1}$ is featured by the iterative relation
$X_{t+1}=\left(1-\beta(t+1)^{-\alpha}\right)X_{t}+(t+1)^{-\nu}.$ (38)
For simplicity, we denote
$C=\frac{1}{\nu-1}\left(\frac{2(\nu-\alpha)}{\beta}\right)^{\frac{\nu-\alpha}{1-\alpha}}.$
then the lemma is equivalent to the claim that
$t^{\nu-\alpha}X_{t}\leq C$ (39)
holds for all positive integer $t$, and can hence be proved by an induction argument. On one hand, we observe that the iterative relation (38) implies
$X_{t+1}<X_{t}+(t+1)^{-\nu},$
hence $X_{t}$ can be bounded by
$X_{t}<\sum_{i=1}^{t}i^{-\nu}<\int_{1}^{\infty}x^{-\nu}\mathrm{d}x=\frac{1}{\nu-1}.$
Therefore, we can take
$t^{\star}=\left(\frac{2(\nu-\alpha)}{\beta}\right)^{\frac{1}{1-\alpha}},$
then for any $t\leq t^{\star}$, it can be guaranteed that
$t^{\nu-\alpha}X_{t}<(t^{\star})^{\nu-\alpha}\cdot\frac{1}{\nu-1}=C.$
On the other hand, assume that (39) holds for an integer $t=k\geq t^{\star}$, then for $t=k+1$, we observe
$t^{\nu-\alpha}X_{t}$
$=(k+1)^{\nu-\alpha}\left[(1-\beta(k+1)^{-\alpha})X_{k}+(k+1)^{-\alpha}\right]$
$\leq(k+1)^{\nu-\alpha}\cdot Ck^{\alpha-\nu}-\beta Ck^{\alpha-\nu}(k+1)^{\nu-2\alpha}+(k+1)^{-\alpha}$
$=C+C\left[\left(\frac{k+1}{k}\right)^{\nu-\alpha}-1\right]-\beta C(k+1)^{-\alpha}\left(\frac{k+1}{k}\right)^{\nu-\alpha}+(k+1)^{-\alpha}$
$=C+C\left(\frac{k+1}{k}\right)^{\nu-\alpha}\left[1-\left(\frac{k}{k+1}\right)^{\nu-\alpha}\right]$
$-\beta C(k+1)^{-\alpha}\left(\frac{k+1}{k}\right)^{\nu-\alpha}+(k+1)^{-\alpha}$
<$\overset{(i)}{<}C+C\left(\frac{k+1}{k}\right)^{\nu-\alpha}\cdot\frac{\nu-\alpha}{k+1}-\beta C(k+1)^{-\alpha}\left(\frac{k+1}{k}\right)^{\nu-\alpha}+(k+1)^{-\alpha}$
<$\overset{(ii)}{<}C-\frac{\beta}{2}C(k+1)^{-\alpha}\left(\frac{k+1}{k}\right)^{\nu-\alpha}+(k+1)^{-\alpha}$
<$\overset{(iii)}{<}C-(k+1)^{-\alpha}\left(\frac{k+1}{k}\right)^{\nu-\alpha}+(k+1)^{-\alpha}<C.$
Throughout the process, we applied the following facts:
$(i):\left(\frac{k}{k+1}\right)^{\nu-\alpha}=\left(1-\frac{1}{k+1}\right)^{\nu-\alpha}\geq 1-\frac{\nu-\alpha}{k+1},\quad\text{since }\nu-\alpha\in(1-\alpha,1];$
$(ii):\frac{\nu-\alpha}{k+1}<\frac{\beta}{2}(k+1)^{-\alpha},\quad\text{since }k>t^{\star};\quad\text{and}$
$(iii):\beta C=\frac{\beta}{\nu-1}\left(\frac{2(\nu-\alpha)}{\beta}\right)^{\frac{\nu-\alpha}{1-\alpha}}>\frac{\beta}{\nu-1}\frac{2(\nu-\alpha)}{\beta}>2,\quad\text{since }\nu>1>\alpha.$
An induction argument then validates (39) for every positive integer $t$.
Proof of Equation (36).
In order to derive an iterative relationship with respect to the left-hand-side of (36), we denote
$X_{t}=\sum_{i=1}^{t}i^{-2\alpha}\prod_{k=i+1}^{t}\left(1-\beta k^{-\alpha}\right)$
for any positive integer $t$. Then by definition, $X_{1}=1$ and the sequence $\{X_{t}\}_{t\geq 1}$ is featured by the iterative relation
$X_{t}=\left(1-\beta t^{-\alpha}\right)X_{t-1}+t^{-2\alpha},\quad\text{for every }t>1.$ (40)
Direct calculation yields
$t^{\alpha}X_{t}$ $=\left(\frac{t}{t-1}\right)^{\alpha}\left(1-\beta t^{-\alpha}\right)[(t-1)^{\alpha}X_{t-1}]+t^{-\alpha}$
$=\left(\frac{t}{t-1}\right)^{\alpha}\left(1-\beta t^{-\alpha}\right)\cdot\frac{1}{\beta}+\left(\frac{t}{t-1}\right)^{\alpha}\left(1-\beta t^{-\alpha}\right)\left[(t-1)^{\alpha}X_{t-1}-\frac{1}{\beta}\right]+t^{-\alpha}$
$=\frac{1}{\beta}+\left[\left(\frac{t}{t-1}\right)^{\alpha}-1\right]\left(\frac{1}{\beta}+t^{-\alpha}\right)$
$+\left(\frac{t}{t-1}\right)^{\alpha}\left(1-\beta t^{-\alpha}\right)\left[(t-1)^{\alpha}X_{t-1}-\frac{1}{\beta}\right].$
Setting $Z_{t}=t^{\alpha}X_{t}-\frac{1}{\beta}$, we immediately obtain
$Z_{t}=\left(\frac{t}{t-1}\right)^{\alpha}\left(1-\beta t^{-\alpha}\right)Z_{t-1}+\left[\left(\frac{t}{t-1}\right)^{\alpha}-1\right]\left(\frac{1}{\beta}+t^{-\alpha}\right).$
Notice that for every $t>1$ and $\alpha<1$,
$\left(\frac{t}{t-1}\right)^{\alpha}-1<\frac{t}{t-1}-1=\frac{1}{t-1}<\frac{2}{t}.$
Hence, the sequence $\{Z_{t}\}_{t\geq 1}$ is upper bounded iteratively by
$Z_{t}<\left(\frac{t}{t-1}\right)^{\alpha}\left(1-\beta t^{-\alpha}\right)Z_{t-1}+\frac{2}{t}\left(\frac{1}{\beta}+t^{-\alpha}\right).$
By induction, this directly implies
$Z_{t}<\underset{I_{1}}{\underbrace{t^{\alpha}\prod_{k=1}^{t}\left(1-\beta k^{-\alpha}\right)Z_{1}}}+\underset{I_{2}}{\underbrace{2t^{\alpha}\sum_{i=1}^{t}\left(\frac{1}{\beta}i^{-\alpha-1}+i^{-2\alpha-1}\right)\prod_{k=i+1}^{t}(1-\beta k^{-\alpha})}},$
where $Z_{1}=1^{\alpha}X_{1}-\frac{1}{\beta}=1-\frac{1}{\beta}$. The proof now boils down to showing that both $I_{1}$ and $I_{2}$ converges to 00 as $t\to\infty$. On one hand, $I_{1}$ is bounded by
$t^{\alpha}\exp\left(-\beta\sum_{k=1}^{t}k^{-\alpha}\right)<\left(\frac{1}{\beta}\right)^{\frac{1}{1-\alpha}}t^{\alpha-1}.$
On the other hand, the convergence of $I_{2}$ is guaranteed by (35):
$I_{2}$ $\leq\frac{4t^{\alpha}}{\beta}\sum_{i=1}^{t}i^{-\alpha-1}\prod_{k=i+1}^{t}(1-\beta k^{-\alpha})$
$\leq\frac{2}{\alpha}\left(\frac{2}{\beta}\right)^{\frac{2-\alpha}{1-\alpha}}t^{\alpha-1}.$
Combining these results, we obtain
$t^{\alpha}X_{t}-\frac{1}{\beta}=Z_{t}<5\left(\frac{2}{\beta}\right)^{\frac{2-\alpha}{1-\alpha}}t^{\alpha-1},$
which directly implies (36).
Proof of Equation (37).
For every positive interger $t$, we begin by identifying the index $i$ that maximizes
$i^{-\alpha}\prod_{k=i+1}^{t}\left(1-\beta k^{-\alpha}\right):=X_{i}^{t}.$
Consider the ratio of two consecutive terms:
$\frac{X_{i+1}}{X_{i}}$ $=\frac{(i+1)^{-\alpha}}{i^{-\alpha}(1-\beta(i+1)^{-\alpha})}$
$=\frac{i^{\alpha}}{(i+1)^{\alpha}-\beta},\quad\text{for any }1\leq i<t.$
Equivalently, we observe
$\frac{X_{i}}{X_{i+1}}=1+i^{-\alpha}\left[(i+1)^{\alpha}-i^{\alpha}-\beta\right].$
Since $(i+1)^{\alpha}-i^{\alpha}\searrow 0$ for any $\alpha\in(\frac{1}{2},1)$, we can conclude that when
$2^{\alpha}-1-\beta<0,$
the sequence ${X_{i}^{t}}_{1\leq i\leq t}$ is monotonically increasing; otherwise, there exists a constant $i^{\star}$ such that $\{X_{i}^{t}\}_{1\leq i\leq t}$ decreases when $i\leq i^{\star}$ and increases when $i>i^{\star}$. Either way, the maximum over $i$ is reached at either $i=1$ or $i=t$. That is,
$\max_{1\leq i\leq t}X_{i}^{t}=\max\left\{\prod_{k=2}^{t}\left(1-\beta k^{-\alpha}\right),t^{-\alpha}\right\}$
where the first term on the right hand side is bounded above by (34). The equation follows naturally. ∎
Lemma A.5.
For every $\beta\in(0,1)$, $\alpha\in(\frac{1}{2},1)$ and sufficiently large $T$, it can be guaranteed that
$t^{-\alpha}\sum_{j=t}^{T}\prod_{k=t+1}^{j}(1-\beta k^{-\alpha})<3\left(\frac{2}{\beta}\right)^{\frac{1}{1-\alpha}},\quad\text{and}$ (41)
$t^{-\alpha}\sum_{j=t}^{\infty}\prod_{k=t+1}^{j}(1-\beta k^{-\alpha})-\frac{1}{\beta}\lesssim\left(\frac{1}{\beta}\right)^{\frac{1}{1-\alpha}}\Gamma\left(\frac{1}{1-\alpha}\right)t^{\alpha-1}.$ (42)
Proof: We address the two inequalities in order.
Proof of (41).
For simplicity, we denote
$t^{\star}=\left(\frac{2\alpha}{\beta}\right)^{\frac{1}{1-\alpha}},$
$X_{t}=\sum_{j=t}^{T}\prod_{k=t+1}^{j}\left(1-\beta k^{-\alpha}\right),\quad\text{and}$
$Y_{t}=t^{-\alpha}X_{t}.$
By definition, we observe that $X_{T}=1$, $Y_{T}=T^{-\alpha}$ and that the sequences $\{X_{t}\}_{1\leq t\leq T}$ and $\{Y_{t}\}_{1\leq t\leq T}$ is featured by the iterative relation
$X_{t}=1+\left[1-\beta(t+1)^{-\alpha}\right]X_{t+1}$ (43)
$Y_{t}=t^{-\alpha}+\left(\frac{t+1}{t}\right)^{\alpha}\left[1-\beta(t+1)^{-\alpha}\right]Y_{t+1}.$ (44)
In what follows, we consider the bound on $Y_{t}$ for both $t\geq t^{\star}$ and $1\leq t<t^{\star}$.
Bounding $Y_{t}$ for $t\geq t^{\star}$.
Without loss of generality, we assume that $T>t^{\star}$; the bound on $Y_{t}$ for $t\geq t^{\star}$ is shown by an inverse induction from $t=T$ to $t=t^{\star}$. Namely, we observe that $Y_{T}=T^{-\alpha}<1<\frac{2}{\beta}$; if $Y_{t+1}<\frac{2}{\beta}$ holds for an integer $t\in(t^{\star},T)$, then the iterative relation (44) implies
$Y_{t}$ $<t^{-\alpha}+\left(\frac{t+1}{t}\right)^{\alpha}\left[1-\beta(t+1)^{-\alpha}\right]\cdot\frac{2}{\beta}$
$<\frac{2}{\beta}+t^{-\alpha}\left\{\left[(t+1)^{\alpha}-t^{\alpha}\right]\frac{2}{\beta}-1\right\}.$
Hence, it boils down to showing
$(t+1)^{\alpha}-t^{\alpha}\leq\frac{\beta}{2},$
for which we observe
$(t+1)^{\alpha}-t^{\alpha}<\alpha t^{\alpha-1}\leq\alpha\left(t^{\star}\right)^{\alpha-1}=\frac{2}{\beta}.$
In this way, we prove that $Y_{t}\leq\frac{2}{\beta}$ for every $t\geq t^{\star}$.
Bounding $Y_{t}$ for $1\leq t<t^{\star}$.
For every positive integer $t$ satisfying $1\leq t<t^{\star}$, the iterative relation (43)
$Y_{t}=t^{-\alpha}X_{t}<X_{t}<X_{t^{\star}}+(t^{\star}-t)$ $<X_{t^{\star}}+t^{\star}$
$=(t^{\star})^{\alpha}Y_{t^{\star}}+t^{\star}$
$<\left(\frac{2\alpha}{\beta}\right)^{\frac{\alpha}{1-\alpha}}\frac{2}{\beta}+\left(\frac{2\alpha}{\beta}\right)^{\frac{1}{1-\alpha}}$
$=\left(\frac{2\alpha}{\beta}\right)^{\frac{1}{1-\alpha}}(\frac{1}{\alpha}+1).$
Proof of (42).
We firstly expand the left-hand-side of (42) as
$t^{-\alpha}\sum_{j=t}^{\infty}\prod_{k=t+1}^{j}(1-\beta k^{-\alpha})$
$=\sum_{j=t}^{\infty}(j+1)^{-\alpha}\prod_{k=t+1}^{j}(1-\beta k^{-\alpha})+\sum_{j=t}^{\infty}(t^{-\alpha}-(j+1)^{-\alpha})\prod_{k=t+1}^{j}(1-\beta k^{-\alpha}).$ (45)
Here, the first term is bounded by applying the telescoping technique
$\sum_{j=t}^{\infty}(j+1)^{-\alpha}\prod_{k=t+1}^{j}(1-\beta k^{-\alpha})$ $=\frac{1}{\beta}\sum_{j=t}^{\infty}\left(\prod_{k=t+1}^{j}-\prod_{k=t+1}^{j+1}\right)(1-\beta k^{-\alpha})$
$=\frac{1}{\beta}\left(1-\prod_{k=t+1}^{\infty}(1-\beta k^{-\alpha})\right)=\frac{1}{\beta}.$
Therefore, the proof boils down to showing that the second term is bounded by
$\sum_{j=t}^{\infty}((j+1)^{-\alpha}-t^{-\alpha})\prod_{k=t+1}^{j}(1-\beta k^{-\alpha})=O(t^{\alpha-1}).$ (46)
Towards this end, we denote
$m_{t}^{j}=\sum_{k=t}^{j}k^{-\alpha}.$ (47)
In terms of $m_{t}^{j}$, the left-hand-side of (46) is bounded by
$\sum_{j=t}^{\infty}((j+1)^{-\alpha}-t^{-\alpha})\prod_{k=t+1}^{j}(1-\beta k^{-\alpha})$
$\asymp\sum_{j=t}^{\infty}((j+1)^{-\alpha}-t^{-\alpha})\exp\left(-\beta m_{t}^{j}\right).$
Furthermore, we observe that on one hand, $m_{t}^{j}$ is bounded by
$m_{t}^{j}=\sum_{k=t}^{j}k^{-\alpha}>\int_{t}^{j+1}x^{-\alpha}\mathrm{d}x=\frac{(j+1)^{1-\alpha}-t^{1-\alpha}}{1-\alpha};$ (48)
on the other hand, the term $(j+1)^{-\alpha}-t^{-\alpha}$ can be bounded by
$t^{-\alpha}-(j+1)^{-\alpha}$ $=(j+1)^{-\alpha}\left[\left(\frac{j+1}{t}\right)^{\alpha}-1\right]$
$=(m_{t}^{j+1}-m_{t}^{j})\cdot\left[\left(\frac{j+1}{t}\right)^{\alpha}-1\right].$
In what follows, we aim to bound the last term in terms of $m_{t}^{j}$. Specifically, since
$\left(\frac{j+1}{t}\right)^{1-\alpha}-1<\frac{(1-\alpha)m_{t}^{j}}{t^{1-\alpha}},$
we obtain
$\left(\frac{j+1}{t}\right)^{\alpha}-1$ $<\left[\frac{(1-\alpha)m_{t}^{j}}{t^{1-\alpha}}+1\right]^{\frac{\alpha}{1-\alpha}}-1$
$<t^{\alpha-1}\left((1-\alpha)m_{t}^{j}+1\right)^{\frac{\alpha}{1-\alpha}}.$
Therefore, the left-hand-side of (46) is bounded by
$\sum_{j=t}^{\infty}((j+1)^{-\alpha}-t^{-\alpha})\prod_{k=t+1}^{j}(1-\beta k^{-\alpha})$
$<\sum_{j=t}^{\infty}(m_{t}^{j+1}-m_{t}^{j})t^{\alpha-1}\left((1-\alpha)m_{t}^{j}+1\right)^{\frac{\alpha}{1-\alpha}}\exp\left(\beta t^{-\alpha}-\beta m_{t}^{j}\right)$
$<t^{\alpha-1}\sum_{j=t}^{\infty}\left((1-\alpha)m_{t}^{j}+1\right)^{\frac{\alpha}{1-\alpha}}\exp\left(-\beta m_{t}^{j}\right)(m_{t}^{j+1}-m_{t}^{j})$
$\asymp t^{\alpha-1}\int_{0}^{\infty}\left((1-\alpha)m+1\right)^{\frac{\alpha}{1-\alpha}}\exp(-\beta m)\mathrm{d}m=O(t^{\alpha-1});$
the integral on the last line is further bounded by
$\int_{0}^{\infty}\left((1-\alpha)m+1\right)^{\frac{\alpha}{1-\alpha}}\exp(-\beta m)\mathrm{d}m$
$<\int_{0}^{\infty}\left(m^{\frac{\alpha}{1-\alpha}}+2^{\frac{\alpha}{1-\alpha}}\right)\exp(-\beta m)\mathrm{d}m$
$=2^{\frac{\alpha}{1-\alpha}}\frac{1}{\beta}+\left(\frac{1}{\beta}\right)^{\frac{1}{1-\alpha}}\Gamma\left(\frac{1}{1-\alpha}\right)\lesssim\left(\frac{1}{\beta}\right)^{\frac{1}{1-\alpha}}\Gamma\left(\frac{1}{1-\alpha}\right).$
The bound (42) follows naturally. ∎
Lemma A.6.
Let $\eta_{t}=\eta_{0}t^{-\alpha}$ with $\alpha\in(\frac{1}{2},1)$ and $0\leq\eta_{0}\leq\frac{1+\gamma}{2\lambda_{\Sigma}}$. Recall that $\bm{Q}_{t}$ is defined as in (19). Then for any positive integer $\ell<T$, the following relation hold:
$\sum_{t=1}^{T}(\bm{Q}_{t}-\bm{A}^{-1})=-\bm{A}^{-1}\sum_{j=1}^{T}\prod_{k=1}^{j}(\bm{I}-\eta_{k}\bm{A}),$ (49a)
Furthermore, there exist a sequence of matrix $\{\bm{S}_{t}\}_{1\leq t\leq T}$, along with a constant $\widetilde{C}$, depending on $\alpha,\eta_{0}$ and $\gamma$, such that the difference between $\bm{Q}_{t}$ and $\bm{A}^{-1}$ can be represented as
$\bm{Q}_{t}-\bm{A}^{-1}=\bm{S}_{t}-\bm{A}^{-1}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A}),$ (49b)
in which $\bm{S}_{t}$ is independent of $T$, and its norm is bounded by
$\|\bm{S}_{t}\|\lesssim\eta_{0}\Gamma\left(\frac{1}{1-\alpha}\right)\left(\frac{2}{(1-\gamma)\lambda_{0}\eta_{0}}\right)^{\frac{1}{1-\alpha}}t^{\alpha-1}.$ (49c)
Proof: We address these equations and bounds separately.
Proof of (49a)
By definition, the sum of $Q_{t}$ for $1\leq t\leq T$ can be expressed as
$\sum_{t=1}^{T}\bm{Q}_{t}=\sum_{t=1}^{T}\eta_{t}\sum_{j=t}^{T}\prod_{k=t+1}^{j}(\bm{I}-\eta_{k}\bm{A}).$
By switching the order of summation we obtain
$\sum_{t=1}^{T}\bm{Q}_{t}$ $=\sum_{j=1}^{T}\sum_{t=1}^{j}\eta_{t}\prod_{k=t+1}^{j}(\bm{I}-\eta_{k}\bm{A})$
$=\sum_{j=1}^{T}\sum_{t=1}^{j}\bm{A}^{-1}\left(\prod_{k=t+1}^{j}-\prod_{k=t}^{j}\right)(\bm{I}-\eta_{k}\bm{A})$
$=\bm{A}^{-1}\sum_{j=1}^{T}\left[\bm{I}-\prod_{k=1}^{j}(\bm{I}-\eta_{k}\bm{A})\right],$
which directly implies (49a).
Construction of the decomposition (49b).
The definition of $\bm{Q}_{t}$ (19) directly implies that on one hand,
$\bm{Q}_{t}$ $=\eta_{t}\sum_{j=t}^{T}\prod_{k=t+1}^{j}(\bm{I}-\eta_{k}\bm{A})=\eta_{t}\bm{I}+\eta_{t}\sum_{j=t+1}^{T}\prod_{k=t+1}^{j}(\bm{I}-\eta_{k}\bm{A}),$
and that
$(\bm{I}-\eta_{t}\bm{A})\bm{Q}_{t}=\eta_{t}\sum_{j=t}^{T}\prod_{k=t}^{j}(\bm{I}-\eta_{k}\bm{A})$ $=\eta_{t}\sum_{j=t}^{T-1}\prod_{k=t}^{j}(\bm{I}-\eta_{k}\bm{A})+\eta_{t}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A})$
$=\eta_{t}\sum_{j=t+1}^{T-1}\prod_{k=t}^{j-1}(\bm{I}-\eta_{k}\bm{A})+\eta_{t}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A})$
On the other hand,
$\prod_{k=t+1}^{j}(\bm{I}-\eta_{k}\bm{A})-\prod_{k=t}^{j-1}(\bm{I}-\eta_{k}\bm{A})$
$=\left[\prod_{k=t+1}^{j-1}(\bm{I}-\eta_{k}\bm{A})\right](\bm{I}-\eta_{j}\bm{A})-(\bm{I}-\eta_{t}\bm{A})\left[\prod_{k=t+1}^{j-1}(\bm{I}-\eta_{k}\bm{A})\right]$
$=\prod_{k=t+1}^{j-1}(\bm{I}-\eta_{k}\bm{A})-\eta_{j}\left[\prod_{k=t+1}^{j-1}(\bm{I}-\eta_{k}\bm{A})\right]\bm{A}-\prod_{k=t+1}^{j-1}(\bm{I}-\eta_{k}\bm{A})+\eta_{t}\bm{A}\prod_{k=t+1}^{j-1}(\bm{I}-\eta_{k}\bm{A})$
$=\eta_{t}\bm{A}\prod_{k=t+1}^{j-1}(\bm{I}-\eta_{k}\bm{A})-\eta_{j}\prod_{k=t+1}^{j-1}(\bm{I}-\eta_{k}\bm{A})\bm{A}$
$=(\eta_{t}-\eta_{j})\bm{A}\prod_{k=t+1}^{j-1}(I-\eta_{k}\bm{A}).$
Therefore, by taking a difference between $\bm{Q}_{t}$ and $(\bm{I}-\eta_{t}\bm{A})\bm{Q}_{t}$, we obtain
$\eta_{t}\bm{A}\bm{Q}_{t}$ $=\bm{Q}_{t}-(\bm{I}-\eta_{t}\bm{A})\bm{Q}_{t}$
$=\eta_{t}\bm{I}+\eta_{t}\sum_{j=t+1}^{T-1}\left[\prod_{k=t+1}^{j}(\bm{I}-\eta_{k}\bm{A})-\prod_{k=t}^{j-1}(\bm{I}-\eta_{k}\bm{A})\right]-\eta_{t}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A})$
$=\eta_{t}\bm{I}+\eta_{t}\sum_{j=t+1}^{T-1}(\eta_{t}-\eta_{j})\bm{A}\prod_{k=t+1}^{j-1}(\bm{I}-\eta_{k}\bm{A})-\eta_{t}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A}).$
Consequently, the difference between $\bm{Q}_{t}$ and $\bm{A}^{-1}$ can be expressed as
$\bm{Q}_{t}-\bm{A}^{-1}=\underset{\bm{S}_{t}}{\underbrace{\sum_{j=t+1}^{T-1}(\eta_{t}-\eta_{j})\prod_{k=t+1}^{j-1}(\bm{I}-\eta_{k}\bm{A})}}-\bm{A}^{-1}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A}),$ (50)
which takes the form of (49b).
Proof of (49c).
Lemma A.3 implies that
$\left\|\prod_{k=t+1}^{j-1}(\bm{I}-\eta_{k}\bm{A})\right\|$ $\leq\prod_{k=t+1}^{j-1}\left(1-\frac{1}{2}\eta_{k}(1-\gamma)\lambda_{0}\right)$
$\leq\prod_{k=t+1}^{j}\exp\left(-\frac{1}{2}\eta_{k}(1-\gamma)\lambda_{0}\right)$
$=\exp\left(-\frac{1}{2}(1-\gamma)\lambda_{0}\sum_{k=t+1}^{j-1}\eta_{k}\right).$
Recalling the definition of $m_{t}^{j}$ (47),(48), we can bound the norm $\bm{S}_{t}$ by
$\|\bm{S}_{t}\|_{2}$ $\leq\sum_{j=t}^{\infty}(\eta_{t}-\eta_{j})\left\|\prod_{k=t+1}^{j-1}(\bm{I}-\eta_{k}\bm{A})\right\|_{2}$
$\leq\sum_{j=t}^{\infty}(\eta_{t}-\eta_{j})\exp\left(-\frac{1}{2}(1-\gamma)\lambda_{0}\sum_{k=t+1}^{j-1}\eta_{k}\right)$
$=\sum_{j=t}^{\infty}\eta_{j}\left[\left(\frac{j}{t}\right)^{\alpha}-1\right]\exp\left(-\frac{1}{2}(1-\gamma)\lambda_{0}\eta_{0}m_{t}^{j}\right).$
The right-hand-side can then be bounded by a similar argument to that of Lemma A.5. ∎
The following Lemma features the compressive rate of the matrix $\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A})$.
Lemma A.7.
Under the condition 12, for every positive definite matrix $\bm{Y}\in\mathbb{R}^{d\times d}$, it can be guarnateed that
$T^{-\alpha}\sum_{t=1}^{T}\left(\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A})\bm{Y}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})\right)=\bm{X}+\bm{O}(T^{\alpha-1}),$ (51)
where $\bm{X}$ is the solution to the Lyapunov equation
$\eta_{0}(\bm{A}\bm{X}+\bm{XA}^{\top})=\bm{Y}.$ (52)
Proof: For simplicity, denote the left-hand-side of (51) as $\bm{X}_{T}$. It can be easily verified that the sequence $\{\bm{X}_{T}\}_{T\geq 1}$ is featured by the iterative relation
$\bm{X}_{T}$ $=\left(\frac{T-1}{T}\right)^{\alpha}(\bm{I}-\eta_{T}\bm{A})\bm{X}_{T-1}(\bm{I}-\eta_{T}\bm{A}^{\top})$
$+T^{-\alpha}(\bm{I}-\eta_{T}\bm{A})\bm{Y}(\bm{I}-\eta_{T}\bm{A}^{\top}).$
Direct calculation yields
$\bm{X}_{T}-\bm{X}$ $=\left(\frac{T-1}{T}\right)^{\alpha}(\bm{I}-\eta_{T}\bm{A})(\bm{X}_{T-1}-\bm{X})(\bm{I}-\eta_{T}\bm{A}^{\top})$
$+\left(\frac{T-1}{T}\right)^{\alpha}(\bm{I}-\eta_{T}\bm{A})\bm{X}(\bm{I}-\eta_{T}\bm{A}^{\top})-\bm{X}$
$+T^{-\alpha}(\bm{I}-\eta_{T}\bm{A})\bm{Y}(\bm{I}-\eta_{T}\bm{A}^{\top})$
$=\left(\frac{T-1}{T}\right)^{\alpha}(\bm{I}-\eta_{T}\bm{A})(\bm{X}_{T-1}-\bm{X})(\bm{I}-\eta_{T}\bm{A}^{\top})$
$+T^{-\alpha}[\eta_{0}(\bm{AX}+\bm{XA}^{\top})-\bm{Y}]+\frac{\bm{X}}{T}+O(T^{-2\alpha}).$
Notice that the first term on the last line is $\bm{0}$ according to the definition of $\bm{X}$; therefore by iteration, the difference between $\bm{X}_{T}$ and $\bm{X}$ can be represented by
$\bm{X}_{T}-\bm{X}$ $=T^{-\alpha}\sum_{i=1}^{T}i^{\alpha-1}\prod_{k=i+1}^{T}(\bm{I}-\eta_{k}\bm{A})\bm{X}\prod_{k=i+1}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})$
$-T^{-\alpha}\prod_{k=1}^{T}(\bm{I}-\eta_{k}\bm{A})\bm{X}\prod_{k=1}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})+O(T^{-\alpha});$
By letting $\beta=\frac{1-\gamma}{2}\lambda_{0}\eta_{0}$, LemmaA.3 inplies the following bound on the norm of the first term on the right-hand-side:
$\left\|T^{-\alpha}\sum_{i=1}^{T}i^{\alpha-1}\prod_{k=i+1}^{T}(\bm{I}-\eta_{k}\bm{A})\bm{X}\prod_{k=i+1}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})\right\|$
$<T^{\alpha-1}\sum_{i=1}^{T}i^{-\alpha}\prod_{k=1}^{T}(1-\beta k^{-\alpha})\|\bm{X}\|$
$=T^{\alpha-1}\cdot\frac{1}{\beta}\sum_{i=1}^{T}\left(\prod_{k=i}^{T}-\prod_{k=i+1}^{T}(1-\beta k^{-\alpha})\right)\|\bm{X}\|<\frac{\|\bm{X}\|}{\beta}T^{\alpha-1}.$
The two terms on the second line are clearly both $\bm{o}(T^{\alpha-1})$; hence,
$\|\bm{X}_{T}-\bm{X}\|=\bm{O}(T^{\alpha-1}).$
Lemma A.8.
Let $\eta_{t}=\eta_{0}t^{-\alpha}$ with $\alpha\in(\frac{1}{2},1)$ and $0\leq\eta\leq\frac{1+\gamma}{2\lambda_{\Sigma}}$. For $\bm{Q}_{t}$ defined as in (19), it can be guaranteed that
$\|\bm{Q}_{t}\|\leq 3\eta_{0}^{-\frac{\alpha}{1-\alpha}}\left(\frac{4\alpha}{\lambda_{0}(1-\gamma)}\right)^{\frac{1}{1-\alpha}},\quad\forall t\in[T];$ (53)
$\|\bm{A}\bm{Q}_{t}\|\leq 2+\eta_{0}\left(\frac{1}{\beta}\right)^{\frac{1}{1-\alpha}}\Gamma\left(\frac{1}{1-\alpha}\right)t^{\alpha-1};\quad\text{and}$ (54)
$\frac{1}{T}\sum_{t=1}^{T}\|\bm{Q}_{t}-\bm{A}^{-1}\|^{2}\leq\frac{1}{\lambda_{0}(1-\gamma)}T^{\alpha-1}+\widetilde{C}T^{2\alpha-2}$ (55)
where $\beta=\frac{1-\gamma}{2}\lambda_{0}\eta_{0}$ and $\widetilde{C}$ depends on $\alpha,\eta_{0},\lambda_{0},\gamma$.
Proof: According to Lemma A.3, since $0<\eta_{t}\leq\eta_{0}\leq\frac{1+\gamma}{2\lambda_{\Sigma}}$ for all $t$, the norm of $\bm{Q}_{t}$ is bounded by
$\|\bm{Q}_{t}\|_{2}$ $\leq\eta_{t}\sum_{j=t}^{T}\prod_{k=t+1}^{j}\left(1-\frac{1}{2}\lambda_{0}(1-\gamma)\eta_{k}\right)$
$=\eta_{0}t^{-\alpha}\cdot\sum_{j=t}^{T}\prod_{k=t+1}^{j}\left(1-\frac{1}{2}\lambda_{0}(1-\gamma)\eta_{0}k^{-\alpha}\right).$
The bound (53) follows immediately by taking $\beta=\frac{1-\gamma}{2}\lambda_{0}\eta_{0}$ in (41). For the bound (54), observe due to (49b) that
$\bm{AQ}_{t}-\bm{I}=\bm{AS}_{t}-\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A}).$
Therefore, the triangle inequality and (49c) directly imply
$\|\bm{AQ}_{t}\|\leq\|\bm{I}\|+\|\bm{AS}_{t}\|+\left\|\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A})\right\|$ $\leq 1+\eta_{0}\left(\frac{1}{\beta}\right)^{\frac{1}{1-\alpha}}\Gamma\left(\frac{1}{1-\alpha}\right)t^{\alpha-1}+1$
$=2+\eta_{0}\left(\frac{1}{\beta}\right)^{\frac{1}{1-\alpha}}\Gamma\left(\frac{1}{1-\alpha}\right)t^{\alpha-1}.$
Furthermore, it can be easily verified that
$\|\bm{AQ}_{t}-\bm{I}\|^{2}\leq\|\bm{AS}_{t}\|^{2}+\left\|\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A})\right\|^{2}$ $\leq\widetilde{C}^{\prime}t^{2\alpha-2}+\prod_{k=t}^{T}(1-\beta k^{-\alpha})$
$<\widetilde{C}^{\prime}t^{2\alpha-2}+T^{\alpha}(t-1)^{-\alpha}\prod_{k=t}^{T}(1-\beta k^{-\alpha})$
$=\widetilde{C}^{\prime}t^{2\alpha-2}+T^{\alpha}\frac{1}{\beta}\left(\prod_{k=t}^{T}-\prod_{k=t-1}^{T}\right)(1-\beta k^{-\alpha});$
The bound (55) follows by taking an average from $t=1$ to $t=T$ and applying the telescoping method. ∎
The decomposition (49b), combined with the bounds (49c) Theorem A.7, implies the following bound on the difference between $\overline{\bm{\Lambda}}_{T}$ and $\bm{\Lambda}^{\star}$.
Theorem A.9.
Let $\overline{\bm{\Lambda}}_{T}$ and $\bm{\Lambda}^{\star}$ be defined as in (18) and (15) respectively. Define $\bm{X}(\bm{\Lambda}^{\star})$ as the unique solution to the Lyapunov equation
$\eta_{0}(\bm{AX+XA}^{\top})=\bm{\Lambda}^{\star};$ (56)
then as $T\to\infty$, the difference between $\overline{\bm{\Lambda}_{T}}$ and $\bm{\Lambda}^{\star}$ is featured by
$T^{\alpha-1}\bm{X}(\bm{\Lambda}^{\star})-\bm{O}(T^{2\alpha-2})\preceq\overline{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star}\preceq T^{\alpha-1}\bm{X}(\bm{\Lambda}^{\star})+\bm{O}(T^{2\alpha-2})$
Proof: By definition, the difference bewteen $\overline{\bm{\Lambda}}_{T}$ and $\bm{\Lambda}^{\star}$ can be expressed as
$\overline{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star}$ $=\frac{1}{T}\sum_{t=1}^{T}(\bm{Q}_{t}\bm{\Gamma}\bm{Q}_{t}^{\top}-\bm{A}^{-1}\bm{\Gamma}\bm{A}^{-\top})$
$=\underset{I_{1}}{\underbrace{\frac{1}{T}\sum_{t=1}^{T}(\bm{Q}_{t}-\bm{A}^{-1})\bm{\Gamma}\bm{A}^{-\top}+\frac{1}{T}\sum_{t=1}^{T}\bm{A}^{-1}\bm{\Gamma}(\bm{Q}_{t}-\bm{A}^{-1})^{\top}}}$
$+\underset{I_{2}}{\underbrace{\frac{1}{T}\sum_{t=1}^{T}(\bm{Q}_{t}-\bm{A}^{-1})\bm{\Gamma}(\bm{Q}_{t}-\bm{A}^{-1})^{\top}}}.$ (57)
We now proceed to bound the norms of $I_{1}$ and $I_{2}$ respectively.
Bounding the norm of $I_{1}$.
Notice that the first term can be expressed by
$\frac{1}{T}\sum_{t=1}^{T}(\bm{Q}_{t}-\bm{A}^{-1})\bm{\Gamma}\bm{A}^{-\top}$ $=-\frac{1}{T}\bm{A}^{-1}\sum_{j=1}^{T}\prod_{k=1}^{j}(\bm{I}-\eta_{k}\bm{A})\cdot\bm{\Gamma}\bm{A}^{-\top}$
$=-\frac{1}{T}\sum_{j=1}^{T}\prod_{k=1}^{j}(\bm{I}-\eta_{k}\bm{A})\cdot\bm{A}^{-1}\bm{\Gamma}\bm{A}^{-\top}$
$=-\frac{1}{T}\sum_{j=1}^{T}\prod_{k=1}^{j}(\bm{I}-\eta_{k}\bm{A})\cdot\bm{\Lambda}^{\star}.$ (58)
where we applied Equation (49a) and the interchangability of the production. Similarly, the second term can be expressed by
$\bm{A}^{-1}\bm{\Gamma}\frac{1}{T}\sum_{t=1}^{T}(\bm{Q}_{t}-\bm{A}^{-1})^{\top}=-\bm{\Lambda}^{\star}\cdot\frac{1}{T}\sum_{j=1}^{T}\prod_{k=1}^{j}(\bm{I}-\eta_{k}\bm{A})^{\top}.$
Lemma A.3 directly implies that the norm of these two terms are bounded by
$\left\|\frac{1}{T}\sum_{t=1}^{T}(\bm{Q}_{t}-\bm{A}^{-1})\bm{\Gamma}\bm{A}^{-\top}\right\|$ $\leq\frac{1}{T}\|\bm{\Lambda}^{\star}\|\sum_{j=1}^{T}\prod_{k=1}^{j}\left(1-\frac{1-\gamma}{2}\lambda_{0}\eta_{0}k^{-\alpha}\right).$
For simplicity, we denote $\beta=\frac{1-\gamma}{2}\lambda_{0}\eta_{0}$; it can then be seen that the summand on the right-hand-side can be bounded by
$\sum_{j=1}^{T}\prod_{k=1}^{j}(1-\beta k^{-\alpha})$ $<\sum_{j=1}^{T}\exp\left(\beta\sum_{k=1}^{j}k^{-\alpha}\right)$
$=\sum_{j=1}^{T}\exp\left(-\beta m_{1}^{j}\right)$ (59)
$\asymp\sum_{j=1}^{T}\exp(-\beta m_{1}^{j})(m_{1}^{j+1}-m_{1}^{j})((1-\alpha)m_{1}^{j})^{\frac{\alpha}{1-\alpha}}$
$\lesssim(1-\alpha)^{\frac{\alpha}{1-\alpha}}\int_{0}^{\infty}m^{\frac{\alpha}{1-\alpha}}\exp(-\beta m)\mathrm{d}m$
$=\beta^{-\frac{1}{1-\alpha}}(1-\alpha)^{\frac{\alpha}{1-\alpha}}\Gamma\left(\frac{1}{1-\alpha}\right),$ (60)
a constant independent of $T$. Therefore, $I_{1}$ is bounded above by $\bm{O}(T^{-1})$.
Decomposing $I_{2}$.
The term $I_{2}$ can be further expressed by invoking the decomposition (49b). Specifically, we observe
$I_{2}$ $=\underset{I_{21}}{\underbrace{\frac{1}{T}\sum_{t=1}^{T}\bm{S}_{t}\bm{\Gamma}\bm{S}_{t}^{\top}}}+\underset{I_{22}}{\underbrace{\frac{1}{T}\sum_{t=1}^{T}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A})\bm{\Gamma}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})}}$
($-\underset{I_{23}}{\underbrace{\left(\frac{1}{T}\sum_{t=1}^{T}\bm{S}_{t}\bm{\Gamma}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})+\frac{1}{T}\sum_{t=1}^{T}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A})\bm{\Gamma}\bm{S}_{t}^{\top}\right)}}.$
Since $\|\bm{S}_{t}\|=O(T^{\alpha-1})$, it is easy to verify that $I_{21}=\bm{O}(T^{2\alpha-2})$; meanwhile, theorem A.7 indicates that $I_{22}=T^{\alpha-1}\bm{X}(\bm{\Lambda}^{\star})+O(T^{2\alpha-2})$. As for the term $I_{23}$, we observe that
$\left\|\sum_{t=1}^{T}\bm{S}_{t}\bm{\Gamma}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})\right\|$ $\lesssim\sum_{t=1}^{T}t^{\alpha-1}\prod_{k=t}^{T}(1-\beta k^{-\alpha})$
$<T^{2\alpha-1}\sum_{t=1}^{T}t^{\alpha-1}\prod_{k=t+1}^{T}(1-\beta k^{-\alpha})$
$=T^{2\alpha-1}\cdot\frac{1}{\beta}\sum_{t=1}^{T}\left(\prod_{k=t+1}^{T}-\prod_{k=t}^{T}\right)(1-\beta k^{-\alpha})$
$<\frac{T^{2\alpha-1}}{\beta}.$
Consequently, it can be guaranteed that $I_{23}=\bm{O}(T^{2\alpha-2})$. In combintation, we have shown that
$T^{\alpha-1}\bm{X}(\bm{\Lambda}^{\star})-\bm{O}(T^{2\alpha-2})\preceq\frac{1}{T}\sum_{t=1}^{T}(\bm{Q}_{t}-\bm{A}^{-1})\bm{\Gamma}(\bm{Q}_{t}-\bm{A}^{-1})^{\top}\preceq T^{\alpha-1}\bm{X}(\bm{\Lambda}^{\star})+\bm{O}(T^{2\alpha-2});$
the theorem follows naturally by combining this with the bound on $I_{1}$.∎
The following Lemma controls the minimum eigenvalue of $\bm{A}\overline{\bm{\Lambda}}_{T}\bm{A}^{\top}$.
Lemma A.10.
With $\bm{A}$ defined as (6a) and $\overline{\bm{\Lambda}}_{T}$ defined as in (18), when
$T\geq 4\left(\frac{2}{(1-\gamma)\lambda_{0}\eta_{0}}\right)^{\frac{1}{1-\alpha}}(1-\alpha)^{\frac{\alpha}{1-\alpha}}\Gamma(\frac{1}{1-\alpha})\mathsf{cond}(\Gamma),$ (61)
it can be guaranteed that
$\lambda_{\min}(\bm{A}\bm{\Lambda}_{T}\bm{A}^{\top})\geq\frac{1}{2}\lambda_{\min}(\bm{\Gamma})=\frac{1}{2\|\bm{\Gamma}^{-1}\|}.$
Proof: We firstly decompose $\bm{A}\bm{\Lambda}_{T}\bm{A}^{\top}$ as
$\bm{A}\bm{\Lambda}_{T}\bm{A}^{\top}$ $=\bm{A}\bm{\Lambda}^{\star}\bm{A}^{\top}+\bm{A}(\bm{\Lambda}_{T}-\bm{\Lambda}^{\star})\bm{A}^{\top}$
$=\bm{\Gamma}+\bm{A}(\bm{\Lambda}_{T}-\bm{\Lambda}^{\star})\bm{A}^{\top};$
here, (A.1) indicates that the difference $(\bm{\Lambda}_{T}-\bm{\Lambda}^{\star})$ is bounded above by
$\bm{\Lambda}_{T}-\bm{\Lambda}^{\star}$ $\succeq\frac{1}{T}\sum_{t=1}^{T}\left[(\bm{Q}_{t}-\bm{A}^{-1})\bm{\Gamma}\bm{A}^{-\top}+\bm{A}^{-1}\bm{\Gamma}(\bm{Q}_{t}-\bm{A}^{-1})^{\top}\right].$
Furthermore, (49a) reveals that
$\bm{A}(\bm{\Lambda}_{T}-\bm{\Lambda}^{\star})\bm{A}^{\top}$
$=\frac{1}{T}\bm{A}\sum_{t=1}^{T}(\bm{Q}_{t}-\bm{A}^{-1})\bm{\Gamma}\bm{A}^{-\top}\bm{A}^{\top}+\frac{1}{T}\bm{A}\bm{A}^{-1}\sum_{t=1}^{T}\bm{\Gamma}(\bm{Q}_{t}-\bm{A}^{-1})^{\top}\bm{A}^{\top}$
$=-\frac{1}{T}\bm{A}\bm{A}^{-1}\sum_{j=1}^{T}\prod_{k=1}^{j}(\bm{I}-\eta_{k}\bm{A})\bm{\Gamma}-\frac{1}{T}\bm{\Gamma}\sum_{j=1}^{T}\prod_{k=1}^{j}(\bm{I}-\eta_{k}\bm{A})^{\top}\bm{A}^{-\top}\bm{A}^{\top}$
$=-\frac{1}{T}\sum_{j=1}^{T}\prod_{k=1}^{j}(\bm{I}-\eta_{k}\bm{A})\bm{\Gamma}-\frac{1}{T}\bm{\Gamma}\sum_{j=1}^{T}\prod_{k=1}^{j}(\bm{I}-\eta_{k}\bm{A})^{\top}.$
It has been revealed in (A.1) that
$\left\|\sum_{j=1}^{T}\prod_{k=1}^{j}(\bm{I}-\eta_{k}\bm{A})\right\|\leq\left(\frac{2}{(1-\gamma)\lambda_{0}\eta_{0}}\right)^{\frac{1}{1-\alpha}}(1-\alpha)^{\frac{\alpha}{1-\alpha}}\Gamma(\frac{1}{1-\alpha}).$
Consequently, the triangle inequality yields
$\lambda_{\min}(\bm{A}\bm{\Lambda}_{T}\bm{A}^{\top})$ $\geq\lambda_{\min}(\bm{\Gamma})-\left\|\frac{1}{T}\sum_{j=1}^{T}\prod_{k=1}^{j}(\bm{I}-\eta_{k}\bm{A})\bm{\Gamma}+\frac{1}{T}\bm{\Gamma}\sum_{j=1}^{T}\prod_{k=1}^{j}(\bm{I}-\eta_{k}\bm{A})^{\top}\right\|$
$\geq\lambda_{\min}(\bm{\Gamma})-\frac{2}{T}\left(\frac{2}{(1-\gamma)\lambda_{0}\eta_{0}}\right)^{\frac{1}{1-\alpha}}(1-\alpha)^{\frac{\alpha}{1-\alpha}}\Gamma(\frac{1}{1-\alpha})\|\bm{\Gamma}\|.$
The lemma follows immediately by applying the condition (61). ∎
A.2 Gaussian approximation
This section is concerned with Gaussian approximation of random variables. We will firstly introduce some Berry-Esseen bounds and then illustrate some properties of the multi-dimensional Gaussian distribution.
A.2.1 Berry-Esseen bounds
The following theorem, proved by Shao and Zhang, (2021), gives a Berry-Esseen bound for non-linear statistics which would be crucial for our analysis of the non-asymptotic distribution of the TD estimation error.
Theorem A.11.
(Theorem 2.1 of Shao and Zhang, (2021)) Let $\bm{X}_{1},...,\bm{X}_{n}$ be independent random variables, and let $\bm{T}(\bm{X}_{1},...,\bm{X}_{n})$ be a $d$-dimensional statistic that can be represented as
$\bm{T}=\bm{W}+\bm{D},$
where $\bm{W}$ is the sum of independent random variables
$\bm{W}=\sum_{i=1}^{n}\bm{\xi}_{i}=\sum_{i=1}^{n}\bm{h}_{i}(\bm{X}_{i})$
for Borel measurable functions $\{\bm{h}_{i}\}_{1\leq i\leq n}$ that satisfies
$\mathbb{E}[\bm{\xi}_{i}]=0,\text{ for each }1\leq i\leq n\text{ and }\sum_{i=1}^{n}\mathbb{E}[\bm{\xi}_{i}\bm{\xi}_{i}^{\top}]=\bm{\Lambda}.$
Further assume that $\lambda_{\min}(\bm{\Lambda})\geq\sigma>0$. Let $\bm{z}\sim N(0,\bm{I}_{d})$. Then for any convex set $\mathcal{A}\subseteq\mathbb{R}^{d}$,
$d_{\mathsf{C}}(\bm{T},\bm{\Lambda}^{\frac{1}{2}}\bm{z})\leq 259\sqrt{d}\sigma^{-\frac{3}{2}}\cdot\sum_{i=1}^{n}\mathbb{E}\|\bm{\xi}_{i}\|_{2}^{3}+2\sigma^{-1}\mathbb{E}\{\|\bm{W}\|_{2}\Delta\}+2\sigma^{-1}\sum_{i=1}^{n}\mathbb{E}\{\|\bm{\xi}_{i}\|_{2}|\Delta-\Delta^{(i)}|\},$ (62)
for any random variables $\Delta$ and $\{\Delta^{(i)}\}_{1\leq i\leq n}$ such that $\Delta\geq\|D\|_{2}$ and $\Delta^{(i)}$ is independent of $\bm{X}_{i}$.
A.2.2 Properties of multivariate Gaussian distributions
The following theorem proposed by Devroye et al., (2018) gives an upper bound for the total variation (TV) distance between two Gaussian random variables with same means and different covariance matrixes.
Theorem A.12 (Devroye et al., (2018)).
Let $\bm{\Lambda}_{1},\bm{\Lambda}_{2}\in\mathbb{S}^{d\times d}$ be two positive definite matrices, and $\bm{\mu}$ be any vector in $\mathbb{R}^{d}$. Then the TV distance between Gaussian distirbutions $\mathcal{N}(\bm{\mu},\bm{\Lambda}_{1})$ and $\mathcal{N}(\bm{\mu},\bm{\Lambda}_{2})$ is bounded by
$\min\left\{1,\frac{1}{100}\left\|\bm{\Lambda}_{1}^{-1/2}\bm{\Lambda}_{2}\bm{\Lambda}_{1}^{-1/2}-\bm{I}_{d}\right\|_{\mathsf{F}}\right\}$
$\leq d_{\mathsf{TV}}(\mathcal{N}(\bm{\mu},\bm{\Lambda}_{1}),\mathcal{N}(\bm{\mu},\bm{\Lambda}_{2}))$
$\leq\frac{3}{2}\left\|\bm{\Lambda}_{1}^{-1/2}\bm{\Lambda}_{2}\bm{\Lambda}_{1}^{-1/2}-\bm{I}_{d}\right\|_{\mathsf{F}}.$
In the case of $\bm{\Lambda}_{1}\succ\bm{\Lambda}_{2}$, the following proposition further specifies that the TV distance and the distance on convex sets are the same bewteen $\mathcal{N}(\bm{0},\bm{\Lambda}_{1})$ and $\mathcal{N}(\bm{0},\bm{\Lambda}_{2})$.
Proposition A.13.
Let $\bm{\Lambda}_{1},\bm{\Lambda}_{2}\in\mathbb{S}^{d\times d}$ be two positive definite matrices, and $\bm{\mu}$ be any vector in $\mathbb{R}^{d}$. When $\bm{\Lambda}_{1}\succ\bm{\Lambda_{2}}$, it can be guaranteed that
$d_{\mathsf{C}}(\mathcal{N}(\bm{\mu},\bm{\Lambda}_{1}),\mathcal{N}(\bm{\mu},\bm{\Lambda}_{2}))=d_{\mathsf{TV}}(\mathcal{N}(\bm{\mu},\bm{\Lambda}_{1}),\mathcal{N}(\bm{\mu},\bm{\Lambda}_{2})).$
Proof: Without loss of generality, let $\bm{\mu}=\bm{0}$. Define $\mathcal{C}$ as
$\mathcal{C}:=\{\bm{x}\in\mathbb{R}^{d}\mid p_{1}(\bm{x})<p_{2}(\bm{x})\},$
where $p_{1}(\cdot)$ and $p_{2}(\cdot)$ denote the probability density functions of $\mathcal{N}(\bm{0},\bm{\Lambda}_{1})$ and $\mathcal{N}(\bm{0},\bm{\Lambda}_{2})$ respectively; we aim to show that $\mathcal{C}$ is a convex set. Notice that the ratio between these two density functions is featured by
$\frac{p_{1}(\bm{x})}{p_{2}(\bm{x})}$ $=\frac{\frac{1}{\sqrt{\mathsf{det}(\bm{\Lambda_{1}})}}\exp(-\bm{x}^{\top}\bm{\Lambda_{1}}^{-1}\bm{x})}{\frac{1}{\sqrt{\mathsf{det}(\bm{\Lambda_{1}})}}\exp(-\bm{x}^{\top}\bm{\Lambda_{2}}^{-1}\bm{x})}$
$=\sqrt{\frac{\mathsf{det}(\bm{\Lambda}_{2})}{\mathsf{det}(\bm{\Lambda_{1}})}}\exp(\bm{x}^{\top}(\bm{\Lambda_{2}}^{-1}-\bm{\Lambda_{1}}^{-1})\bm{x}).$
Therefore by definition, the set $\mathcal{C}$ is characterized by
$\mathcal{C}=\left\{\bm{x}\in\mathbb{R}^{d}\bigg{|}\bm{x}^{\top}(\bm{\Lambda_{2}}^{-1}-\bm{\Lambda_{1}}^{-1})\bm{x}<\frac{1}{2}\log\frac{\mathsf{det}(\bm{\Lambda}_{1})}{\mathsf{det}(\bm{\Lambda}_{2})}\right\}.$
Since $\bm{\Lambda_{1}}\succ\bm{\Lambda}_{2}$, the matrix $(\bm{\Lambda_{2}}^{-1}-\bm{\Lambda_{1}}^{-1})$ must be positive definite; consequently, $\mathcal{C}$ is a $d$-dimensional hyper-ellipsoid and is therefore convex. With $\mathcal{C}$ being convex, it is seen that
$\left|\mathbb{P}_{z\sim\mathcal{N}(\bm{0},\bm{\Lambda}_{1})}(\bm{z}\in\mathcal{C})-\mathbb{P}_{z\sim\mathcal{N}(\bm{0},\bm{\Lambda}_{2})}(\bm{z}\in\mathcal{C})\right|$
$\leq\sup_{\mathcal{A}\in\mathscr{C}_{d}}\left|\mathbb{P}_{z\sim\mathcal{N}(\bm{\mu},\bm{\Lambda}_{1})}(\bm{z}\in\mathcal{A})-\mathbb{P}_{z\sim\mathcal{N}(\bm{\mu},\bm{\Lambda}_{2})}(\bm{z}\in\mathcal{A})\right|$
$=d_{\mathsf{C}}(\mathcal{N}(\bm{0},\bm{\Lambda}_{1}),\mathcal{N}(\bm{0},\bm{\Lambda}_{2})),$
while in the meantime
$\left|\mathbb{P}_{z\sim\mathcal{N}(\bm{\mu},\bm{\Lambda}_{1})}(\bm{z}\in\mathcal{C})-\mathbb{P}_{z\sim\mathcal{N}(\bm{\mu},\bm{\Lambda}_{2})}(\bm{z}\in\mathcal{C})\right|$ $=d_{\mathsf{TV}}(\mathcal{N}(\bm{\mu},\bm{\Lambda}_{1}),\mathcal{N}(\bm{\mu},\bm{\Lambda}_{2}))$
$\geq d_{\mathsf{C}}(\mathcal{N}(\bm{0},\bm{\Lambda}_{1}),\mathcal{N}(\bm{0},\bm{\Lambda}_{2})).$
Combining these two inequalitites completes the proof. ∎
A.3 Concentration inequalities
The following Azuma’s inequality for matrix martingales is shown in Tropp, (2011).
Theorem A.14.
Let $\{\bm{X}_{k}\}$ be a random sequence of matrices in $\mathbb{S}^{d\times d}$, and $\{\bm{A}_{k}\}$ be a sequence of fixed matrices in $\mathbb{S}^{d\times d}$. Assume that $\mathbb{E}_{k-1}[\bm{X}_{k}]=\bm{0}$ and $\bm{X}_{k}^{2}\preceq\bm{A}_{k}^{2}$ almost surely. Define
$\sigma^{2}:=\left\|\sum_{k}\bm{A}_{k}^{2}\right\|,$
then for all $\varepsilon>0$,
$\mathbb{P}\left(\left\|\sum_{k}\bm{X}_{k}\right\|\geq\varepsilon\right)\leq d\exp\left(-\frac{\varepsilon^{2}}{8\sigma^{2}}\right).$
Specifically, for vector-valued martingales, the following proposition is shown by Kallenberg and Sztencel, (1991).
Proposition A.15.
Let $\{\bm{x}_{i}\}_{i\geq 1}$ be a martingale in $\mathbb{R}^{d}$. Then there exists a martingale $\{\bm{y}_{i}\}_{i\geq 1}$ in $\mathbb{R}^{2}$ such that for every $i>1$, $\|\bm{y}_{i}\|_{2}$ has the same distribution as $\|\bm{x}_{i}\|_{2}$ and $\|\bm{y}_{i}-\bm{y}_{i-1}\|_{2}$ has the same distribution as $\|\bm{x}_{i}-\bm{x}_{i-1}\|_{2}$.
The following Azuma’s inequality on vector-valued martingales comes as a direct corollary.
Corollary A.16.
Let $\{\bm{x}_{i}\}_{i\geq 1}$ be a martingale in $\mathbb{R}^{d}$, and let $W_{\max}$ be a positive constant that satisfies
$W_{\max}\geq\sum_{i=1}^{t}\|\bm{x}_{i}-\bm{x}_{i-1}\|_{2}^{2}\quad\text{almost surely.}$.
Then it can be guaranteed with probability at least $1-\delta$ that
$\|\bm{x}_{t}\|_{2}\leq 2\sqrt{2W_{\max}\log\frac{3}{\delta}}.$
Appendix B Proof of main results
B.1 High-probability convergence guarantee for the original TD estimation error
In order to derive the high-probability convergence guarantee for $\overline{\bm{\Delta}}_{T}$, we firstly propose the following result on the time-uniform convergence of the original TD estimator $\bm{\Delta}_{t}$, under condition 12.
Theorem B.1.
Consider the TD algorithm (9) with independent samples and decaying stepsizes as in condition 12. For any $\delta\in(0,1)$, there exists $\eta_{0}>0$, such that with probability at least $1-\delta$, the norm of $\bm{\Delta}_{t}$ is bounded by
$\|\bm{\Delta}_{t}\|_{2}$ $\lesssim\frac{\eta_{0}}{\sqrt{2\alpha-1}}\left(\frac{(1-\gamma)\lambda_{0}\eta_{0}}{4\alpha}\right)^{-\frac{\alpha}{2(1-\alpha)}}\sqrt{\log\frac{9T}{\delta}}(2\|\bm{\theta}^{\star}\|_{2}+1)t^{-\frac{\alpha}{2}}$ (63)
simultaneously for all $t\in[T]$.
Proof: Under the TD update rule (9), we represent $\bm{\Delta}_{t}$ as
$\bm{\Delta}_{t}=\bm{\theta}_{t}-\bm{\theta}^{\star}$ $=(\bm{\theta}_{t-1}-\eta_{t}(\bm{A}_{t}\bm{\theta}_{t-1}-\bm{b}_{t}))-\bm{\theta}^{\star}$
$=\bm{\Delta}_{t-1}-\eta_{t}(\bm{A}\bm{\theta}_{t-1}-\bm{b}+\bm{\zeta}_{t})$
$=(\bm{I}-\eta_{t}\bm{A})\bm{\Delta}_{t-1}-\eta_{t}\bm{\zeta}_{t},$
where $\bm{\zeta}_{t}$ is defined as
$\bm{\zeta}_{t}=(\bm{A}_{t}-\bm{A})\bm{\theta}_{t-1}-(\bm{b}_{t}-\bm{b}).$ (64)
Therefore by induction, $\bm{\Delta}_{t}$ can be expressed as a weighted sum of $\{\bm{\zeta}_{i}\}_{0\leq i<t}$, namely
$\bm{\Delta}_{t}$ $=\prod_{k=1}^{t}(\bm{I}-\eta_{k}\bm{A})\Delta_{0}-\sum_{i=1}^{t}\bm{R}_{i}^{t}\bm{\zeta}_{i}$
$=\prod_{k=1}^{t}(\bm{I}-\eta_{k}\bm{A})\Delta_{0}-\sum_{i=1}^{t}\bm{R}_{i}^{t}(\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i})-\sum_{i=1}^{t}\bm{R}_{i}^{t}(\bm{A}_{i}-\bm{A})\bm{\Delta}_{i-1}$ (65)
in which the matrices $\{\bm{R}_{i}^{t}\}_{1\leq i\leq t}$ are defined as
$\bm{R}_{i}^{t}=\eta_{i}\prod_{k=i+1}^{t}(\bm{I}-\eta_{k}\bm{A}).$ (66)
For simplicity, in this proof we deonte
$\beta=\frac{1-\gamma}{2}\lambda_{0}\eta_{0},\quad R=\frac{5}{2}\|\bm{\theta}^{\star}\|_{2}+\frac{1}{4},\quad\text{and}$
$C=\frac{2\alpha-1}{512\eta_{0}^{2}}\left(\frac{\beta}{2\alpha}\right)^{\frac{\alpha}{1-\alpha}}.$
For any given $\delta$, we define $t^{\star}$ as a function of $\delta$, speicifically
$t^{\star}(\delta)=\inf\left\{t\in\mathbb{N}^{+}:\int_{t}^{\infty}\exp(-Cx^{\alpha})\mathrm{d}x\leq\frac{\delta}{9}\right\},$ (67)
and assume that
$16\eta_{0}\sqrt{\frac{2}{2\alpha-1}}\sqrt{\log\frac{9t^{\star}}{\delta}}\leq 1$ (68)
Furthermore, we denote, for simplicity,
$\mathcal{H}_{i}=\{\max_{1\leq j\leq i}\|\bm{\Delta}_{j}\|\leq R\},\quad\text{and}\quad\widetilde{\bm{\zeta}}_{i}=\bm{\zeta}_{i}\mathds{1}(\mathcal{H}_{i-1}).$
It can then be easily verified that
$\|\widetilde{\bm{\zeta}}_{i}\|_{2}\leq 4R+2\|\bm{\theta}^{\star}\|_{2}+1.$
In what follows, we prove the desired result in the following 4 steps:
  1. 1.
    Showing the existence of $t^{\star}(\delta)$ and featuring its relationship with $\delta$;
  2. 2.
    Showing that $\mathbb{P}(\mathcal{H}_{t^{\star}})\geq 1-\frac{1}{3}\delta$;
  3. 3.
    Showing that $\mathbb{P}(\mathcal{H}_{\infty})\geq 1-\frac{2}{3}\delta$;
  4. 4.
    Showing the high-probability bound (63).
Step 1: the existence of $t^{\star}(\delta)$.
By definition, $t^{\star}$ depends on the integral
$\int_{t}^{\infty}\exp(-Cx^{\alpha})\mathrm{d}x,$
which we handle with a change of variable
$u=x^{\alpha}.$
It is easy to verify that
$\mathrm{d}x=\frac{1}{\alpha}u^{\frac{1-\alpha}{\alpha}}\mathrm{d}u,$
and hence
$\int_{t}^{\infty}\exp(-Cx^{\alpha})\mathrm{d}x$ $=\frac{1}{\alpha}\int_{t^{\alpha}}^{\infty}u^{\frac{1-\alpha}{\alpha}}\exp(-Cu)\mathrm{d}u.$
since $\alpha>0.5$, it can be guaranteed that $u^{\frac{1-\alpha}{\alpha}}<u$, for any $u>t^{\alpha}\geq 1$. Hence, this integral is bounded by
$\int_{t}^{\infty}\exp(-Cx^{\alpha})\mathrm{d}x$ $<\frac{1}{\alpha}\int_{t^{\alpha}}^{\infty}u\exp(-Cu)\mathrm{d}u$
$=\frac{1}{\alpha C^{2}}(Cu+1)\exp(-Cu).$
Consequently, in order that
$\int_{t}^{\infty}\exp(-Cx^{\alpha})\mathrm{d}x\leq\frac{\delta}{9},$
it suffices to let
$(Ct^{\alpha}+1)\exp(-Ct^{\alpha})\leq\frac{\alpha C^{2}\delta}{9};$
It is easy to verify that this condition is satisfied by taking
$t=\left[\frac{2}{C}\left(\log\frac{9}{\alpha C^{2}\delta}\vee\frac{5}{2}\right)\right]^{\frac{1}{\alpha}}.$
In other words, $t^{\star}(\delta)$ is a well-defined function of $\delta$ that scales as $O\left(\log^{\alpha}\frac{1}{\delta}\right)$ as $\delta\to 0$.
Step 2: Bounding $\mathbb{P}(\mathcal{H}_{t^{\star}})$.
We aim to prove, through an induction argument, that
$\mathbb{P}(\mathcal{H}_{t})\geq 1-\frac{t}{t^{\star}}\cdot\frac{\delta}{3}.$ (69)
Since $\bm{\theta}_{0}=0$, the norm of $\bm{\Delta}_{0}$ is equal to that of $\bm{\theta}^{\star}$, and therefore $\mathbb{P}(\mathcal{H}_{0})=1$ is trivial. When (69) holds for $t-1$, we consider the probability of $\mathcal{H}_{t}$. We firstly observe that
$\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\widetilde{\bm{\zeta}}_{i}\right\|_{2}^{2}$ $\leq\sum_{i=1}^{t}\|\bm{R}_{i}^{t}\|^{2}\|\widetilde{\bm{\zeta}}_{i}\|_{2}^{2}$
$<\eta_{0}^{2}(4R+2\|\bm{\theta}^{\star}\|_{2}+1)^{2}\underset{W_{t}}{\underbrace{\sum_{i=1}^{t}i^{-2\alpha}\prod_{k=1}^{t}(1-\beta k^{-\alpha})}}.$
The term $W_{t}$ is bounded by $\sum_{i=1}^{t}i^{-2\alpha}<\frac{1}{2\alpha-1}$. Therefore, the vectors Azuma’s inequality (Corollary A.16) shows that, with probability at least $1-\frac{\delta}{3t^{\star}}$,
$\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\widetilde{\bm{\zeta}}_{i}\right\|_{2}$ $<2\eta_{0}(4R+2\|\bm{\theta}^{\star}\|_{2}+1)\sqrt{\frac{2}{2\alpha-1}\log\frac{9t^{\star}}{\delta}}$
$\leq\frac{1}{8}(4R+2\|\bm{\theta}^{\star}\|_{2}+1),$
where we invoked the condition (68) in the second inequality. The triangle inequality and a union bound argument yield that
$\|\bm{\Delta}_{t}\|$ $\leq\left\|\prod_{k=1}^{t}(\bm{I}-\eta_{k}\bm{A})\bm{\Delta}_{0}\right\|_{2}+\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\bm{\zeta}_{i}\right\|_{2}$
$=\left\|\prod_{k=1}^{t}(\bm{I}-\eta_{k}\bm{A})\bm{\Delta}_{0}\right\|_{2}+\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\widetilde{\bm{\zeta}}_{i}\right\|_{2}$
$<\|\bm{\theta}^{\star}\|_{2}+\frac{1}{8}(4R+2\|\bm{\theta}^{\star}\|_{2}+1)=R,$
with probability at least $\mathbb{P}(\mathcal{H}_{t-1})-\frac{\delta}{3t^{\star}}>1-\frac{t}{t^{\star}}\cdot\frac{\delta}{3}$. The induction assumption then guarantees $\mathbb{P}(\mathcal{H}_{t^{\star}})\geq 1-\frac{\delta}{3}$.
Step 3: Bounding $\mathbb{P}(\mathcal{H}_{\infty})$.
When $t>t^{\star}$, we invoke a tighter bound on $W_{t}$ to bound the norm of $\bm{\Delta}_{t}$. Namely, Lemma A.4 indicates that
$W_{t}\leq\frac{1}{2\alpha-1}\left(\frac{2\alpha}{\beta}\right)^{\frac{\alpha}{1-\alpha}}t^{-\alpha}.$
Furthermore, define
$\delta_{t}=3\exp\left(-\frac{2\alpha-1}{512\eta_{0}^{2}}\left(\frac{\beta}{2\alpha}\right)^{\frac{\alpha}{1-\alpha}}t^{\alpha}\right)=3\exp(-Ct^{\alpha});$
it can then be seen that with probability at least $1-\delta_{t}$,
$\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\widetilde{\bm{\zeta}}_{i}\right\|_{2}$ $\leq 2\eta_{0}(4R+2\|\bm{\theta}^{\star}\|_{2}+1)\sqrt{2W_{t}\log\frac{3}{\delta_{t}}}$
$\leq 2\eta_{0}(4R+2\|\bm{\theta}^{\star}\|_{2}+1)\cdot\frac{1}{16}=\frac{1}{2}R+\frac{1}{4}(2\|\bm{\theta}^{\star}\|_{2}+1).$
With a similar reasoning as the first step, the triangle inequality then indicates
$\mathbb{P}(\mathcal{H}_{t})\geq\mathbb{P}(\mathcal{H}_{t-1})-\delta_{t}.$
Hence, the probability of $\mathcal{H}_{\infty}$ is bounded by
$\mathbb{P}(\mathcal{H}_{\infty})$ $\geq\mathbb{P}(\mathcal{H}_{t^{\star}})-\sum_{t=t^{\star}+1}^{\infty}\delta_{t}$
$\geq\left(1-\frac{\delta}{3}\right)-\sum_{t=t^{\star}+1}^{\infty}3\exp(-Ct^{\alpha})$
$\geq\left(1-\frac{\delta}{3}\right)-\int_{t^{\star}}^{\infty}3\exp(-Cx^{\alpha})\mathrm{d}x$
$\geq\left(1-\frac{\delta}{3}\right)-\frac{\delta}{3}=1-\frac{2}{3}\delta,$
where we applied the definition of $t^{\star}$ at the last inequality.
Step 4: Refining the bound on $\|\bm{\Delta}_{t}\|_{2}$.
With $\mathbb{P}(\mathcal{H}_{\infty})$ being bounded, we now proceed to refining the bound on $\|\bm{\Delta}_{t}\|_{2}$ for any $t\in\mathbb{N}$. Specifically, we observe, due to Lemma A.4, that for every $t\in[T]$,
$\left\|\prod_{k=1}^{t}(\bm{I}-\eta_{k}\bm{A})\bm{\Delta}_{0}\right\|_{2}\leq\exp\left(-\frac{\beta}{1-\alpha}t^{1-\alpha}\right)\|\bm{\Delta}_{0}\|_{2}\leq\left(\frac{\alpha}{2e\beta}\right)^{\frac{\alpha}{2(1-\alpha)}}\|\bm{\theta}^{\star}\|_{2}t^{-\frac{\alpha}{2}}.$
Meanwhile, the vector Azuma’s inequality (Corollary A.16) implies with probability at least $1-\frac{\delta}{3T}$,
$\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\widetilde{\bm{\zeta}}_{i}\right\|_{2}\leq 2\eta_{0}\left(\frac{2\alpha}{\beta}\right)^{\frac{\alpha}{2(1-\alpha)}}\sqrt{\frac{2}{2\alpha-1}\log\frac{9T}{\delta}}(4R+2\|\bm{\theta}^{\star}\|_{2}+1)t^{-\frac{\alpha}{2}}.$
The bound (63) then follows by invoking the triangle inequality and a union bound, and plugging in the definition of $R$ and $\beta$. ∎
B.2 Proof of Theorem 3.1
Similar to the proof of Theorem B.1, we firstly expand $\overline{\bm{\Delta}}_{T}$ into three terms: a non-random term capturing the error introduced by the intialization $\bm{\theta}_{0}$, a martingale, and a sum of $i.i.d.$ vectors. Specifically, (B.1) implies
$\overline{\bm{\Delta}}_{T}$ $=\frac{1}{T}\sum_{t=1}^{T}\left(\prod_{k=1}^{t}(\bm{I}-\eta_{k}\bm{A})\Delta_{0}-\sum_{i=1}^{t}\bm{R}_{i}^{t}(\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i})-\sum_{i=1}^{t}\bm{R}_{i}^{t}(\bm{A}_{i}-\bm{A})\bm{\Delta}_{i-1}\right)$
$=\frac{1}{T}\sum_{t=1}^{T}\prod_{k=1}^{t}(\bm{I}-\eta_{k}\bm{A})\Delta_{0}-\frac{1}{T}\sum_{t=1}^{T}\sum_{i=1}^{t}\bm{R}_{i}^{t}(\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i})-\frac{1}{T}\sum_{t=1}^{T}\sum_{i=1}^{t}\bm{R}_{i}^{t}(\bm{A}_{i}-\bm{A})\bm{\Delta}_{i-1}$
$=\frac{1}{T}\sum_{t=1}^{T}\prod_{k=1}^{t}(\bm{I}-\eta_{k}\bm{A})\Delta_{0}-\frac{1}{T}\sum_{i=1}^{T}\sum_{t=i}^{T}\bm{R}_{i}^{t}(\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i})-\frac{1}{T}\sum_{i=1}^{T}\sum_{t=i}^{T}\bm{R}_{i}^{t}(\bm{A}_{i}-\bm{A})\bm{\Delta}_{i-1},$
where we change the order of summation in the last equation. Hence, the definition of $\bm{R}_{i}^{t}$ (66) and $\bm{Q}_{t}$ (19) indicates that
$\sum_{t=1}^{T}\prod_{k=1}^{t}(\bm{I}-\eta_{k}\bm{A})=\frac{\bm{Q}_{0}}{\eta_{0}},\quad\text{and}$
$\sum_{t=i}^{T}\bm{R}_{i}^{t}=\sum_{t=i}^{T}\left(\eta_{i}\prod_{k=i+1}^{t}(\bm{I}-\eta_{k}\bm{A})\right)=\eta_{i}\sum_{t=i}^{T}\prod_{k=i+1}^{t}(\bm{I}-\eta_{k}\bm{A})=\bm{Q}_{i}.$
Consequently, $\overline{\bm{\Delta}}_{T}$ can be expressed as
$\overline{\bm{\Delta}}_{T}=\underset{I_{1}}{\underbrace{\frac{1}{T\eta_{0}}\bm{Q}_{0}\bm{\Delta}_{0}}}-\underset{I_{2}}{\underbrace{\frac{1}{T}\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i})}}-\underset{I_{3}}{\underbrace{\frac{1}{T}\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}-\bm{A})\bm{\Delta}_{i-1}}}.$ (70)
Here, $I_{1}$ is a non-random term that captures the error introduced by $\bm{\Delta}_{0}$, $I_{2}$ is the sum of $i.i.d.$ random vectors, and $I_{3}$ is a martingale. We address the bounds on these terms in the following paragraphs in order. Through this proof, we denote, for simplicity, $\beta$ as
$\beta:=\frac{1-\gamma}{2}\lambda_{0}\eta_{0}$
Bounding the norm of $I_{1}$.
The norm of $I_{1}$ is bounded by the universal control over $\|\bm{Q}_{t}\|$ as indicated by Lemma A.8. Specifically,
$\left\|\frac{1}{T\eta_{0}}\bm{Q}_{0}\bm{\Delta}_{0}\right\|_{2}\leq\frac{1}{T\eta_{0}}\|\bm{Q}_{0}\|\|\bm{\Delta}_{0}\|\leq\frac{3}{T}\left(\frac{2}{\beta}\right)^{\frac{1}{1-\alpha}}\|\bm{\Delta}_{0}\|_{2}.$ (71)
Bounding the norm of $I_{2}$.
The term $I_{2}$ is a sum of $i.i.d.$ random vectors, and can hence be bounded by invoking the matrix Berstein’s inequality. Specifically, define
$W_{\max}:=\sum_{i=1}^{T}\mathbb{E}\left\|\frac{1}{T}\bm{Q}_{i}(\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i})\right\|_{2}^{2},\quad\text{and}$
$B_{\max}:=\max_{1\leq i\leq T}\sup\left\|\frac{1}{T}\bm{Q}_{i}(\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i})\right\|_{2}.$
It can then be guaranteed with probability at least $1-\delta$ that
$\left\|\frac{1}{T}\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i})\right\|_{2}\leq 2\sqrt{W_{\max}\log\frac{2d}{\delta}}+\frac{4}{3}B_{\max}\log\frac{2d}{\delta}.$
Furthermore, the terms $W_{\max}$ and $B_{\max}$ can be controlled by
$W_{\max}$ $=\frac{1}{T^{2}}\sum_{i=1}^{T}\mathbb{E}\|\bm{Q}_{i}(\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i})\|_{2}^{2}$
$=\frac{1}{T^{2}}\sum_{i=1}^{T}\mathbb{E}\left[(\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i})^{\top}\bm{Q}_{i}^{\top}\bm{Q}_{i}(\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i})\right]$
$=\frac{1}{T^{2}}\sum_{i=1}^{T}\mathbb{E}\left[\mathsf{Tr}\left(\bm{Q}_{i}(\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i})(\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i})^{\top}\bm{Q}_{i}^{\top}\right)\right]$
$=\frac{1}{T^{2}}\sum_{i=1}^{T}\mathsf{Tr}\left(\bm{Q}_{i}\mathbb{E}\left[(\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i})(\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i})^{\top}\right]\bm{Q}_{i}^{\top}\right)$
$=\frac{1}{T}\mathsf{Tr}(\overline{\bm{\Lambda}}_{T});$
$\text{and}\qquad B_{\max}$ $\leq\frac{1}{T}\sup_{1\leq i\leq T}\|\bm{Q}_{i}\|\sup\|\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i}\|_{2}\leq\frac{3\eta_{0}}{T}\left(\frac{2}{\beta}\right)^{\frac{1}{1-\alpha}}(2\|\bm{\theta}^{\star}\|_{2}+1).$
Hence, the norm of $I_{2}$ is bounded by
$\left\|\frac{1}{T}\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i})\right\|_{2}\leq 2\sqrt{\frac{\mathsf{Tr}(\overline{\bm{\Lambda}}_{T})}{T}\log\frac{6d}{\delta}}+\frac{4\eta_{0}}{T}\left(\frac{2}{\beta}\right)^{\frac{1}{1-\alpha}}(2\|\bm{\theta}^{\star}\|_{2}+1)\log\frac{6d}{\delta}.$ (72)
with probability at least $1-\frac{\delta}{3}$.
Bounding the norm of $I_{3}$.
The term $I_{3}$ is a martingale, to which the matrix Azuma’s inequality is applicable. Towards this end, we need the universal bound on $\{\|\bm{\Delta}_{i-1}\|_{2}\}_{1\leq i\leq T}$, for which Theorem B.1 comes in handy. Specifically, under the condition (68), it can be guaranteed with probability at least $1-\frac{\delta}{3}$ that
$\left\|\bm{\Delta}_{t}\right\|_{2}\lesssim\eta_{0}\left(\frac{2\alpha}{\beta}\right)^{\frac{\alpha}{2(1-\alpha)}}\sqrt{\frac{2}{2\alpha-1}\log\frac{27T}{\delta}}(2\|\bm{\theta}^{\star}\|_{2}+1)t^{-\frac{\alpha}{2}},\quad\forall t\in[T].$
For simplicity, we denote
$R=\eta_{0}\left(\frac{2\alpha}{\beta}\right)^{\frac{\alpha}{2(1-\alpha)}}\sqrt{\frac{2}{2\alpha-1}\log\frac{27T}{\delta}}(2\|\bm{\theta}^{\star}\|_{2}+1),$
$\mathcal{H}_{t}=\left\{\max_{1\leq j\leq t}j^{\frac{\alpha}{2}}\|\bm{\Delta}_{j}\|_{2}\leq R\right\},\quad\text{and}$
$\widetilde{\bm{\Delta}}_{t}=\bm{\Delta}_{t}\mathds{1}(\mathcal{H}_{t}).$
Then by definition, the probability of $\mathcal{H}_{t}$ is bounded by
$\mathbb{P}(\mathcal{H}_{1})\geq\mathbb{P}(\mathcal{H}_{2})\geq...\geq\mathbb{P}(\mathcal{H}_{T})\geq 1-\frac{\delta}{3}.$ (73)
Meanwhile, the vector Azuma’s inequality can be invoked to obtain that with probability at least $1-\frac{\delta}{3}$,
$\left\|\frac{1}{T}\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}-\bm{A})\widetilde{\bm{\Delta}}_{i-1}\right\|_{2}\leq 2\sqrt{2W_{\max}\log\frac{9}{\delta}},$
where
$W_{\max}=\sup\frac{1}{T^{2}}\sum_{i=1}^{T}\|\bm{Q}_{i}(\bm{A}_{i}-\bm{A})\widetilde{\bm{\Delta}}_{i-1}\|_{2}^{2}$ $\lesssim\frac{1}{T^{2}}3\eta_{0}^{2}\left(\frac{2}{\beta}\right)^{\frac{1}{1-\alpha}}\cdot 4R^{2}\sum_{i=1}^{T}i^{-\alpha}$
$\asymp\frac{12\eta_{0}^{2}}{1-\alpha}\left(\frac{2}{\beta}\right)^{\frac{1}{1-\alpha}}R^{2}T^{-1-\alpha}.$
In other words, it can be guaranteed with probability at least $1-\frac{\delta}{3}$ that
$\left\|\frac{1}{T}\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}-\bm{A})\widetilde{\bm{\Delta}}_{i-1}\right\|_{2}$ $\lesssim\eta_{0}\left(\frac{2}{\beta}\right)^{\frac{1}{2(1-\alpha)}}R\sqrt{\frac{1}{1-\alpha}\log\frac{9}{\delta}}$
$\lesssim\eta_{0}^{2}\left(\frac{2}{\beta}\right)^{\frac{2+\alpha}{2(1-\alpha)}}(6\|\bm{\theta}^{\star}\|_{2}+1)\sqrt{\frac{1}{(1-\alpha)(2\alpha-1)}}T^{-\frac{\alpha+1}{2}}\log\frac{27T}{\delta}.$ (74)
Combining (70), (71), (72), (73) and (B.2) with the triangle inequality and a union bound argument, we obtain with probability $1-\delta$ that
$\|\overline{\bm{\Delta}}_{T}\|_{2}$ $\leq\frac{3}{T}\left(\frac{2}{\beta}\right)^{\frac{1}{1-\alpha}}\|\bm{\Delta}_{0}\|_{2}$
$+2\sqrt{\frac{\mathsf{Tr}(\overline{\bm{\Lambda}}_{T})}{T}\log\frac{6d}{\delta}}+\frac{4\eta_{0}}{T}\left(\frac{2}{\beta}\right)^{\frac{1}{1-\alpha}}(2\|\bm{\theta}^{\star}\|_{2}+1)\log\frac{6d}{\delta}$
$+\eta_{0}^{2}\left(\frac{2}{\beta}\right)^{\frac{2+\alpha}{2(1-\alpha)}}(6\|\bm{\theta}^{\star}\|_{2}+1)\sqrt{\frac{1}{(1-\alpha)(2\alpha-1)}}T^{-\frac{\alpha+1}{2}}\log\frac{27T}{\delta}.$
Theorem 3.1 follows by taking
$\widetilde{C}$ $=\eta_{0}^{2}\left(\frac{2}{\beta}\right)^{\frac{2+\alpha}{2(1-\alpha)}}(6\|\bm{\theta}^{\star}\|_{2}+1)\sqrt{\frac{1}{(1-\alpha)(2\alpha-1)}}.$ (75)
B.3 Proof of Theorem 3.2
This proof is based on Theorem A.11, the Berry-Esseen bound for multivariate non-linear statistics. Throughout the proof, we will firstly define corresponding statistics, and then bound the terms on the right-hand-side of (62).
Defining $\bm{W}$,$\bm{D}$ and other statistics.
Recall from (70) that the averaged TD error can be represented as
$\overline{\bm{\Delta}}_{T}=\frac{1}{T\eta_{0}}\bm{Q}_{0}\bm{\Delta}_{0}-\frac{1}{T}\sum_{i=1}^{T-1}\bm{Q}_{i}\bm{\zeta}_{i},$
where $\bm{Q}_{t}$ is defined as in (19). Therefore, the statistics $\bm{W}$ and $\bm{D}$ can be constructed by
$\sqrt{T}\bm{A}\overline{\bm{\Delta}}_{T}$ $=\frac{\bm{A}}{\sqrt{T}\eta_{0}}\bm{Q}_{0}\bm{\Delta}_{0}-\frac{\bm{A}}{\sqrt{T}}\sum_{i=1}^{T}\bm{Q}_{i}\bm{\zeta}_{i}$
$=\underset{\bm{W}}{\underbrace{\sum_{i=1}^{T}\bm{\xi}_{i}}}+\underset{\bm{D}}{\underbrace{\left(\frac{\bm{A}}{\sqrt{T}\eta_{0}}\bm{Q}_{0}\Delta_{0}+\frac{\bm{A}}{\sqrt{T}}\sum_{i=1}^{T}\bm{Q}_{i}\bm{\omega}_{i}\right)}},$ (76)
in which we define $\bm{\xi}_{i}$ and $\bm{\omega}_{i}$ as
$\bm{\xi}_{i}=\frac{1}{\sqrt{T}}\bm{A}\bm{Q}_{i}(\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i});$ (77)
$\bm{\omega}_{i}=(\bm{A}_{i}-\bm{A})\Delta_{i-1}.$ (78)
Notice that since $\{(s_{i},s_{i}^{\prime})\}_{i=1}^{T}$ are independent, $\{(\bm{A}_{i},\bm{b}_{i})\}_{i=1}^{T}$ are also independent. By definition, it is easy to verify that $\mathbb{E}[\bm{\xi}_{i}]=0$ and that
$\sum_{i=1}^{T}\mathbb{E}[\bm{\xi}_{i}\bm{\xi}_{i}^{\top}]$ $=\frac{1}{T}\bm{A}\left\{\sum_{i=1}^{T}\bm{Q}_{i}\mathbb{E}[(\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i})(\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i})^{\top}]\bm{Q}_{i}^{\top}\right\}\bm{A}^{\top}$
$=\bm{A}\left\{\frac{1}{T}\sum_{i=1}^{T}\bm{Q}_{i}\bm{\Gamma}\bm{Q}_{i}^{\top}\right\}\bm{A}^{\top}=\bm{A}\overline{\bm{\Lambda}}_{T}\bm{A}^{\top}.$
With the minimum eigenvalue bounded below in Lemma A.10 by
$\lambda_{\min}(\bm{A}\overline{\bm{\Lambda}}_{T}\bm{A}^{\top})\geq\frac{1}{2\|\bm{\Gamma}^{-1}\|},$
Theorem A.11 guarantees that
$d_{\mathsf{C}}(\bm{A}\overline{\bm{\Delta}}_{T},(\bm{A}\overline{\bm{\Lambda}}_{T}\bm{A}^{\top})^{\frac{1}{2}}\bm{z})$ $\lesssim\underset{I_{1}}{\underbrace{\sqrt{d}\|\bm{\Gamma}^{-\frac{3}{2}}\|\cdot\sum_{i=1}^{n}\mathbb{E}\|\bm{\xi}_{i}\|_{2}^{3}}}+\underset{I_{2}}{\underbrace{\|\bm{\Gamma}^{-1}\|\mathbb{E}\{\|\bm{W}\|_{2}\Delta\}}}+\underset{I_{3}}{\underbrace{\|\bm{\Gamma}^{-1}\|\sum_{i=1}^{n}\mathbb{E}\{\|\bm{\xi}_{i}\|_{2}|\Delta-\Delta^{(i)}|\}}},$ (79)
where $\Delta=\|\bm{D}\|$ and $\Delta^{(i)}$ is any random variable independent of $(\bm{A}_{i},\bm{b}_{i})$. Now, the proof boils down to bounding the three terms on the right-hand-side of (79).
Bounding $I_{1}$.
We firstly observe that
$\|\bm{\xi}_{i}\|_{2}^{3}$ $\leq T^{-3/2}\|\bm{AQ}_{i}\|^{3}\|\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i}\|_{2}^{3}$
$\lesssim T^{-3/2}\cdot\left(2+\eta_{0}\left(\frac{1}{\beta}\right)^{\frac{1}{1-\alpha}}\Gamma\left(\frac{1}{1-\alpha}\right)i^{\alpha-1}\right)^{3}\cdot\|\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i}\|_{2}^{3},$
in which $\|\bm{AQ}_{i}\|$ is bounded by Lemma A.8. Furthermore, since $\max_{s\in\mathcal{S}}\|\bm{\phi}(s)\|_{2}\leq 1$ and $\max_{s\in\mathcal{S}}r(s)\leq 1$, it is easy to verify that
$\|\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i}\|_{2}<2\|\bm{\theta}^{\star}\|_{2}+1.$
Therefore, the term $I_{1}$ is bounded by
$\sqrt{d}\sum_{i=1}^{T}\mathbb{E}\|\bm{\xi}_{i}\|_{2}^{3}\lesssim\sqrt{d}\cdot\left(\|\bm{\Gamma}^{-1}\|\right)^{3/2}\left(2\|\bm{\theta}^{\star}\|_{2}+1\right)^{3}\left(\eta_{0}\left(\frac{1}{\beta}\right)^{\frac{1}{1-\alpha}}\Gamma\left(\frac{1}{1-\alpha}\right)\right)^{3}T^{-\frac{1}{2}}=O(T^{-\frac{1}{2}}).$ (80)
Bounding $I_{2}$.
First of all, we define $\Delta=\|\bm{D}\|_{2}$. By Cauchy-Schwartz inequality, the second term of (79) is upper bounded by
$\mathbb{E}\left\{\|\bm{W}\|_{2}\Delta\right\}\leq\sqrt{\mathbb{E}\|\bm{W}\|_{2}^{2}\mathbb{E}\Delta^{2}}.$
Regarding $\mathbb{E}\|\bm{W}\|_{2}^{2}$, notice that
$\mathbb{E}\|\bm{W}\|_{2}^{2}$ $=\mathbb{E}[\mathsf{Tr}(\bm{W}\bm{W}^{\top})]=\mathsf{Tr}(\bm{A}\overline{\bm{\Lambda}}_{T}\bm{A}^{\top})=\mathsf{Tr}(\bm{\Gamma})+O(T^{\alpha-1}),$
where the last identity is a consequence of Theorem A.9. In detail,
$\bm{A}\overline{\bm{\Lambda}}_{T}\bm{A}^{\top}-\bm{\Gamma}=\bm{A}(\overline{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star})\bm{A}^{\top}=T^{\alpha-1}\bm{X}(\bm{\Lambda}^{\star})\bm{A}^{\top}+\bm{O}(T^{2\alpha-2})=\bm{O}(T^{\alpha-1}),$
where $\bm{X}(\bm{\Lambda}^{\star})$ is defined as in (56).
As for the expectation of $|\Delta|^{2}$, observe that
$\mathbb{E}\Delta^{2}$ $\leq 2\mathbb{E}\left\|\frac{1}{\sqrt{T}\eta_{0}}\bm{AQ}_{0}\bm{\Delta}_{0}\right\|_{2}^{2}+2\mathbb{E}\left\|\frac{1}{\sqrt{T}}\sum_{i=1}^{T}\bm{AQ}_{i}\bm{\omega}_{i}\right\|_{2}^{2}$
$\leq\frac{2}{T\eta_{0}^{2}}\|\bm{AQ}_{0}\|^{2}\|\bm{\Delta}_{0}\|_{2}^{2}+\frac{2}{T}\sum_{i=1}^{T}\|\bm{AQ}_{i}\|^{2}\mathbb{E}\|\bm{\omega}_{i}\|_{2}^{2}$
$\lesssim\frac{1}{T\eta_{0}^{2}}\cdot\left(\frac{2}{(1-\gamma)\lambda_{0}\eta_{0}}\right)^{\frac{2}{1-\alpha}}\|\bm{\Delta}_{0}\|_{2}^{2}$
$+\frac{1}{T}\sum_{i=1}^{T}(2+O(i^{\alpha-1}))^{2}\mathbb{E}\|(\bm{A}_{i}-\bm{A})\|_{2}^{2}\mathbb{E}\|\bm{\Delta}_{i-1}\|_{2}^{2}$
$\lesssim\frac{1}{T}\sum_{i=1}^{T}\mathbb{E}\|\bm{\Delta}_{i-1}\|_{2}^{2}+O\left(\frac{1}{T}\right).$
The first line is the inequality $\|\bm{x}+\bm{y}\|_{2}^{2}\leq 2\|\bm{x}\|_{2}^{2}+2\|\bm{y}\|_{2}^{2}$, for any vectors $\bm{x}$ and $\bm{y}$; in the third line, we applied the upper bounds for $\|\bm{AQ}_{t}\|$ established in Lemma A.8. For the term $\mathbb{E}\|\bm{\Delta}_{i-1}\|_{2}^{2}$, the following lemma comes in handy.
Lemma B.2.
Under the TD iteration (9) with $i.i.d.$ samples and decaying stepsizes $\eta_{t}=\eta_{0}t^{-\alpha}$ for $\alpha\in(\frac{1}{2},1)$ and $\eta_{0}\leq\frac{1}{2}$, it can be guaranteed that
$\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}\leq\frac{3\eta_{0}\mathsf{Tr}(\bm{\Gamma})}{(1-\gamma)\lambda_{0}}t^{-\alpha}+O(t^{-1}).$
holds for any $t>0$.
Proof: See Appendix C.1. ∎
As an immediate result of Lemma B.2, the last term is bounded by
$\sum_{i=1}^{T}\mathbb{E}\|\bm{\Delta}_{i}\|_{2}^{2}$ $\leq\sum_{i=1}^{T}\frac{3\eta_{0}}{(1-\gamma)\lambda_{0}}\mathsf{Tr}(\bm{\Gamma})i^{-\alpha}+O(i^{-1})$
$=\frac{3\eta_{0}}{(1-\alpha)(1-\gamma)\lambda_{0}}(2\|\bm{\theta}^{\star}\|_{2}+1)^{2}T^{1-\alpha}+O(\log T).$
Therefore, $\mathbb{E}\Delta^{2}$ can be bounded by
$\mathbb{E}\Delta^{2}$ $\lesssim\frac{1}{T}\sum_{i=1}^{T}\mathbb{E}\|\bm{\Delta}_{i-1}\|_{2}^{2}+O\left(\frac{1}{T}\right)$
$\leq\frac{1}{T}\left(\frac{3\eta_{0}}{(1-\alpha)(1-\gamma)\lambda_{0}}\mathsf{Tr}(\bm{\Gamma})T^{1-\alpha}+O(\log T)\right)+O\left(\frac{1}{T}\right)$
$=\frac{3\eta_{0}\mathsf{Tr}(\bm{\Gamma})}{(1-\alpha)(1-\gamma)\lambda_{0}}T^{-\alpha}+O\left(\frac{\log T}{T}\right)$
So in conclusion, the second term of (79) can be bounded by
$\mathbb{E}\{\|W\|\Delta\}\leq\sqrt{\mathbb{E}\|W\|_{2}^{2}\mathbb{E}\Delta^{2}}$ $\lesssim\sqrt{\frac{3\eta_{0}}{(1-\alpha)(1-\gamma)\lambda_{0}}}\mathsf{Tr}(\bm{\Gamma})T^{-\frac{\alpha}{2}}+O\left(\sqrt{\frac{\log T}{T}}\right).$ (81)
Bounding $I_{3}$.
The third term of (79) involves random variables $\{\Delta^{(i)}\}_{0\leq i\leq T}$ such that $\Delta^{(i)}$ is independent of $(s_{i},s_{i}^{\prime})$. For this purpose, we define $\{(\check{s}_{i},\check{s}_{i}^{\prime})\}_{i=1}^{n}$ as independent copies of $\{(s_{i},s_{i}^{\prime})\}$ and let $\check{\bm{A}}_{i},\check{\bm{b}}_{i}$ be generated by $(\check{s}_{i},\check{s}_{i}^{\prime})$. Consider the following iterations for any $i$:
$\bm{\theta}_{j}^{(i)}=\begin{cases}\bm{\theta}_{j},\text{ if }j<i;\\ \bm{\theta}_{j-1}-\eta_{j}(\check{\bm{A}}_{j}\bm{\theta}_{j-1}-\check{\bm{b}}_{j}),\text{ if }j=i;\\ \bm{\theta}_{j-1}^{(i)}-\eta_{j}(\bm{A}_{j}\bm{\theta}_{j-1}^{(i)}-\bm{b}_{j}),\text{ if }j>i.\end{cases}$ (82)
It is easy to verify that for all $j\in[T]$, $\bm{\theta}_{j}^{(i)}$ is independent of $(s_{i},s_{i}^{\prime})$. Therefore, with the definitions of
$\bm{\omega}_{j}^{(i)}=\begin{cases}(\check{\bm{A}}_{j}-\bm{A})(\bm{\theta}_{j-1}^{(i)}-\bm{\theta}^{\star}),\quad\text{ if }\quad j=i;\\ (\bm{A}_{j}-\bm{A})(\bm{\theta}_{j-1}^{(i)}-\bm{\theta}^{\star}),\quad\text{otherwise.}\end{cases}$ (83)
and
$\Delta^{(i)}=\left\|\frac{1}{\sqrt{T}\eta_{0}}\bm{AQ}_{0}\bm{\Delta}_{0}-\frac{1}{\sqrt{T}}\sum_{j=1}^{T}\bm{AQ}_{j}\bm{\omega}_{j}^{(i)}\right\|_{2},$
It can be guaranteed that $\Delta^{(i)}$ is independent of $(s_{i},s_{i}^{\prime})$. The proof now boils down to bounding the difference between $\Delta$ and $\Delta^{(i)}$, for which the triangle inequality implies
$|\Delta-\Delta^{(i)}|\leq\left\|\frac{1}{\sqrt{T}}\sum_{j=1}^{T}\bm{AQ}_{j}(\bm{\omega}_{j}-\bm{\omega}_{j}^{(i)})\right\|_{2}$ $\leq\frac{1}{\sqrt{T}}\left\|\sum_{j=i}^{T}\bm{AQ}_{j}(\bm{\omega}_{j}-\bm{\omega}_{j}^{(i)})\right\|_{2}.$
By squaring and taking expectations on both sides, we observe that
$\mathbb{E}|\Delta-\Delta^{(i)}|^{2}$ $\leq\frac{1}{T}\mathbb{E}\left\|\sum_{j=i}^{T}\bm{AQ}_{j}(\bm{\omega}_{j}-\bm{\omega}_{j}^{(i)})\right\|_{2}^{2}$
$=\frac{1}{T}\mathbb{E}\sum_{j=i}^{T}\left\|\bm{AQ}_{j}(\bm{\omega}_{j}-\bm{\omega}_{j}^{(i)})\right\|_{2}^{2}+\frac{2}{T}\mathbb{E}\left\{\sum_{i\leq j<j^{\prime}<T}\left[\bm{AQ}_{j}\left(\bm{\omega}_{j}-\bm{\omega}_{j}^{(i)})\right)\right]^{\top}\left[\bm{AQ}_{j^{\prime}}\left(\bm{\omega}_{j^{\prime}}-\bm{\omega}_{j^{\prime}}^{(i)}\right)\right]\right\}.$ (84)
For the last term on the right-hand-side of (B.3), recall from the definitions of $\bm{\omega}_{j}$ (78) and $\bm{\omega}_{j}^{(i)}$ (83), for any $j>i$,
$\bm{\omega}_{j}-\bm{\omega}_{j}^{(i)}$ $=\left[(\bm{A}_{j}-\bm{A})(\bm{\theta}_{j-1}-\bm{\theta}^{\star})\right]-\left[(\bm{A}_{j}-\bm{A})\left(\bm{\theta}_{j-1}^{(i)}-\bm{\theta}^{\star}\right)\right]$
$=(\bm{A}_{j}-\bm{A})\left(\bm{\theta}_{j-1}-\bm{\theta}_{j-1}^{(i)}\right)$
Hence, for $i\leq j<j^{\prime}<T$, the law of total expectation implies
$\mathbb{E}\left\{\left[\bm{AQ}_{j}\left(\bm{\omega}_{j}-\bm{\omega}_{j}^{(i)})\right)\right]^{\top}\left[\bm{AQ}_{j^{\prime}}\left(\bm{\omega}_{j^{\prime}}-\bm{\omega}_{j^{\prime}}^{(i)}\right)\right]\right\}$
$=\mathbb{E}\mathbb{E}_{j^{\prime}-1}\left\{\left[\bm{AQ}_{j}\left(\bm{\omega}_{j}-\bm{\omega}_{j}^{(i)}\right)\right]^{\top}\left[\bm{AQ}_{j^{\prime}}\left(\bm{\omega}_{j^{\prime}}-\bm{\omega}_{j^{\prime}}^{(i)}\right)\right]\right\}$
$=\mathbb{E}\left\{\left[\bm{AQ}_{j}\left(\bm{\omega}_{j}-\bm{\omega}_{j}^{(i)}\right)\right]^{\top}\left[\bm{AQ}_{j^{\prime}}\mathbb{E}_{j^{\prime}-1}\left(\bm{\omega}_{j^{\prime}}-\bm{\omega}_{j^{\prime}}^{(i)}\right)\right]\right\}$
$=\mathbb{E}\left\{\left[\bm{AQ}_{j}\left(\bm{\omega}_{j}-\bm{\omega}_{j}^{(i)}\right)\right]^{\top}\left[\bm{AQ}_{j^{\prime}}\mathbb{E}_{j^{\prime}-1}\left[(\bm{A}_{j^{\prime}}-\bm{A})\left(\bm{\theta}_{j^{\prime}-1}-\bm{\theta}_{j^{\prime}-1}^{(i)}\right)\right]\right]\right\}=0,$
since $\mathbb{E}_{j^{\prime}-1}[\bm{A}_{j^{\prime}}-\bm{A}]=\bm{0}$. This effectively means that the last term on the right-hand-side of (B.3) is 00. Therefore, the difference bewteen $\Delta$ and $\Delta^{(i)}$ is bounded by
$\mathbb{E}|\Delta-\Delta^{(i)}|^{2}$ $\leq\frac{1}{T}\mathbb{E}\sum_{j=i}^{T}\left\|\bm{AQ}_{j}(\bm{\omega}_{j}-\bm{\omega}_{j}^{(i)})\right\|_{2}^{2}$
$\leq\frac{1}{T}\sum_{j=i}^{T}\|\bm{AQ}_{j}\|^{2}\mathbb{E}\left\|\bm{\omega}_{j}-\bm{\omega}_{j}^{(i)}\right\|_{2}^{2}$
$\lesssim\frac{1}{T}\cdot\sum_{j=i}^{T}(2+O(j^{\alpha-1}))\mathbb{E}\left\|\bm{\omega}_{j}-\bm{\omega}_{j}^{(i)}\right\|_{2}^{2},$ (85)
where in the last line we applied Lemma A.8 to bound the norm of $\bm{Q}_{j}$. The proof now boils down to bounding the expected difference between $\bm{\omega}_{j}$ and $\bm{\omega}_{j}^{(i)}$, for which we recall from definition that
$\bm{\omega}_{j}-\bm{\omega}_{j}^{(i)}$ =$=\begin{cases}(\bm{A}_{i}-\check{\bm{A}}_{i})(\bm{\theta}_{i-1}-\bm{\theta}^{\star}),\quad\text{for}\quad j=i;\\ (\bm{A}_{j}-\bm{A})\left(\bm{\theta}_{j-1}-\bm{\theta}_{j-1}^{(i)}\right),\quad\text{for}\quad j>i.\end{cases}$
The norm of $(\bm{\theta}_{i-1}-\bm{\theta}^{\star})$ can be bounded by Lemma B.2; for the norm of $\left(\bm{\theta}_{j-1}-\bm{\theta}_{j-1}^{(i)}\right)$, the following proposition comes in handy.
Proposition B.3.
With $\bm{\theta}_{j}^{(i)}$ defined in (82),it can be guaranteed that for every $i\in[T]$,
$\mathbb{E}\|\bm{\theta}_{i}-\bm{\theta}_{i}^{(i)}\|_{2}^{2}\leq 2\eta_{i}^{2}\mathsf{Tr}(\bm{\Gamma})+2\eta_{i}^{2}\mathbb{E}\|\bm{\Delta}_{i-1}\|_{2}^{2}.$ (86a)
Furthermore, for every $j>i$,
$\mathbb{E}\|\bm{\theta}_{j}-\bm{\theta}_{j}^{(i)}\|_{2}^{2}$ $\leq\prod_{k=i+1}^{j}\left(1-\frac{1}{2}(1-\gamma)\lambda_{0}\eta_{k}\right)\mathbb{E}\|\bm{\theta}_{i}-\bm{\theta}_{i}^{(i)}\|_{2}^{2}.$ (86b)
Proof: See appendix C.2. ∎
For $j=i$, direct calculation yields
$\mathbb{E}\|\bm{\omega}_{i}-\bm{\omega}_{i}^{(i)}\|_{2}^{2}$ $=\mathbb{E}\|(\bm{A}_{i}-\check{\bm{A}}_{i})(\bm{\theta}_{i-1}-\bm{\theta}^{\star})\|_{2}^{2}$
$\lesssim\mathbb{E}\|\bm{\Delta}_{i-1}\|_{2}^{2};$ (87)
and similarly, for $j>i$,
$\mathbb{E}\|\bm{\omega}_{j}-\bm{\omega}_{j}^{(i)}\|_{2}^{2}$ $=\mathbb{E}\left\|(\bm{A}_{j}-\bm{A})\left(\bm{\theta}_{j-1}-\bm{\theta}_{j-1}^{(i)}\right)\right\|_{2}^{2}$
$\lesssim\mathbb{E}\left\|\bm{\theta}_{j-1}-\bm{\theta}_{j-1}^{(i)}\right\|_{2}^{2}.$ (88)
By combining (B.3), (B.3),(B.3) with (86b), we obtain
$\mathbb{E}|\Delta-\Delta^{(i)}|^{2}$ $\leq\frac{1}{T}\left((2+O(i^{\alpha-1}))\mathbb{E}\|\bm{\omega}_{i}-\bm{\omega}_{i}^{(i)}\|_{2}^{2}+\sum_{j=i+1}^{T}(2+O(j^{\alpha-1}))\mathbb{E}\|\bm{\omega}_{j}-\bm{\omega}_{j}^{(i)}\|_{2}^{2}\right)$
$\lesssim\frac{1}{T}\cdot(2+O(i^{\alpha-1}))\underset{I_{1}}{\underbrace{\mathbb{E}\|\bm{\omega}_{i}-\bm{\omega}_{i}^{(i)}\|_{2}^{2}}}$
$+\frac{1}{T}\underset{I_{2}}{\underbrace{\sum_{j=i+1}^{T}(2+O(j^{\alpha-1}))\prod_{k=i+1}^{j}\left(1-\frac{1-\gamma}{2}\lambda_{0}\eta_{k}\right)\mathbb{E}\|\bm{\theta}_{i}-\bm{\theta}_{i}^{(i)}\|_{2}^{2}}}.$
With $I_{1}$ bounded by (B.3), the proof now boils down to bounding $I_{2}$. Notice that (86a) implies
$\sum_{j=i+1}^{T}\prod_{k=i+1}^{j}\left(1-\frac{1-\gamma}{2}\lambda_{0}\eta_{k}\right)\mathbb{E}\|\bm{\theta}_{i}-\bm{\theta}_{i}^{(i)}\|_{2}^{2}$
$\lesssim\left[\mathsf{Tr}(\bm{\Gamma})+2\mathbb{E}\|\bm{\Delta}_{i-1}\|_{2}^{2}\right]\eta_{0}^{2}i^{-2\alpha}\sum_{j=i+1}^{T}\prod_{k=i+1}^{j}\left(1-\frac{1-\gamma}{2}\lambda_{0}\eta_{k}\right),$
where we can invoke (42) in Lemma A.5 to obtain
$i^{-2\alpha}\sum_{j=i+1}^{T}\prod_{k=i+1}^{j}\left(1-\frac{1-\gamma}{2}\lambda_{0}\eta_{k}\right)<\frac{2}{(1-\gamma)\lambda_{0}\eta_{0}}i^{-\alpha}+O(i^{-1}).$
Combining these results with Lemma B.2, we establish the following bound for $I_{1}+I_{2}$:
$I_{1}+I_{2}\lesssim\frac{\eta_{0}}{(1-\gamma)\lambda_{0}}\mathsf{Tr}(\bm{\Gamma})i^{-\alpha}+O(i^{-1}).$
Hence,
$\mathbb{E}|\Delta-\Delta^{(i)}|^{2}\lesssim\frac{1}{T}\frac{\eta_{0}}{(1-\gamma)\lambda_{0}}\mathsf{Tr}(\bm{\Gamma})i^{-\alpha}+O(i^{-1}).$
Therefore, by Cauchy-Schwartz inequality,
$\left(\mathbb{E}\sum_{i=1}^{T}\|\bm{\xi}_{i}\||\Delta-\Delta^{(i)}|\right)^{2}\leq\left(\sum_{i=1}^{T}\mathbb{E}\|\bm{\xi}_{i}\|_{2}^{2}\right)\left(\sum_{i=1}^{T}\mathbb{E}|\Delta-\Delta^{(i)}|^{2}\right),$ (89)
in which the two terms on the right-hand-side can be bounded respectively by
$\sum_{i=1}^{T}\mathbb{E}\|\bm{\xi}_{i}\|_{2}^{2}$ $=\sum_{i=1}^{T}\mathbb{E}\mathsf{Tr}[\bm{\xi}_{i}\bm{\xi}_{i}^{\top}]=\mathsf{Tr}(\bm{A}\overline{\bm{\Lambda}}_{T}\bm{A}^{\top})=\mathsf{Tr}(\bm{\Gamma})+O(T^{\alpha-1})$ (90)
and
$\sum_{i=1}^{T}\mathbb{E}|\Delta-\Delta^{(i)}|^{2}$ $\lesssim\frac{1}{T}\frac{\eta_{0}}{(1-\gamma)\lambda_{0}}\mathsf{Tr}(\bm{\Gamma})\sum_{i=1}^{T}\left(i^{-\alpha}+O(i^{-1})\right)$
$=\frac{\eta_{0}\mathsf{Tr}(\bm{\Gamma})}{(1-\alpha)(1-\gamma)\lambda_{0}}T^{-\alpha}+O\left(\frac{\log T}{T}\right).$ (91)
A direct combination of (89), (90) and (B.3) yields the following bound for the third term of (79):
$\mathbb{E}\sum_{i=1}^{T}\|\bm{\xi}_{i}\||\Delta-\Delta^{(i)}|\lesssim\sqrt{\frac{\eta_{0}}{(1-\alpha)(1-\gamma)\lambda_{0}}}\mathsf{Tr}(\bm{\Gamma})T^{-\frac{\alpha}{2}}+O\left(\sqrt{\frac{\log T}{T}}\right).$ (92)
Completing the proof.
A direct combination of (79), (80), (81) and (92) yields
$\left|\mathbb{P}(\sqrt{T}\bm{A}\overline{\bm{\Delta}}_{T}\in\mathcal{A})-\mathbb{P}((\bm{A}\overline{\bm{\Lambda}}_{T}\bm{A}^{\top})^{\frac{1}{2}}\bm{z}\in\mathcal{A})\right|\leq\sqrt{\frac{\eta_{0}}{(1-\alpha)(1-\gamma)\lambda_{0}}}\|\bm{\Gamma}^{-1}\|\mathsf{Tr}(\bm{\Gamma})T^{-\frac{\alpha}{2}}+O\left(\sqrt{\frac{\log T}{T}}\right).$
Furthermore, for any convex set $\mathcal{A}\in\mathbb{R}^{d}$, the set
$\mathcal{A}^{\prime}=\{\bm{A}\bm{x}:\bm{x}\in\mathcal{A}\}$
must also be convex; by definition, $\mathcal{A}$ and $\mathcal{A}^{\prime}$ are related by
$\mathbb{P}(\sqrt{T}\overline{\bm{\Lambda}}_{T}^{-1/2}\overline{\bm{\Delta}}_{T}\in\mathcal{A})=\mathbb{P}(\sqrt{T}\overline{\bm{\Delta}}_{T}\in\mathcal{A}^{\prime}),\quad\text{and}$
$\mathbb{P}((\bm{A}\overline{\bm{\Lambda}}_{T}\bm{A}^{\top})^{\frac{1}{2}}\bm{z}\in\mathcal{A})=\mathbb{P}(\overline{\bm{\Lambda}}_{T}^{1/2}\bm{z}\in\mathcal{A}^{\prime}).$
Therefore, for any convex set $\mathcal{A}\in\mathbb{R}^{d}$, it can be guaranteed that
$\left|\mathbb{P}(\sqrt{T}\overline{\bm{\Delta}}_{T}\in\mathcal{A})-\mathbb{P}(\overline{\bm{\Lambda}}_{T}^{1/2}\bm{z}\in\mathcal{A})\right|\leq\sqrt{\frac{\eta_{0}}{(1-\alpha)(1-\gamma)\lambda_{0}}}\|\bm{\Gamma}^{-1}\|\mathsf{Tr}(\bm{\Gamma})T^{-\frac{\alpha}{2}}+O\left(\sqrt{\frac{\log T}{T}}\right).$
Taking supremum over $\mathcal{A}$ completes the proof. ∎
B.4 Proof of Theorem 3.3
For any Borel set $\mathcal{A}\subset\mathbb{R}^{d}$, define $\mathcal{A}^{\prime}=\{\bm{Ax}:\bm{x}\in\mathbb{R}^{d}\}$. By definition, when $\bm{z}$ is the $d$-dimensional standard Gaussian distribution,
$\left|\mathbb{P}(\overline{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}\in\mathcal{A})-\mathbb{P}({\bm{\Lambda}}^{\star\frac{1}{2}}\bm{z}\in\mathcal{A})\right|=\left|\mathbb{P}(\bm{A}\overline{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}\in\mathcal{A}^{\prime})-\mathbb{P}(\bm{A}{\bm{\Lambda}}^{\star\frac{1}{2}}\bm{z}\in\mathcal{A}^{\prime})\right|.$
Since $\bm{A}$ is non-singular, the map $\mathcal{A}\mapsto\mathcal{A}^{\prime}$ is a bijection. Hence by taking supremum over $\mathcal{A}$ and $\mathcal{A}^{\prime}$ on both sides of the equation, we obtain
$d_{\mathsf{TV}}(\overline{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z},{\bm{\Lambda}}^{\star\frac{1}{2}}\bm{z})=d_{\mathsf{TV}}(\bm{A}\overline{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z},\bm{A}{\bm{\Lambda}}^{\star\frac{1}{2}}\bm{z});$
In other words,
$d_{\mathsf{TV}}(\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T}),\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))=d_{\mathsf{TV}}(\mathcal{N}(\bm{0},\bm{A}\overline{\bm{\Lambda}}_{T}\bm{A}^{\top}),\mathcal{N}(\bm{0},\bm{\Gamma}));$
Therefore, the proof boils down to bounding the TV distance between $\mathcal{N}(\bm{0},\bm{A}\overline{\bm{\Lambda}}_{T}\bm{A}^{\top})$ and $\mathcal{N}(\bm{0},\bm{\Gamma})$. Towards this end, Theorem A.12 indicates that this TV distance converges at the same rate of
$\left\|\bm{\Gamma}^{-\frac{1}{2}}\bm{A}\overline{\bm{\Lambda}}_{T}\bm{A}^{\top}\bm{\Gamma}^{-\frac{1}{2}}-\bm{I}_{d}\right\|_{\mathsf{F}}=\left\|\bm{\Gamma}^{-\frac{1}{2}}\bm{A}(\overline{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star})\bm{A}^{\top}\bm{\Gamma}^{-\frac{1}{2}}\right\|_{\mathsf{F}}.$
Here, $\overline{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star}$ featured by Theorem A.9 as
$\overline{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star}=T^{\alpha-1}\bm{X}+\bm{O}(T^{2\alpha-2}),$
where $\bm{X}$ satisfies the Lyapunov equation
$\eta_{0}(\bm{AX}+\bm{XA}^{\top})=\bm{\Lambda}^{\star};$
Therefore it is easy to verify that
$\bm{\Gamma}^{-\frac{1}{2}}\bm{A}(\overline{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star})\bm{A}^{\top}\bm{\Gamma}^{-\frac{1}{2}}=T^{\alpha-1}\bm{Y}+\bm{O}(T^{2\alpha-2}),$
in which
$\eta_{0}(\bm{\Gamma}^{-\frac{1}{2}}\bm{A}\bm{\Gamma}^{\frac{1}{2}}\bm{Y}+\bm{Y}\bm{\Gamma}^{\frac{1}{2}}\bm{A}^{\top}\bm{\Gamma}^{-\frac{1}{2}})=\bm{I}.$
Let $\lambda$ be any eigenvalue of $\bm{Y}$ and $\bm{x}$ be its corresponding eigenvector with $\|\bm{x}\|_{2}=1$. Then by definition,
11 11 $=\eta_{0}\bm{x}^{\top}(\bm{\Gamma}^{-\frac{1}{2}}\bm{A}\bm{\Gamma}^{\frac{1}{2}}\bm{Y}+\bm{Y}\bm{\Gamma}^{\frac{1}{2}}\bm{A}^{\top}\bm{\Gamma}^{-\frac{1}{2}})\bm{x}$
$=\eta_{0}\lambda\bm{x}^{\top}(\bm{\Gamma}^{-\frac{1}{2}}\bm{A}\bm{\Gamma}^{\frac{1}{2}}+\bm{\Gamma}^{\frac{1}{2}}\bm{A}^{\top}\bm{\Gamma}^{-\frac{1}{2}})\bm{x}$
$\geq\eta_{0}\lambda\cdot\lambda_{\min}(\bm{\Gamma}^{-\frac{1}{2}}\bm{A}\bm{\Gamma}^{\frac{1}{2}}+\bm{\Gamma}^{\frac{1}{2}}\bm{A}^{\top}\bm{\Gamma}^{-\frac{1}{2}})$
$\geq\frac{2(1-\gamma)\lambda_{0}\eta_{0}}{\mathsf{cond}(\bm{\Gamma}^{\frac{1}{2}})}\lambda.$
Therefore, the Frobenius norm of $\bm{Y}$ is bounded by
$\|\bm{Y}\|_{\mathsf{F}}\leq\sqrt{d}\lambda_{\max}(\bm{Y})\leq\frac{\sqrt{d\mathsf{cond}(\bm{\Gamma}^{\frac{1}{2}})}}{2(1-\gamma)\lambda_{0}\eta_{0}}T^{\alpha-1}.$
Hence the distance between $\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T})$ and $\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})$ is bounded by
$d_{\mathsf{C}}(\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T}),\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))$ $\leq\|\bm{Y}\|_{\mathsf{F}}T^{\alpha-1}+\|\bm{O}(T^{2\alpha-2})\|_{\mathsf{F}}=\frac{\sqrt{d\mathsf{cond}(\bm{\Gamma})}}{2(1-\gamma)\lambda_{0}\eta_{0}}T^{\alpha-1}+O(T^{2\alpha-2}).$
B.4.1 Tightness of Theorem 3.3.
As is indicated by Lemma A.9, the difference between $\overline{\bm{\Lambda}}_{T}$ and $\bm{\Lambda}^{\star}$ is featured by
$\overline{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star}\geq\bm{O}(T^{\alpha-1})$
when $T$ is sufficiently large. Therefore, since $\overline{\bm{\Lambda}}_{T}\succ\bm{\Lambda}^{\star}$, Proposition A.13 guarantees that
$d_{\mathsf{C}}(\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T}),\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))=d_{\mathsf{TV}}(\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T}),\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})).$
Meanwhile, Theorem A.12 indicates the TV distance between $\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T})$ and $\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))$ converges at the same rate, i.e.
$\left\|\bm{\Lambda}^{\star-\frac{1}{2}}\overline{\bm{\Lambda}}_{T}\bm{\Lambda}^{\star-\frac{1}{2}}-\bm{I}\right\|_{\mathsf{F}}=\left\|\bm{\Lambda}^{\star-\frac{1}{2}}(\overline{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star})\bm{\Lambda}^{\star-\frac{1}{2}}\right\|_{\mathsf{F}}=O(T^{\alpha-1}).$
Consequently, the $O(T^{\alpha-1})$ rate by which $\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T})$ converges to $\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})$ cannot be further improved.
B.5 Proof of Theorem 3.4
Applying the same reasoning as in the proof of Theorem 3.3,
$d_{\mathsf{TV}}(\mathcal{N}(\bm{0},\widehat{\bm{\Lambda}}_{T}),\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))$ $=d_{\mathsf{TV}}(\mathcal{N}(\bm{0},\bm{A}\widehat{\bm{\Lambda}}_{T}\bm{A}^{\top}),\mathcal{N}(\bm{0},\bm{\Gamma}))$
$\lesssim\|\bm{\Gamma}^{-\frac{1}{2}}\bm{A}(\widehat{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star})\bm{A}^{\top}\bm{\Gamma}^{-\frac{1}{2}}\|_{\mathsf{F}}$
$\leq\|\bm{\Gamma}^{-1}\|\|\bm{A}(\widehat{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star})\bm{A}^{\top}\|_{\mathsf{F}}$
In order to bound the right-hand-side of the inequality above, we firstly present the difference between $\widehat{\bm{\Lambda}}$ and $\bm{\Lambda}^{\star}$ as
$\widehat{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star}$ $=\overline{\bm{A}}^{-1}\widehat{\bm{\Gamma}}_{T}\overline{\bm{A}}^{-\top}-\bm{A}\bm{\Gamma}\bm{A}$
$=\overline{\bm{A}}^{-1}(\widehat{\bm{\Gamma}}_{T}-\bm{\Gamma})\overline{\bm{A}}^{-\top}+\left[\overline{\bm{A}}^{-1}\bm{\Gamma}\overline{\bm{A}}^{-\top}-\bm{A}^{-1}\bm{\Gamma}\bm{A}^{-\top}\right].$ (93)
The proof now boils down to bounding the estimation error of $\overline{\bm{A}}^{-1}$ and $\widehat{\bm{\Gamma}}$, for which the following lemmas come in handy.
Lemma B.4.
Let $\overline{\bm{A}}$ be defined as in (20) with $\{(s_{t},s_{t}^{\prime})\}_{0\leq t\leq T}$ being $i.i.d.$ samples. It can then be guaranteed with probability at least $1-\delta$ that
$\|(\overline{\bm{A}}^{-1}-\bm{A}^{-1})\bm{A}\|_{\mathsf{F}}\leq\frac{4\sqrt{2}}{\lambda_{0}(1-\gamma)\sqrt{T}}\sqrt{\log\frac{3}{\delta}},$
with the proviso that
$T\geq\frac{128}{\lambda_{0}(1-\gamma)}\log\frac{3}{\delta}.$ (94)
Proof: See Appendix C.3. ∎
Lemma B.5.
Let $\bm{\Gamma}$ be defined as in (16) and $\widehat{\bm{\Gamma}}_{T}$ be defined as in (21). Then the difference between $\widehat{\bm{\Gamma}}_{T}$ and $\bm{\Gamma}$ can be bounded by
$\left\|\widehat{\bm{\Gamma}}_{T}-\bm{\Gamma}\right\|_{\mathsf{F}}\leq\frac{\sqrt{\mathsf{Tr}(\bm{\Gamma})}}{\lambda_{0}(1-\gamma)}(2\|\bm{\theta}^{\star}\|_{2}+1)\|\bm{\Gamma}^{-1}\|T^{-\frac{1}{3}}+o(T^{-\frac{1}{3}}),$
with probability at least $1-2T^{-\frac{1}{3}}-\frac{\mathsf{Tr}(\bm{\Gamma})}{\lambda_{0}(1-\gamma)}T^{-\frac{1}{2}}$.
Proof: See Appendix C.4. ∎
Based on Lemmas B.4 and B.5, we now consider the estimation error of $\widehat{\bm{\Lambda}}_{T}$. By triangle inequality, (B.5) implies that
$\left\|\bm{A}(\widehat{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star})\bm{A}^{\top}\right\|_{\mathsf{F}}\leq\underset{I_{1}}{\underbrace{\|\overline{\bm{A}}^{-1}\bm{A}\|^{2}\|\widehat{\bm{\Gamma}}_{T}-\bm{\Gamma}\|_{\mathsf{F}}}}+\underset{I_{2}}{\underbrace{\left\|\bm{A}\overline{\bm{A}}^{-1}\bm{\Gamma}\overline{\bm{A}}^{-\top}\bm{A}^{\top}-\bm{\Gamma}\right\|_{\mathsf{F}}}}.$
In what follows, we bound $I_{1}$ and $I_{2}$ in order.
Bounding $I_{1}$.
The norm of $\widehat{\bm{\Gamma}}_{T}-\bm{\Gamma}$ is bounded by Lemma B.5, while the norm of $\bm{A}\overline{\bm{A}}^{-1}$ is bounded by the fact that
$\|\bm{A}\overline{\bm{A}}^{-1}\|\leq\|\bm{I}\|+\|\bm{A}(\overline{\bm{A}}^{-1}-\bm{A}^{-1})\|\leq 3$
with probability at least $1-\delta$, according to Lemma B.4 under condition (94). In combination, $I_{1}$ can be bounded by
$I_{1}\leq\left(\frac{\sqrt{\mathsf{Tr}(\bm{\Gamma})}}{\lambda_{0}(1-\gamma)}+(2\|\bm{\theta}^{\star}\|_{2}+1)\right)(2\|\bm{\theta}^{\star}\|_{2}+1)\sqrt{\frac{\log T}{T}}+o\left(\sqrt{\frac{\log T}{T}}\right)$ (95)
with probability at least $1-2T^{-\frac{1}{2}}$.
Bounding $I_{2}$.
The norm of $I_{2}$ is bounded by estimation error of $\overline{\bm{A}}^{-1}$. Namely, we further decompose $I_{2}$ by triangle inequality as
$I_{2}$ $\leq\|(\bm{A}\overline{\bm{A}}^{-1}-\bm{I})\bm{\Gamma}\|_{\mathsf{F}}+\|\bm{\Gamma}(\overline{\bm{A}}^{-\top}\bm{A}^{\top}-\bm{I})\|_{\mathsf{F}}+\|(\overline{\bm{A}}^{-1}-\bm{A}^{-1})\bm{\Gamma}(\overline{\bm{A}}^{-\top}-\bm{A}^{-\top})\|_{\mathsf{F}}$
$\leq 2\|\bm{\Gamma}\|\|\bm{A}(\overline{\bm{A}}^{-1}-\bm{A}^{-1})\|_{\mathsf{F}}+\|\bm{\Gamma}\|\|\bm{A}(\overline{\bm{A}}^{-1}-\bm{A}^{-1})\|_{\mathsf{F}}^{2},$
where we can apply Lemma B.4 to conclude that
$I_{2}\lesssim\frac{\|\bm{\Gamma}\|}{\lambda_{0}(1-\gamma)}\sqrt{\frac{\log T}{T}}.$ (96)
with probability at least $1-T^{-\frac{1}{2}}$. Theorem 3.4 follows by combining (95) and (96).
B.6 Proof of Theorem 3.5
We approach the theorem in two steps: first measuring the difference between the distribution of $\sqrt{T}\overline{\bm{\Delta}}_{T}$ and $\mathcal{N}(\bm{0},\widehat{\bm{\Lambda}}_{T})$, and then bounding the probability that $\bm{\theta}$ is within the region $\mathcal{C}_{\delta}$.
Step 1: Measuring the difference between the distribution of $\sqrt{T}\overline{\bm{\Delta}}_{T}$ and $\mathcal{N}(\bm{0},\widehat{\bm{\Lambda}}_{T})$.
We firstly observe that by triangle inequality, for any convex set $\mathcal{A}\subset\mathbb{R}^{d}$ and a $d$-dimensional standard Gaussian random variable $\bm{z}$,
$\left|\mathbb{P}(\sqrt{T}\overline{\bm{\Delta}}_{T}\in\mathcal{A})-\mathbb{P}(\widehat{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}\in\mathcal{A})\right|$ |($\leq\underset{I_{1}}{\underbrace{\left|\mathbb{P}(\sqrt{T}\overline{\bm{\Delta}}_{T}\in\mathcal{A})-\mathbb{P}(\overline{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}\in\mathcal{A})\right|}}$
+|($\quad+\underset{I_{2}}{\underbrace{\left|\mathbb{P}(\overline{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}\in\mathcal{A})-\mathbb{P}({\bm{\Lambda}}^{\star\frac{1}{2}}\bm{z}\in\mathcal{A})\right|}}+\underset{I_{3}}{\underbrace{\left|\mathbb{P}({\bm{\Lambda}}^{\star\frac{1}{2}}\bm{z}\in\mathcal{A})-\mathbb{P}(\widehat{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}\in\mathcal{A})\right|}}.$
Here, $I_{1}$ and $I_{2}$ measures the difference between the non-asymptotic and asymptotic distributions of $\overline{\bm{\Delta}}_{T}$, and are bounded by Theorems 3.2 and 3.3 respectively. Specifically,
$\left|\mathbb{P}(\sqrt{T}\overline{\bm{\Delta}}_{T}\in\mathcal{A})-\mathbb{P}({\bm{\Lambda}}^{\star\frac{1}{2}}\bm{z}\in\mathcal{A})\right|$
$\lesssim\sqrt{\frac{\eta_{0}}{(1-\alpha)(1-\gamma)\lambda_{0}}}\mathsf{Tr}(\bm{\Gamma})\|\bm{\Gamma}^{-1}\|T^{-\frac{\alpha}{2}}+\frac{\sqrt{d\mathsf{cond}(\bm{\Gamma})}}{2(1-\gamma)\lambda_{0}\eta_{0}}T^{\alpha-1}+O(T^{2\alpha-2}+T^{-\frac{1}{2}}).$ (97)
where $\bm{z}$ is a $d$-dimensional standard Gaussian random variable independent of the sample trajectory $\{(s_{t},s_{t}^{\prime})\}_{t\geq 0}$. Hence, the proof boils down to bounding $I_{3}$, which measures between the difference the centered Gaussian distribution with variance matrices $\widehat{\bm{\Lambda}}_{T}$ and $\bm{\Lambda}^{\star}$. Towards this end, we use $\mathcal{E}$ to denote the event of
$d_{\mathsf{TV}}(\mathcal{N}(\bm{0},\widehat{\bm{\Lambda}}_{T}),\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))$
$\lesssim\left(\frac{\sqrt{\mathsf{Tr}(\bm{\Gamma})}}{\lambda_{0}(1-\gamma)}+(2\|\bm{\theta}^{\star}\|_{2}+1)\right)(2\|\bm{\theta}^{\star}\|_{2}+1)\|\bm{\Gamma}^{-1}\|\sqrt{\frac{\log T}{T}}+o\left(\sqrt{\frac{\log T}{T}}\right).$
it has been illustrated in Theorem 3.4 that
$\mathbb{P}(\mathcal{E})\geq 1-3T^{-\frac{1}{2}};$
meanwhile by definition,
$\left|\mathbb{P}\left({\bm{\Lambda}}^{\star\frac{1}{2}}\bm{z}\in\mathcal{A}\right)-\mathbb{P}\left(\widehat{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}\in\mathcal{A}\mid\mathcal{E}\right)\right|$
$\lesssim\left(\frac{\sqrt{\mathsf{Tr}(\bm{\Gamma})}}{\lambda_{0}(1-\gamma)}+(2\|\bm{\theta}^{\star}\|_{2}+1)\right)(2\|\bm{\theta}^{\star}\|_{2}+1)\|\bm{\Gamma}^{-1}\|\sqrt{\frac{\log T}{T}}+o\left(\sqrt{\frac{\log T}{T}}\right).$
Therefore, since $\bm{z}$ is independent of the sample trajectory, the law of total probability implies
$\left|\mathbb{P}\left(\widehat{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}\in\mathcal{A}\right)-\mathbb{P}\left(\widehat{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}\in\mathcal{A}\mid\mathcal{E}\right)\right|$ $=\left|\mathbb{P}\left(\widehat{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}\in\mathcal{A}\mid\mathcal{E}\right)(\mathbb{P}(\mathcal{E})-1)+\mathbb{P}\left(\widehat{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}\in\mathcal{A}\mid\mathcal{E}^{\complement}\right)\mathbb{P}\left(\mathcal{E}^{\complement}\right)\right|$
$=\mathbb{P}(\mathcal{E}^{\complement})\cdot\left|\mathbb{P}\left(\widehat{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}\in\mathcal{A}\mid\mathcal{E}\right)-\mathbb{P}\left(\widehat{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}\in\mathcal{A}\mid\mathcal{E}^{\complement}\right)\right|$
$\leq\mathbb{P}(\mathcal{E}^{\complement})\leq 3T^{-\frac{1}{2}}.$
Consequently, the triangle inequality directly implies that
$\left|\mathbb{P}\left({\bm{\Lambda}}^{\star\frac{1}{2}}\bm{z}\in\mathcal{A}\right)-\mathbb{P}\left(\widehat{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}\in\mathcal{A}\right)\right|$
$\leq\left|\mathbb{P}\left({\bm{\Lambda}}^{\star\frac{1}{2}}\bm{z}\in\mathcal{A}\right)-\mathbb{P}\left(\widehat{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}\in\mathcal{A}\mid\mathcal{E}\right)\right|+\left|\mathbb{P}\left(\widehat{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}\in\mathcal{A}\right)-\mathbb{P}\left(\widehat{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}\in\mathcal{A}\mid\mathcal{E}\right)\right|$
$\leq\left(\frac{\sqrt{\mathsf{Tr}(\bm{\Gamma})}}{\lambda_{0}(1-\gamma)}+(2\|\bm{\theta}^{\star}\|_{2}+1)\right)(2\|\bm{\theta}^{\star}\|_{2}+1)\|\bm{\Gamma}^{-1}\|\sqrt{\frac{\log T}{T}}+o\left(\sqrt{\frac{\log T}{T}}\right)+3T^{-\frac{1}{2}}$
$=\left(\frac{\sqrt{\mathsf{Tr}(\bm{\Gamma})}}{\lambda_{0}(1-\gamma)}+(2\|\bm{\theta}^{\star}\|_{2}+1)\right)(2\|\bm{\theta}^{\star}\|_{2}+1)\|\bm{\Gamma}^{-1}\|\sqrt{\frac{\log T}{T}}+o\left(\sqrt{\frac{\log T}{T}}\right).$ (98)
Combining (B.6), (B.6) and taking the supremum of $\mathcal{A}$ over all convex sets, we obtain
$d_{\mathsf{C}}\left(\sqrt{T}\overline{\bm{\Delta}}_{T},\mathcal{N}(\bm{0},\widehat{\bm{\Lambda}}_{T})\right)\lesssim\frac{\sqrt{d\mathsf{cond}(\bm{\Gamma})}}{2(1-\gamma)\lambda_{0}\eta_{0}}T^{-\frac{1}{3}}+\frac{\sqrt{\mathsf{Tr}(\bm{\Gamma})}}{\lambda_{0}(1-\gamma)}(2\|\bm{\theta}^{\star}\|_{2}+1)\|\bm{\Gamma}^{-1}\|T^{-\frac{1}{3}}+o(T^{-\frac{1}{3}}).$ (99)
Step 2: Bounding $\mathbb{P}(\bm{\theta}^{\star}\in\mathcal{C}_{\delta})$.
Define $\mathcal{C}_{\delta}^{\prime}$ as
$\mathcal{C}_{\delta}^{\prime}:=\left\{y=\sqrt{T}(\overline{\bm{\theta}}_{T}-\bm{x}):\bm{x}\in\mathcal{C}_{\delta}\right\}.$
Then since $\mathcal{C}_{\delta}$ is convex, it is easy to verify that $\mathcal{C}_{\delta}^{\prime}$ is also convex. Meanwhile on one hand,
$\mathbb{P}(\bm{\theta}^{\star}\in\mathcal{C}_{\delta})=\mathbb{P}\left(\sqrt{T}(\overline{\bm{\theta}}_{T}-\bm{\theta}^{\star})\in\mathcal{C}_{\delta}^{\prime}\right);$
On the other hand,
$\mathbb{P}_{\bm{z}\sim\mathcal{N}(\bm{0},\bm{I}_{d})}\left(\overline{\bm{\theta}}_{T}-\frac{\widehat{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}}{\sqrt{T}}\in\mathcal{C}_{\delta}\bigg{|}\overline{\bm{\theta}}_{T},\widehat{\bm{\Lambda}}_{T}\right)=\mathbb{P}_{\bm{z}\sim\mathcal{N}(\bm{0},\bm{I}_{d})}\left(\widehat{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}\in\mathcal{C}_{\delta}^{\prime}\right).$
As a direct consequence, (99) and the triangle inequality indicate that
$\mathbb{P}(\bm{\theta}^{\star}\in\mathcal{C}_{\delta})$ $=\mathbb{P}\left(\sqrt{T}(\overline{\bm{\theta}}_{T}-\bm{\theta}^{\star})\in\mathcal{C}_{\delta}^{\prime}\right)$
$\geq\mathbb{P}_{\bm{z}\sim\mathcal{N}(\bm{0},\bm{I}_{d})}\left(\widehat{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}\in\mathcal{C}_{\delta}^{\prime}\right)-d_{\mathsf{C}}\left(\sqrt{T}\overline{\bm{\Delta}}_{T},\mathcal{N}(\bm{0},\widehat{\bm{\Lambda}}_{T})\right)$
$=\mathbb{P}_{\bm{z}\sim\mathcal{N}(\bm{0},\bm{I}_{d})}\left(\overline{\bm{\theta}}_{T}-\frac{\widehat{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}}{\sqrt{T}}\in\mathcal{C}_{\delta}\bigg{|}\overline{\bm{\theta}}_{T},\widehat{\bm{\Lambda}}_{T}\right)-d_{\mathsf{C}}\left(\sqrt{T}\overline{\bm{\Delta}}_{T},\mathcal{N}(\bm{0},\widehat{\bm{\Lambda}}_{T})\right)$
$\geq 1-\delta-\frac{\sqrt{\mathsf{Tr}(\bm{\Gamma})}}{\lambda_{0}(1-\gamma)}(2\|\bm{\theta}^{\star}\|_{2}+1)\|\bm{\Gamma}^{-1}\|T^{-\frac{1}{3}}-\frac{\sqrt{d\mathsf{cond}(\bm{\Gamma})}}{2(1-\gamma)\lambda_{0}\eta_{0}}T^{-\frac{1}{3}}-o(T^{-\frac{1}{3}}).$
Appendix C Proof of supportive lemmas and propositions
C.1 Proof of Lemma B.2
The lemma is proved by developing an iterative relation for the series $\{\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}\}$. By definition,
$\bm{\Delta}_{t}$ $=(\bm{\theta}_{t-1}-\bm{\theta}^{\star})-\eta_{t}(\bm{A}_{t}\bm{\theta}_{t-1}-\bm{b}_{t})$
$=(\bm{\theta}_{t-1}-\bm{\theta}^{\star})-\eta_{t}\bm{A}_{t}(\bm{\theta}_{t-1}-\bm{\theta}^{\star})-\eta_{t}(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})$
$=(\bm{I}-\eta_{t}\bm{A}_{t})\bm{\Delta}_{t-1}-\eta_{t}(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t}).$
Therefore, under $i.i.d.$ samples, we can relate $\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}$ with $\mathbb{E}\|\bm{\Delta}_{t-1}\|_{2}^{2}$ by
$\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}$ $=\mathbb{E}\|(\bm{I}-\eta_{t}\bm{A}_{t})\bm{\Delta}_{t-1}\|_{2}^{2}+\eta_{t}\mathbb{E}\langle(\bm{I}-\eta_{t}\bm{A}_{t})\bm{\Delta}_{t-1},\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t}\rangle+\eta_{t}^{2}\mathbb{E}\|\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t}\|_{2}^{2}$
$=\mathbb{E}\|(\bm{I}-\eta_{t}\bm{A}_{t})\bm{\Delta}_{t-1}\|_{2}^{2}+\eta_{t}^{2}\mathbb{E}\langle\bm{A}_{t}\bm{\Delta}_{t-1},\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t}\rangle+\eta_{t}^{2}\mathbb{E}\|\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t}\|_{2}^{2}$
$=\mathbb{E}\|(\bm{I}-\eta_{t}\bm{A}_{t})\bm{\Delta}_{t-1}\|_{2}^{2}+\eta_{t}^{2}\left(\frac{1}{2}\mathbb{E}\|\bm{A}_{t}(\bm{\theta}_{t-1}-\bm{\theta}^{\star})\|+\mathbb{E}\|\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t}\|\right)$
$+\eta_{t}^{2}\mathbb{E}\|\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t}\|_{2}^{2}$
$=\|\bm{\Delta}_{t-1}\|_{2}^{2}-\eta_{t}\bm{\Delta}_{t-1}^{\top}(\bm{A}+\bm{A}^{\top})\bm{\Delta}_{t-1}$
$+\frac{3}{2}\eta_{t}^{2}\bm{\Delta}_{t-1}^{\top}\mathbb{E}[\bm{A}_{t}^{\top}\bm{A}_{t}]\bm{\Delta}_{t-1}+\frac{3}{2}\eta_{t}^{2}\mathbb{E}\|\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t}\|_{2}^{2}$
$\leq\|\bm{\Delta}_{t-1}\|_{2}^{2}-(\eta_{t}-\frac{3}{2}\eta_{t}^{2})\bm{\Delta}_{t-1}^{\top}(\bm{A}+\bm{A}^{\top})\bm{\Delta}_{t-1}+\frac{3}{2}\eta_{t}^{2}\mathbb{E}\|\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t}\|_{2}^{2}.$
Notice that we applied Lemma A.3 in the last inequality. Now, since the stepsize $\eta_{t}$ satisfies $\eta_{t}\leq\eta_{0}\leq\frac{1}{2}$, Lemma A.3 implies
$\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}$ $\leq\mathbb{E}\|\bm{\Delta}_{t-1}\|_{2}^{2}-\frac{1}{4}\eta_{t}\mathbb{E}[(\bm{\theta}_{t-1}-\bm{\theta}^{\star})^{\top}(\bm{A}+\bm{A}^{\top})(\bm{\theta}_{t-1}-\bm{\theta}^{\star})]+\frac{3}{2}\eta_{t}^{2}\mathbb{E}\|\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t}\|_{2}^{2}$
$<\left(1-\frac{1-\gamma}{2}\lambda_{0}\eta_{t}\right)\mathbb{E}\|\bm{\Delta}_{t-1}\|_{2}^{2}+\frac{3}{2}\eta_{t}^{2}\mathsf{Tr}(\bm{\Gamma}).$
Therefore by induction, $\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}$ is bounded by
$\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}<\prod_{i=1}^{t}\left(1-\frac{1-\gamma}{2}\lambda_{0}\eta_{i}\right)\|\bm{\Delta}_{0}\|_{2}^{2}+\frac{3}{2}\sum_{i=1}^{n}\eta_{i}^{2}\prod_{j=i+1}^{n}\left(1-\frac{1-\gamma}{2}\lambda_{0}\eta_{j}\right)\mathsf{Tr}(\bm{\Gamma}).$ (100)
The proof now boils down to bounding the two terms on the right hand side of (100). For the first term, we observe
$\prod_{i=1}^{t}\left(1-\frac{1-\gamma}{2}\lambda_{0}\eta_{i}\right)$ $<\exp\left(\frac{1-\gamma}{2}\lambda_{0}\eta_{0}\sum_{i=1}^{t}i^{-\alpha}\right)$
$\lesssim\exp\left(-\frac{1-\gamma}{2(1-\alpha)}\lambda_{0}\eta_{0}t^{1-\alpha}\right)=o(t^{-1});$ (101)
for the second term, we can apply Lemma A.4 with $\beta=\frac{1-\gamma}{2}\lambda_{0}\eta_{0}$ to show that
$\sum_{i=1}^{n}\eta_{i}^{2}\prod_{j=i+1}^{n}\left(1-\frac{1-\gamma}{2}\lambda_{0}\eta_{j}\right)=\frac{2\eta_{0}}{(1-\gamma)\lambda_{0}}t^{-\alpha}+O(t^{-1}).$ (102)
The Lemma follows immediately by plugging (C.1) and (102) into (100).
C.2 Proof of Proposition B.3
We address the two properties in order.
Proof of Equation (86a).
Recall from (9) and (82) that for any $i\in[T]$,
$\bm{\theta}_{i}=\bm{\theta}_{i-1}-\eta_{i}(\bm{A}_{i}\bm{\theta}_{i-1}-\bm{b}_{i});\quad\text{and}$
$\bm{\theta}_{i}^{(i)}=\bm{\theta}_{i-1}-\eta_{i}(\check{\bm{A}}_{i}\bm{\theta}_{i-1}-\check{\bm{b}}_{i}).$
Therefore, the difference between $\bm{\theta}_{i}$ and $\bm{\theta}_{i}^{(i)}$ can be bounded by
$\mathbb{E}\|\bm{\theta}_{i}-\bm{\theta}_{i}^{(i)}\|_{2}^{2}$ $\leq\eta_{i}^{2}\mathbb{E}\|(\bm{A}_{i}-\check{\bm{A}}_{i})\bm{\theta}_{i-1}-(\bm{b}_{i}-\check{\bm{b}}_{i})\|_{2}^{2}$
$=\eta_{i}^{2}\mathbb{E}\|(\bm{A}_{i}-\bm{A})\bm{\theta}^{\star}+(\bm{A}_{i}-\bm{A})\bm{\Delta}_{i-1}-(\bm{b}_{i}-\bm{b})\|_{2}^{2}$
$\leq\eta_{i}^{2}(2\mathbb{E}\|(\bm{A}_{i}-\bm{A})\bm{\Delta}_{i-1}\|_{2}^{2}+2\mathbb{E}\|(\bm{A}_{i}-\bm{A})\bm{\theta}^{\star}-(\bm{b}_{i}-\bm{b})\|_{2}^{2})$
$\leq 2\eta_{i}^{2}\mathsf{Tr}(\bm{\Gamma})+2\eta_{i}^{2}\mathbb{E}\|\bm{\Delta}_{i-1}\|_{2}^{2}.$
Proof of Equation (86b).
Recall from (9) and (82) that
$\bm{\theta}_{j}=\bm{\theta}_{j-1}-\eta_{j}(\bm{A}_{j}\bm{\theta}_{j-1}-\bm{b}_{j});$
$\bm{\theta}_{j}^{(i)}=\bm{\theta}_{j-1}^{(i)}-\eta_{j}(\bm{A}_{j}\bm{\theta}_{j-1}^{(i)}-\bm{b}_{j}).$
Therefore, the difference between $\bm{\theta}_{j}$ and $\bm{\theta}_{j}^{(i)}$ can be expressed as
$\mathbb{E}\|\bm{\theta}_{j}-\bm{\theta}_{j}^{(i)}\|_{2}^{2}$ $=\mathbb{E}\|(\bm{I}-\eta_{j}\bm{A}_{j})(\bm{\theta}_{j-1}-\bm{\theta}_{j-1}^{(i)})\|_{2}^{2}$
$\leq\left(1-\frac{1}{2}(1-\gamma)\lambda_{0}\eta_{j}\right)\mathbb{E}\|\bm{\theta}_{j-1}-\bm{\theta}_{j-1}^{(i)}\|_{2}^{2},$
where the inequality follows from Lemma A.3. By induction, we obtain
$\mathbb{E}\|\bm{\theta}_{j}-\bm{\theta}_{j}^{(i)}\|_{2}^{2}$ $\leq\prod_{k=i+1}^{j}\left(1-\frac{1}{2}(1-\gamma)\lambda_{0}\eta_{k}\right)\mathbb{E}\|\bm{\theta}_{i}-\bm{\theta}_{i}^{(i)}\|_{2}^{2}.$
C.3 Proof of Lemma B.4
We firstly observe that since $\|\bm{\phi}(s)\|_{2}\leq 1$ for all $s\in\mathcal{S}$, it can be guaranteed that
$\|\textbf{vec}(\bm{A}_{t})\|_{2}^{2}=\|\bm{A}_{t}\|_{\mathsf{F}}^{2}$ $=\mathsf{Tr}(\bm{A}_{t}^{\top}\bm{A}_{t})$
$=\mathsf{Tr}\left([\bm{\phi}(s_{t})-\gamma\bm{\phi}(s_{t}^{\prime})]\bm{\phi}(s_{t})^{\top}\bm{\phi}(s_{t})[\bm{\phi}(s_{t})-\gamma\bm{\phi}(s_{t}^{\prime})]^{\top}\right)$
$\leq\mathsf{Tr}\left([\bm{\phi}(s_{t})-\gamma\bm{\phi}(s_{t}^{\prime})][\bm{\phi}(s_{t})-\gamma\bm{\phi}(s_{t}^{\prime})]^{\top}\right)<4$
holds uniformly for all $0\leq t\leq T$. Hence the vector Azuma’s inequality A.16 directly implies that
$\|\overline{\bm{A}}-\bm{A}\|_{\mathsf{F}}$ $=\left\|\mathbf{vec}(\overline{\bm{A}}-\bm{A})\right\|\leq\frac{4\sqrt{2}}{\sqrt{T}}\sqrt{\log\frac{3}{\delta}}$
with probability at least $1-\delta$. In what follows, we aim to transform this error bound to the error bound for $\overline{\bm{A}}^{-1}$. Towards this end, we observe that condition 94 guarantees with probability at least $1-\delta$,
$\left\|\bm{A}^{-1}(\overline{\bm{A}}-\bm{A})\right\|$ $\leq\|\bm{A}^{-1}\|\|\overline{\bm{A}}-\bm{A}\|$
$\leq\lambda_{0}^{-1}(1-\gamma)^{-1}\frac{4\sqrt{2}}{\sqrt{T}}\sqrt{\log\frac{3}{\delta}}<\frac{1}{2}$
where the last inequality is guaranteed by the condition (94). Hence, the condition of Lemma A.2 is satisfied with probability at least $1-\delta$ by taking $\bm{Y}=\bm{A}^{-1}\overline{\bm{A}}$ and $\bm{X}=\bm{I}$. Consequently, the difference between $\overline{\bm{A}}^{-1}\bm{A}$ and $\bm{I}$ is bounded by
$\|(\overline{\bm{A}}^{-1}-\bm{A}^{-1})\bm{A}\|_{\mathsf{F}}\leq 2\|(\overline{\bm{A}}-\bm{A})\bm{A}^{-1}\|_{\mathsf{F}}$ $\leq\frac{4\sqrt{2}}{\sqrt{T}}\sqrt{\log\frac{3}{\delta}}\|\bm{A}^{-1}\|$
$\leq\frac{4\sqrt{2}}{\lambda_{0}(1-\gamma)\sqrt{T}}\sqrt{\log\frac{3}{\delta}}.$
C.4 Proof of Lemma B.5
We firstly represent $\widehat{\bm{\Gamma}}_{T}-\bm{\Gamma}$ as
$\widehat{\bm{\Gamma}}_{T}-\bm{\Gamma}$ $=\underset{I_{1}}{\underbrace{\frac{1}{T}\sum_{t=1}^{T}\left[(\bm{A}_{t}\overline{\bm{\theta}}_{T}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\theta}}_{T}-\bm{b}_{t})^{\top}-(\bm{A}_{t}{\bm{\theta}}^{\star}-\bm{b}_{t})(\bm{A}_{t}{\bm{\theta}^{\star}}-\bm{b}_{t})^{\top}\right]}}$
$+\underset{I_{2}}{\underbrace{\frac{1}{T}\sum_{t=1}^{T}\left[(\bm{A}_{t}{\bm{\theta}}^{\star}-\bm{b}_{t})(\bm{A}_{t}{\bm{\theta}^{\star}}-\bm{b}_{t})^{\top}]-\mathbb{E}[(\bm{A}_{t}{\bm{\theta}}^{\star}-\bm{b}_{t})(\bm{A}_{t}{\bm{\theta}^{\star}}-\bm{b}_{t})^{\top}\right]}},$
where $I_{1}$ is mainly controlled by the estimation error of $\overline{\bm{\theta}}_{T}$, and $I_{2}$ only involves the randomness of $\{(\bm{A}_{t},\bm{b}_{t})\}$. We now control the norms of these two terms in order.
Bounding the Frobenius norm of $I_{1}$.
By definition, $I_{1}$ can be further decomposed as
$I_{1}$ $=\frac{1}{T}\sum_{t=1}^{T}\left[(\bm{A}_{t}\overline{\bm{\theta}}_{T}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\theta}}_{T}-\bm{b}_{t})^{\top}-(\bm{A}_{t}{\bm{\theta}}^{\star}-\bm{b}_{t})(\bm{A}_{t}{\bm{\theta}^{\star}}-\bm{b}_{t})^{\top}\right]$
$=\underset{I_{11}}{\underbrace{\frac{1}{T}\sum_{t=1}^{T}\left[(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}\right]}}+\underset{I_{12}}{\underbrace{\frac{1}{T}\sum_{t=1}^{T}\left[(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}+(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}\right]}}.$
By triangle inequality, the Frobenius norm of $I_{11}$ is bounded by
$\left\|\frac{1}{T}\sum_{t=1}^{T}\left[(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}\right]\right\|_{\mathsf{F}}$ $\leq\frac{1}{T}\sum_{t=1}^{T}\left\|(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}\right\|_{\mathsf{F}}$
$\leq\frac{1}{T}\sum_{t=1}^{T}\sqrt{\mathsf{Tr}\left((\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}\right)}$
$\leq\frac{1}{T}\sum_{t=1}^{T}4\|\overline{\bm{\Delta}}_{T}\|_{2}^{2}=4\|\overline{\bm{\Delta}}_{T}\|_{2}^{2}.$ (103)
Since
$\mathbb{E}\|\overline{\bm{\Delta}}_{T}\|_{2}^{2}$ $=\frac{1}{T}\mathsf{Tr}(\overline{\bm{\Lambda}}_{T})+\widetilde{C}T^{-\alpha-1}=\frac{1}{T}\mathsf{Tr}(\bm{\Lambda}^{\star})+o(T^{-1})\leq\frac{\mathsf{Tr}(\bm{\Gamma})}{\lambda_{0}^{2}(1-\gamma)^{2}T}+o(T^{-1}),$
the Markov’s inequality indicates
$\left\|\frac{1}{T}\sum_{t=1}^{T}\left[(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}\right]\right\|_{\mathsf{F}}\leq\frac{\sqrt{\mathsf{Tr}(\bm{\Gamma})}}{\lambda_{0}(1-\gamma)\sqrt{T}}+o(T^{-\frac{1}{2}})$ (104)
with probability at least $1-\frac{\sqrt{\mathsf{Tr}(\bm{\Gamma})}}{\lambda_{0}(1-\gamma)\sqrt{T}}$.
Meanwhile, the norm of $I_{12}$ is bounded by
$\left\|\frac{1}{T}\sum_{t=1}^{T}\left[(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}+(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}\right]\right\|_{\mathsf{F}}$
$\quad\leq\frac{1}{T}\sum_{t=1}^{T}\left\|(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}+(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}\right\|_{\mathsf{F}}$
$\quad\leq\frac{2}{T}\sum_{t=1}^{T}\left\|(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}\right\|_{\mathsf{F}}$
$\quad\leq\frac{2}{T}\sum_{t=1}^{T}\sqrt{\mathsf{Tr}\left((\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}\right)}$
$\quad\leq\frac{4}{T}\sum_{t=1}^{T}4(2\|\bm{\theta}^{\star}\|_{2}+1)\|\overline{\bm{\Delta}}_{T}\|_{2}$
$\quad=4(2\|\bm{\theta}^{\star}\|_{2}+1)\|\overline{\bm{\Delta}}_{T}\|_{2}.$ (105)
Therefore, the Chebyshev’s inequality indicates
$\left\|\frac{1}{T}\sum_{t=1}^{T}\left[(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}+(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}\right]\right\|_{\mathsf{F}}$
$\leq 4(2\|\bm{\theta}^{\star}\|_{2}+1)\frac{\sqrt{\mathsf{Tr}(\bm{\Gamma})}}{\lambda_{0}(1-\gamma)}T^{-\frac{1}{3}}+o(T^{-\frac{1}{3}})$ (106)
with probability at least $1-T^{-\frac{1}{3}}$.
Hence by combining (104) and (C.4), we obtain
$\|I_{1}\|_{\mathsf{F}}\lesssim 4(2\|\bm{\theta}^{\star}\|_{2}+1)\frac{\sqrt{\mathsf{Tr}(\bm{\Gamma})}}{\lambda_{0}(1-\gamma)}T^{-\frac{1}{3}}+o(T^{-\frac{1}{3}})$ (107)
with probability at least $1-T^{-\frac{1}{3}}-\frac{\sqrt{\mathsf{Tr}(\bm{\Gamma})}}{\lambda_{0}(1-\gamma)\sqrt{T}}$.
Bounding the Frobenius norm of $I_{2}$.
Notice that $I_{2}$ only involves the randomness of $i.i.d.$ samples $\{(\bm{A}_{t},\bm{b}_{t})\}$, and
$\left\|\textbf{vec}\left((\bm{A}_{t}{\bm{\theta}}^{\star}-\bm{b}_{t})(\bm{A}_{t}{\bm{\theta}^{\star}}-\bm{b}_{t})^{\top}\right)\right\|_{2}$
$=\left\|(\bm{A}_{t}{\bm{\theta}}^{\star}-\bm{b}_{t})(\bm{A}_{t}{\bm{\theta}^{\star}}-\bm{b}_{t})^{\top}\right\|_{\mathsf{F}}$
$=\sqrt{\mathsf{Tr}((\bm{A}_{t}{\bm{\theta}}^{\star}-\bm{b}_{t})(\bm{A}_{t}{\bm{\theta}^{\star}}-\bm{b}_{t})^{\top}(\bm{A}_{t}{\bm{\theta}}^{\star}-\bm{b}_{t})(\bm{A}_{t}{\bm{\theta}^{\star}}-\bm{b}_{t})^{\top})}$
$<(2\|\bm{\theta}^{\star}\|_{2}+1)^{2}$
holds uniformly for all $t\in[T]$. Consequently, the Frobenius norm of $I_{2}$ can be bounded by directly applying the vector Azuma’s inequality: for every $\delta\in(0,1)$,
$\left\|I_{2}\right\|_{\mathsf{F}}\leq\frac{2\sqrt{2}}{\sqrt{T}}(2\|\bm{\theta}^{\star}\|_{2}+1)^{2}\sqrt{\log\frac{3}{\delta}}$ (108)
with probability at least $1-\delta$. Lemma B.5 follows by combining (107) and (108) with $\delta=T^{-\frac{1}{3}}$.∎
Appendix D Experimental details
D.1 The Markov decision process and the feature map
Given a finite state space $\mathcal{S}$, define a sequence of MDP $\{\mathcal{M}_{\bm{q}}\}$ indexed by $\bm{q}\in\mathcal{Q}\subset\{q_{+},q_{-}\}^{d-1}$ where for each $\bm{q}$, the transition kernel equals to
$P_{\bm{q}}(s^{\prime}\mid s)$
=$=\left\{\begin{array}[]{lcc}q_{s}\mathds{1}(s^{\prime}=s)+\frac{1-q_{s}}{|\mathcal{S}|-d+1}\mathds{1}(s^{\prime}\geq d)&\text{for}&s<d;\\[5.69046pt] \frac{\gamma}{|\mathcal{S}|-d+1}\mathds{1}(s^{\prime}\geq d)+\frac{1-q_{s^{\prime}}}{d-1}\mathds{1}(s^{\prime}<d)&\text{for}&s\geq d.\end{array}\right.$ (111)
and the reward function equals to $r(s)=\mathds{1}(s\geq d)$.
Here, for each $i\in[d-1]$, $q_{i}$ is taken to be either $q_{+}$ or $q_{-}$ where
$q_{+}:=\gamma+(1-\gamma)^{2}\varepsilon,\qquad\text{and}\qquad q_{-}:=\gamma-(1-\gamma)^{2}\varepsilon.$
We further impose the constraint that the number of $q_{+}$’s and $q_{-}$’s in $\bm{q}$ are the same, namely,
$\sum_{s=1}^{d-1}\mathds{1}(q_{s}=q_{+})=\sum_{s=1}^{d-1}\mathds{1}(q_{s}=q_{-})=(d-1)/2.$ (112)
Here without loss of generality, assume $d$ is an odd number. With these definitions in place, it can be easily verified that the stationary distribution for $\bm{P}$ obeys
$\mu(s)=\left\{\begin{array}[]{lcc}\frac{1}{2(d-1)}&\text{for}&s<d;\\[5.69046pt] \frac{1}{2(|\mathcal{S}|-d+1)}&\text{for}&s\geq d.\end{array}\right.$ (115)
Moreover, suppose the feature map is taken to be
$\bm{\phi}(s)=\bm{e}_{s\wedge d}\in\mathbb{R}^{d},$
then one can further verify that
$\theta^{\star}(d)$ $=V^{\star}(s)=\frac{1}{1-\gamma^{2}-\sum_{i=1}^{d-1}\frac{\gamma^{2}(1-q_{i})^{2}}{(d-1)(1-\gamma q_{i})}},$
$\theta^{\star}(i)$ $=V^{\star}(i)=\frac{\gamma(1-q_{i})}{1-\gamma q_{i}}V^{\star}(s),\text{ for }s\geq d\text{ and }i<d.$ (116)
From the expressions above, we remark that, the values of $q$ and $V^{\star}(s)$ with $s\geq d$ are fixed for all $\bm{q}\in\mathcal{Q}$ which is ensured by the construction (112). It is shown in Theorem 2 of Li et al., 2024b that the sequence $\{\mathcal{M}_{\bm{q}}\}_{q\in\mathcal{Q}}$ is hard to distinguish for a considerably large set $\mathcal{Q}$.
D.2 Parameter specification
In our experiments, we fix the size of the state space to $\|\mathcal{S}\|=10$, the discount factor to be $\gamma=0.2$, and let $\varepsilon=0.01$. For the TD learning algorithm, we set the initial stepsize to be $\eta_{0}=5$, and total number of iterations to be $T=10^{5}$. Data is recorded every $100$ iterations. In order to show the role played by the stepsize decaying speed $\alpha$, we fix $d=3$ and $\alpha$ iterate through the group $(\frac{1}{2},\frac{2}{3},\frac{3}{4},\frac{4}{5})$; and in order to illustrate the influence of the dimension of the feature map, we fix $\alpha=\frac{2}{3}$ and let $d$ iterate through the group of $(3,5,7,9)$. For every given $d$, we set $\bm{q}$ to be
$q_{i}=\begin{cases}q_{+},\quad\text{for }1\leq i\leq\frac{d-1}{2};\\ q_{-},\quad\text{for }\frac{d+1}{2}\leq i\leq d.\end{cases}$
We verify our theoretical results through the following:
  1. 1.
    In order to illustrate the high-probability convergence of the TD estimation error as indicated by Theorem 3.1, we measure the $95\%$ quantile of the $L_{2}$ norm of $\overline{\bm{\Delta}}_{T}$ in $10,000$ independent trials;
  2. 2.
    In order to illustrate the Berry-Esseen bound as indicated by Theorems 3.2 and 3.3, we measure the difference between the empirical and asymptotic distributions of $\overline{\bm{\Delta}}_{T}$ by
    $d(\sqrt{T}\overline{\bm{\Delta}}_{T},\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})):=\max_{1\leq j\leq d}\sup_{x\in\mathbb{R}}\left|\mathbb{P}(\sqrt{T}(\overline{\bm{\Delta}}_{T})_{j}\leq x)-\mathbb{P}((\bm{\Lambda}^{\star\frac{1}{2}}\bm{z})_{j}\leq x)\right|,$
    averaged among $10,000$ independent trials. Besides, as an intuitive demonstration, Figure 1 in the introduction section records the empirical and asymptotic distributions of $(\overline{\bm{\Delta}}_{T})_{1}$ when $d=3$ and $\alpha=\frac{2}{3}$.
  3. 3.
    In order to illustrate the efficiency of our variance esitmator $\widehat{\bm{\Lambda}}_{T}$, we measure the averaged Frobenius norm of the difference between $\widehat{\bm{\Lambda}}_{T}$ and $\bm{\Lambda}^{\star}$ among $100$ independent trials. As is indicated by A.12, this is a direct proxy for the TV distance between the zero-mean Gaussian distributions with the two corresponding variance matrices.
  4. 4.
    In order to validate our statistical inference procedure, we run Algorithm 1 on the MDP with $d=3$ using $\alpha=\frac{2}{3}$, construct both individual and simultaneous confidence intervals for $\bm{\theta}^{\star}$, and measure the coverage rates of these confidence intervals in $1000$ independent trials. Results are shown in section 4.2.
D.3 Full experimental results
D.3.1 High-probability convergence guarantee for $\overline{\bm{\Delta}}_{T}$
Figure 5 shows the decay of the $95\%$ quantile of the $L_{2}$ norm of $\overline{\bm{\Delta}}_{T}$; the slope of this log-log plot verifies our theoretical claim in Theorem 3.1 that $\|\overline{\bm{\Delta}}_{T}\|_{2}$ converges by the rate of $O(T^{-\frac{1}{2}})$ with high probability. Plot (a) shows that the convergence rate is similar for different choices of $\alpha>\frac{1}{2}$ when $T$ is large, while plot(b) shows that the convergence is slower when $d$ is larger. Both are in accordance with our conclusion in Theorem (1) that $\|\overline{\bm{\Delta}}_{T}\|_{2}$ converges by the rate of
$O\left(\sqrt{\frac{\mathsf{Tr}(\bm{\Lambda}^{\star})}{T}\log\frac{d}{\delta}}\right)$
with probability at least $1-\delta$.
Refer to caption Refer to caption
Figure 5: The decay of the $95\%$ quantile of $\|\overline{\bm{\Delta}}_{T}\|_{2}^{2}$ when (a) $d=3$, for different choices of $\alpha$; and (b) $\alpha=\frac{2}{3}$, for different choices of $d$.
D.3.2 Berry-Esseen bound
Figure 6 illustrates how the finite-sample distribution of $\sqrt{T}\overline{\bm{\Delta}}_{T}$ converges to its asymptotic Gaussian distribution $\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})$. Notably, plot (a), which is a replication of Figure 2(b) in Section 4, indicates that the convergence is slower when $\alpha=\frac{1}{2}$. This directly contradicts the suggestion by Samsonov et al., (2024) to choose $\alpha=\frac{1}{2}$ in order to yield a faster in-distribution convergence of the TD estimation error. In fact, with $d=3$ fixed and other choices of $\alpha$, the convergence rate are similar when $T$ is large. Meanwhile, plot (b) shows the convergence generally is slower when $d$ is larger.
Refer to caption Refer to caption
Figure 6: The decay of $d(\sqrt{T}\overline{\bm{\Delta}}_{T},\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))$, when (a) $d=3$, for different choices of $\alpha$; and (b) $\alpha=\frac{2}{3}$, for different choices of $d$.
D.3.3 Estimator of the covariance matrix
Figure 7 illustrates the rate by which our variance estimator $\widehat{\bm{\Lambda}}_{T}$ converges to the asymptotic variance $\bm{\Lambda}^{\star}$. Plot (a) shows that the convergence rate is similar for different choices of $\alpha$ when the number of interations $T$ is sufficiently large; plot(b) indicates that the convergence is slower when $d$ is larger, by the nearly parallel curves in the log-log plot demonstrate that the convergence rate with respect to $T$ is similar among different choices of $d$.
Refer to caption Refer to caption
Figure 7: The decay of the difference bewteen estimated and asymptotic covariance matrices, measured in squared Frobenius norm, when (a) $d=3$, for different choices of $\alpha$; and (b) $\alpha=\frac{2}{3}$, for different choices of $d$.

### v3

@@
-Statistical Inference for Temporal Difference Learning with Linear Function Approximation
+Statistical Inference for Policy Evaluation with Temporal Difference Learning
 Abstract
-Statistical inference with finite-sample validity for the value function of a given policy in Markov decision processes (MDPs) is crucial for ensuring the reliability of reinforcement learning.
-Temporal Difference (TD) learning, arguably the most widely used algorithm for policy evaluation, serves as a natural framework for this purpose.
-In this paper, we study the consistency properties of TD learning with Polyak-Ruppert averaging and linear function approximation, and obtain three significant improvements over existing results.
-First, we derive a novel sharp high-dimensional probability convergence guarantee that depends explicitly on the asymptotic variance and holds under weak conditions.
-We further establish refined high-dimensional Berry-Esseen bounds over the class of convex sets that guarantee faster rates than those in the literature.
-Finally, we propose a plug-in estimator for the asymptotic covariance matrix, designed for efficient online computation.
-These results enable the construction of confidence regions and simultaneous confidence intervals for the linear parameters of the value function, with guaranteed finite-sample coverage.
+We investigate the statistical properties of Temporal Difference (TD) learning with Polyak-Ruppert averaging, arguably one of the most widely used algorithms in reinforcement learning, for the task of estimating the parameters of the optimal linear approximation to the value function.
+We make three significant contributions that improve the current state-of-the-art results: (i) we derive sharper high probability convergence guarantee that depend explicitly on the asymptotic variance and hold under weaker conditions than those normally assumed; (ii) we establish refined high-dimensional Berry-Esseen bounds over the class of convex sets, achieving faster rates than those previously established in the literature, and (iii) we propose and analyze a novel, computationally efficient online plug-in estimator of the asymptotic covariance matrix.
+These results enable the construction of confidence regions and simultaneous confidence intervals for the linear parameters of the value function approximation, with guaranteed finite-sample coverage.
 We demonstrate the applicability of our theoretical findings through numerical experiments.
 1 Introduction
-Statistical inference tasks, such as constructing confidence regions or simultaneous confidence intervals, are often addressed by deriving distributional theory such as central limit theorems (CLTs) for the estimator in use.
-Due to the high dimensionality of modern science and engineering applications, there has been a surge of interests in deriving convergence results that are valid in a finite-sample manner.
-In Reinforcement Learning (RL), a discipline that underpins many recent machine learning breakthroughs (Sutton and Barto, (2018); Agarwal et al., (2019)) one central question is to evaluate with confidence the quality of a given policy, measured by its value function.
-As RL is often modeled as decision making in Markov decision processes (MDPs), the task of statistical inference needs to accommodate the online nature of the sampling mechanism.
-Temporal Difference (TD) learning (Sutton, (1988)) is arguably the most widely used algorithm designed for this purpose.
-TD learning, which is an instance of stochastic approximation (SA) method (Robbins and Monro, (1951)), approximates the value function of a given policy in an iterative manner.
-Towards understanding the non-asymptotic convergence rate of TD to the target value function, extensive recent efforts have been made (see, e.g.
-Nemirovski et al., (2009); Rakhlin et al., (2012); Li et al., 2024a ; Bhandari et al., (2021); Dalal et al., (2018); Khamaru et al., (2021); Srikant and Ying, (2019); Lakshminarayanan and Szepesvari, (2018); Mou et al., (2020); Durmus et al., (2022); Patil et al., (2023); Li et al., 2024b ; Samsonov et al., (2023) and references therein) and it is known that for the tabular case, where both state and action spaces are assumed to be finite, the TD algorithm is minimax optimal (Li et al., 2024a ).
-Another thread of work considers the asymptotic performance of TD where it has been shown that with Polyak-Ruppert averaging (Polyak and Juditsky, (1992); Ruppert, (1998)), TD converges to a Gaussian distribution in the large sample limit (Fort, (2015); Mou et al., (2020); Li et al., (2023)).
-However, these two types of results remain insufficient for conducting statistical inference that is valid in the finite-sample fashion.
-Figure 1: Emergence of the first entry of TD estimation error on an MDP with a 3-dimensional feature map.
-To close this gap, in this paper, based on the TD learning algorithm, we design confidence regions and simultaneous confidence intervals for the linear coefficients of the value function when the it is approximated by a linear combination of features.
-To achieve this goal, we first provide a novel sharp high-dimensional probability convergence guarantee for the TD iterates, and follow up with refined high-dimensional Berry-Esseen bounds over the class of convex sets.
-Finally, we propose a plug-in estimator for the asymptotic covariance matrix, improving upon other alternatives proposed by recent literature (Chen et al., (2020), White and White, (2010), Hanna et al., (2017), Hao et al., (2021), Ramprasad et al., (2022)) in both computational and statistical efficiency.
-We summarize our main contributions in details in the following section.
-While we are writing this paper, we become aware of a great recent paper by Samsonov et al., (2024), which considers similar problems to ours and derives a Berry-Esseen bound for linear SA with Polyak-Ruppert averaging.
-It remains unclear whether their general framework gives tight results when specialized to TD learning.
-We offer more detailed comparisons between our results and theirs in the following sections.
-1.1 Contributions
-Focusing on TD learning with linear function approximation, polynomial-decaying stepsize, Polyak-Ruppert averaging and $i.i.d.$ samples, this paper makes the following theoretical and methodological contributions.
+The feasibility and accuracy of statistical inference tasks, such as hypothesis testing or constructing confidence regions or simultaneous confidence intervals, are traditionally predicated on using estimators that are not only consistent but also satisfy some type of distributional approximation.
+In most cases, this is accomplished via Central Limit Theorems (CLTs), which ensure that the procedures in use admit a Gaussian approximation.
+Due to the high dimensionality of modern scientific and engineering applications, there has been a surge of interest in establishing new high-dimensional Berry-Esseen bounds, CLTs that hold in high dimensions and have a finite-sample guarantee.
+While efforts in this area have focused on Berry-Esseen bounds for normalized sums, few results are available to handle more complicated scenarios, where estimators are computed in a nonlinear, iterative manner.
+In this paper, we investigate one such scenario and study in detail the consistency, Gaussian approximations, and statistical inference for the output of the Temporal Difference (TD) learning algorithm for policy evaluation with linear function approximation in Reinforcement Learning (RL).
+A critical task in RL is to evaluate with confidence the quality of a given policy, measured by its value function.
+As RL models are based on Markov Decision Processes (MDPs), statistical inference must accommodate the online nature of the sampling mechanism.
+TD learning (see, e.g., Sutton, 1988), an iterative procedure that can be framed as an instance of stochastic approximation (SA), is arguably the most widely used and effective algorithm for policy evaluation in RL.
+The last several years have witnessed a flurry of activities within the statistical machine learning community to derive finite sample consistency rates of TD iterates, leading to deep results and a much improved understanding of the performance and limitations TD learning.
+See, e.g., Khamaru et al.
+(2020); Li et al.
+(2024); Bhandari et al.
+(2021); Dalal et al.
+(2018); Srikant and Ying (2019); Lakshminarayanan and Szepesvari (2018); Mou et al.
+(2020); Durmus et al.
+(2022); Li et al.
+(2024); Samsonov et al.
+(2024); Prashanth et al.
+(2013); Borkar et al.
+(2025) for a partial and growing list of relevant references.
+Though for the tabular case (finite state and action spaces) the TD algorithm is known to be minimax rate optimal Li et al.
+(2024), its properties – and, in particular, its sample complexity – in more complex and high-dimensional settings have not been fully understood.
+paper stepsize convergence guarantee constraints on $t,T,\delta$
+[…]
+Li et al. (2024) $\eta_{t}=\eta$ $\|\overline{\bm{\Delta}}_{T}\|_{2}\leq O\left(T^{-\frac{1}{2}}\log^{\frac{1}{2}}\frac{T}{\delta}\right)$ $\eta\sqrt{\log\frac{T}{\delta}}\leq\tilde{C}$
+Samsonov et al. (2024) $\eta_{t}=\eta$ $\|\overline{\bm{\Delta}}_{T}\|_{2}\leq O\left(T^{-\frac{1}{2}}\log\frac{T}{\delta}\right)$ $\eta\log\frac{T}{\delta}\leq\tilde{C}$
+Theorem 3.1 $\eta_{t}=\eta_{0}t^{-\alpha},\alpha\in\left(\frac{1}{2},1\right)$ $\|\overline{\bm{\Delta}}_{T}\|_{2}\leq O\left(T^{-\frac{1}{2}}\sqrt{\log\frac{d}{\delta}}\right)$ $\eta_{0}\sqrt{\log\frac{1}{\delta\eta_{0}}}\leq\tilde{C}$
+Table 1: Comparisons between our results on TD convergence rate with prior work.
+The quantity $\tilde{C}$, whose value may change at each occurrence, may depend on problem-related parameters other than $T$ and $\delta$.
+In a more specialized line of work, concerned with establishing distributional approximations of the TD iterates, it has been shown that, using Polyak-Ruppert averaging (Polyak and Juditsky, 1992; Ruppert, 1998), the TD estimation error converges to a Gaussian distribution in the large sample limit; see, e.g., Fort (2015); Mou et al.
+(2020); Li et al.
+(2023).
+While these results provide a formal justification for the validity of statistical inference, the guarantees provided in those contributions are non-quantitative, in the sense that they are of asymptotic nature and hold only in fixed dimensions.
+As a result, they do not adequately capture the dependence on the sample size and other problem-related quantities, most notably the dimensionality, all of which affect the accuracy of statistical inference.
+In this paper, we partially address this gap.
+We focus on TD learning with Polyak-Ruppert averaging for estimating the coefficients of the best linear approximation of the value function, assuming i.i.d.
+samples.
+We make the following theoretical and methodological contributions.
   1. 1.
-    Improved high-probability guarantees. We show that, for an appropriate choice of the initial stepsize, the $L_{2}$ norm of the estimation error is of order
-    $O\left(\sqrt{\frac{\mathsf{Tr}(\bm{\Lambda}^{\star})}{T}\log\frac{d}{\delta}}\right),$
+    Improved high-probability guarantees. We show that, for an appropriate choice of the initial stepsize, the Euclidean norm of the estimation error for the TD learning algorithm is of order
+    $O\left(\sqrt{\frac{\mathsf{Tr}(\bm{\Lambda}^{\star})}{T}\log\frac{1}{\delta}}\right),$
     with probability at least $1-\delta$, where $T$ denotes the number of iterations, $\bm{\Lambda}^{\star}$ represents the asymptotic variance matrix of the estimation error, and $\mathsf{Tr}(\bm{\Lambda}^{\star})$ its trace.
 To the best of our knowledge, this is the first high-probability convergence guarantee for this problem that matches the asymptotic variance, up to logarithm factors, without requiring a projection step to control the iterative trajectory.
-Table 1 compares our high-probability convergence guarantees with the latest results in the literature.
+Furthermore, our assumptions are weaker than those used in the literature.
+Table 1 compares our high-probability convergence guarantees with the latest available results.
   2. 2.
     Sharper Berry-Esseen bounds.
-We derive a new high-dimensional Berry-Esseen bound for the estimation error over the class of convex sets, achieving a $O(T^{-\frac{1}{3}})$ rate as the function of the number of iterations.
-Our results improve upon the sharpest rate to date, or order $O(T^{-\frac{1}{4}})$, due to Samsonov et al., (2024).
+We derive a new high-dimensional Berry-Esseen bound for the estimation error over the class of convex sets and using an estimated variance, achieving an $O(T^{-\frac{1}{3}})$ rate as a function of the number of iterations, which appears to be the fastest in the literature.
+While completing this work, we became aware of the recent related paper by Samsonov et al.
+(2024), which, under the same settings, derived a high-dimensional Berry-Esseen bound for linear SA with Polyak-Ruppert averaging, obtaining a slower rate, of order $O(T^{-\frac{1}{4}})$.
+See Section 3.2 for detailed comparisons.
   3. 3.
-    Novel variance estimator.
-We develop a computationally efficient online algorithm to construct an estimator $\widehat{\bm{\Lambda}}_{T}$ of the asymptotic variance $\bm{\Lambda}^{\star}$, and establish a converge rate for $\widehat{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star}$ of order $O(T^{-\frac{1}{3}})$ in Frobenius norm.
-We also verify our theoretical results by numerical simulations.
-As an intuitive illustration, Figure 1 shows how the empirical distribution of TD estimation error in $1,000$ independent trials (as shown in blue histograms), scaled by the square-root of the number of iterations $T$, converges to its asymptotic Gaussian distribution (as shown in red curves).
-1.2 Notation
+    Novel, computationally efficient variance estimator.
+We develop a computationally efficient online algorithm to construct an estimator $\widehat{\bm{\Lambda}}_{T}$ of the asymptotic variance $\bm{\Lambda}^{\star}$.
+We also show that the total variation distance between the (computable) distribution $\mathcal{N}(\bm{0},\widehat{\bm{\Lambda}}_{T})$ and the asymptotic distribution $\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})$ vanishes with high probability at a rate $O(T^{-\frac{1}{3}})$.
+Figure 1: Emergence of the first entry of TD estimation error on an MDP with a 3-dimensional feature map.
+We verify our theoretical findings with some numerical simulations, discussed in the appendix.
+As an illustration, Figure 1 shows how the empirical distribution of TD estimation error in $1,000$ independent trials (as shown in blue histograms), scaled by the square-root of the number of iterations $T$, converges to its asymptotic Gaussian distribution (shown by the red curves).
+The i.i.d.
+condition.
+Our results are derived under the simplifying but standard assumption of i.i.d.
+data, which has been widely used in the recent literature on statistical analyses of TD learning, SA and stochastic gradient descent (see, e.g.
+Sheshukova et al., 2025).
+Analogous results, obtained using different techniques and yielding slower rates, have been recently obtained for the more challenging case of Markovian data: see Srikant (2024) and Wu et al.
+(2025).
+Notation.
+We use boldface small letters to denote vectors and capital letters to denote matrices.
+For a vector $\bm{x}$, $\|\bm{x}\|_{2}$ incates its Euclidean norm; for a matrix $\bm{M}$, we use $\mathsf{Tr}(\bm{M})$, $\mathsf{det}(\bm{M})$, $\|\bm{M}\|_{\mathsf{F}}$ and $\|\bm{M}\|$ to denote its trace, determinant, Frobenius and spectral norm, respectively.
+For a symmetric matrix $\bm{M}\in\mathbb{R}^{d\times d}$, we use $\lambda_{\max}(\bm{M})=\lambda_{1}(\bm{M})\geq\ldots\lambda_{d}(\bm{M})=\lambda_{\min}(\bm{M})$ to denote its eigenvalues and $\mathrm{cond}(\bm{M})$ to indicate its condition number.
+For a vector $\bm{x}\in\mathbb{R}^{d}$ and a symmetric matrix $\bm{M}\in\mathbb{R}^{d\times d}$, we let $\|\bm{x}\|_{\bm{M}}=\sqrt{\bm{x}^{\top}\bm{Mx}}$.
+We will use the notation $\tilde{C}$ to express a quantity independent of $T$, the number of iterations/sample size, but possibly dependent on other problem-related parameters.
+Similarly, for positive sequences $\{f_{t}\}_{t\in\mathbb{N}}$ and $\{g_{t}\}_{t\in\mathbb{N}}$, we write $f_{t}\lesssim g_{t}$, or $f_{t}=O(g_{t})$, to signify that $f_{t}\leq\tilde{C}g_{t}$, for some $C>0$ independent of $t$.
+The convex distance between two probability distributions on $\mathbb{R}^{d}$, say $P$ and $Q$, is defined as
+$d_{\mathsf{C}}(P,Q)=\sup_{\mathcal{A}\in\mathscr{C}_{d}}|P(\mathcal{A})-Q(\mathcal{A})|,$
+where $\mathscr{C}_{d}$ is the class of all convex subsets of $\mathbb{R}^{d}$.
+We use $d_{\mathsf{TV}}(P,Q)$ to denote instead the total variation (TV) distance between $P$ and $Q$.
+Notice that $d_{\mathsf{C}}(P,Q)\leq d_{\mathsf{TV}}(P,Q)$.
+2 Background
+Markov decision process.
+An infinite-horizon Markov Decision Process (MDP) (Bertsekas, 2017; Sutton and Barto, 2018) is the tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},P,r,\gamma)$ where $\mathcal{S}$ and $\mathcal{A}$ represent the state space of the environment and the action space available to an agent, respectively, both assumed for simplicity to be finite (though our results hold for countable spaces).
+The function $P:\mathcal{S}\times\mathcal{A}\mapsto\Delta(\mathcal{S})$ is a transition kernel that associates to any state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$ the probability distribution $P(\cdot\mid s,a)$ of the next state of the environment when the agent executes action $a$ while the environment is in state $s$.
+The reward function $r:\mathcal{S}\times\mathcal{A}\mapsto[0,1]$ measures the immediate reward to the agent after taking action $a$ at state $s$; finally, $\gamma\in[0,1)$ is the discount factor, used to assess the impact of future rewards.
+In choosing their actions, the agent follows a randomized action selection rule, or policy, represented as a (time-invariant) conditional probability distribution $\pi:\mathcal{S}\mapsto\Delta(\mathcal{A})$ over the action space given the state of the environment.
+For any policy $\pi$, its associated reward function $r:\mathcal{S}\mapsto[0,1]$ is the expected reward of the next action selected according to the policy $\pi$, i.e.
+$r(s):=\mathbb{E}_{a\sim\pi(\cdot\mid s)}[r(s,a)],$ for $s\in\mathcal{S}.$ Any policy $\pi$ induces a transition kernel $P$, representing the probability distribution of the next state given the current state $s$ and the policy $\pi$, given by $P(s^{\prime}\mid s)=\mathbb{E}_{a\sim\pi(\cdot\mid s)}P(s^{\prime}\mid s,a)$ for $s,s^{\prime}\in\mathcal{S}.$ We assume throughout that there exists a unique stationary distribution $\mu$ for the transition matrix ${P}$; that is, $\mu(i)=\mathbb{P}_{s\sim\mu,s^{\prime}\sim P(\cdot\mid s)}(s^{\prime}=i)$, $i\in\mathcal{S}$.
+The value function and the Bellman equation.
+The worth of a policy $\pi$ is measured by its value function $s\in\mathcal{S}\mapsto V(s):=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\,\big{|}\,s_{0}=s\right]$ returning, for each initial state, the expected value of all cumulative future rewards, appropriately discounted.
+A fundamental, defining feature of the value function is that it arises as the unique solution to the the Bellman equation, a functional fixed-point equation of the form
+$V(s)=r(s)+\mathbb{E}_{s^{\prime}\sim P(\cdot\mid s)}V(s^{\prime})=:\mathcal{T}V(s).$ (1)
+where $\mathcal{T}$ is known as Bellman’s operator.
+Linear policy approximation and the projected Bellman equation.
+In most practical scenarios, the value function is very complex and impossible to compute due to the prohibitively large size of the state space $\mathcal{S}$ A common way to overcome this issue is to consider a parametric approximation of $V$ through a linear function, which is both computationally tractable and amenable to statistical estimation.
+In detail, for a given feature map $\bm{\phi}:\mathcal{S}\mapsto\mathbb{R}^{d}$, mapping each state in $\mathcal{S}$ to a centered feature vector in $\mathbb{R}^{d}$, where $d$ is typically small(er than $|\mathcal{S}|$), a linear value function approximation is given by
+$s\in\mathcal{S}\mapsto V_{\bm{\theta}}(s):={\bm{\phi}}(s)^{\top}{\bm{\theta}},$ (2)
+where ${\bm{\theta}}\in\mathbb{R}^{d}$ is a vector of linear coefficients.
+Throughout this paper, we assume that the Eculidean norm of the features are uniformly bounded, i.e., $\|\bm{\phi}(s)\|_{2}\leq 1$ for all $s\in\mathcal{S}$.
+The best approximating linear function is the one parametrized by the coefficient $\bm{\theta}^{\star}$ solving the projected Bellman equation (Tsitsiklis and Van Roy, 1997) $V_{\bm{\theta}^{\star}}=\Pi\mathcal{T}V_{\bm{\theta}^{\star}},$ with $\Pi$ the orthogonal projection onto the subspace of functions that can be represented by linear combinations of the features.
+See, e.g., Duan et al.
+(2021).
+In turn, this (unique!) solution is given by
+$\bm{\theta}^{\star}:=\bm{A}^{-1}\bm{b},$ (3)
+where
+$\bm{A}:=\mathop{\mathbb{E}}\limits_{s\sim\mu,s^{\prime}\sim P(\cdot\mid s)}\left[\bm{\phi}(s)\left(\bm{\phi}(s)-\gamma\bm{\phi}(s^{\prime})\right)^{\top}\right]\in\mathbb{R}^{d\times d}\quad\text{and}\quad\bm{b}:=\mathop{\mathbb{E}}\limits_{s\sim\mu}\left[\bm{\phi}(s)r(s)\right]\in\mathbb{R}^{d}.$ (4)
+Throughout, we let
+$\bm{\Sigma}:=\mathbb{E}_{s\sim\mu}[\bm{\phi}(s)\bm{\phi}^{\top}(s)],\quad\lambda_{0}:=\lambda_{\min}(\bm{\Sigma}),\quad\text{and}\quad\lambda_{\Sigma}:=\lambda_{\max}(\bm{\Sigma})$ (5)
+be the feature covariance matrix and its smallest and largest eigenvalues, respectively.
+The Temporal Difference (TD) learning with Polyak-Ruppert averaging.
+While in most real-world RL scenarios the state-action-reward tuples follow a Markov decision process, many of the most recent theoretical contributions – including those delivering the sharpest guarantees listed in Table 1 and Samsonov et al.
+(2024) – have made the simplifying but technically very convenient assumption of $i.i.d.$ data.
+Here we follow the same convention and analyze the properties of the TD learning algorithm applied to a sequence of $T$ $i.i.d.$ state-action-reward tuples of the form
+$\{(s_{t},a_{t},s_{t}^{\prime},r_{t})\}_{t=1}^{T},$ (6)
+where $(s_{1},\ldots,s_{T})\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mu$, the stationary distribution, and, for eact $t$, $a_{t}\sim\pi(s_{t}),r_{t}=r(s_{t},a_{t})$ and $s_{t}^{\prime}\sim P(\cdot\mid s_{t},a_{t})$.
+The TD learning algorithm is a stochastic approximation method to find in an iterative manner the solution (3).
+In detail, for a sequence of pre-selected step sizes $\{\eta_{t}\}_{t\geq 1},$ TD proceeds using the update rule
+$\bm{\theta}_{t}=\bm{\theta}_{t-1}-\eta_{t}(\bm{A}_{t}\bm{\theta}_{t-1}-\bm{b}_{t}),$ (7)
+where
+$\bm{A}_{t}:=\bm{\phi}(s_{t})\left(\bm{\phi}(s_{t})-\gamma\bm{\phi}(s_{t}^{\prime})\right)^{\top}\quad\text{and }\quad\bm{b}_{t}:=r_{t}\bm{\phi}(s_{t}).$ (8)
+Rather than the final TD iterate $\bm{\theta}_{T}$, to estimate $\bm{\theta}^{\star}$ it is often preferred to use instead the average of all the iterates
+$\overline{\bm{\theta}}_{T}$ $=\frac{1}{T}\sum_{t=1}^{T}\bm{\theta}_{t}.$ (9)
+This estimator is an instance of the Polyak-Ruppert averaging scheme (Polyak and Juditsky, 1992; Ruppert, 1998), a stochastic approximation methodology that provably reduces the variability and accelerates convergence.
+Furthermore, with a proper choice of the stepsizes, the averaged iterates converge to a Gaussian distribution, thus enabling statistical inference.
+To date, the best Berry-Esseen bounds rates for stochastic approximation methods and TD learning with Polyak-Ruppert averaging are due to Samsonov et al.
+(2024), which in turn are based on general non-linear Berry-Esseeen bounds by Shao and Zhang (2022).
+In this paper we improve on these results by deriving novel and sharper high-probability and Berry-Esseen bounds for the estimation error of the TD algorithm
+$\overline{\bm{\Delta}}_{T}=\overline{\bm{\theta}}_{T}-\bm{\theta}^{\star}.$ (10)
+We conclude by mentioning that the choice of learning rates $\{\eta_{t}\}_{t\geq 0}$ in the updates (7) plays a critical role in determining the statistical properties of (both the original and) the averaged TD estimator (9).
+Here we assume a polynomial-decaying stepsizes schedule $\eta_{t}=\eta_{0}t^{-\alpha}$ for $\alpha\in(1/2,1)$ and $\eta_{0}<1/2$.
+3 Results
+It is known (Polyak and Juditsky, 1992; Fort, 2015) that the TD estimation error of (10) satisfies
+[…]
+where the asymtoptic covariance matrix $\bm{\Lambda}^{\star}$ is given by
+$\bm{\Lambda}^{\star}=\bm{A}^{-1}\bm{\Gamma}\bm{A}^{-\top},\quad\text{with}\quad\bm{\Gamma}=\mathbb{E}[(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}].$ (12)
+Thus, in the fixed-dimensional setting, the above result implies the rate
+$\|\overline{\bm{\Delta}}_{T}\|_{2}=O_{P}\left(\sqrt{\frac{\mathsf{Tr}(\bm{\Lambda}^{\star})}{T}}\right),$ (13)
+which may serve as a benchmark.
+Our main goal is to demonstrate how the TD algorithm can be used to construct non-asymptotic confidence regions and simultaneous confidence intervals for $\bm{\theta}^{*}$ in high dimensions.
+Towards this goal, in the subsequent sections we will establish a high-probability convergence guarantee for $\overline{\bm{\Delta}}_{T}$, develop high-dimensional Berry-Esseen bounds over the class of convex sets, and finally build confidence regions/intervals for $\bm{\theta}^{\star}$ using the estimator $\overline{\bm{\theta}}_{T}$.
+3.1 High-probability bounds
+We start with a high-dimensional, high-probability bound on the size of the TD estimation error.
+Theorem 3.1.
+Consider the TD algorithm with Polyak-Ruppert averaging (7), independent samples and decaying stepsizes $\eta_{t}=\eta_{0}t^{-\alpha}$ for $\alpha\in(\frac{1}{2},1)$.
+For any $\delta\in(0,1)$, there exists $\eta_{0}=\eta_{0}(\delta)>0$, such that
+$\|\overline{\bm{\Delta}}_{T}\|_{2}$ $\lesssim\sqrt{\frac{\mathsf{Tr}(\bm{\Lambda}^{\star})}{T}\log\frac{1}{\delta}}+\tilde{C}T^{-\frac{\alpha+1}{2}}\log\frac{T}{\delta}$
+with probability at least $1-\delta$, where $\tilde{C}$ is a problem-related quantity independent of $T$ and $\delta$.
+The proof can be found in Appendix B.1, where the value of $\tilde{C}$ is given in equation (78).
+Throughout the paper, for the sake of readability, we have opted to express unwieldy problem-dependent quantities that do not directly depend on $T$ in implicit form; their values can be tracked in the proofs.
+Accordingly, when writing our finite-sample bounds, we will first present the leading terms in $T$.
+In a nutshell, Theorem 3.1 states that $\overline{\bm{\theta}}_{T}$ converges to $\bm{\theta}^{*}$ (in Euclidean distance) at the rate of $O(\sqrt{{\mathsf{Tr}(\bm{\Lambda}^{\star})}/{T}})$ up to a log factor.
+It is worth comparing the high-probability sample complexity bound of Theorem 3.1 with those in the literature.
+To the best of our knowledge, the tightest results were attained by Dalal et al.
+(2018) for inverse-decaying stepsizes and Li et al.
+(2024); Samsonov et al.
+(2024) for constant stepsize as summarized in Table 1.
+Theorem 3.1 is the first result to achieve tight high-probability convergence guarantees up to log factors using polynomial-decaying stepsizes.
+Another subtle but consequential difference concerns the restriction on the step size.
+For our bound to hold, the choice of the initial stepsize $\eta_{0}$ depends on $\delta$; specifically, $\eta_{0}$ is required to be sufficiently small such that $\eta_{0}\sqrt{\log{1}/(\delta\eta_{0}})\leq\tilde{C}$, where $\tilde{C}$ is independent of $T$ and $\delta$ but depends on the other parameters of interest; its specific form is shown in (70) and (71) in the Appendix.
+Though not ideal, this condition is necessary for a Chernoff-type convergence guarantee; we refer to Appendix B.1.3 for a detailed discussion.
+Nonetheless, our stepsize requirements are considerably weaker than those imposed by Samsonov et al.
+(2024) and Li et al.
+(2024), which depend on both $\delta$ and $T$; see Table 1.
+3.2 High-dimensional Berry-Esseen bounds
+In our next result, we derive novel high-dimensional Berry-Esseen bounds for the TD learning estimator, certifying that, uniformly over the class of $d$-dimensional convex sets, $\sqrt{T}\overline{\bm{\Delta}}_{T}$ is close in distribution to a centered Gaussian random vector with the optimal asymptotic covariance (12), thus closely mimicking the classical, fixed-dimensional CLT guarantee in (11).
+We reiterate that, to the best of our knowledge, the resulting rates are the sharpest to date.
+Theorem 3.2.
+Consider the TD algorithm with Polyak-Ruppert averaging (7), independent samples and decaying stepsizes $\eta_{t}=\eta_{0}t^{-\alpha}$ for $\alpha\in(\frac{1}{2},1)$ and $\eta_{0}<1/2$. Then,
+$d_{\mathsf{C}}(\sqrt{T}\overline{\bm{\Delta}}_{T},\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))$ $\lesssim\sqrt{\frac{\eta_{0}}{(1-\alpha)(1-\gamma)\lambda_{0}}}\mathsf{Tr}(\bm{\Gamma})\|\bm{\Gamma}^{-1}\|T^{-\frac{\alpha}{2}}+\frac{\sqrt{d\mathsf{cond}(\bm{\Gamma})}}{(1-\gamma)\lambda_{0}\eta_{0}}T^{\alpha-1}$
+$+\tilde{C}\;T^{-\frac{1}{2}}+\tilde{C}^{\prime}\;T^{2\alpha-2}$
+where $\tilde{C},\tilde{C}^{\prime}$ are problem-related quantities independent of $T$.
+On the final rate.
+By choosing $\alpha=2/3$, we obtain a rate in $T$ of order $O(T^{-1/3})$.
+In fact, in Theorem B.3 in the Appendix we establish the more refined result that the Berry-Esseen bound (in the convex distance) between the distribution of $\sqrt{T}\overline{\bm{\Delta}}_{T}$ and the Gaussian distribution with (approximately) matching first two moments is of order $\sqrt{\frac{\eta_{0}}{(1-\alpha)(1-\gamma)\lambda_{0}}}\mathsf{Tr}(\bm{\Gamma})\|\bm{\Gamma}^{-1}\|T^{-\frac{\alpha}{2}}$.
+As $\alpha\to 1$, this rate approaches the optimal rate of $1/\sqrt{T}$.
+The additional term $\frac{\sqrt{d\mathsf{cond}(\bm{\Gamma})}}{(1-\gamma)\lambda_{0}\eta_{0}}T^{\alpha-1}$ in the above bound, which is ultimately the reason for the final rate $O(T^{-1/3})$, originates from replacing an approximation to the finite sample variance of $\sqrt{T}\overline{\bm{\Delta}}_{T}$ with its asymptotic variance $\bm{\Lambda}^{\star}$.
+Comparison with Samsonov et al.
+(2024).
+Prior to this work, the best Berry-Esseen bounds for stochastic approximation with polynomial-decaying stepsizes and Polyak-Ruppert averaging are those obtained by Samsonov et al.
+(2024), thus covering the settings considered here.
+Specifically, the authors show in their Theorem 3 that when $\eta_{0}\leq\min\{\lambda_{0}(1-\gamma),\frac{1-\gamma}{(1+\gamma)^{2}}\}$ and $\alpha\in[\frac{1}{2},1)$,
+$d_{\mathsf{C}}(\sqrt{T}\overline{\bm{\Delta}}_{T},\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\lesssim O(T^{-\frac{\alpha}{2}}+T^{\frac{\alpha-1}{2}}).$
+Choosing $\alpha=\frac{1}{2}$ leads to a convergence rate of order $O(T^{-\frac{1}{4}})$.
+In contrast, Theorem 3.2 shows that, by selecting $\alpha=\frac{2}{3}$, we obtain the faster rate $O(T^{-\frac{1}{3}})$.
+As shown in Appendix B.2.3, this rate is tight for $\alpha\in(\frac{2}{3},1)$.
+See also Figure 2 below, showing the output of a numerical experiment demonstrating that the choice of $\alpha=\frac{2}{3}$ leads to a smaller approximation error.
+Though worked out independently, the proofs of our Theorem 3.2 and of Theorem 3 of Samsonov et al.
+(2024) bear similarities, as they both rely on the general Berry-Esseen bounds for non-linear statistics developed by Shao and Zhang (2022).
+We obtain sharper rates by a more targeted analysis of the TD settings, while, in comparison, the results of Samsonov et al.
+(2024) apply more broadly to SA problems.
+Specifically we (i) adopt a different decomposition of the TD estimation error, (ii) handle in a more refined way certain terms appearing in the bound of Theorem 3.1 of Shao and Zhang (2022) [^2] Specifically, the term $I_{3}$ in Appendix B.2.1.
+and (iii) carry out a more specialized analysis of the properties of the matrices $\bm{A}_{t}$ and $\bm{A}$ (see, e.g., Lemma A.4 in the Appendix) originating from TD learning, which has allowed us to tighten problem-related coefficient, as well as to work with a substantially larger intial stepsize $\eta_{0}$.
+This non-trivial sharpening may be potentially generalized to analyze stochastic gradient algorithms applied to non-linear functions, as, e.g., in Theorem 3.4 of Shao and Zhang (2022).
+3.3 Online estimation of the covariance matrix
+In the last section, we obtained finite sample bounds on the convex distance between $\sqrt{T}\overline{\bm{\Delta}}_{T}$ and its Gaussian approximation, which features an unknown variance.
+To construct confidence regions for $\bm{\theta}^{\star}$, it is necessary to estimate the covariance matrix $\bm{\Lambda}^{\star}$.
+Two types of approaches have been recently proposed: plug-in estimation (Chen et al., 2020) and bootstraping (White and White, 2010; Hanna et al., 2017; Hao et al., 2021; Ramprasad et al., 2022; Zhu et al., 2023; Samsonov et al., 2024).
+Here we propose to use a computationally feasible online plug-in estimator for $\bm{\Lambda}^{\star}$ with superior computational efficiency than existing methodologies and establish its non-asymptotic rate of convergence.
+We will use
+$\overline{\bm{A}}_{T}=\frac{1}{T}\sum_{t=1}^{T}\bm{A}_{t},\quad\text{and}\quad\widehat{\bm{\Gamma}}_{T}=\frac{1}{T}\sum_{t=1}^{T}(\bm{A}_{t}\overline{\bm{\theta}}_{T}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\theta}}_{T}-\bm{b}_{t})^{\top},$ (14)
+as estimators for $\bm{A}$ and $\bm{\Gamma}$ respectively, and set
+$\widehat{\bm{\Lambda}}_{T}=\overline{\bm{A}}_{T}^{-1}\widehat{\bm{\Gamma}}_{T}\overline{\bm{A}}_{T}^{-\top}$ (15)
+as the final estimator of $\bm{\Lambda}^{\star}$.
+Crucially, because of the linear nature of the problem, the estimator $\widehat{\bm{\Lambda}}_{T}$ can be evaluated in an online manner.
+See Algorithm 1 in Appendix D.1 for a detailed description of the procedure, which has time and space complexities of order $O(Td^{4})$ and $O(d^{4})$, respectively.
+The estimator $\widehat{\bm{\Lambda}}_{T}$ also enjoys good statistical properties.
+This is made precise by the following Gaussian comparison result, whose proof can be found in Appendix B.3, showing that the unknown asymptotic variance $\bm{\Lambda}^{\star}$ in the Berry-Esseen bound of Theorem B.4 can be replaced by $\widehat{\bm{\Lambda}}_{T}$ without compromising the accuracy of the approximation (and without affecting the rate in $T$).
+Theorem 3.3.
+Consider the TD algorithm (7) with independent samples and decaying stepsizes $\eta_{t}=\eta_{0}t^{-\alpha}$ for all $t>0$, any $\alpha\in(1/2,1)$ and $\eta_{0}<1/2$ , and let $\widehat{\bm{\Lambda}}_{T}$, $\bm{\Lambda}^{\star}$ be defined as in (15), (12) respectively.
+Then,
+$d_{\mathsf{TV}}(\mathcal{N}(\bm{0},\widehat{\bm{\Lambda}}_{T}),\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\lesssim\frac{\sqrt{\mathsf{Tr}(\bm{\Gamma})}}{\lambda_{0}(1-\gamma)}(\|\bm{\theta}^{\star}\|_{2}+1)\|\bm{\Gamma}^{-1}\|T^{-\frac{1}{3}}+o(T^{-\frac{1}{3}}).$
+with probability at least $1-3T^{-\frac{1}{3}}-\frac{\sqrt{\mathsf{Tr}(\bm{\Gamma})}}{\lambda_{0}(1-\gamma)\sqrt{T}}$.
+Above, the term $o(T^{-\frac{1}{3}})$ depends on all the parameters of the problem (and, in particular, to $\alpha$) but, as a function of $T$ alone, is of smaller order than $T^{-\frac{1}{3}}$.
+See Appendix B.3 for the proof and details.
+Comparison with Chen et al.
+(2020).
+In different settings, Chen et al.
+(2020) considered the problem of carrying out statistical inference for the true model parameters using stochastic gradient descent and assuming a strongly convex and smooth population loss function.
+In their work, they consider the slightly different plug-in estimator $\widehat{\bm{\Lambda}}_{T}^{\prime}=\overline{\bm{A}}_{T}^{-1}\widehat{\bm{\Gamma}}_{T}^{\prime}\overline{\bm{A}}_{T}^{\top}$, where $\widehat{\bm{\Gamma}}_{T}^{\prime}:=\frac{1}{T}\sum_{t=1}^{T}(\bm{A}_{t}{\bm{\theta}}_{t}-\bm{b}_{t})(\bm{A}_{t}{\bm{\theta}}_{t}-\bm{b}_{t})^{\top}$.
+The authors proved that the estimation error of $\widehat{\bm{\Lambda}}_{T}^{\prime}$ is bounded in expectation as
+$\mathbb{E}\|\widehat{\bm{\Lambda}}_{T}^{\prime}-\bm{\Lambda}^{\star}\|\leq\|\bm{\Gamma}\|\cdot(\widetilde{C}_{1}T^{-\frac{\alpha}{2}}+\widetilde{C}_{2}T^{-\alpha}),$ (16)
+for problem-related constants $\widetilde{C}_{1}$ and $\widetilde{C}_{2}$.
+It is possible to modify the proof of Theorem 3.3 to obtain a guarantee in expectation for our problem as well, albeit in the TV metric.
+Specifically, we can show that
+$\mathbb{E}\left[d_{\mathsf{TV}}(\mathcal{N}(\bm{0},\widehat{\bm{\Lambda}}_{T}),\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\right]$ $\lesssim\widetilde{C}\|\bm{\Gamma}\|T^{-\frac{1}{2}}$ (17)
+where $\widetilde{C}$ depends on $\lambda_{0},\gamma,\eta_{0},\alpha$ and $\bm{\theta}^{\star}$.
+Comparing (16) with (17), we have sharpened the rate from $O\left(T^{-\frac{\alpha}{2}}\right)$ to $O\left(T^{-\frac{1}{2}}\right)$, independent of the choice of $\alpha$.
+It is noteworthy that the online Algorithm 1 takes advantage of the linearity of the problem; in general, potentially non-linear settings, replacing $\bm{\theta}_{t}$ in $\widehat{\bm{\Gamma}}_{T}^{\prime}$ by $\overline{\bm{\theta}}_{T}$ in $\widehat{\bm{\Gamma}}_{T}$ would require storing all the samples throughout the iterative trajectory, resulting in a larger space complexity.
+Comparison with Samsonov et al. (2024).
+Leveraging the multiplier-bootstrap technique proposed by Zhu et al.
+(2023), Samsonov et al.
+(2024) constructs a randomized estimator $\overline{\bm{\theta}}_{T}^{(b)}$ based on the samples $\mathcal{S}_{T}:=\{s_{t},s_{t}^{\prime}\}_{t=1}^{T}$.
+Their Theorems 4 and 5 imply that, in our notation,
+$d_{\mathsf{C}}(\sqrt{T}(\overline{\bm{\theta}}_{T}^{(b)}-\bm{\theta}^{\star})|\mathcal{S}_{T},\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\lesssim\frac{\|\bm{\Gamma}^{-1}\|}{[\lambda_{0}(1-\gamma)]^{\frac{3}{2}}}(2\|\bm{\theta}^{\star}\|_{2}+1)^{2}\frac{\log T}{T^{\frac{1}{4}}}+o(T^{-\frac{1}{4}})$ (18)
+Compared with this bootstrap-based procedure, our methodology is advantageous in the following ways: (i) the upper bound in Theorem 3.3 vanishes, as a function of $T$, at the rate of $O(T^{-\frac{1}{3}})$, which is faster than the $O(T^{-\frac{1}{4}}\log T)$ rate from the bound (18).
+(ii) The leading term in the upper bound in Theorem 3.3 entails a smaller problem-related coefficient than that of (18).
+(iii) Perhaps most importantly, in order to estimate the bootstrap distribution $\sqrt{T}(\overline{\bm{\theta}}_{T}^{(b)}-\bm{\theta}^{\star})|\mathcal{S}_{T}$ with an $O(T^{-\frac{1}{4}})$ error rate, it is necessary to store the sample trajectory $\mathcal{S}_{T}$ and generate $O(T^{\frac{1}{2}})$ perturbed trajectories.
+Consequently, the bootstrap procedure has a time complexity of order $O(T^{\frac{3}{2}}d^{2})$ and a space complexity of order $O(T)$.
+When $d\ll T$, this amounts to a much higher computational burden than our online Algorithm 1.
+3.4 Statistical inference
+Combining the results from Sections 3.2 and 3.3, we can now propose a general methodology for constructing convex confidence sets for $\bm{\theta}^{\star}$ with guaranteed finite sample coverage.
+Based on the TD estimator $\overline{\bm{\theta}}_{T}$ (9) and the variance estimator $\widehat{\bm{\Lambda}}_{T}$ (15), and given any tolerance level $\delta\in(0,1)$, it is enough to pick any convex set $\mathcal{C}_{\delta}$ that has probability at least $1-\delta$ under the (random!) Gaussian distribution $\mathcal{N}(\overline{\bm{\theta}}_{T},\frac{\widehat{\bm{\Lambda}}_{T}}{T})$.
+That is, ssattisfying, the condition
+$\mathbb{P}_{\bm{\theta}\sim\mathcal{N}(\overline{\bm{\theta}}_{T},\frac{\widehat{\bm{\Lambda}}_{T}}{T})}\left(\bm{\theta}\in\mathcal{C}_{\delta}\bigg{|}\overline{\bm{\theta}}_{T},\widehat{\bm{\Lambda}}_{T}\right)\geq 1-\delta.$ (19)
+Natural choices of $\mathcal{C}_{\delta}$ are (i) hyperctangles, leading to simultaneous confidence intervals of the form
+$\mathcal{C}_{\delta}=\prod_{j=1}^{d}[(\overline{\bm{\theta}}_{T})_{j}-R/\sqrt{T},(\overline{\bm{\theta}}_{T})_{j}+R/\sqrt{T}],$ (20)
+where $R>0$ is the quantile of the $\ell_{\infty}$ norm of a centered Gaussian random vectors with covariance $(\widehat{\bm{\Lambda}}_{T}/{T})$, which can be calculated via simulations; and (ii) ellipsoids or balls, for example
+$\mathcal{C}_{\delta}=\left\{\bm{\theta}:\sqrt{T}\|\bm{\theta}-\overline{\bm{\theta}}_{T}\|_{\widehat{\bm{\Lambda}}_{T}^{-1}}^{2}\leq R\right\},$ (21)
+where $R$ is the $1-\delta$ quantile of the $\chi^{2}(d)$ distribution.
+The next result, whose proof is in Appendix B.4, derives a coverage rate for $\mathcal{C}_{\delta}$, which converges to the nominal level of $1-\delta$ at a rate of order $O(T^{-1/3})$, as a function of $T$ alone.
+Theorem 3.4.
+Consider TD with Polyak-Ruppert averaging (7) with independent samples and decaying stepsizes $\eta_{t}=\eta_{0}t^{-\alpha}$, for all $t>0$ with $\alpha=2/3$ and $\eta_{0}<1/2$.For a given significance level $\delta\in(0,1)$, it holds that
+$\mathbb{P}(\bm{\theta}^{\star}\in\mathcal{C}_{\delta})$ $\geq 1-\delta-\frac{\sqrt{\mathsf{Tr}(\bm{\Gamma})}}{\lambda_{0}(1-\gamma)}(2\|\bm{\theta}^{\star}\|_{2}+1)\|\bm{\Gamma}^{-1}\|T^{-\frac{1}{3}}-\frac{\sqrt{d\mathsf{cond}(\bm{\Gamma})}}{2(1-\gamma)\lambda_{0}\eta_{0}}T^{-\frac{1}{3}}-o(T^{-\frac{1}{3}}),$
+uniformly over all convex sets $\mathcal{C}_{\delta}\subset\mathbb{R}^{d}$ satisfying (19).
+We highlight two important features of the above result.
+First, the coverage guarantee holds uniformly over any choice of the convex set $\mathcal{C}_{\delta}$ meeting condition (19), including the common choices like (20) and (21).
+Secondly, the size (i.e., volume) of the confidence sets vanishes at the rate of $O(T^{-1/2})$, keeping all the other variable constant, suggesting that, in this regard, there is little or no room for improvements, at least in terms of rates.
+4 Simulations
+We illustrate our theoretical results with some numerical experiments.
+[^3] The code for our experiments can be found at https://www.dropbox.com/scl/fo/tbw256dlx4gzxzgb4t3hq/AOUaiLFj79iSnc29dFTPpvQ?rlkey=nsgwlcbfx0b73uebwv8xlovag&st=1axbxaa1&dl=0 In order to test our proposed framework for statistical inference, we consider a member of a family of MDPs that are designed to be difficult to distinguish among each other.
+The specifics of the MDP are the same as described in Li et al.
+(2024), and we refer readers to Appendix D for the construction details.
+With $10,000$ independent trials, we examine the high-probability convergence of the $\ell_{2}$ norm of TD estimation errors, the rate by which they converge to their asymptotic Gaussian distributions, and the precision of our estimator for the covariance matrix.
+Figure 2 reveals that the empirical distribution of TD estimation error converges to the asymptotic Gaussian distribution slowly when $\alpha$ is set to ${1}/{2}$ as suggested by Samsonov et al.
+(2024).
+Notice that, for computational simplicity, the difference bewteen the empirical distribution of $\sqrt{T}\overline{\bm{\Delta}}_{T}$ and the asymptotic distribution $\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})$ is measured by
+$d(\sqrt{T}\overline{\bm{\Delta}}_{T},\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})):=\max_{1\leq j\leq d}\sup_{x\in\mathbb{R}}\left|\mathbb{P}(\sqrt{T}(\overline{\bm{\Delta}}_{T})_{j}\leq x)-\mathbb{P}((\bm{\Lambda}^{\star\frac{1}{2}}\bm{z})_{j}\leq x)\right|$
+instead of the convex distance we invoked in the theoretical results.
+Figure 2: Distance between empirical and asymptotic distributions of the TD estimation error.
+5 Discussion
+We develop high-probability convergence guarantees and Berry-Esseen bounds for TD learning with linear function approximation and $i.i.d.$ samples in a non-asymptotic manner.
+We also propose an efficient way to construct a covariance estimator and derive its consistency rates.
+These results lead to a practicable methodology for constructing convex confidence sets for the target parameter, with its size vanishing at a parametric rate $O(T^{-1/2})$.
+An important direction for future research is to investigate whether the order $O(T^{-1/3})$ for the convergence speed of coverage rate in Theorem 3.4 are tight.
+Another interesting extension would be to derive specialized Berry-Esseen bounds for the sub-class of hyper-rectangles, where one would expect to find a better dependence on the dimension $d$ of the feature space, possibly of poly-logarithmic, as opposed to polynomial, order.
+We believe our analysis and results can be adapted or extended to cover other iterative algorithms, such as stochastic gradient descent and SA methods.
+Recent analyses (see, e.g., Sheshukova et al., 2025; Samsonov et al., 2024) have demonstrated Berry-Esseen bounds for the convex distance (using an estimated variance and assuming i.i.d.
+data), as well as valid bootstrap schemes.
+Acknowledgments and Disclosure of Funding
+W.
+Wu and A.
+Rinaldo acknowledge partial support from the NIH under Grant R01NS121913.
+Gen Li was supported in part by the Chinese University of Hong Kong Direct Grant for Research.
+Y.
+Wei was supported in part by the NSF grants CCF-2106778, CCF-2418156 and CAREER award DMS-2143215.
+The authors thank Yang Peng for pointing us to Theorem A.19.
+Appendix A Preliminaries
+In this section, we summarize some preliminary facts that would be useful in the proof of the theoretical results in this paper.
+A.1 Useful algebraic facts
+Supplementary notation.
 For any complex number $z\in\mathbb{C}$, we use $\mathsf{Re}(z)$ to denote its real part.
-Throughout the paper, we use boldface small letters to denote vectors and capital letters to denote matrices.
-For any vector $\bm{x}$, we use $\|\bm{x}\|_{2}$ to denote its $L_{2}$ norm; for any matrix $\bm{M}$, we use $\mathsf{Tr}(\bm{M})$ to denote its trace, $\mathsf{det}(\bm{M})$ to denote its determinant, $\mathsf{cond}(\bm{M})$ to denote its conditional number, $\|\bm{M}\|$ to denote its spectral norm (i.e.
-the largest singular value), and $\|\bm{M}\|_{\mathsf{F}}$ to denote its Frobenius norm, i.e.
-$\|\bm{M}\|_{\mathsf{F}}=\sqrt{\mathsf{Tr}(\bm{M}^{\top}\bm{M})}$.
-For any symmetric matrix $\bm{M}\in\mathbb{R}^{d\times d}$, we use $\lambda_{\max}(\bm{M})=\lambda_{1}(\bm{M})\geq\lambda_{2}(\bm{M})\geq...\geq\lambda_{d}(\bm{M})=\lambda_{\min}(\bm{M})$ to denote its eigenvalues.
 If a matrix $\bm{M}$ is positive (semi-)definite, i.e.
@@
 With subscripts, it can be represented as $(\bm{X}\otimes\bm{Y})_{(i-1)m+j,(k-1)p+\ell}=X_{ik}Y_{j\ell}$.
-For sequences $\{f_{t}\}_{t\in\mathbb{N}}$ and $\{g_{t}\}_{t\in\mathbb{N}}$ of numbers, we write $f_{t}\lesssim g_{t}$, or $f_{t}=O(g_{t})$, to rsignify that there exists a quantity $C>0$ independent of $t$ such that $f_{t}\leq Cg_{t}$ for all $t$.
-We will use the notatin $\widetilde{C}$ to express a quantity independent of $T$, the number of iterations/sample size, but possibly dependent on other problem-related parameters.
-We will measure the difference between two probability distributions on $\mathbb{R}^{d}$ using the convex distance, the maximal distance convex sets.
-Specifically, for any two probability distributions $P$ and $Q$, we set
-$d_{\mathsf{C}}(P,Q)=\sup_{\mathcal{A}\in\mathscr{C}_{d}}|P(\mathcal{A})-Q(\mathcal{A})|.$
-where $\mathscr{C}_{d}$ is the class of all convex set in $\mathbb{R}^{d}$. We use $d_{\mathsf{TV}(P,Q)}$ to denote the Total Variation (TV) distance between $P$ and $Q$.
-paper stepsize convergence guarantee constraint on $t,T,\delta$
-[…]
-Li et al., 2024b $\eta_{t}=\eta$ $\|\overline{\bm{\Delta}}_{T}\|_{2}\leq O\left(T^{-\frac{1}{2}}\log^{\frac{1}{2}}\frac{T}{\delta}\right)$ $\eta\sqrt{\log\frac{T}{\delta}}\leq\widetilde{C}$
-Samsonov et al., (2023) $\eta_{t}=\eta$ $\|\overline{\bm{\Delta}}_{T}\|_{2}\leq O\left(T^{-\frac{1}{2}}\log\frac{T}{\delta}\right)$ $\eta\log\frac{T}{\delta}\leq\widetilde{C}$
-This work, Theorem 3.1 $\eta_{t}=\eta_{0}t^{-\alpha},\alpha\in\left(\frac{1}{2},1\right)$ $\|\overline{\bm{\Delta}}_{T}\|_{2}\leq O\left(T^{-\frac{1}{2}}\sqrt{\log\frac{d}{\delta}}\right)$ $\eta_{0}\sqrt{\log\frac{1}{\delta\eta_{0}}}\leq\widetilde{C}$
-Table 1: Comparisons between our results on convergence rates for TD with prior work that derived with high probability control.
-The quantity $\widetilde{C}$, whose value may change at each occurrence, may depend on problem-related parameters other than $T$ and $\delta$.
-2 Background
-Markov decision process.
-An infinite-horizon Markov Decision Process (MDP) (Bertsekas,, 2017; Sutton and Barto,, 2018) can be represented mathematically by the tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},P,r,\gamma)$.
-Here, $\mathcal{S}$ and $\mathcal{A}$ represent the state space of the environment and the action space available to an agent respectively, both assumed finite throughout this paper; the function $P:\mathcal{S}\times\mathcal{A}\mapsto\Delta(\mathcal{S})$ is a transition kernel that associates to any state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$ the probability distribution $P(\cdot\mid s,a)$ of the next state of the environment when the agent executes action $a$ while the environment is in state $s$; the reward function $r:\mathcal{S}\times\mathcal{A}\mapsto[0,1]$ measures the immediate reward to the agent after taking action $a$ at state $s$; finally, $\gamma\in(0,1]$ is the discount factor, used to assess the impact of future rewards.
-In choosing their actions, the agent follows a randomized action selection rule, or policy, represented as a (time-invariant) conditional probability distribution $\pi:\mathcal{S}\mapsto\Delta(\mathcal{A})$ over the action space given the state of the environment.
-For any policy $\pi$, its associated reward function $r^{\pi}:\mathcal{S}\mapsto[0,1]$ is the expected reward of the next action selected according to the policy $\pi$, i.e.
-$r^{\pi}(s):=\mathbb{E}_{a\sim\pi(\cdot\mid s)}[r(s,a)],\quad s\in\mathcal{S}.$
-For readability, we will omit the superscript $\pi$ and use $r(\cdot)$ to denote $r^{\pi}(\cdot)$.
-Any policy $\pi$ induces a transition kernel $P^{\pi}:\mathcal{S}\mapsto\Delta(\mathcal{S})$, representing the probability distribution of the next state, given the current state $s$ and the policy $\pi$.
-Formally,
-$P^{\pi}(s^{\prime}\mid s)=\mathbb{E}_{a\sim\pi(\cdot\mid s)}P(s^{\prime}\mid s,a),\quad s^{\prime}\in\mathcal{S}.$
-We assume throughout that there exists a unique stationary distribution $\mu^{\pi}\in\Delta(\mathcal{S})$ for the transition matrix ${P}^{\pi}$, assigning positive probability to each state. That is,
-$0<\mu^{\pi}(i)=\mathbb{P}_{s\sim\mu^{\pi},s^{\prime}\sim P^{\pi}(\cdot\mid s)}(s^{\prime}=i),\quad\forall i\in\mathcal{S}.$
-The value function and the Bellman equation.
-The worth of a policy $\pi$ is measured through its value function $V^{\pi}:\mathcal{S}\mapsto\mathbb{R}_{+}$ returning, for each initial state, the expected value of all cumulative future rewards, appropriately discounted, when policy $\pi$ is executed under the MDP $\mathcal{M}$.
-Formally,
-$V^{\pi}(s):=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\,\big{|}\,s_{0}=s\right],\quad s\in\mathcal{S}.$ (1)
-A fundamental, defining feature of the value function is that it satisfies a certain self-consistency property, known as the Bellman equation, a functional fixed-point equation of the form
-$V^{\pi}(s)=r(s)+\mathbb{E}_{s^{\prime}\sim P^{\pi}(\cdot\mid s)}V^{\pi}(s^{\prime})=:\mathcal{T}V^{\pi}(s),$ (2)
-where $\mathcal{T}$ is the Bellman’s operator.
-Indeed, the value function is the unique solution to the Bellman equation.
-With a slight abuse of notation, we shall omit the superscript $\pi$ in $V^{\pi}$ and abbreviate it as $V$ as the target policy is considered fixed.
-Linear policy approximation and the projected Bellman equation.
-In most practical scenarios, though finite, the state space $\mathcal{S}$ is prohibitively large.
-As a result, the value function is often very complex and impractical to compute.
-To overcome this issue, it is customary to consider a parametric approximation of $V$ through a linear function, which is computationally tractable and amenable to statistical estimation (see, e.g.
-Tsitsiklis and Van Roy, (1997); Dalal et al., (2018); Bhandari et al., (2021); Srikant and Ying, (2019); Lakshminarayanan and Szepesvari, (2018)) In detail, for a given feature map $\bm{\phi}:\mathcal{S}\mapsto\mathbb{R}^{d}$, mapping each state in $\mathcal{S}$ to a feature vector in $\mathbb{R}^{d}$, where $d$ is usually smaller than $|\mathcal{S}|$, the value function at $s$, $V(s)$, is expproximated by
-$V_{\bm{\theta}}(s):={\bm{\phi}}(s)^{\top}{\bm{\theta}},\qquad\quad s\in\mathcal{S},$ (3)
-where ${\bm{\theta}}\in\mathbb{R}^{d}$ is a vector of linear coefficients.
-Computing the optimal linear approximation to $V$ can be cast as a least-squares minimization problem over $\mathbb{R}^{d}$ of finding a coefficient vector $\bm{\theta}$ minimizing the expected difference between $V_{\bm{\theta}}$ and its Bellman projection $\mathcal{T}V_{\bm{\theta}}$.
-That is,
-$\bm{\theta}^{\star}:=\underset{\bm{\theta}\in\mathbb{R}^{d}}{\text{argmin }}\mathbb{E}_{s\sim\mu}(V_{\bm{\theta}}(s)-\mathcal{T}V_{\bm{\theta}}(s))^{2},$ (4)
-where we recall that $\mu$ that denote the stationary distribution of $P^{\pi}$. It turns out that the problem (4) admits a unique solution (Tsitsiklis and Van Roy,, 1997), given by
-$\bm{\theta}^{\star}:=\bm{A}^{-1}\bm{b},$ (5)
-where
-$\bm{A}$ $:=\mathop{\mathbb{E}}\limits_{s\sim\mu,s^{\prime}\sim P(\cdot\mid s)}\left[\bm{\phi}(s)\left(\bm{\phi}(s)-\gamma\bm{\phi}(s^{\prime})\right)^{\top}\right]\in\mathbb{R}^{d\times d},$ (6a)
-[…]
-For notional simplicity, we further introduce the notation $\bm{\Sigma}$ as feature covariance matrix
-$\bm{\Sigma}:=\mathbb{E}_{s\sim\mu}[\bm{\phi}(s)\bm{\phi}^{\top}(s)],$ (7)
-and use $\lambda_{\Sigma}$ and $\lambda_{0}$ to denote its largest and smallest eigenvalue, namely
-$\lambda_{0}:=\lambda_{\min}(\bm{\Sigma}),\quad\text{and}\quad\lambda_{\Sigma}:=\lambda_{\max}(\bm{\Sigma}).$ (8)
-The Temporal Difference (TD) learning algorithm.
-To estimate the optimal vector $\bm{\theta}^{*}$ given by (4) we use the observe data, consisting of $T$ triplets of (state, action, reward),
-$\{(s_{0},a_{0},r_{0}),(s_{1},a_{1},r_{1}),\ldots(s_{T},a_{T},r_{T})\},$
-where, for each $t=0,1,\ldots,T$, $a_{t}\sim\pi(\cdot\mid s_{t})$, $r_{t}=r(s_{t},a_{t})$ and $s_{t+1}\sim P(\cdot\mid s_{t},a_{t})$.
-The TD learning algorithm is a stochastic approximation algorithm that iteratively approximates the solution to the linear system $\bm{A}\bm{\theta}=\bm{b}$.
-Specifically, given a sequence of pre-selected step size $\{\eta_{t}\}_{t\geq 1},$ TD proceeds using the updating rule
-$\bm{\theta}_{t}$ $=\bm{\theta}_{t-1}-\eta_{t}(\bm{A}_{t}\bm{\theta}_{t-1}-\bm{b}_{t}),$ (9a)
-where
-$\bm{A}_{t}$ $:=\bm{\phi}(s_{t-1})\left(\bm{\phi}(s_{t-1})-\gamma\bm{\phi}(s_{t})\right)^{\top}$ (9b)
-and $\text{and }\qquad\bm{b}_{t}$ $:=\bm{\phi}(s_{t-1})r_{t-1}.$ (9c)
-After a total number of $T$ iterations, the final iteration $\bm{\theta}_{T}$ is returned as the estimator for $\bm{\theta}^{\star}$.
-Polyak-Ruppert averaging.
-The Polyak-Ruppert averaging (Polyak and Juditsky,, 1992; Ruppert,, 1998) is a stochastic approximation methodology known to reduce the variability and accelerate the convergence of algorithms by averaging across the iterates.
-In addition, with a proper choice of stepsize, the averaged iterates converge to a Gaussian distribution asymptotically, which enables valid statistical inference that holds with large samples.
-When applied to the TD algorithm, Polyak-Ruppert averaging returns an estimator of $\bm{\theta}^{*}$ of the form
-$\overline{\bm{\theta}}_{T}$ $=\frac{1}{T}\sum_{t=1}^{T}\bm{\theta}_{t},$ (10)
-with associated estimation error
-$\overline{\bm{\Delta}}_{T}=\overline{\bm{\theta}}_{T}-\bm{\theta}^{\star}.$ (11)
-TD with Polyak-Ruppert averaging has been shown to deliver significant improvements, both of computational and statistical nature.
-While there is a recent strand of non-asymptotic results established for stochastic approximation (e.g.
-Mou et al., (2021); Moulines and Bach, (2011); Nemirovski et al., (2009); Lan, (2012)), the current results on Berry-Esseen bounds tailored to policy evaluation with linear function approximation still remain inadequate.
-In the work, we derive improved consistency and Gaussian approximation rates for $\overline{\bm{\Delta}}_{T}$.
-It is important to mention that the choice of learning rates $\{\eta_{t}\}_{t\geq 0}$ in the update (9) plays a critical role in determining the statistical properties of both the original and the averaged TD estimator, including their convergence rates and asymptotic distributions.
-This work focuses on the typical choice — the polynomial-decaying stepsizes — where $\eta_{t}=\eta_{0}t^{-\alpha}$ for $\alpha\in(1/2,1)$ and $\eta_{0}$ satisfying
-$\eta_{0}<{1}/{(2\lambda_{\Sigma})}.$ (12)
-3 Main results
-While in most real-world RL scenarios the state-action-reward tuples are generated following a Markov decision process, there are applications where $i.i.d.$ samples become available, such as in game playing and autonomous driving.
-Furthermore, it is illuminating to first understand the statistical properties of this simpler setting before considering the more complicated setting of Markov samples.
-In this paper, like in most of the RL literature, we will make this simplifying assumption.
-Specifically, we will analyze the performance of the TD learning algorithm with Polyak-Ruppert averaging assuming a sample consisting of $T$ $i.i.d.$ state-action-reward tuples of the form
-$\{(s_{t},a_{t},s_{t}^{\prime},r_{t})\}_{t=1}^{T},$ (13)
-where $(s_{1},\ldots,s_{T})\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mu$ and, for eact $t$, $a_{t}\sim\pi(s_{t}),r_{t}=r(s_{t},a_{t})$ and $s_{t}^{\prime}\sim P(\cdot\mid s_{t},a_{t})$.
-Correspondingly, $\bm{A}_{t}$ and $\bm{b}_{t}$ are defined as
-$\bm{A}_{t}$ $=\bm{\phi}(s_{t})\left(\bm{\phi}(s_{t})-\gamma\bm{\phi}(s_{t}^{\prime})\right)^{\top},$
-$\text{and}\qquad\bm{b}_{t}$ $=r_{t}\bm{\phi}(s_{t}).$
-In this setting, with polynomial-decaying stepsizes, it has been shown (Polyak and Juditsky, (1992); Borkar, (2008); Fort, (2015)) that the estimation error satisfies
-[…]
-where the covariance matrix $\bm{\Lambda}^{\star}$ is given as
-$\bm{\Lambda}^{\star}=\bm{A}^{-1}\bm{\Gamma}\bm{A}^{-\top},\quad\text{with}$ (15)
-$\bm{\Gamma}=\mathbb{E}[(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}].$ (16)
-Thus, in the fixed-dimensional setting, the above result implies the rate
-$\|\overline{\bm{\Delta}}_{T}\|_{2}=O_{P}\left(\sqrt{\frac{\mathsf{Tr}(\bm{\Lambda}^{\star})}{T}}\right),$ (17)
-which may serve as a comparison term.
-Our main goal is to demonstrate how the TD algorithm can be used to construct confidence regions and simultaneous confidence intervals for $\bm{\theta}^{*}$, the linear parameters of the best value function linear approximation (see (5)).
-Towards this goal, we shall establish a high-probability convergence guarantee for $\overline{\bm{\Delta}}_{T}$, develop high-dimensional Berry-Esseen bounds over the class of convex sets, and finally build confidence regions/intervals for $\bm{\theta}^{\star}$ using the estimator $\overline{\bm{\theta}}_{T}$.
-These topics will be addressed in order.
-3.1 High-probability convergence guarantees
-We start by controlling the convergence rate of the TD iterates with Polyak-Ruppert averaging in high probability. Before stating our main results, let us first introduce additional quantities
-$\overline{\bm{\Lambda}}_{T}=\frac{1}{T}\sum_{t=1}^{T}\bm{Q}_{t}\bm{\Gamma}\bm{Q}_{t}^{\top},$ (18)
-where $\bm{\Gamma}$ is defined in (16) an, for each $t$, we set
-$\bm{Q}_{t}=\eta_{t}\sum_{j=t}^{T}\prod_{k=t+1}^{j}(\bm{I}-\eta_{k}\bm{A}).$ (19)
-The matrix $\overline{\bm{\Lambda}}_{T}$ plays the role of a non-asymptotic covariance matrix for $\overline{\bm{\theta}}_{T}$.
-Theorem 3.1.
-Consider TD with Polyak-Ruppert averaging (9) with independent samples and decaying stepsizes $\eta_{t}=\eta_{0}t^{-\alpha}$ for $\alpha\in(\frac{1}{2},1)$.
-For any $\delta\in(0,1)$, there exists $\eta_{0}=\eta_{0}(\delta)>0$, such that
-$\|\overline{\bm{\Delta}}_{T}\|_{2}$ $\lesssim\sqrt{\frac{\mathsf{Tr}(\overline{\bm{\Lambda}}_{T})}{T}\log\frac{d}{\delta}}+\widetilde{C}T^{-\frac{\alpha+1}{2}}\log\frac{T}{\delta}$
-with probability at least $1-\delta$, where $\widetilde{C}$ is a problem-related quantity independent of $T$ and $\delta$.
-The proof can be found in Appendix B.2, where the value of $\widetilde{C}$ is given in equation (75).
-Throughout the paper, for the sake of readability, we have opted to express unwieldy problem-dependent quantities that do not directly depend on $T$ in implicit form; their values can be tracked in the proofs.
-Accordingly, when writing our finite-sample bounds, we will first present the leading terms in $T$.
-In a nutshell, Theorem 3.1 states that $\overline{\bm{\theta}}_{T}$ converges to $\bm{\theta}^{*}$ in $\ell_{2}$ error at a rate $O(\sqrt{{\mathsf{Tr}(\overline{\bm{\Lambda}}_{T})}/{T}})$ up to log factor.
-It is worth comparing the high-probability sample complexity bound of Theorem 3.1 with those in the literature.
-To the best of our knowledge, the tightest results were attained by Dalal et al., (2018) for inverse-decaying stepsizes and Li et al., 2024b ; Samsonov et al., (2023) for constant stepsize as summarized in Table 1.
-With polynomial-decaying stepsizes, Theorem 3.1 is the first of its kind to achieve tight high-probability convergence guarantees up to log factors.
-Another subtle but consequential difference concerns the restriction on the step size.
-For our bound to hold, the choice of the initial stepsize $\eta_{0}$ depends on $\delta$, and is required to be sufficiently small to satisfy $\eta_{0}\sqrt{\log{1}/(\delta\eta_{0}})\leq\widetilde{C}$, where $\widetilde{C}$ is independent of $T$ and $\delta$ but may depend on the other parameters of interest.
-Though not ideal, this condition is a considerable weakening of the stepsize requirements imposed by Samsonov et al., (2023) and Li et al., 2024b , which depend on both $\delta$ and $T$; see Table 1.
-The bound of Theorem 3.1 can be made dependent on the asymptotic variance $\bm{\Lambda}^{\star}$ without affecting the rate.
-Indeed, as we will see below, the difference between $\overline{\bm{\Lambda}}_{T}$ and $\bm{\Lambda}^{\star}$ converges at the rate of $O(T^{\alpha-1})$.
-Thus, the dominant terms in Theorem 3.1 match the convergence rate (17) implied by the asymptotic distribution up to a logarithm term.
-3.2 High-dimensional Berry-Esseen bounds
-Next we derive high-dimensional Berry-Esseen bounds for TD with Polyak-Ruppert averaging.
-We first characterize the convex distance between $\sqrt{T}\overline{\bm{\Delta}}_{T}$ and $\overline{\bm{\Lambda}}_{T}^{1/2}\bm{z}$, where $\bm{z}$ is a $d$-dimensional standard gaussian.
-The proof of this result is included in Appendix B.3.
-Theorem 3.2.
-Consider the TD with Polyak-Ruppert averaging with independent samples and decaying stepsizes $\eta_{t}=\eta_{0}t^{-\alpha}$ for $\alpha\in(\frac{1}{2},1)$. Then,
-$d_{\mathsf{C}}(\sqrt{T}\overline{\bm{\Delta}}_{T},\overline{\bm{\Lambda}}_{T}^{1/2}\bm{z})\lesssim\sqrt{\frac{\eta_{0}}{(1-\alpha)(1-\gamma)\lambda_{0}}}\mathsf{Tr}(\bm{\Gamma})\|\bm{\Gamma}^{-1}\|T^{-\frac{\alpha}{2}}+\widetilde{C}\;T^{-\frac{1}{2}}\log T,$
-where $\widetilde{C}$ is a problem related quantity indepedent of $T$.
-To relate this result with the classical CLT (as in (14)), we will use the following comparison bound for the total variation distance between $\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T})$ and $\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})$; see Appendix B.4 for the proof.
-Theorem 3.3.
-Let $\bm{\Lambda}^{\star}$ and $\overline{\bm{\Lambda}}_{T}$ be defined as in (15) and (18), respectively, with the stepsizes $\eta_{t}=\eta_{0}t^{-\alpha}$, for all $t$ and any $\alpha\in(\frac{1}{2},1)$.
-Then,
-$d_{\mathsf{C}}(\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T}),\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\leq d_{\mathsf{TV}}(\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T}),\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\lesssim\frac{\sqrt{d\mathsf{cond}(\bm{\Gamma})}}{(1-\gamma)\lambda_{0}\eta_{0}}T^{\alpha-1}+\widetilde{C}\;T^{2\alpha-2},$
-where $\widetilde{C}$ is a problem related quantity indepedent of $T$.
-Comparison with existing results.
-The latest and sharpest Berry-Esseen bounds for stochastic approximation under condition (12) are du to Samsonov et al., (2024). Specifically, the authors show that
-$d_{\mathsf{C}}(\sqrt{T}\overline{\bm{\Delta}}_{T},{\bm{\Lambda}}^{\star 1/2}\bm{z})\leq\widetilde{C}\;T^{\frac{\alpha-1}{2}},$
-where $\widetilde{C}$ depend on other problem parameters.
-Upon choosing $\alpha=\frac{1}{2}$, they obtain a convergence rate of order $O(T^{-\frac{1}{4}})$.
-In contrast, the Berry-Esseen bounds of Theorem 3.2 and the comparison bound of Theorem 3.3 indicate that, selecting the optimal $\alpha=\frac{2}{3}$ leads to the improved rate
-$d_{\mathsf{C}}(\sqrt{T}\overline{\bm{\Delta}}_{T},\bm{\Lambda}^{\star 1/2}\bm{z})\leq\widetilde{C}\;T^{-\frac{1}{3}}.$
-The proof of Theorem 3.2 relies on Theorem 3.1 of Shao and Zhang, (2021), a high-dimensional Berry-Esseen bound for non-linear statistics, which is also used by Samsonov et al., (2024) to establish their main result about Gaussian approximations for the TD error, Theorem 3.4 therein.
-We are able to improve their the rate of convergence from $O(T^{\frac{1}{2}-\alpha})$ to $O(T^{-\frac{\alpha}{2}})$ by carefully handling some of the quantities appearing in the bound of Theorem 3.1 of Shao and Zhang, (2021) [^2] Specifically, the term $I_{3}$ of Equation (79) in Appendix B.3..
-This sharpening is non-trivial and may be of independent interest, as it may be potentially generalized to analyze stochastic gradient algorithms applied to non-linear functions (as of Theorem 3.4 of Shao and Zhang, (2021)).
-We leave this extension to future work.
-3.3 Online estimation of the covariance matrix
-Thus far, we have provided high-dimensional Berry-Esseen bounds to quantify the non-asymptotic discrepancy between $\overline{\bm{\Delta}}_{T}$ and its Gaussian limit.
-In order to invoke these results for constructing confidence region or confidence intervals for $\bm{\theta}^{\star}$, it is necessary to estimate the covariance matrix $\bm{\Lambda}^{\star}$.
-Two types of approaches have been proposed recently, namely plug-in estimation (Chen et al., (2020); Zhu et al., (2023); Li et al., (2022)) and bootstraping (White and White, (2010); Hanna et al., (2017); Hao et al., (2021); Ramprasad et al., (2022); Fang et al., (2018); Samsonov et al., (2024)); nontheless, the statistical and computational efficiency of these estimations can both be further enhanced.
-In this section, we introduce a computationally feasible online plug-in estimator for $\bm{\Lambda}^{\star}$, and establish its non-asymptotic rate of convergence. We will use
-$\overline{\bm{A}}_{T}=\frac{1}{T}\sum_{t=1}^{T}\bm{A}_{t};$ (20)
-$\widehat{\bm{\Gamma}}_{T}=\frac{1}{T}\sum_{t=1}^{T}(\bm{A}_{t}\overline{\bm{\theta}}_{T}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\theta}}_{T}-\bm{b}_{t})^{\top},$ (21)
-as estimators for $\bm{A}$ and $\bm{\Gamma}$ respectively, and set
-$\widehat{\bm{\Lambda}}_{T}=\overline{\bm{A}}_{T}^{-1}\widehat{\bm{\Gamma}}_{T}\overline{\bm{A}}_{T}^{-\top}$ (22)
-as the final estimator of $\bm{\Lambda}^{\star}$.
-Crucially, because of the linear nature of the problem, the estimator $\widehat{\bm{\Lambda}}_{T}$ can be evaluated in an online manner.
-Specifically, the basic properties of Kronecker products indicate
-$\textbf{vec}\left(\bm{A}_{t}\overline{\bm{\theta}}_{T}\overline{\bm{\theta}}_{T}^{\top}\bm{A}_{t}\right)=(\bm{A}_{t}\otimes\bm{A}_{t})(\overline{\bm{\theta}}_{T}\otimes\overline{\bm{\theta}}_{T});$
-$\textbf{vec}\left(\bm{A}_{t}\overline{\bm{\theta}}_{T}\overline{\bm{b}}_{t}^{\top}\right)=(\bm{A}_{t}\otimes\bm{b}_{t})\overline{\bm{\theta}}_{T};$
-$\textbf{vec}\left(\bm{b}_{t}\bm{\overline{\theta}}_{T}^{\top}\bm{A}_{t}^{\top}\right)=(\bm{b}_{t}\otimes\bm{A}_{t})\overline{\bm{\theta}}_{T};$
-$\textbf{vec}\left(\bm{b}_{t}\bm{b}_{t}^{\top}\right)=\bm{b}_{t}\otimes\bm{b}_{t}.$
-Therefore, the vectorization of $\widehat{\bm{\Gamma}}_{T}$ can be represented as
-$\textbf{vec}(\widehat{\bm{\Gamma}}_{T})$ $=\frac{1}{T}\sum_{t=1}^{T}\textbf{vec}\left(\bm{A}_{t}\overline{\bm{\theta}}_{T}\overline{\bm{\theta}}_{T}^{\top}\bm{A}_{t}-\bm{A}_{t}\overline{\bm{\theta}}_{T}\overline{\bm{b}}_{t}^{\top}-\bm{b}_{t}\bm{\overline{\theta}}_{T}^{\top}\bm{A}_{t}^{\top}+\bm{b}_{t}\bm{b}_{t}^{\top}\right)$
-$=\frac{1}{T}\sum_{t=1}^{T}\left((\bm{A}_{t}\otimes\bm{A}_{t})(\overline{\bm{\theta}}_{T}\otimes\overline{\bm{\theta}}_{T})-(\bm{A}_{t}\otimes\bm{b}_{t})\overline{\bm{\theta}}_{T}-(\bm{b}_{t}\otimes\bm{A}_{t})\overline{\bm{\theta}}_{T}+\bm{b}_{t}\otimes\bm{b}_{t}\right)$
-=($=\underset{I_{1}}{\underbrace{\left(\frac{1}{T}\sum_{t=1}^{T}\bm{A}_{t}\otimes\bm{A}_{t}\right)}}(\overline{\bm{\theta}}_{T}\otimes\overline{\bm{\theta}}_{T})-\underset{I_{2}}{\underbrace{\left(\frac{1}{T}\sum_{t=1}^{T}[\bm{A}_{t}\otimes\bm{b}_{t}+\bm{b}_{t}\otimes\bm{A}_{t}]\right)}}\overline{\bm{\theta}}_{T}$
-$+\underset{I_{3}}{\underbrace{\frac{1}{T}\sum_{t=1}^{T}\bm{b}_{t}\otimes\bm{b}_{t}}},$
-where $I_{1}$, $I_{2}$ and $I_{3}$ are sample means and can hence be calculated using online updates.
-Based on this, we propose Algorithm 1 for the calculation $\widehat{\bm{\Lambda}}_{T}$ in an online manner.
-Since Algorithm 1 iterates from $t=0$ to $t=T-1$ in a single loop and does not store the samples for each iteration, it has a space complexity is $O(Td^{4})$ and a sapce complexity of $O(d^{4})$.
-Here, the $d^{4}$ term stems from calculating the Kronecker product $\bm{A}\otimes\bm{A}$.
-Not only is the estimator $\widehat{\bm{\Lambda}}_{T}$ computationally feasible, but it also enjoys good statistical properties.
-This is made precise by the following Gaussian comparison result, whose proof is in Appendix B.5, showing that the unknown asymptotic variance $\bm{\Lambda}^{\star}$ in the Berry-Esseen bound of Theorem 3.3 can be replaced by $\widehat{\bm{\Lambda}}_{T}$ without compromising the accuracy of the approximation.
-Theorem 3.4.
-Consider the TD algorithm (9) with independent samples and decaying stepsizes $\eta_{t}=\eta_{0}t^{-\alpha}$ for all $t$ and any $\alpha\in(1/2,1)$, and let $\widehat{\bm{\Lambda}}_{T}$, $\bm{\Lambda}^{\star}$ be defined as in (22), (15) respectively.
-Then,
-$d_{\mathsf{TV}}(\mathcal{N}(\bm{0},\widehat{\bm{\Lambda}}_{T}),\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\lesssim\frac{\sqrt{\mathsf{Tr}(\bm{\Gamma})}}{\lambda_{0}(1-\gamma)}(\|\bm{\theta}^{\star}\|_{2}+1)\|\bm{\Gamma}^{-1}\|T^{-\frac{1}{3}}+o(T^{-\frac{1}{3}}).$
-with probability at least $1-3T^{-\frac{1}{3}}-\frac{\sqrt{\mathsf{Tr}(\bm{\Gamma})}}{\lambda_{0}(1-\gamma)\sqrt{T}}$.
-Above, the term $o(T^{-\frac{1}{3}})$ depends on all the parameters of the problem (and, in particular, to $\alpha$) but, as a function of $T$ alone, is of smaller order than $T^{-\frac{1}{3}}$.
-See Appendix B.5 for the proof and details.
-Comparison with Chen et al., (2020).
-In different settings, Chen et al., (2020) considered the problem of carrying out statistical inference for the true model parameters using stochastic gradient descent and assuming a strongly convex and smooth population loss function.
-In their work, a plug-in estimator of the form $\widehat{\bm{\Lambda}}_{T}^{\prime}=\overline{\bm{A}}^{-1}\widehat{\bm{\Gamma}}_{T}^{\prime}\overline{\bm{A}}^{\top}$, where
-$\widehat{\bm{\Gamma}}_{T}^{\prime}:=\frac{1}{T}\sum_{t=1}^{T}(\bm{A}_{t}{\bm{\theta}}_{t}-\bm{b}_{t})(\bm{A}_{t}{\bm{\theta}}_{t}-\bm{b}_{t})^{\top},$ (23)
-is proposed, which differs slightly from ours. The authors proved that the estimation error of $\widehat{\bm{\Lambda}}_{T}^{\prime}$ is bounded in expectatin by
-$\mathbb{E}\|\widehat{\bm{\Lambda}}_{T}^{\prime}-\bm{\Lambda}^{\star}\|\leq\|\bm{\Gamma}\|\cdot(\widetilde{C}_{1}T^{-\frac{\alpha}{2}}+\widetilde{C}_{2}T^{-\alpha}),$ (24)
-for problem-related constants $\widetilde{C}_{1}$ and $\widetilde{C}_{2}$.
-It is possible to modify the proof of Theorem 3.4 to obtain a guarantee in expectation for our problem as well, albeit in the TV metric.
-Specifically, we can show that
-$\mathbb{E}\left[d_{\mathsf{TV}}(\mathcal{N}(\bm{0},\widehat{\bm{\Lambda}}_{T}),\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\right]$ $\lesssim\widetilde{C}\|\bm{\Gamma}\|T^{-\frac{1}{2}}$ (25)
-where $\widetilde{C}$ depends on $\lambda_{0},\gamma,\eta_{0},\alpha,\bm{\theta}^{\star}$ and $d$.
-Comparing (24) with (25), we have tightened the bound from $O\left(T^{-\frac{\alpha}{2}}\right)$ to $O\left(T^{-\frac{1}{2}}\right)$, independent of $\alpha\in(1/2,1)$.
-This is a significant improvement, at least in our setting, not only because of the different metric but because a root-$T$ rate based on (24) would require $\alpha=1$, which is not permissible in our problem.
-It is noteworthy that the online algorithm 1 takes advantage of the linearity of the problem; in general, potentially non-linear settings, replacing $\bm{\theta}_{t}$ in (23) by $\overline{\bm{\theta}}_{T}$ in (21) would require storing all the samples throughout the iterative trajectory, resulting in a larger space complexity.
-Comparison with Samsonov et al., (2024).
-Using a multiplier-bootstrap technique, Samsonov et al., (2024) constructs a random estimator based on the samples $\mathcal{S}_{T}:=\{s_{t},s_{t}^{\prime}\}_{t=1}^{T}$, denoted as $\overline{\bm{\theta}}_{T}^{(b)}$.Theorem 4 and 5 in Samsonov et al., (2024)) imply that
-$d_{\mathsf{C}}(\sqrt{T}(\overline{\bm{\theta}}_{T}^{(b)}-\bm{\theta}^{\star})|\mathcal{S}_{T},\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\lesssim\frac{\|\bm{\Gamma}^{-1}\|}{[\lambda_{0}(1-\gamma)]^{\frac{3}{2}}}(2\|\bm{\theta}^{\star}\|_{2}+1)^{2}\frac{\log T}{T^{\frac{1}{4}}}$ (26)
-Compared with the bootstrap-based procedure of Samsonov et al., (2024), our plug-in estimator $\widehat{\bm{\Lambda}}_{T}$ differ in three fundamental ways.
-(i) The upper bound in Theorem 3.4 vanishes, as a function of $T$, at the rate of $O(T^{-\frac{1}{3}})$, which is faster than the $O(T^{-\frac{1}{4}}\log T)$ convergence rate obtained from the bound (26).
-(ii) The leading term in the upper bound in Theorem 3.4 entails a smaller problem-related coefficient than that of (26).
-(iii) Perhaps most importantly, in order to estimate the bootstrap distribution $\sqrt{T}(\overline{\bm{\theta}}_{T}^{(b)}-\bm{\theta}^{\star})|\mathcal{S}_{T}$ with an $O(T^{-\frac{1}{4}})$ error rate, it is necessary to store the sample trajectory $\mathcal{S}_{T}$ and generate $O(T^{\frac{1}{2}})$ perturbed trajectories.
-Consequently, the bootstrap procedure has a time complexity of order $O(T^{\frac{3}{2}}d^{2})$ and a space complexity of order $O(T)$.
-When $d\ll T$, this results in a significant computational border compared to our online procedure of Algorithm 1.
-3.4 Statistical inference for $\bm{\theta}^{\star}$
-By combining the results from Sections 3.2 and 3.3, we obtain a general methodology for constructing convex confidence sets for $\bm{\theta}^{\star}$ with guaranteed finite sample coverage, as shown in our final result, whose proof is in Appendix B.6.
-Theorem 3.5.
-Consider TD with Polyak-Ruppert averaging (9) with independent samples and decaying stepsizes $\eta_{t}=\eta_{0}t^{-\alpha}$, for all $t$ and with $\alpha=2/3$.For a given significance level $\delta\in(0,1)$, it holds that
-$\mathbb{P}(\bm{\theta}^{\star}\in\mathcal{C}_{\delta})$ $\geq 1-\delta-\frac{\sqrt{\mathsf{Tr}(\bm{\Gamma})}}{\lambda_{0}(1-\gamma)}(2\|\bm{\theta}^{\star}\|_{2}+1)\|\bm{\Gamma}^{-1}\|T^{-\frac{1}{3}}-\frac{\sqrt{d\mathsf{cond}(\bm{\Gamma})}}{2(1-\gamma)\lambda_{0}\eta_{0}}T^{-\frac{1}{3}}-o(T^{-\frac{1}{3}}),$
-uniformly over all convex sets $\mathcal{C}_{\delta}\subset\mathbb{R}^{d}$ satisfying
-$\mathbb{P}_{\bm{z}\sim\mathcal{N}(\bm{0},\bm{I}_{d})}\left(\overline{\bm{\theta}}_{T}-\frac{\widehat{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}}{\sqrt{T}}\in\mathcal{C}_{\delta}\bigg{|}\overline{\bm{\theta}}_{T},\widehat{\bm{\Lambda}}_{T}\right)\geq 1-\delta,$ (27)
-where the variance matrix estimator $\widehat{\bm{\Lambda}}_{T}$ is given in (22).
-It is important to highlight that the above coverage guarantee holds uniformly over any choice of the convex set $\mathcal{C}_{\delta}$ meeting condition (27).
-In particular, this class includes all hyper-rectangles (cartesian products of univariate intervals), returning simultaneous confidence intervals that are especially useful due to their simplicity and interpretability.
-The construction of $\mathcal{C}_{\delta}$ requires computing appropriate quantiles of the $\mathcal{N}(\overline{\bm{\theta}}_{T},\widehat{\bm{\Lambda}}_{T})$ distribution, conditionally on $\overline{\bm{\theta}}_{T}$ and $\widehat{\bm{\Lambda}}_{T}$, depending on the choice of convex set.
-For the case of hyper-rectangles, this task is reduced to finding the quantile of the maximum of Gaussian variables.
-Input: MDP transition kernel $\mathcal{P}$, stationary distribution $\mu$, discount factor $\gamma$, feature matrix $\bm{\Phi}$, initial stepsize $\eta_{0}$, stepsize decaying parameter $\alpha\in(0.5,1)$, sample size $T$
-1 Initialize $\bm{\theta}=\bm{0}$, $\overline{\bm{\theta}}=\bm{0}$, $\overline{\bm{A}}=\bm{0}$, $\overline{\bm{AA}}=\bm{0}$, $\overline{\bm{Ab}}=\bm{0}$, $\overline{\bm{bb}}=\bm{0}$;
-2 for $t=0:(T-1)$ do
-3       sample $s\sim\mu$, $s^{\prime}\sim\mathcal{P}(\cdot\mid s)$;
-4       compute
-$\bm{A}\leftarrow\bm{\phi}(s)[\bm{\phi}(s)-\gamma\bm{\phi}(s^{\prime})];$
-$\bm{b}\leftarrow r(s)\bm{\phi}(s);$
-5       perform TD update
-$\bm{\theta}\leftarrow\bm{\theta}-\eta_{0}t^{-\alpha}(\bm{A}\bm{\theta}-\bm{b});$
-6       perform online updates
-$\overline{\bm{\theta}}\leftarrow\overline{\bm{\theta}}+\frac{1}{t}(\bm{\theta}-\overline{\bm{\theta}});$
-$\overline{\bm{A}}\leftarrow\overline{\bm{A}}+\frac{1}{t}(\bm{A}-\overline{\bm{A}});$
-$\overline{\bm{AA}}\leftarrow\overline{\bm{AA}}+\frac{1}{t}(\bm{A}\otimes\bm{A}-\overline{\bm{AA}});$
-$\overline{\bm{Ab}}\leftarrow\overline{\bm{Ab}}+\frac{1}{t}(\bm{A}\otimes\bm{b}+\bm{b}\otimes\bm{A}-\overline{\bm{Ab}});$
-$\overline{\bm{bb}}\leftarrow\overline{\bm{bb}}+\frac{1}{t}(\bm{b}\otimes\bm{b}-\overline{\bm{bb}});$
-7compute
-$\widehat{\bm{\Gamma}}\leftarrow\overline{\bm{AA}}(\overline{\bm{\theta}}\otimes\overline{\bm{\theta}})-2\overline{\bm{Ab}}(\overline{\bm{\theta}})+\overline{\bm{bb}};$
-8 reshape
-$\widehat{\bm{\Gamma}}\leftarrow\mathsf{reshape}(\widehat{\bm{\Gamma}},(d\times d))$
-compute
-$\widehat{\bm{\Lambda}}\leftarrow\overline{\bm{A}}^{-1}\widehat{\bm{\Gamma}}\overline{\bm{A}}^{-\top}$
-Output: averaged TD estimator $\overline{\bm{\theta}}$, covariance matrix estimator $\widehat{\bm{\Lambda}}$
-Algorithm 1 Online algorithm for computing $\overline{\bm{\theta}}_{T}$ and $\widehat{\bm{\Lambda}}_{T}$
-4 Numerical experiments
-We illustrate our theoretical results with some numerical experiments.
-In order to test our proposed framework for statistical inference, we consider a member of a family of MDPs that are designed to be difficult to distinguish among each other.
-The specifics of the MDP follow the precedent of Li et al., 2024b and we refer readers to Appendix D for the construction details.
-We examine the high-probability convergence of the norm of TD errors, the rate by which they converge to their asymptotic Gaussian distributions, the precision of our estimator for the covariance matrix, and the validity of our statistical inference procedure.
-While the full details and results of the experiments are postponed to Appendix D, we demonstrate some important findings in the simulations that verify our theoretical conclusions.
-4.1 Uncertainty quantification with different choices of $\alpha$
-In order to explore the role played by the parameter $\alpha$ that decides the decaying speed of the stepsizes, we run TD iterates using different choices of $\alpha$ on the MDP with feature dimension $d=3$, record the $95\%$ quantile of $\|\overline{\bm{\Delta}}_{T}\|_{2}^{2}$ and measure the difference between the empirical and asymptotic distributions of $\overline{\bm{\Delta}}_{T}$  [^3] For computational simplicity, this difference between distributions is not measured by the distance on convex sets as we used in Theorems 3.2 and 3.3.
-Details are showin in Appendix D.
-.
-Figure 2(a) shows the decay of the $95\%$ quantile of the $L_{2}$ norm of $\overline{\bm{\Delta}}_{T}$; the slope of this log-log plot verifies our theoretical claim in Theorem 3.1 that $\|\overline{\bm{\Delta}}_{T}\|_{2}$ converges by the rate of $O(T^{-\frac{1}{2}})$ with high probability.
-Notably, the convergence rate is similar for different choices of $\alpha>\frac{1}{2}$ when $T$ is large.
-Figure 2 (b) reveals that the empirical distribution of TD estimation error converges to the asymptotic Gaussian distribution slowly when $\alpha$ is set to ${1}/{2}$ as suggested by Samsonov et al., (2024).
-Both plots appear to indicate that choosing $\alpha=\frac{1}{2}$ yields a slower converging rate both in $\ell_{2}$ norm and in distribution.
-Refer to caption Refer to caption
-Figure 2: (a) $95\%$ quantile of $\|\overline{\bm{\Delta}}_{T}\|_{2}^{2}$, and (b) distance between empirical and asymptotic distributions of the TD estimation error $\overline{\bm{\Delta}}_{T}$, for an MDP with $d=3$ with different choices of $\alpha$.
-4.2 Confidence intervals
-Furthermore, we validate our statistical inference framework on the MDP with $d=3$ and TD learning with $\alpha=\frac{2}{3}$.
-We run 1000 independent trials of Algorithm 1, and at each iteration $T$, we construct both the individual and the simultaneous confidence intervals.
-Specifically, the individual confidence intervals for the three entries of $\bm{\theta}^{\star}$ are constructed by
-$\mathcal{C}_{j}=\left[\overline{\bm{\theta}}_{T}-\Phi(1-\delta/2)\cdot\sqrt{\frac{(\widehat{\bm{\Lambda}}_{T})_{jj}}{T}},\overline{\bm{\theta}}_{T}+\Phi(1-\delta/2)\cdot\sqrt{\frac{(\widehat{\bm{\Lambda}}_{T})_{jj}}{T}}\right],\quad\text{for}\quad j=1,2,3.$
-Here, $\Phi(1-\delta/2)$ is the $1-\frac{\delta}{2}$ quantile of the standard Gaussian distribution.
-In our experiment, we take $\delta=0.05$ and therefore $\Phi(1-\delta/2)=0.96$.
-Figure 3 illustrates the emergence of $\overline{\bm{\theta}}_{T}$ and $\{\mathcal{C}_{j}\}_{j=1,2,3}$ in one single trial.
-As $T$ grows larger, the three entries of the point estimate $\overline{\bm{\theta}}_{T}$ (blue curves) converges to the corresponding entries of the optimal coefficients $\bm{\theta}^{\star}$ (black dotted lines), while the individual confidence intervals $\{\mathcal{C}_{j}\}_{j=1,2,3}$ (blue shades) also concentrate around $\bm{\theta}^{\star}$.
-Figure 3: The emergence of the point estimate $\overline{\bm{\theta}}_{T}$ (blue curves) and individual confidence intervals (blue shades) on the three entries of the optimal coefficients $\bm{\theta}^{\star}$ (black dotted lines) for the MDP with $d=3$, using TD with Polyak-Ruppert averaging, $\alpha=\frac{2}{3}$
-Meanwhile, the simultaneous confidence intervals are constructed as
-$\mathcal{C}=\{\bm{\theta}:\|\bm{\theta}-\overline{\bm{\theta}}_{T}\|_{\infty}\leq x^{\star}\},\quad\text{where}$
-$x^{\star}:=\sup\left\{x:\mathbb{P}_{z\sim\mathcal{N}(\bm{0},\bm{I}_{3})}\left(\left\|\widehat{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}\right\|_{\infty}\leq x\right)\leq 1-\delta\right\}.$
-Essentially, $x^{\star}$ is the $1-\delta$ quantile of the $\ell_{\infty}$ norm of the $\mathcal{N}(\bm{0},\widehat{\bm{\Lambda}}_{T})$ random variable.
-Calculating this quantity is a non-trivial task, so we use $1000$ independent simulations from $\mathcal{N}(\bm{0},\widehat{\bm{\Lambda}}_{T})$ and take the $1-\delta$ quantile of the $\ell_{\infty}$ norm of the samples as an estimate.
-Figure 4 shows the coverage rates of the independent and simultaneous confidence intervals, defined as the percentage of times when the optimal coefficient $\bm{\theta}^{\star}$ (or its entries) are covered by the confidence intervals in the $1000$ independent trials.
-As is indicated by the plot, the coverage rates all approach the target confidence level of $1-\delta=95\%$ when $T$ grows large.
-Figure 4: The emergence coverage rates of independent and simultaneous confidence intervals on the three entries of the optimal coefficients $\bm{\theta}^{\star}$ of the MDP with $d=3$, using TD with Polyak-Ruppert averaging, $\alpha=\frac{2}{3}$.
-Black dotted line represents the target confidence level of $95\%$; blue, orange and green curves represent the coverage rates of the individual confidence intervals $\mathcal{C}_{1}$, $\mathcal{C}_{2}$ and $\mathcal{C}_{3}$ correspondingly; the red curve represents the coverage rate of the confidence region (simultaneous confidence intervals) $\mathcal{C}$.
-5 Discussion
-We develop high-probability convergence gaurantees and Berry-Esseen bounds for TD learning with linear function approximation and $i.i.d.$ samples in a non-asymptotic manner.
-We also propose an efficient way to construct a covariance estimator, thus paving the way for statistical inference on policy evaluation.
-One may naturally be interested in whether and how our techniques can be generalized to the more realistic scenario of Markov samples, where the TD estimation error is also gauranteed to be asymptotically normal (Fort, (2015)).
-On this front, the answer is multifaceted.
-For the high-probability convergence guarantee, our induction-based technique used in the proof of Theorem 3.1 may be generalized to Markovian samples under appropriate assumptions on the mixing time of the Markov chain.
-As for the CLT, a cornerstone result invoked in the proof of Theorem 3.2 holds only with $i.i.d.$ samples, and a new framework is necessary for generalizing the Berry-Esseen bound to Markovian samples.
-Furthermore, it would be of interest to derive specialized Berry-Esseen bounds for the sub-class of hyper-rectangles.
-In this setting, we expect to find a better dependence on the dimension $d$ of the feature space, possibly of poly-logarithmic, as opposed to polynmial, order.
-Finally, the plug-in estimator $\widehat{\bm{\Lambda}}_{T}$ may be generalized to a batch-mean estimator.
-Its efficiency, however, needs to be further scrutinized, as the variance estimation problem for samples generated from a Markov chain remains still largely unsolved.
-Acknowledgements
-W.
-Wu and A.Rinaldo are supported in part by NIH under Grant R01 NS121913.
-Gen Li is supported in part by the Chinese University of Hong Kong Direct Grant for Research.
-Y.
-Wei is supported in part by the NSF grants DMS-2147546/2015447, CAREER award DMS-2143215, CCF-2106778, CCF-2418156 and the Google Research Scholar Award.
-Appendix A Preliminary facts
-In this section, we summarize some preliminary facts that would be useful in the proof of the theoretical results in this paper.
-A.1 Useful algebriac facts
-The following lemma captures some basic properties in linear algebra regarding the relaitionship between the Frobenius and operator norms of matrices.
+For a matrix sequence $\{\bm{F}_{t}\}_{t\in\mathbb{N}}$, a number sequence $\{f_{t}\}_{t\in\mathbb{N}}$ and a fixed matrix $\bm{X}$, if there exists a constant $\widetilde{C}$ independent of $t$ and $\bm{X}$ such that
+$\|\bm{F}_{t}\|\leq\widetilde{C}f_{t}\|\bm{X}\|,\quad\text{and}\quad\mathsf{Tr}(\bm{F}_{t})\leq\widetilde{C}f_{t}\mathsf{Tr}(\bm{X})$ (22)
+hold simultaneously, we denote for simplicity $\bm{F}_{t}=O(f_{t})\bm{X}$.
+The following lemma captures some basic properties in linear algebra regarding the relationship between Frobenius and operator norms of matrices.
 Lemma A.1.
@@
   1. 1.
-    For any sequence of matrices $\bm{X}_{1},\bm{X}_{2},\ldots,\bm{X}_{n}\in\mathbb{R}^{d}$:
+    For any sequence of matrices $\bm{X}_{1},\bm{X}_{2},\ldots,\bm{X}_{n}\in\mathbb{R}^{d\times d}$:
     $\left\|\prod_{i=1}^{n}\bm{X}_{i}\right\|_{\mathsf{F}}\leq\|\bm{X}_{j}\|_{\mathsf{F}}\prod_{\begin{subarray}{c}1\leq i\leq n\\ i\neq j\end{subarray}}\|\bm{X}_{i}\|.$ (23)
   2. 2.
-    For any matrix $\bm{X}\in\mathbb{R}^{d}$,
+    For any matrix $\bm{X}\in\mathbb{R}^{d\times d}$,
     $\|\bm{X}\|_{\mathsf{F}}\leq\sqrt{d}\|\bm{X}\|.$ (24)
-The following lemma, introduced as Lemma 4.1 in Chen et al., (2020), bounds the difference between inverse matrices and would be useful for our analysis of the plug-in variance estimator.
 Lemma A.2.
-Let $\bm{X},\bm{Y}$ be two invertible matrices satifying $\|\bm{X}^{-1}(\bm{Y}-\bm{X})\|\leq\frac{1}{2}$. Then the difference between $\bm{Y}^{-1}$ and $\bm{X}^{-1}$ can be bounded by
+For any matrix $\bm{X}\in\mathbb{R}^{d\times d}$ and a symmetric matrix $\bm{Y}\in\mathbb{S}^{d\times d}$, if $\bm{Y}\succeq\bm{0}$, then
+$\mathsf{Tr}(\bm{XY})=\mathsf{Tr}(\bm{YX})\leq\|\bm{X}\|\mathsf{Tr}(\bm{Y}).$
+Lemma A.2 directly imply that for any matrix sequence $\{\bm{F}_{t}\}_{t\in\mathbb{N}}$ on $\mathbb{R}^{d\times d}$ and a fixed matrix $\bm{X}\succeq\bm{0}$, it can be guaranteed that
+$\bm{F}_{t}\bm{X}=O(\|\bm{F}_{t}\|)\bm{X},\quad\text{and}\quad\bm{X}\bm{F}_{t}=O(\|\bm{F}_{t}\|)\bm{X}.$ (25)
+The following lemma, due to Chen et al. (2020)[Lemma 4.1], bounds the difference between the inverse matrices and would be useful for our analysis of the plug-in variance estimator.
+Lemma A.3.
+Let $\bm{X},\bm{Y}$ be two invertible matrices that satisfy $\|\bm{X}^{-1}(\bm{Y}-\bm{X})\|\leq\frac{1}{2}$. Then the difference between $\bm{Y}^{-1}$ and $\bm{X}^{-1}$ can be bounded by
 $\left\|\bm{Y}^{-1}-\bm{X}^{-1}\right\|\leq 2\|\bm{Y}-\bm{X}\|\|\bm{X}^{-2}\|.$
-The following lemma reveals several critical algebriac features of the matrix $\bm{A}$.
+The following lemma reveals several critical algebraic features of the matrix $\bm{A}$.
 Lemma A.4.
-Let $\bm{A}_{t},\bm{A}$ be defined as in (9b) and (6a) respectively, and $\bm{\Sigma}$, $\lambda_{0}$ and $\lambda_{\Sigma}$ be defined as in (7) and (8). Then the following features hold:
+Let $\bm{A}_{t},\bm{A}$ be defined as in (8) and (4) respectively, and $\bm{\Sigma}$, $\lambda_{0}$ and $\lambda_{\Sigma}$ be defined as in (5).Then the following features hold:
 $2(1-\gamma)\bm{\Sigma}\preceq\bm{A}+\bm{A}^{\top}\preceq 2(1+\gamma)\bm{\Sigma};$ (26)
 $\min_{1\leq i\leq d}\mathsf{Re}(\lambda_{i}(\bm{A}))\geq(1-\gamma)\lambda_{0};$ (27)
-$\mathbb{E}[\bm{A}_{t}^{\top}\bm{A}_{t}]\leq\bm{A}+\bm{A}^{\top};$ (32)
-$\left\|\bm{I}-\eta\bm{A}\right\|\leq 1-\frac{1-\gamma}{2}\lambda_{0}\eta,\quad\forall\eta\in\left(0,\frac{1+\gamma}{2\lambda_{\Sigma}}\right).$ (33)
+$\mathbb{E}[\bm{A}_{t}^{\top}\bm{A}_{t}]\preceq\bm{A}+\bm{A}^{\top};$ (28)
+$\bm{A}^{\top}\bm{A}\preceq\lambda_{\Sigma}(\bm{A}+\bm{A}^{\top});$ (29)
+$\left\|\bm{I}-\eta\bm{A}\right\|\leq 1-\frac{1-\gamma}{2}\lambda_{0}\eta,\quad\forall\eta\in\left(0,\frac{1}{2\lambda_{\Sigma}}\right);$ (30)
+$\|\bm{A}^{-1}\|\leq\frac{1}{\lambda_{0}(1-\gamma)}.$ (31)
 Proof: We address these properties in order.
@@
 By Cauchy-Schwartz inequality, the second term is bounded by
-$-\left(\bm{\phi}(s)\bm{\phi}(s)^{\top}+\bm{\phi}(s^{\prime})\bm{\phi}(s^{\prime})^{\top}\right)$
-$\preceq\bm{\phi}(s)\bm{\phi}(s^{\prime})^{\top}+\bm{\phi}(s^{\prime})\bm{\phi}(s)^{\top}$
-$\preceq\bm{\phi}(s)\bm{\phi}(s)^{\top}+\bm{\phi}(s^{\prime})\bm{\phi}(s^{\prime})^{\top};$
+$-\left(\bm{\phi}(s)\bm{\phi}(s)^{\top}+\bm{\phi}(s^{\prime})\bm{\phi}(s^{\prime})^{\top}\right)\preceq\bm{\phi}(s)\bm{\phi}(s^{\prime})^{\top}+\bm{\phi}(s^{\prime})\bm{\phi}(s)^{\top}\preceq\bm{\phi}(s)\bm{\phi}(s)^{\top}+\bm{\phi}(s^{\prime})\bm{\phi}(s^{\prime})^{\top};$
 Hence, $\bm{A}+\bm{A}^{\top}$ is bounded above by
@@
 $=\mathbb{E}\|\bm{\phi}(s_{t})\|_{2}^{2}[(\bm{\phi}(s_{t})-\gamma\bm{\phi}(s_{t}^{\prime}))(\bm{\phi}(s_{t})-\gamma\bm{\phi}(s_{t}^{\prime}))^{\top}]$
-$\leq\max_{s\in\mathcal{S}}\|\bm{\phi}(s)\|_{2}^{2}\mathbb{E}[(\bm{\phi}(s_{t})-\gamma\bm{\phi}(s_{t}^{\prime}))(\bm{\phi}(s_{t})-\gamma\bm{\phi}(s_{t}^{\prime}))^{\top}]$
-$\leq\mathbb{E}[\bm{\phi}(s_{t})\bm{\phi}(s_{t})^{\top}]-\gamma\mathbb{E}[\bm{\phi}(s_{t})\bm{\phi}(s_{t}^{\prime})^{\top}+\bm{\phi}(s_{t}^{\prime})\bm{\phi}(s_{t})^{\top}]+\gamma^{2}\mathbb{E}[\bm{\phi}(s_{t}^{\prime})\bm{\phi}(s_{t}^{\prime})^{\top}]$
-$\leq 2\mathbb{E}[\bm{\phi}(s_{t})\bm{\phi}(s_{t})^{\top}]-\gamma\mathbb{E}[\bm{\phi}(s_{t})\bm{\phi}(s_{t}^{\prime})^{\top}+\bm{\phi}(s_{t}^{\prime})\bm{\phi}(s_{t})^{\top}]$
+$\preceq\max_{s\in\mathcal{S}}\|\bm{\phi}(s)\|_{2}^{2}\mathbb{E}[(\bm{\phi}(s_{t})-\gamma\bm{\phi}(s_{t}^{\prime}))(\bm{\phi}(s_{t})-\gamma\bm{\phi}(s_{t}^{\prime}))^{\top}]$
+$\preceq\mathbb{E}[\bm{\phi}(s_{t})\bm{\phi}(s_{t})^{\top}]-\gamma\mathbb{E}[\bm{\phi}(s_{t})\bm{\phi}(s_{t}^{\prime})^{\top}+\bm{\phi}(s_{t}^{\prime})\bm{\phi}(s_{t})^{\top}]+\gamma^{2}\mathbb{E}[\bm{\phi}(s_{t}^{\prime})\bm{\phi}(s_{t}^{\prime})^{\top}]$
+$\preceq 2\mathbb{E}[\bm{\phi}(s_{t})\bm{\phi}(s_{t})^{\top}]-\gamma\mathbb{E}[\bm{\phi}(s_{t})\bm{\phi}(s_{t}^{\prime})^{\top}+\bm{\phi}(s_{t}^{\prime})\bm{\phi}(s_{t})^{\top}]$
 $=\bm{A}+\bm{A}^{\top},$
-Notice that the inequality on the fourth line follows from the assumption that $\max_{s\in\mathcal{S}}\|\bm{\phi}(s)\|_{2}\leq 1$ and the inequality on the fifth line follows from the fact that since $s_{t}$ and $s_{t}^{\prime}$ and both drawn from the stationary distribution, $\mathbb{E}[\phi(s_{t}^{\prime})\phi(s_{t}^{\prime})^{\top}]=\mathbb{E}[\phi(s_{t})\phi(s_{t})^{\top}]$.
+Notice that the inequality on the fourth line follows from the assumption that $\max_{s\in\mathcal{S}}\|\bm{\phi}(s)\|_{2}\leq 1$ and the inequality on the fifth line follows from the fact that since $s_{t}$ and $s_{t}^{\prime}$ and both drawn from the stationary distribution, $\mathbb{E}[\bm{\phi}(s_{t}^{\prime})\bm{\phi}(s_{t}^{\prime})^{\top}]=\mathbb{E}[\bm{\phi}(s_{t})\bm{\phi}(s_{t})^{\top}]$.
+Proof of Equation (29).
+It suffices to show that for any $\bm{x}\in\mathbb{R}^{d}$,
+$\bm{x}^{\top}\bm{A}^{\top}\bm{Ax}\leq\lambda_{\Sigma}\bm{x}^{\top}(\bm{A}+\bm{A}^{\top})\bm{x}.$
+By definition, the left hand side of this inequality is featured by
+$\bm{x}^{\top}\bm{A}^{\top}\bm{Ax}=\|\bm{Ax}\|_{2}^{2}=\left\|\mathbb{E}[\bm{\phi}(s)(\bm{\phi}(s)-\gamma\bm{\phi}(s^{\prime}))^{\top}\bm{x}]\right\|_{2}^{2},$
+where the expectation is taken over the distribution $s\sim\mu,s^{\prime}\sim P(\cdot\mid s)$. For simplicity, denote
+$f(s)=\mathbb{E}_{s^{\prime}\sim P(\cdot\mid s)}[(\bm{\phi}(s)-\gamma\bm{\phi}(s^{\prime}))^{\top}\bm{x}];$
+then the law of total expectations yields
+$\bm{x}^{\top}\bm{A}^{\top}\bm{Ax}=\left\|\mathbb{E}[f(s)\bm{\phi}(s)]\right\|_{2}^{2}.$
+In what follows, we aim to show that
+$\left\|\mathbb{E}[f(s)\bm{\phi}(s)]\right\|_{2}^{2}\leq\lambda_{\Sigma}\mathbb{E}[f^{2}(s)].$ (32)
+Observe that for any vector $\bm{y}\in\mathbb{R}^{d}$, it can always be guaranteed that
+$\mathbb{E}[f(s)-\bm{\phi}^{\top}(s)\bm{y}]^{2}\geq 0.$
+A direct expansion yields
+00 0 $\leq\mathbb{E}[f^{2}(s)]-2\mathbb{E}[f(s)\bm{\phi}^{\top}(s)\bm{y}]+\bm{y}^{\top}\mathbb{E}[\bm{\phi}(s)\bm{\phi}(s)^{\top}]\bm{y}$
+$=\mathbb{E}[f^{2}(s)]-2\mathbb{E}[f(s)\bm{\phi}^{\top}(s)\bm{y}]+\bm{y}^{\top}\bm{\Sigma}\bm{y}$
+$\leq\mathbb{E}[f^{2}(s)]-2\mathbb{E}[f(s)\bm{\phi}^{\top}(s)\bm{y}]+\lambda_{\Sigma}\|\bm{y}\|_{2}^{2},\quad\forall\bm{y}\in\mathbb{R}^{d}.$
+By taking
+$y=\frac{\mathbb{E}[f(s)\bm{\phi}(s)]}{\lambda_{\Sigma}},$
+we obtain
+$0\leq\mathbb{E}[f^{2}(s)]-\frac{\left\|\mathbb{E}[f(s)\bm{\phi}(s)]\right\|_{2}^{2}}{\lambda_{\Sigma}},$
+and (32) follows immediately. In other words, it can be guaranteed for every $\bm{x}\in\mathbb{R}^{d}$ that
+$\bm{x}^{\top}\bm{A}^{\top}\bm{Ax}\leq\lambda_{\Sigma}\mathbb{E}[f^{2}(s)].$
+and we now proceed to bounding $\mathbb{E}[f^{2}(s)]$. The Jensen’s inequality implies
+$\mathbb{E}[f^{2}(s)]$ $=\mathbb{E}_{s\sim\mu}\left\{\bm{x}^{\top}\mathbb{E}_{s^{\prime}\sim P(\cdot\mid s)}[\bm{\phi}(s)-\gamma\bm{\phi}(s^{\prime})]\mathbb{E}_{s^{\prime}\sim P(\cdot\mid s)}[\bm{\phi}(s)-\gamma\bm{\phi}(s^{\prime})]^{\top}\bm{x}\right\}$
+$\leq\mathbb{E}_{s\sim\mu}\left\{\bm{x}^{\top}\mathbb{E}_{s^{\prime}\sim P(\cdot\mid s)}[\bm{\phi}(s)-\gamma\bm{\phi}(s^{\prime})][\bm{\phi}(s)-\gamma\bm{\phi}(s^{\prime})]^{\top}\bm{x}\right\}$
+$=\bm{x}^{\top}\mathbb{E}\left\{[\bm{\phi}(s)-\gamma\bm{\phi}(s^{\prime})][\bm{\phi}(s)-\gamma\bm{\phi}(s^{\prime})]^{\top}\right\}\bm{x}$
+$\leq\bm{x}^{\top}(\bm{A}+\bm{A}^{\top})\bm{x},$
+where the last inequality follows from the same logic as in the proof of Equation (28). Equation (29) follows immediately.
 Proof of Equation (30).
 We firstly notice that
-$(\bm{I}-\eta\bm{A})^{\top}(\bm{I}-\eta\bm{A})=\bm{I}-\eta(\bm{A}+\bm{A}^{\top})+\eta^{2}\bm{A}^{\top}\bm{A},$
-where $\bm{A}^{\top}\bm{A}$ can be bounded above by
-$\bm{A}^{\top}\bm{A}$ $\preceq\bm{A}^{\top}\bm{A}+\bm{A}\bm{A}^{\top}\preceq\frac{1}{2}(\bm{A}+\bm{A}^{\top})^{2}\preceq(1+\gamma)\lambda_{\Sigma}(\bm{A}+\bm{A}^{\top}),$
-where we applied Equation (30) in the last inequality. Hence,
-$(\bm{I}-\eta\bm{A})^{\top}(\bm{I}-\eta\bm{A})\preceq\bm{I}-(\eta-(1+\gamma)\lambda_{\Sigma}\eta^{2})(\bm{A}+\bm{A}^{\top}).$
-For any $\eta\in(0,\frac{1+\gamma}{2\lambda_{\Sigma}})$, this upper bound is controlled by
-$(\bm{I}-\eta\bm{A})^{\top}(\bm{I}-\eta\bm{A})\preceq\bm{I}-\frac{\eta}{2}(\bm{A}+\bm{A}^{\top})\preceq(1-(1-\gamma)\lambda_{0})\bm{I}.$
+$(\bm{I}-\eta\bm{A})^{\top}(\bm{I}-\eta\bm{A})=\bm{I}-\eta(\bm{A}+\bm{A}^{\top})+\eta^{2}\bm{A}^{\top}\bm{A}.$
+By invoking (29), we obtain
+$(\bm{I}-\eta\bm{A})^{\top}(\bm{I}-\eta\bm{A})\preceq\bm{I}-(\eta-\lambda_{\Sigma}\eta^{2})(\bm{A}+\bm{A}^{\top}).$
+For any $\eta\in(0,\frac{1}{2\lambda_{\Sigma}})$, this upper bound is controlled by
+$(\bm{I}-\eta\bm{A})^{\top}(\bm{I}-\eta\bm{A})\preceq\bm{I}-\frac{\eta}{2}(\bm{A}+\bm{A}^{\top})\preceq(1-(1-\gamma)\lambda_{0}\eta)\bm{I}.$
 Hence, the spectral norm of $\bm{I}-\eta\bm{A}$ is controlled by
-$\|\bm{I}-\eta\bm{A}\|\leq\sqrt{1-(1-\gamma)\lambda_{0}}\leq 1-\frac{1-\gamma}{2}\lambda_{0}\eta.$
+$\|\bm{I}-\eta\bm{A}\|\leq\sqrt{1-(1-\gamma)\lambda_{0}\eta}\leq 1-\frac{1-\gamma}{2}\lambda_{0}\eta.$
+Proof of Equation (31).
+For any unit vector $\bm{x}\in\mathbb{R}^{d}$, we observe
+$\bm{x}^{\top}\bm{\Sigma}^{-\frac{1}{2}}\bm{A}\bm{\Sigma}^{-\frac{1}{2}}\bm{x}$ $=\bm{x}^{\top}\bm{\Sigma}^{-\frac{1}{2}}\mathbb{E}[\bm{\phi}(s)\bm{\phi}^{\top}(s)]\bm{\Sigma}^{-\frac{1}{2}}\bm{x}-\gamma\bm{x}^{\top}\bm{\Sigma}^{-\frac{1}{2}}\mathbb{E}[\bm{\phi}(s)\bm{\phi}^{\top}(s^{\prime})]\bm{\Sigma}^{-\frac{1}{2}}\bm{x}$
+$=\bm{x}^{\top}\bm{\Sigma}^{-\frac{1}{2}}\bm{\Sigma}\bm{\Sigma}^{-\frac{1}{2}}\bm{x}-\gamma\bm{x}^{\top}\bm{\Sigma}^{-\frac{1}{2}}\mathbb{E}[\bm{\phi}(s)\bm{\phi}^{\top}(s^{\prime})]\bm{\Sigma}^{-\frac{1}{2}}\bm{x}$
+$=1-\gamma\mathbb{E}[(\bm{x}^{\top}\bm{\Sigma}^{-\frac{1}{2}}\bm{\phi}(s))(\bm{x}^{\top}\bm{\Sigma}^{-\frac{1}{2}}\bm{\phi}(s^{\prime}))]$
+$\leq 1-\gamma\sqrt{\mathbb{E}[(\bm{x}^{\top}\bm{\Sigma}^{-\frac{1}{2}}\bm{\phi}(s))^{2}]\mathbb{E}[(\bm{x}^{\top}\bm{\Sigma}^{-\frac{1}{2}}\bm{\phi}(s^{\prime}))^{2}]}$
+$=1-\gamma\sqrt{(\bm{x}^{\top}\bm{\Sigma}^{-\frac{1}{2}}\mathbb{E}[\bm{\phi}(s)\bm{\phi}^{\top}(s)]\bm{\Sigma}^{-\frac{1}{2}}\bm{x})\cdot(\bm{x}^{\top}\bm{\Sigma}^{-\frac{1}{2}}\mathbb{E}[\bm{\phi}(s^{\prime})\bm{\phi}^{\top}(s^{\prime})]\bm{\Sigma}^{-\frac{1}{2}}\bm{x})}$
+$=1-\gamma\bm{x}^{\top}\bm{\Sigma}^{-\frac{1}{2}}\bm{\Sigma}\bm{\Sigma}^{-\frac{1}{2}}\bm{x}=1-\gamma,$
+where the fourth line is an application of the Cauchy-Schwartz inequality. Consequently, for any unit vector $\bm{x}\in\mathbb{R}^{d}$, it can be guaranteed that
+$\|\bm{\Sigma}^{-\frac{1}{2}}\bm{A}\bm{\Sigma}^{-\frac{1}{2}}\bm{x}\|_{2}$ $=\|\bm{x}\|_{2}\|\bm{\Sigma}^{-\frac{1}{2}}\bm{A}\bm{\Sigma}^{-\frac{1}{2}}\bm{x}\|_{2}$
+$\geq\bm{x}^{\top}\bm{\Sigma}^{-\frac{1}{2}}\bm{A}\bm{\Sigma}^{-\frac{1}{2}}\bm{x}\geq 1-\gamma,$
+where the inequality, again, is an application of the Cauchy-Schwartz inequality.
+This essentially means that $\|\bm{\Sigma}^{\frac{1}{2}}\bm{A}^{-1}\bm{\Sigma}^{\frac{1}{2}}\|\leq\frac{1}{1-\gamma}$, and therefore
+$\|\bm{A}^{-1}\|\leq\frac{1}{1-\gamma}\frac{1}{\lambda_{\min}(\bm{\Sigma})}=\frac{1}{\lambda_{0}(1-\gamma)}.$
+Lemma A.4 implies the following property regarding a Lyapunov equation.
+Lemma A.5.
+For any matrix $\bm{E}\succeq\bm{0}$, there exists a unique positive definite matrix $\bm{X}\succeq\bm{0}$, such that
+$\bm{AX}+\bm{XA}^{\top}=\bm{E}.$
+Furthermore, it can be guaranteed that
+$\|\bm{X}\|\leq\frac{1}{2(1-\gamma)\lambda_{0}}\|\bm{E}\|,\quad\text{and}\quad\mathsf{Tr}(\bm{X})\leq\frac{1}{2(1-\gamma)\lambda_{0}}\mathsf{Tr}(\bm{E}).$
+Proof: The existence of $\bm{X}$ follows from the basic property of Lyapunov equation and (27); we focus on bounding the norm and trace of $\bm{X}$.
+Bounding the operator norm of $\bm{X}$.
+Let $\lambda=\lambda_{\max}(\bm{X})$ and $\bm{x}$ denote a corresponding eigen vector with $\|\bm{x}\|_{2}=1$. By definition, $\bm{Xx}=\lambda\bm{x}$, and therefore
+$\|\bm{E}\|\geq\bm{x}^{\top}\bm{Ex}=\bm{x}^{\top}(\bm{AX}+\bm{XA}^{\top})\bm{x}=\bm{x}^{\top}(\bm{A}+\bm{A}^{\top})\bm{x}\geq[2(1-\gamma)\lambda_{0}]\cdot\lambda,$
+where the last inequality follows from (26). The bound on $\|\bm{X}\|$ follows naturally.
+Bounding the norm of $\bm{X}$
+. Since $\bm{X}\succeq\bm{0}$, direct calculation yields
+$\mathsf{Tr}(\bm{E})$ $=\mathsf{Tr}(\bm{AX}+\bm{XA}^{\top})$
+$=\mathsf{Tr}(\bm{X}^{\frac{1}{2}}\bm{AX}^{\frac{1}{2}}+\bm{X}^{\frac{1}{2}}\bm{A}^{\top}\bm{X}^{\frac{1}{2}})$
+$=\mathsf{Tr}(\bm{X}^{\frac{1}{2}}(\bm{A}+\bm{A}^{\top})\bm{X}^{\frac{1}{2}})$
+$\geq\mathsf{Tr}(\bm{X}^{\frac{1}{2}}\cdot 2(1-\gamma)\lambda_{0}\cdot\bm{X}^{\frac{1}{2}})=2(1-\gamma)\lambda_{0}\mathsf{Tr}(\bm{X}),$
+where we invoked the cyclability of trace and (26) through the process. ∎
 The following lemma further features the compressive power of the matrix $\prod_{k=1}^{t}(\bm{I}-\eta_{k}\bm{A})$.
@@
 Proof of Equation (36).
-For every positive interger $t$, we begin by identifying the index $i$ that maximizes
+For every positive integer $t$, we begin by identifying the index $i$ that maximizes
 $i^{-\alpha}\prod_{k=i+1}^{t}\left(1-\beta k^{-\alpha}\right):=X_{i}^{t}.$
@@
 Here, the first term is bounded by applying the telescoping technique
-$\sum_{j=t}^{\infty}(j+1)^{-\alpha}\prod_{k=t+1}^{j}(1-\beta k^{-\alpha})$ $=\frac{1}{\beta}\sum_{j=t}^{\infty}\left(\prod_{k=t+1}^{j}-\prod_{k=t+1}^{j+1}\right)(1-\beta k^{-\alpha})$
-$=\frac{1}{\beta}\left(1-\prod_{k=t+1}^{\infty}(1-\beta k^{-\alpha})\right)=\frac{1}{\beta}.$
+$\sum_{j=t}^{\infty}(j+1)^{-\alpha}\prod_{k=t+1}^{j}(1-\beta k^{-\alpha})$
+$=\sum_{j=t}^{\infty}\frac{1}{\beta}\left[1-(1-\beta(j+1)^{-\alpha})\right]\prod_{k=t+1}^{j}(1-\beta k^{-\alpha})$
+$=\frac{1}{\beta}\sum_{j=t}^{\infty}\left(\prod_{k=t+1}^{j}(1-\beta k^{-\alpha})-\prod_{k=t+1}^{j+1}(1-\beta k^{-\alpha})\right)$
+$=\frac{1}{\beta}\left(\sum_{j=t}^{\infty}\prod_{k=t+1}^{j}(1-\beta k^{-\alpha})-\sum_{j=t}^{\infty}\prod_{k=t+1}^{j+1}(1-\beta k^{-\alpha})\right)$
+$=\frac{1}{\beta}\left[\left(\prod_{k=t+1}^{t}(1-\beta k^{-\alpha})+\sum_{j=t+1}^{\infty}\prod_{k=t+1}^{j}(1-\beta k^{-\alpha})\right)-\sum_{j=t}^{\infty}\prod_{k=t+1}^{j+1}(1-\beta k^{-\alpha})\right]$
+=$\overset{(i)}{=}\frac{1}{\beta}\left(1+\sum_{i=t}^{\infty}\prod_{k=t+1}^{i+1}(1-\beta k^{-\alpha})-\sum_{j=t}^{\infty}\prod_{k=t+1}^{j+1}(1-\beta k^{-\alpha})\right)=\frac{1}{\beta}.$ (45)
+Notice that in (i) we invoked the convention
+$\prod_{k=t+1}^{t}(1-\beta k^{-\alpha})=1,$
+and a change of variable $i=j-1$.
 Therefore, the proof boils down to showing that the second term is bounded by
@@
 The bound (41) follows naturally. ∎
+Next, we introduce additional quantities
+$\overline{\bm{\Lambda}}_{T}=\frac{1}{T}\sum_{t=1}^{T}\bm{Q}_{t}\bm{\Gamma}\bm{Q}_{t}^{\top},$ (49)
+where $\bm{\Gamma}$ is defined in (12) and, for each $t$, we set
+$\bm{Q}_{t}=\eta_{t}\sum_{j=t}^{T}\prod_{k=t+1}^{j}(\bm{I}-\eta_{k}\bm{A}).$ (50)
+The matrix $\overline{\bm{\Lambda}}_{T}$ plays the role of a non-asymptotic covariance matrix for $\overline{\bm{\theta}}_{T}$.
+The following lemmas feature the properties of $\bm{Q}_{t}$ and $\overline{\bm{\Lambda}}_{T}$.
 Lemma A.8.
-Let $\eta_{t}=\eta_{0}t^{-\alpha}$ with $\alpha\in(\frac{1}{2},1)$ and $0\leq\eta_{0}\leq\frac{1+\gamma}{2\lambda_{\Sigma}}$.
+Let $\eta_{t}=\eta_{0}t^{-\alpha}$ with $\alpha\in(\frac{1}{2},1)$ and $0\leq\eta_{0}\leq\frac{1}{2\lambda_{\Sigma}}$.
 Recall that $\bm{Q}_{t}$ is defined as in (50).
@@
 Proof of (51a)
-By definition, the sum of $Q_{t}$ for $1\leq t\leq T$ can be expressed as
-$\sum_{t=1}^{T}\bm{Q}_{t}=\sum_{t=1}^{T}\eta_{t}\sum_{j=t}^{T}\prod_{k=t+1}^{j}(\bm{I}-\eta_{k}\bm{A}).$
-By switching the order of summation we obtain
+By definition, the sum of $\bm{Q}_{t}$ for $1\leq t\leq T$ can be expressed as
+$\sum_{t=1}^{T}\bm{Q}_{t}$ $=\sum_{t=1}^{T}\eta_{t}\sum_{j=t}^{T}\prod_{k=t+1}^{j}(\bm{I}-\eta_{k}\bm{A})$
+$=\sum_{j=1}^{T}\sum_{t=1}^{j}\eta_{t}\prod_{k=t+1}^{j}(\bm{I}-\eta_{k}\bm{A}),$
+where the second line follows by switching the order of summation. Here, the summand can be represented as
+$\eta_{t}\prod_{k=t+1}^{j}(\bm{I}-\eta_{k}\bm{A})$ $=\bm{A}^{-1}[\bm{I}-(\bm{I}-\eta_{t}\bm{A})]\prod_{k=t+1}^{j}(\bm{I}-\eta_{k}\bm{A})$
+$=\bm{A}^{-1}\left[\prod_{k=t+1}^{j}(\bm{I}-\eta_{k}\bm{A})-\prod_{k=t}^{j}(\bm{I}-\eta_{k}\bm{A})\right].$ (52)
+Notice that summing the difference in the bracket from $t=1$ to $t=j$ would cancel all the terms in the middle and only the first and last terms would remain. Therefore,
 $\sum_{t=1}^{T}\bm{Q}_{t}$ $=\sum_{j=1}^{T}\sum_{t=1}^{j}\eta_{t}\prod_{k=t+1}^{j}(\bm{I}-\eta_{k}\bm{A})$
-$=\sum_{j=1}^{T}\sum_{t=1}^{j}\bm{A}^{-1}\left(\prod_{k=t+1}^{j}-\prod_{k=t}^{j}\right)(\bm{I}-\eta_{k}\bm{A})$
+$=\sum_{j=1}^{T}\sum_{t=1}^{j}\bm{A}^{-1}\left[\prod_{k=t+1}^{j}(\bm{I}-\eta_{k}\bm{A})-\prod_{k=t}^{j}(\bm{I}-\eta_{k}\bm{A})\right]$
 $=\bm{A}^{-1}\sum_{j=1}^{T}\left[\bm{I}-\prod_{k=1}^{j}(\bm{I}-\eta_{k}\bm{A})\right],$
@@
 $=\eta_{t}\bm{A}\prod_{k=t+1}^{j-1}(\bm{I}-\eta_{k}\bm{A})-\eta_{j}\prod_{k=t+1}^{j-1}(\bm{I}-\eta_{k}\bm{A})\bm{A}$
-$=(\eta_{t}-\eta_{j})\bm{A}\prod_{k=t+1}^{j-1}(I-\eta_{k}\bm{A}).$
+$=(\eta_{t}-\eta_{j})\bm{A}\prod_{k=t+1}^{j-1}(\bm{I}-\eta_{k}\bm{A}).$
 Therefore, by taking a difference between $\bm{Q}_{t}$ and $(\bm{I}-\eta_{t}\bm{A})\bm{Q}_{t}$, we obtain
@@
 Lemma A.9.
-Under the condition 12, for every positive definite matrix $\bm{Y}\in\mathbb{R}^{d\times d}$, it can be guarnateed that
-$T^{-\alpha}\sum_{t=1}^{T}\left(\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A})\bm{Y}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})\right)=\bm{X}+\bm{O}(T^{\alpha-1}),$ (51)
-where $\bm{X}$ is the solution to the Lyapunov equation
-$\eta_{0}(\bm{A}\bm{X}+\bm{XA}^{\top})=\bm{Y}.$ (52)
-Proof: For simplicity, denote the left-hand-side of (51) as $\bm{X}_{T}$. It can be easily verified that the sequence $\{\bm{X}_{T}\}_{T\geq 1}$ is featured by the iterative relation
-$\bm{X}_{T}$ $=\left(\frac{T-1}{T}\right)^{\alpha}(\bm{I}-\eta_{T}\bm{A})\bm{X}_{T-1}(\bm{I}-\eta_{T}\bm{A}^{\top})$
-$+T^{-\alpha}(\bm{I}-\eta_{T}\bm{A})\bm{Y}(\bm{I}-\eta_{T}\bm{A}^{\top}).$
-Direct calculation yields
-$\bm{X}_{T}-\bm{X}$ $=\left(\frac{T-1}{T}\right)^{\alpha}(\bm{I}-\eta_{T}\bm{A})(\bm{X}_{T-1}-\bm{X})(\bm{I}-\eta_{T}\bm{A}^{\top})$
-$+\left(\frac{T-1}{T}\right)^{\alpha}(\bm{I}-\eta_{T}\bm{A})\bm{X}(\bm{I}-\eta_{T}\bm{A}^{\top})-\bm{X}$
-$+T^{-\alpha}(\bm{I}-\eta_{T}\bm{A})\bm{Y}(\bm{I}-\eta_{T}\bm{A}^{\top})$
-$=\left(\frac{T-1}{T}\right)^{\alpha}(\bm{I}-\eta_{T}\bm{A})(\bm{X}_{T-1}-\bm{X})(\bm{I}-\eta_{T}\bm{A}^{\top})$
-$+T^{-\alpha}[\eta_{0}(\bm{AX}+\bm{XA}^{\top})-\bm{Y}]+\frac{\bm{X}}{T}+O(T^{-2\alpha}).$
-Notice that the first term on the last line is $\bm{0}$ according to the definition of $\bm{X}$; therefore by iteration, the difference between $\bm{X}_{T}$ and $\bm{X}$ can be represented by
-$\bm{X}_{T}-\bm{X}$ $=T^{-\alpha}\sum_{i=1}^{T}i^{\alpha-1}\prod_{k=i+1}^{T}(\bm{I}-\eta_{k}\bm{A})\bm{X}\prod_{k=i+1}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})$
-$-T^{-\alpha}\prod_{k=1}^{T}(\bm{I}-\eta_{k}\bm{A})\bm{X}\prod_{k=1}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})+O(T^{-\alpha});$
-By letting $\beta=\frac{1-\gamma}{2}\lambda_{0}\eta_{0}$, LemmaA.3 inplies the following bound on the norm of the first term on the right-hand-side:
-$\left\|T^{-\alpha}\sum_{i=1}^{T}i^{\alpha-1}\prod_{k=i+1}^{T}(\bm{I}-\eta_{k}\bm{A})\bm{X}\prod_{k=i+1}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})\right\|$
-$<T^{\alpha-1}\sum_{i=1}^{T}i^{-\alpha}\prod_{k=1}^{T}(1-\beta k^{-\alpha})\|\bm{X}\|$
-$=T^{\alpha-1}\cdot\frac{1}{\beta}\sum_{i=1}^{T}\left(\prod_{k=i}^{T}-\prod_{k=i+1}^{T}(1-\beta k^{-\alpha})\right)\|\bm{X}\|<\frac{\|\bm{X}\|}{\beta}T^{\alpha-1}.$
-The two terms on the second line are clearly both $\bm{o}(T^{\alpha-1})$; hence,
-$\|\bm{X}_{T}-\bm{X}\|=\bm{O}(T^{\alpha-1}).$
+When $\eta_{0}<\frac{1}{2\lambda_{\Sigma}}$, for every positive definite matrix $\bm{Y}\in\mathbb{R}^{d\times d}$, it can be guaranteed that
+$T^{-\alpha}\sum_{t=1}^{T}\left(\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A})\bm{Y}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})\right)-\bm{X}\leq O(T^{\alpha-1})\bm{Y}$ (54)
+where $\bm{X}\succeq\bm{0}$ is the solution to the Lyapunov equation
+$\eta_{0}(\bm{A}\bm{X}+\bm{XA}^{\top})=\bm{Y},$ (55)
+and the big $O$ notation for matrix sequences is defined as in (22).
+Proof: For simplicity, we define a sequence of matrices $\{\bm{X}_{i}\}_{i=0}^{T}$ as
+$\bm{X}_{i}=\begin{cases}&i^{-\alpha}\sum_{t=1}^{i}\left(\prod_{k=t}^{i}(\bm{I}-\eta_{k}\bm{A})\bm{Y}\prod_{k=t}^{i}(\bm{I}-\eta_{k}\bm{A}^{\top})\right),\quad\text{for}\quad 1\leq i\leq T;\\ &\bm{0},\quad\text{for}\quad i=0.\end{cases}$
+in this way, the left-hand-side of (54) can be represented by $\bm{X}_{T}-\bm{X}$. It can be easily verified that the sequence $\{\bm{X}_{i}\}_{i=1}^{T}$ is featured by the iterative relation
+$i^{\alpha}\bm{X}_{i}$ $=(i-1)^{\alpha}(\bm{I}-\eta_{i}\bm{A})\bm{X}_{i-1}(\bm{I}-\eta_{i}\bm{A}^{\top})+(\bm{I}-\eta_{i}\bm{A})\bm{Y}(\bm{I}-\eta_{i}\bm{A}^{\top}),\quad\forall 1\leq i\leq T.$
+Direct calculation yields that for every $1\leq i\leq T$,
+$i^{\alpha}(\bm{X}_{i}-\bm{X})$ $=(i-1)^{\alpha}(\bm{I}-\eta_{i}\bm{A})\bm{X}_{i-1}(\bm{I}-\eta_{i}\bm{A}^{\top})+(\bm{I}-\eta_{i}\bm{A})\bm{Y}(\bm{I}-\eta_{i}\bm{A}^{\top})-i^{\alpha}\bm{X}$
+$=(i-1)^{\alpha}(\bm{I}-\eta_{i}\bm{A})(\bm{X}_{i-1}-\bm{X})(\bm{I}-\eta_{i}\bm{A}^{\top})$
+$+(\bm{I}-\eta_{i}\bm{A})\bm{Y}(\bm{I}-\eta_{i}\bm{A}^{\top})-[i^{\alpha}-(i-1)^{\alpha}]\bm{X}-(i-1)^{\alpha}[\eta_{i}(\bm{AX}+\bm{XA}^{\top})+\eta_{i}^{2}\bm{AXA}^{\top}].$
+We now feature the terms on the last line. Firstly, we notice that (55) guarantees
+$\eta_{i}(\bm{AX}+\bm{XA}^{\top})=\eta_{0}i^{-\alpha}(\bm{AX}+\bm{XA}^{\top})=i^{-\alpha}\bm{Y};$
+therefore, a simple reorganization yields
+$(\bm{I}-\eta_{i}\bm{A})\bm{Y}(\bm{I}-\eta_{i}\bm{A}^{\top})-[i^{\alpha}-(i-1)^{\alpha}]\bm{X}-(i-1)^{\alpha}[\eta_{i}(\bm{AX}+\bm{XA}^{\top})+\eta_{i}^{2}\bm{AXA}^{\top}]$
+$=\left[1-\left(\frac{i-1}{i}\right)^{\alpha}\right]\bm{Y}-[i^{\alpha}-(i-1)^{\alpha}]\bm{X}-\eta_{i}(\bm{AY}+\bm{YA}^{\top})+\eta_{i}^{2}\bm{A}(\bm{Y}-(i-1)^{\alpha}\bm{X})\bm{A}^{\top}.$
+For these terms, we observe
+$\left[1-\left(\frac{i-1}{i}\right)^{\alpha}\right]\bm{Y}\preceq\left[1-\left(\frac{i-1}{i}\right)\right]\bm{Y}=\frac{1}{i}\bm{Y};$
+$[i^{\alpha}-(i-1)^{\alpha}]\bm{X}=i^{\alpha}\left[1-\left(\frac{i-1}{i}\right)^{\alpha}\right]\bm{X}\preceq i^{\alpha}\left[1-\left(\frac{i-1}{i}\right)\right]\bm{X}=i^{\alpha-1}\bm{X}=O(i^{\alpha-1})\bm{Y};$
+$\eta_{i}(\bm{AY}+\bm{YA}^{\top})=O(i^{-\alpha})\bm{Y};$
+$\eta_{i}^{2}\bm{A}(\bm{Y}-(i-1)^{\alpha}\bm{X})\bm{A}^{\top}=O(i^{-\alpha})\bm{Y}.$
+Hence, the leading term decreases by the rate of $O(i^{\alpha-1})\bm{Y}$. As a result,
+$i^{\alpha}(\bm{X}_{i}-\bm{X})=(\bm{I}-\eta_{i}\bm{A})[(i-1)^{\alpha}(\bm{X}_{i-1}-\bm{X})](\bm{I}-\eta_{i}\bm{A}^{\top})+O(i^{\alpha-1})\bm{Y}$
+For simplicity, we denote $\bm{Z}_{i}=i^{\alpha}(\bm{X}_{i}-\bm{X})$. The relation above can then be represented as
+$\bm{Z}_{i}=(\bm{I}-\eta_{i}\bm{A})\bm{Z}_{i-1}(\bm{I}-\eta_{i}\bm{A}^{\top})+O(i^{\alpha-1})\bm{Y}.$
+It is easy to verify by iteration that, there exists a quantity $\widetilde{C}$ independent of $T$, such that
+$\sum_{i=1}^{T}i^{\alpha-1}\prod_{k=i+1}^{T}(\bm{I}-\eta_{k}\bm{A})\bm{Y}\prod_{k=i+1}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})$
+$\preceq\bm{Z}_{T}\preceq\sum_{i=1}^{T}i^{\alpha-1}\prod_{k=i+1}^{T}(\bm{I}-\eta_{k}\bm{A})\bm{Y}\prod_{k=i+1}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top});$
+In other words,
+$T^{-\alpha}\sum_{i=1}^{T}i^{\alpha-1}\prod_{k=i+1}^{T}(\bm{I}-\eta_{k}\bm{A})\bm{Y}\prod_{k=i+1}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})$
+$\preceq\bm{X}_{T}-\bm{X}$
+$\preceq T^{-\alpha}\sum_{i=1}^{T}i^{\alpha-1}\prod_{k=i+1}^{T}(\bm{I}-\eta_{k}\bm{A})\bm{Y}\prod_{k=i+1}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top}).$
+By letting $\beta=\frac{1-\gamma}{2}\lambda_{0}\eta_{0}$, Lemma A.4 implies
+$\left\|T^{-\alpha}\sum_{i=1}^{T}i^{\alpha-1}\prod_{k=i+1}^{T}(\bm{I}-\eta_{k}\bm{A})\prod_{k=i+1}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})\right\|$
+$\leq T^{-\alpha}\sum_{i=1}^{T}i^{\alpha-1}\prod_{k=i+1}^{T}\|\bm{I}-\eta_{k}\bm{A}\|\prod_{k=i+1}^{T}\|\bm{I}-\eta_{k}\bm{A}^{\top}\|$
+$\leq T^{-\alpha}\sum_{i=1}^{T}i^{\alpha-1}\prod_{k=i+1}^{T}(1-\beta k^{-\alpha})^{2}$
+$<T^{\alpha-1}\sum_{i=1}^{T}i^{-\alpha}\prod_{k=i+1}^{T}(1-\beta k^{-\alpha})$
+$=T^{\alpha-1}\cdot\frac{1}{\beta}\sum_{i=1}^{T}\left(\prod_{k=i+1}^{T}(1-\beta k^{-\alpha})-\prod_{k=i}^{T}(1-\beta k^{-\alpha})\right)$
+$=\frac{1}{\beta}T^{\alpha-1}\left(\prod_{k=T+1}^{T}(1-\beta k^{-\alpha})-\prod_{k=1}^{T}(1-\beta k^{-\alpha})\right)<\frac{1}{\beta}T^{\alpha-1};$
+notice that in the last two lines we invoked a similar approach to the telescoping method (A.1). Combining this with (25), we obtain
+$T^{-\alpha}\sum_{i=1}^{T}i^{\alpha-1}\prod_{k=i+1}^{T}(\bm{I}-\eta_{k}\bm{A})\bm{Y}\prod_{k=i+1}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})=O(T^{\alpha-1})\bm{Y}.$
+This completes the proof. ∎
 Lemma A.10.
-Let $\eta_{t}=\eta_{0}t^{-\alpha}$ with $\alpha\in(\frac{1}{2},1)$ and $0\leq\eta\leq\frac{1+\gamma}{2\lambda_{\Sigma}}$. For $\bm{Q}_{t}$ defined as in (19), it can be guaranteed that
+Let $\eta_{t}=\eta_{0}t^{-\alpha}$ with $\alpha\in(\frac{1}{2},1)$ and $0\leq\eta\leq\frac{1}{2\lambda_{\Sigma}}$. For $\bm{Q}_{t}$ defined as in (50), it can be guaranteed that
 $\|\bm{Q}_{t}\|\leq 3\eta_{0}^{-\frac{\alpha}{1-\alpha}}\left(\frac{4\alpha}{\lambda_{0}(1-\gamma)}\right)^{\frac{1}{1-\alpha}},\quad\forall t\in[T];$ (56)
@@
 where $\beta=\frac{1-\gamma}{2}\lambda_{0}\eta_{0}$ and $\widetilde{C}$ depends on $\alpha,\eta_{0},\lambda_{0},\gamma$.
-Proof: According to Lemma A.3, since $0<\eta_{t}\leq\eta_{0}\leq\frac{1+\gamma}{2\lambda_{\Sigma}}$ for all $t$, the norm of $\bm{Q}_{t}$ is bounded by
+Proof: According to Lemma A.4, since $0<\eta_{t}\leq\eta_{0}\leq\frac{1}{2\lambda_{\Sigma}}$ for all $t$, the norm of $\bm{Q}_{t}$ is bounded by
 $\|\bm{Q}_{t}\|_{2}$ $\leq\eta_{t}\sum_{j=t}^{T}\prod_{k=t+1}^{j}\left(1-\frac{1}{2}\lambda_{0}(1-\gamma)\eta_{k}\right)$
@@
 Furthermore, it can be easily verified that
-$\|\bm{AQ}_{t}-\bm{I}\|^{2}\leq\|\bm{AS}_{t}\|^{2}+\left\|\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A})\right\|^{2}$ $\leq\widetilde{C}^{\prime}t^{2\alpha-2}+\prod_{k=t}^{T}(1-\beta k^{-\alpha})$
+$\|\bm{AQ}_{t}-\bm{I}\|^{2}$ $\leq\|\bm{AS}_{t}\|^{2}+\left\|\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A})\right\|^{2}$
+$\leq\widetilde{C}^{\prime}t^{2\alpha-2}+\prod_{k=t}^{T}(1-\beta k^{-\alpha})$
 $<\widetilde{C}^{\prime}t^{2\alpha-2}+T^{\alpha}(t-1)^{-\alpha}\prod_{k=t}^{T}(1-\beta k^{-\alpha})$
-$=\widetilde{C}^{\prime}t^{2\alpha-2}+T^{\alpha}\frac{1}{\beta}\left(\prod_{k=t}^{T}-\prod_{k=t-1}^{T}\right)(1-\beta k^{-\alpha});$
-The bound (55) follows by taking an average from $t=1$ to $t=T$ and applying the telescoping method. ∎
+$=\widetilde{C}^{\prime}t^{2\alpha-2}+T^{\alpha}\cdot\frac{1}{\beta}\left(\prod_{k=t}^{T}(1-\beta k^{-\alpha})-\prod_{k=t-1}^{T}(1-\beta k^{-\alpha})\right);$
+The bound (58) follows by taking an average from $t=1$ to $t=T$ and applying the telescoping method as in (A.1). ∎
 The decomposition (51b), combined with the bounds (51c) Theorem A.9, implies the following bound on the difference between $\overline{\bm{\Lambda}}_{T}$ and $\bm{\Lambda}^{\star}$.
@@
 $\eta_{0}(\bm{AX+XA}^{\top})=\bm{\Lambda}^{\star};$ (59)
-then as $T\to\infty$, the difference between $\overline{\bm{\Lambda}_{T}}$ and $\bm{\Lambda}^{\star}$ is featured by
-$T^{\alpha-1}\bm{X}(\bm{\Lambda}^{\star})-\bm{O}(T^{2\alpha-2})\preceq\overline{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star}\preceq T^{\alpha-1}\bm{X}(\bm{\Lambda}^{\star})+\bm{O}(T^{2\alpha-2})$
-Proof: By definition, the difference bewteen $\overline{\bm{\Lambda}}_{T}$ and $\bm{\Lambda}^{\star}$ can be expressed as
+then as $T\to\infty$, the difference between $\overline{\bm{\Lambda}_{T}}$ and $\bm{\Lambda}^{\star}$ satisfies
+$\overline{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star}-T^{\alpha-1}\bm{X}(\bm{\Lambda}^{\star})=O(T^{2\alpha-2})\bm{\Gamma}.$
+Here, the $O$ notation for matrices is defined as in (22).
+Proof: By definition, the difference between $\overline{\bm{\Lambda}}_{T}$ and $\bm{\Lambda}^{\star}$ can be expressed as
 $\overline{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star}$ $=\frac{1}{T}\sum_{t=1}^{T}(\bm{Q}_{t}\bm{\Gamma}\bm{Q}_{t}^{\top}-\bm{A}^{-1}\bm{\Gamma}\bm{A}^{-\top})$
@@
 $=-\frac{1}{T}\sum_{j=1}^{T}\prod_{k=1}^{j}(\bm{I}-\eta_{k}\bm{A})\cdot\bm{\Lambda}^{\star}.$ (61)
-where we applied Equation (49a) and the interchangability of the production. Similarly, the second term can be expressed by
+where we applied Equation (51a) the interchangeability of products. Similarly, the second term can be expressed by
 $\bm{A}^{-1}\bm{\Gamma}\frac{1}{T}\sum_{t=1}^{T}(\bm{Q}_{t}-\bm{A}^{-1})^{\top}=-\bm{\Lambda}^{\star}\cdot\frac{1}{T}\sum_{j=1}^{T}\prod_{k=1}^{j}(\bm{I}-\eta_{k}\bm{A})^{\top}.$
-Lemma A.3 directly implies that the norm of these two terms are bounded by
-$\left\|\frac{1}{T}\sum_{t=1}^{T}(\bm{Q}_{t}-\bm{A}^{-1})\bm{\Gamma}\bm{A}^{-\top}\right\|$ $\leq\frac{1}{T}\|\bm{\Lambda}^{\star}\|\sum_{j=1}^{T}\prod_{k=1}^{j}\left(1-\frac{1-\gamma}{2}\lambda_{0}\eta_{0}k^{-\alpha}\right).$
+Lemma A.4 directly implies that
+$\frac{1}{T}\sum_{j=1}^{T}\prod_{k=1}^{j}(\bm{I}-\eta_{k}\bm{A})\leq\frac{1}{T}\prod_{k=1}^{j}\left(1-\frac{1-\gamma}{2}\lambda_{0}\eta_{0}k^{-\alpha}\right).$
 For simplicity, we denote $\beta=\frac{1-\gamma}{2}\lambda_{0}\eta_{0}$; it can then be seen that the summand on the right-hand-side can be bounded by
@@
 $=\beta^{-\frac{1}{1-\alpha}}(1-\alpha)^{\frac{\alpha}{1-\alpha}}\Gamma\left(\frac{1}{1-\alpha}\right),$ (63)
-a constant independent of $T$. Therefore, $I_{1}$ is bounded above by $\bm{O}(T^{-1})$.
+a constant independent of $T$. Therefore, $I_{1}$ is featured by
+$\frac{1}{T}\sum_{t=1}^{T}(\bm{Q}_{t}-\bm{A}^{-1})\bm{\Gamma}\bm{A}^{-\top}+\frac{1}{T}\sum_{t=1}^{T}\bm{A}^{-1}\bm{\Gamma}(\bm{Q}_{t}-\bm{A}^{-1})^{\top}=O(T^{-1})\bm{\Lambda}^{\star}.$
 Decomposing $I_{2}$.
 The term $I_{2}$ can be further expressed by invoking the decomposition (51b). Specifically, we observe
-$I_{2}$ $=\underset{I_{21}}{\underbrace{\frac{1}{T}\sum_{t=1}^{T}\bm{S}_{t}\bm{\Gamma}\bm{S}_{t}^{\top}}}+\underset{I_{22}}{\underbrace{\frac{1}{T}\sum_{t=1}^{T}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A})\bm{\Gamma}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})}}$
-($-\underset{I_{23}}{\underbrace{\left(\frac{1}{T}\sum_{t=1}^{T}\bm{S}_{t}\bm{\Gamma}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})+\frac{1}{T}\sum_{t=1}^{T}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A})\bm{\Gamma}\bm{S}_{t}^{\top}\right)}}.$
-Since $\|\bm{S}_{t}\|=O(T^{\alpha-1})$, it is easy to verify that $I_{21}=\bm{O}(T^{2\alpha-2})$; meanwhile, theorem A.7 indicates that $I_{22}=T^{\alpha-1}\bm{X}(\bm{\Lambda}^{\star})+O(T^{2\alpha-2})$.
+$I_{2}$ $=\underset{I_{21}}{\underbrace{\frac{1}{T}\sum_{t=1}^{T}\bm{S}_{t}\bm{\Gamma}\bm{S}_{t}^{\top}}}+\underset{I_{22}}{\underbrace{\frac{1}{T}\sum_{t=1}^{T}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A})\bm{\Lambda}^{\star}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})}}$
+($-\underset{I_{23}}{\underbrace{\left(\frac{1}{T}\sum_{t=1}^{T}\bm{S}_{t}\bm{\Gamma}\bm{A}^{-\top}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})+\frac{1}{T}\sum_{t=1}^{T}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A})\bm{A}^{-1}\bm{\Gamma}\bm{S}_{t}^{\top}\right)}}.$
+Since $\|\bm{S}_{t}\|=O(T^{\alpha-1})$, it is easy to verify that $I_{21}=O(T^{2\alpha-2})\bm{\Gamma}$; meanwhile, theorem A.9 indicates that $I_{22}=T^{\alpha-1}\bm{X}(\bm{\Lambda}^{\star})+O(T^{2\alpha-2})\bm{\Lambda}^{\star}$.
 As for the term $I_{23}$, we observe that
-$\left\|\sum_{t=1}^{T}\bm{S}_{t}\bm{\Gamma}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})\right\|$ $\lesssim\sum_{t=1}^{T}t^{\alpha-1}\prod_{k=t}^{T}(1-\beta k^{-\alpha})$
-$<T^{2\alpha-1}\sum_{t=1}^{T}t^{\alpha-1}\prod_{k=t+1}^{T}(1-\beta k^{-\alpha})$
-$=T^{2\alpha-1}\cdot\frac{1}{\beta}\sum_{t=1}^{T}\left(\prod_{k=t+1}^{T}-\prod_{k=t}^{T}\right)(1-\beta k^{-\alpha})$
-$<\frac{T^{2\alpha-1}}{\beta}.$
-Consequently, it can be guaranteed that $I_{23}=\bm{O}(T^{2\alpha-2})$. In combintation, we have shown that
-$T^{\alpha-1}\bm{X}(\bm{\Lambda}^{\star})-\bm{O}(T^{2\alpha-2})\preceq\frac{1}{T}\sum_{t=1}^{T}(\bm{Q}_{t}-\bm{A}^{-1})\bm{\Gamma}(\bm{Q}_{t}-\bm{A}^{-1})^{\top}\preceq T^{\alpha-1}\bm{X}(\bm{\Lambda}^{\star})+\bm{O}(T^{2\alpha-2});$
-the theorem follows naturally by combining this with the bound on $I_{1}$.∎
+$\left\|\sum_{t=1}^{T}\bm{S}_{t}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})\right\|$ $\lesssim\widetilde{C}\sum_{t=1}^{T}t^{\alpha-1}\prod_{k=t}^{T}(1-\beta k^{-\alpha})$
+$<\widetilde{C}T^{2\alpha-1}\sum_{t=1}^{T}t^{\alpha-1}\prod_{k=t+1}^{T}(1-\beta k^{-\alpha})$
+$=\widetilde{C}T^{2\alpha-1}\cdot\frac{1}{\beta}\sum_{t=1}^{T}\left(\prod_{k=t+1}^{T}-\prod_{k=t}^{T}\right)(1-\beta k^{-\alpha})$
+$<\widetilde{C}\frac{T^{2\alpha-1}}{\beta},$
+where $\widetilde{C}$ is a problem-related quantity related to $\eta_{0},\alpha$ and $\gamma$.
+Consequently, it can be guaranteed that $I_{23}=O(T^{2\alpha-2})\bm{\Gamma}$.
+In combintation, we have shown that
+$\overline{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star}$ $=I_{1}+I_{21}+I_{22}+I_{23}$
+$=O(T^{-1})\bm{\Lambda}^{\star}+O(T^{2\alpha-2})\bm{\Gamma}+(T^{\alpha-1}\bm{X}(\bm{\Lambda}^{\star})+O(T^{2\alpha-2})\bm{\Lambda}^{\star})+O(T^{2\alpha-2})\bm{\Gamma}$
+$=T^{\alpha-1}\bm{X}(\bm{\Lambda}^{\star})+O(T^{2\alpha-2})\bm{\Gamma}+O(T^{2\alpha-2})\bm{\Lambda}^{\star};$
+the theorem follows naturally by combining this with the fact that $\bm{\Lambda}^{\star}=O(1)\bm{\Gamma}$, due to (31).∎
 The following Lemma controls the minimum eigenvalue of $\bm{A}\overline{\bm{\Lambda}}_{T}\bm{A}^{\top}$.
 Lemma A.12.
-With $\bm{A}$ defined as (6a) and $\overline{\bm{\Lambda}}_{T}$ defined as in (18), when
+With $\bm{A}$ defined as (4) and $\overline{\bm{\Lambda}}_{T}$ defined as in (49), when
 $T\geq 4\left(\frac{2}{(1-\gamma)\lambda_{0}\eta_{0}}\right)^{\frac{1}{1-\alpha}}(1-\alpha)^{\frac{\alpha}{1-\alpha}}\Gamma(\frac{1}{1-\alpha})\mathsf{cond}(\Gamma),$ (64)
@@
 A.2.1 Berry-Esseen bounds
-The following theorem, proved by Shao and Zhang, (2021), gives a Berry-Esseen bound for non-linear statistics which would be crucial for our analysis of the non-asymptotic distribution of the TD estimation error.
+The following theorem, due to Shao and Zhang (2022), gives a Berry-Esseen bound for non-linear statistics which would be crucial for our analysis of the non-asymptotic distribution of the TD estimation error.
 Theorem A.13.
-(Theorem 2.1 of Shao and Zhang, (2021)) Let $\bm{X}_{1},...,\bm{X}_{n}$ be independent random variables, and let $\bm{T}(\bm{X}_{1},...,\bm{X}_{n})$ be a $d$-dimensional statistic that can be represented as
+(Theorem 2.1 of Shao and Zhang (2022)) Let $\bm{X}_{1},...,\bm{X}_{n}$ be independent random variables, and let $\bm{T}(\bm{X}_{1},...,\bm{X}_{n})$ be a $d$-dimensional statistic that can be represented as
 $\bm{T}=\bm{W}+\bm{D},$
@@
 A.2.2 Properties of multivariate Gaussian distributions
-The following theorem proposed by Devroye et al., (2018) gives an upper bound for the total variation (TV) distance between two Gaussian random variables with same means and different covariance matrixes.
-Theorem A.12 (Devroye et al., (2018)).
+The following theorem by Devroye et al. (2018) gives an upper bound for the total variation (TV) distance between two Gaussian random variables with same means and different covariance matrices.
+Theorem A.14 (Devroye et al. (2018)).
 Let $\bm{\Lambda}_{1},\bm{\Lambda}_{2}\in\mathbb{S}^{d\times d}$ be two positive definite matrices, and $\bm{\mu}$ be any vector in $\mathbb{R}^{d}$.
-Then the TV distance between Gaussian distirbutions $\mathcal{N}(\bm{\mu},\bm{\Lambda}_{1})$ and $\mathcal{N}(\bm{\mu},\bm{\Lambda}_{2})$ is bounded by
+Then the TV distance between Gaussian distributions $\mathcal{N}(\bm{\mu},\bm{\Lambda}_{1})$ and $\mathcal{N}(\bm{\mu},\bm{\Lambda}_{2})$ is bounded by
 $\min\left\{1,\frac{1}{100}\left\|\bm{\Lambda}_{1}^{-1/2}\bm{\Lambda}_{2}\bm{\Lambda}_{1}^{-1/2}-\bm{I}_{d}\right\|_{\mathsf{F}}\right\}$
@@
 $\leq\frac{3}{2}\left\|\bm{\Lambda}_{1}^{-1/2}\bm{\Lambda}_{2}\bm{\Lambda}_{1}^{-1/2}-\bm{I}_{d}\right\|_{\mathsf{F}}.$
-In the case of $\bm{\Lambda}_{1}\succ\bm{\Lambda}_{2}$, the following proposition further specifies that the TV distance and the distance on convex sets are the same bewteen $\mathcal{N}(\bm{0},\bm{\Lambda}_{1})$ and $\mathcal{N}(\bm{0},\bm{\Lambda}_{2})$.
+In the case of $\bm{\Lambda}_{1}\succ\bm{\Lambda}_{2}$, the following proposition further specifies that the TV distance and the distance on convex sets are the same between $\mathcal{N}(\bm{0},\bm{\Lambda}_{1})$ and $\mathcal{N}(\bm{0},\bm{\Lambda}_{2})$.
 Proposition A.15.
@@
 Notice that the ratio between these two density functions is featured by
-$\frac{p_{1}(\bm{x})}{p_{2}(\bm{x})}$ $=\frac{\frac{1}{\sqrt{\mathsf{det}(\bm{\Lambda_{1}})}}\exp(-\bm{x}^{\top}\bm{\Lambda_{1}}^{-1}\bm{x})}{\frac{1}{\sqrt{\mathsf{det}(\bm{\Lambda_{1}})}}\exp(-\bm{x}^{\top}\bm{\Lambda_{2}}^{-1}\bm{x})}$
+$\frac{p_{1}(\bm{x})}{p_{2}(\bm{x})}$ $=\frac{\frac{1}{\sqrt{\mathsf{det}(\bm{\Lambda_{1}})}}\exp(-\bm{x}^{\top}\bm{\Lambda_{1}}^{-1}\bm{x})}{\frac{1}{\sqrt{\mathsf{det}(\bm{\Lambda_{2}})}}\exp(-\bm{x}^{\top}\bm{\Lambda_{2}}^{-1}\bm{x})}$
 $=\sqrt{\frac{\mathsf{det}(\bm{\Lambda}_{2})}{\mathsf{det}(\bm{\Lambda_{1}})}}\exp(\bm{x}^{\top}(\bm{\Lambda_{2}}^{-1}-\bm{\Lambda_{1}}^{-1})\bm{x}).$
@@
 $\geq d_{\mathsf{C}}(\mathcal{N}(\bm{0},\bm{\Lambda}_{1}),\mathcal{N}(\bm{0},\bm{\Lambda}_{2})).$
-Combining these two inequalitites completes the proof. ∎
+Combining these two inequalities completes the proof. ∎
 A.3 Concentration inequalities
-The following Azuma’s inequality for matrix martingales is shown in Tropp, (2011).
+The following Azuma’s inequality for matrix martingales is shown in Tropp (2011).
 Theorem A.16.
@@
 $\mathbb{P}\left(\left\|\sum_{k}\bm{X}_{k}\right\|\geq\varepsilon\right)\leq d\exp\left(-\frac{\varepsilon^{2}}{8\sigma^{2}}\right).$
-Specifically, for vector-valued martingales, the following proposition is shown by Kallenberg and Sztencel, (1991).
+Specifically, for vector-valued martingales, the following proposition is shown by Kallenberg and Sztencel (1991).
 Proposition A.17.
@@
 $\|\bm{x}_{t}\|_{2}\leq 2\sqrt{2W_{\max}\log\frac{3}{\delta}}.$
+The following Freedman’s inequality on Hilbert spaces was shown by Peng et al. (2024) in their Theorem 2.1.
+Theorem A.19.
+Let $\mathcal{X}$ be a Hilbert space, $\{\bm{X}_{i}\}_{i=1}^{n}$ be an $\mathcal{X}$-valued martingale difference sequence adapted to the filtration $\{\mathcal{F}_{i}\}_{i=1}^{n}$, $\bm{Y}_{i}:=\sum_{j=1}^{i}\bm{X}_{j}$ be the corresponding martingale, and $W_{i}=\sum_{j=1}^{i}\mathbb{E}\|\bm{X}_{j}\|^{2}\mid\mathscr{F}_{j-1}$.
+Suppose that $\max_{i\in[n]}\|\bm{X}_{i}\|\leq b$ for some constant $b>0$.
+Then for any $\varepsilon$ and $\sigma>0$, it can be guaranteed that
+$\mathbb{P}\left(\exists k\in[n],\text{ s.t. }\|\bm{Y}_{k}\|\geq\varepsilon\text{ and }W_{k}\leq\sigma^{2}\right)\leq 2\exp\left(-\frac{\varepsilon^{2}/2}{\sigma^{2}+b\varepsilon/3}\right).$
 Appendix B Proof of main results
-B.1 High-probability convergence guarantee for the original TD estimation error
-In order to derive the high-probability convergence guarantee for $\overline{\bm{\Delta}}_{T}$, we firstly propose the following result on the time-uniform convergence of the original TD estimator $\bm{\Delta}_{t}$, under condition 12.
+B.1 Proof of Theorem 3.1
+In order to derive the high-probability convergence guarantee for $\overline{\bm{\Delta}}_{T}$, we firstly propose the following result on the time-uniform convergence of the original TD estimator $\bm{\Delta}_{t}$.
+B.1.1 High-probability convergence guarantee for the original TD estimation error
 Theorem B.1.
-Consider the TD algorithm (9) with independent samples and decaying stepsizes as in condition 12.
+Consider the TD algorithm (7) with independent samples and decaying stepsizes $\eta_{t}=\eta_{0}t^{-\alpha}$, where $\alpha\in(\frac{1}{2},1)$.
 For any $\delta\in(0,1)$, there exists $\eta_{0}>0$, such that with probability at least $1-\delta$, the norm of $\bm{\Delta}_{t}$ is bounded by
@@
 $\bm{R}_{i}^{t}=\eta_{i}\prod_{k=i+1}^{t}(\bm{I}-\eta_{k}\bm{A}).$ (69)
-For simplicity, in this proof we deonte
+For simplicity, in this proof we denote
 $\beta=\frac{1-\gamma}{2}\lambda_{0}\eta_{0},\quad R=\frac{5}{2}\|\bm{\theta}^{\star}\|_{2}+\frac{1}{4},\quad\text{and}$
 $C=\frac{2\alpha-1}{512\eta_{0}^{2}}\left(\frac{\beta}{2\alpha}\right)^{\frac{\alpha}{1-\alpha}}.$
-For any given $\delta$, we define $t^{\star}$ as a function of $\delta$, speicifically
+For any given $\delta$, we define $t^{\star}$ as a function of $\delta$, specifically
 $t^{\star}(\delta)=\inf\left\{t\in\mathbb{N}^{+}:\int_{t}^{\infty}\exp(-Cx^{\alpha})\mathrm{d}x\leq\frac{\delta}{9}\right\},$ (70)
@@
 Furthermore, we denote, for simplicity,
-$\mathcal{H}_{i}=\{\max_{1\leq j\leq i}\|\bm{\Delta}_{j}\|\leq R\},\quad\text{and}\quad\widetilde{\bm{\zeta}}_{i}=\bm{\zeta}_{i}\mathds{1}(\mathcal{H}_{i-1}).$
+$\mathcal{H}_{i}=\{\max_{1\leq j\leq i}\|\bm{\Delta}_{j}\|\leq R\},\quad\text{and}\quad\tilde{\bm{\zeta}}_{i}=\bm{\zeta}_{i}\mathds{1}(\mathcal{H}_{i-1}).$
 It can then be easily verified that
-$\|\widetilde{\bm{\zeta}}_{i}\|_{2}\leq 4R+2\|\bm{\theta}^{\star}\|_{2}+1.$
+$\|\tilde{\bm{\zeta}}_{i}\|_{2}\leq 4R+2\|\bm{\theta}^{\star}\|_{2}+1.$
 In what follows, we prove the desired result in the following 4 steps:
@@
 We firstly observe that
-$\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\widetilde{\bm{\zeta}}_{i}\right\|_{2}^{2}$ $\leq\sum_{i=1}^{t}\|\bm{R}_{i}^{t}\|^{2}\|\widetilde{\bm{\zeta}}_{i}\|_{2}^{2}$
+$\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\tilde{\bm{\zeta}}_{i}\right\|_{2}^{2}$ $\leq\sum_{i=1}^{t}\|\bm{R}_{i}^{t}\|^{2}\|\tilde{\bm{\zeta}}_{i}\|_{2}^{2}$
 $<\eta_{0}^{2}(4R+2\|\bm{\theta}^{\star}\|_{2}+1)^{2}\underset{W_{t}}{\underbrace{\sum_{i=1}^{t}i^{-2\alpha}\prod_{k=1}^{t}(1-\beta k^{-\alpha})}}.$
@@
 Therefore, the vectors Azuma’s inequality (Corollary A.18) shows that, with probability at least $1-\frac{\delta}{3t^{\star}}$,
-$\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\widetilde{\bm{\zeta}}_{i}\right\|_{2}$ $<2\eta_{0}(4R+2\|\bm{\theta}^{\star}\|_{2}+1)\sqrt{\frac{2}{2\alpha-1}\log\frac{9t^{\star}}{\delta}}$
+$\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\tilde{\bm{\zeta}}_{i}\right\|_{2}$ $<2\eta_{0}(4R+2\|\bm{\theta}^{\star}\|_{2}+1)\sqrt{\frac{2}{2\alpha-1}\log\frac{9t^{\star}}{\delta}}$
 $\leq\frac{1}{8}(4R+2\|\bm{\theta}^{\star}\|_{2}+1),$
@@
 $\|\bm{\Delta}_{t}\|$ $\leq\left\|\prod_{k=1}^{t}(\bm{I}-\eta_{k}\bm{A})\bm{\Delta}_{0}\right\|_{2}+\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\bm{\zeta}_{i}\right\|_{2}$
-$=\left\|\prod_{k=1}^{t}(\bm{I}-\eta_{k}\bm{A})\bm{\Delta}_{0}\right\|_{2}+\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\widetilde{\bm{\zeta}}_{i}\right\|_{2}$
+$=\left\|\prod_{k=1}^{t}(\bm{I}-\eta_{k}\bm{A})\bm{\Delta}_{0}\right\|_{2}+\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\tilde{\bm{\zeta}}_{i}\right\|_{2}$
 $<\|\bm{\theta}^{\star}\|_{2}+\frac{1}{8}(4R+2\|\bm{\theta}^{\star}\|_{2}+1)=R,$
@@
 it can then be seen that with probability at least $1-\delta_{t}$,
-$\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\widetilde{\bm{\zeta}}_{i}\right\|_{2}$ $\leq 2\eta_{0}(4R+2\|\bm{\theta}^{\star}\|_{2}+1)\sqrt{2W_{t}\log\frac{3}{\delta_{t}}}$
+$\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\tilde{\bm{\zeta}}_{i}\right\|_{2}$ $\leq 2\eta_{0}(4R+2\|\bm{\theta}^{\star}\|_{2}+1)\sqrt{2W_{t}\log\frac{3}{\delta_{t}}}$
 $\leq 2\eta_{0}(4R+2\|\bm{\theta}^{\star}\|_{2}+1)\cdot\frac{1}{16}=\frac{1}{2}R+\frac{1}{4}(2\|\bm{\theta}^{\star}\|_{2}+1).$
@@
 Meanwhile, the vector Azuma’s inequality (Corollary A.18) implies with probability at least $1-\frac{\delta}{3T}$,
-$\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\widetilde{\bm{\zeta}}_{i}\right\|_{2}\leq 2\eta_{0}\left(\frac{2\alpha}{\beta}\right)^{\frac{\alpha}{2(1-\alpha)}}\sqrt{\frac{2}{2\alpha-1}\log\frac{9T}{\delta}}(4R+2\|\bm{\theta}^{\star}\|_{2}+1)t^{-\frac{\alpha}{2}}.$
+$\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\tilde{\bm{\zeta}}_{i}\right\|_{2}\leq 2\eta_{0}\left(\frac{2\alpha}{\beta}\right)^{\frac{\alpha}{2(1-\alpha)}}\sqrt{\frac{2}{2\alpha-1}\log\frac{9T}{\delta}}(4R+2\|\bm{\theta}^{\star}\|_{2}+1)t^{-\frac{\alpha}{2}}.$
 The bound (66) then follows by invoking the triangle inequality and a union bound, and plugging in the definition of $R$ and $\beta$. ∎
-B.2 Proof of Theorem 3.1
+B.1.2 Completing the proof of Theorem 3.1
 Similar to the proof of Theorem B.1, we firstly expand $\overline{\bm{\Delta}}_{T}$ into three terms: a non-random term capturing the error introduced by the intialization $\bm{\theta}_{0}$, a martingale, and a sum of $i.i.d.$ vectors.
-Specifically, (B.1) implies
+Specifically, (B.1.1) implies
 $\overline{\bm{\Delta}}_{T}$ $=\frac{1}{T}\sum_{t=1}^{T}\left(\prod_{k=1}^{t}(\bm{I}-\eta_{k}\bm{A})\Delta_{0}-\sum_{i=1}^{t}\bm{R}_{i}^{t}(\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i})-\sum_{i=1}^{t}\bm{R}_{i}^{t}(\bm{A}_{i}-\bm{A})\bm{\Delta}_{i-1}\right)$
@@
 Bounding the norm of $I_{2}$.
-The term $I_{2}$ is a sum of $i.i.d.$ random vectors, and can hence be bounded by invoking the matrix Berstein’s inequality. Specifically, define
+The term $I_{2}$ is a sum of $i.i.d.$ random vectors, and can hence be bounded by invoking the Freedman’s inequality on Hilbert spaces (Theorem A.19). Specifically, define
 $W_{\max}:=\sum_{i=1}^{T}\mathbb{E}\left\|\frac{1}{T}\bm{Q}_{i}(\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i})\right\|_{2}^{2},\quad\text{and}$
@@
 It can then be guaranteed with probability at least $1-\delta$ that
-$\left\|\frac{1}{T}\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i})\right\|_{2}\leq 2\sqrt{W_{\max}\log\frac{2d}{\delta}}+\frac{4}{3}B_{\max}\log\frac{2d}{\delta}.$
+$\left\|\frac{1}{T}\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i})\right\|_{2}\leq 2\sqrt{W_{\max}\log\frac{2}{\delta}}+\frac{4}{3}B_{\max}\log\frac{2}{\delta}.$
 Furthermore, the terms $W_{\max}$ and $B_{\max}$ can be controlled by
@@
 Hence, the norm of $I_{2}$ is bounded by
-$\left\|\frac{1}{T}\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i})\right\|_{2}\leq 2\sqrt{\frac{\mathsf{Tr}(\overline{\bm{\Lambda}}_{T})}{T}\log\frac{6d}{\delta}}+\frac{4\eta_{0}}{T}\left(\frac{2}{\beta}\right)^{\frac{1}{1-\alpha}}(2\|\bm{\theta}^{\star}\|_{2}+1)\log\frac{6d}{\delta}.$ (72)
+$\left\|\frac{1}{T}\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i})\right\|_{2}\leq 2\sqrt{\frac{\mathsf{Tr}(\overline{\bm{\Lambda}}_{T})}{T}\log\frac{6}{\delta}}+\frac{4\eta_{0}}{T}\left(\frac{2}{\beta}\right)^{\frac{1}{1-\alpha}}(2\|\bm{\theta}^{\star}\|_{2}+1)\log\frac{6}{\delta}.$ (75)
 with probability at least $1-\frac{\delta}{3}$.
@@
 $\mathcal{H}_{t}=\left\{\max_{1\leq j\leq t}j^{\frac{\alpha}{2}}\|\bm{\Delta}_{j}\|_{2}\leq R\right\},\quad\text{and}$
-$\widetilde{\bm{\Delta}}_{t}=\bm{\Delta}_{t}\mathds{1}(\mathcal{H}_{t}).$
+$\tilde{\bm{\Delta}}_{t}=\bm{\Delta}_{t}\mathds{1}(\mathcal{H}_{t}).$
 Then by definition, the probability of $\mathcal{H}_{t}$ is bounded by
@@
 Meanwhile, the vector Azuma’s inequality can be invoked to obtain that with probability at least $1-\frac{\delta}{3}$,
-$\left\|\frac{1}{T}\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}-\bm{A})\widetilde{\bm{\Delta}}_{i-1}\right\|_{2}\leq 2\sqrt{2W_{\max}\log\frac{9}{\delta}},$
+$\left\|\frac{1}{T}\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}-\bm{A})\tilde{\bm{\Delta}}_{i-1}\right\|_{2}\leq 2\sqrt{2W_{\max}\log\frac{9}{\delta}},$
 where
-$W_{\max}=\sup\frac{1}{T^{2}}\sum_{i=1}^{T}\|\bm{Q}_{i}(\bm{A}_{i}-\bm{A})\widetilde{\bm{\Delta}}_{i-1}\|_{2}^{2}$ $\lesssim\frac{1}{T^{2}}3\eta_{0}^{2}\left(\frac{2}{\beta}\right)^{\frac{1}{1-\alpha}}\cdot 4R^{2}\sum_{i=1}^{T}i^{-\alpha}$
+$W_{\max}=\sup\frac{1}{T^{2}}\sum_{i=1}^{T}\|\bm{Q}_{i}(\bm{A}_{i}-\bm{A})\tilde{\bm{\Delta}}_{i-1}\|_{2}^{2}$ $\lesssim\frac{1}{T^{2}}3\eta_{0}^{2}\left(\frac{2}{\beta}\right)^{\frac{1}{1-\alpha}}\cdot 4R^{2}\sum_{i=1}^{T}i^{-\alpha}$
 $\asymp\frac{12\eta_{0}^{2}}{1-\alpha}\left(\frac{2}{\beta}\right)^{\frac{1}{1-\alpha}}R^{2}T^{-1-\alpha}.$
 In other words, it can be guaranteed with probability at least $1-\frac{\delta}{3}$ that
-$\left\|\frac{1}{T}\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}-\bm{A})\widetilde{\bm{\Delta}}_{i-1}\right\|_{2}$ $\lesssim\eta_{0}\left(\frac{2}{\beta}\right)^{\frac{1}{2(1-\alpha)}}R\sqrt{\frac{1}{1-\alpha}\log\frac{9}{\delta}}$
+$\left\|\frac{1}{T}\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}-\bm{A})\tilde{\bm{\Delta}}_{i-1}\right\|_{2}$ $\lesssim\eta_{0}\left(\frac{2}{\beta}\right)^{\frac{1}{2(1-\alpha)}}R\sqrt{\frac{1}{1-\alpha}\log\frac{9}{\delta}}$
 $\lesssim\eta_{0}^{2}\left(\frac{2}{\beta}\right)^{\frac{2+\alpha}{2(1-\alpha)}}(6\|\bm{\theta}^{\star}\|_{2}+1)\sqrt{\frac{1}{(1-\alpha)(2\alpha-1)}}T^{-\frac{\alpha+1}{2}}\log\frac{27T}{\delta}.$ (77)
-Combining (70), (71), (72), (73) and (B.2) with the triangle inequality and a union bound argument, we obtain with probability $1-\delta$ that
+Combining (73), (74), (75), (76) and (B.1.2) with the triangle inequality and a union bound argument, we obtain with probability $1-\delta$ that
 $\|\overline{\bm{\Delta}}_{T}\|_{2}$ $\leq\frac{3}{T}\left(\frac{2}{\beta}\right)^{\frac{1}{1-\alpha}}\|\bm{\Delta}_{0}\|_{2}$
-$+2\sqrt{\frac{\mathsf{Tr}(\overline{\bm{\Lambda}}_{T})}{T}\log\frac{6d}{\delta}}+\frac{4\eta_{0}}{T}\left(\frac{2}{\beta}\right)^{\frac{1}{1-\alpha}}(2\|\bm{\theta}^{\star}\|_{2}+1)\log\frac{6d}{\delta}$
+$+2\sqrt{\frac{\mathsf{Tr}(\overline{\bm{\Lambda}}_{T})}{T}\log\frac{6}{\delta}}+\frac{4\eta_{0}}{T}\left(\frac{2}{\beta}\right)^{\frac{1}{1-\alpha}}(2\|\bm{\theta}^{\star}\|_{2}+1)\log\frac{6}{\delta}$
 $+\eta_{0}^{2}\left(\frac{2}{\beta}\right)^{\frac{2+\alpha}{2(1-\alpha)}}(6\|\bm{\theta}^{\star}\|_{2}+1)\sqrt{\frac{1}{(1-\alpha)(2\alpha-1)}}T^{-\frac{\alpha+1}{2}}\log\frac{27T}{\delta}.$
 Theorem 3.1 follows by taking
-$\widetilde{C}$ $=\eta_{0}^{2}\left(\frac{2}{\beta}\right)^{\frac{2+\alpha}{2(1-\alpha)}}(6\|\bm{\theta}^{\star}\|_{2}+1)\sqrt{\frac{1}{(1-\alpha)(2\alpha-1)}}.$ (75)
-B.3 Proof of Theorem 3.2
+$\tilde{C}$ $=\eta_{0}^{2}\left(\frac{2}{\beta}\right)^{\frac{2+\alpha}{2(1-\alpha)}}(6\|\bm{\theta}^{\star}\|_{2}+1)\sqrt{\frac{1}{(1-\alpha)(2\alpha-1)}}.$ (78)
+B.1.3 The necessity of the condition (71)
+An important fact to note about Theorem 3.1 is that it relies on choosing an appropriate initial stepsize $\eta_{0}$ dependent of $\delta$.
+The relationship is specified in Equation (71).
+This constraint, albeit only related to $\delta$ in logarithm terms, prevents $\delta$ to converge to 00 for any fixed $\eta_{0}$.
+One might naturally wonder if such a constraint could be relaxed and the convergence of $\overline{\bm{\Delta}}_{T}$ can be guaranteed with high probability for any fixed $\eta_{0}$ and arbitrarily small $\delta$.
+The following proposition shows that the answer is negative, so some form of constraint on $\delta$ is necessary for a Chernoff-style convergence guarantee.
+Proposition B.2.
+Consider the TD algorithm (7) with independent samples and decaying stepsizes $\eta_{t}=\eta_{0}t^{-\alpha}$, where $\eta_{0}$ is fixed and $\alpha\in(\frac{1}{2},1)$.
+There exists an MDP, such that for any bi-variate polynomial function $\mathsf{poly}(\cdot,\cdot)$, there exists $\delta\in(0,1)$ and $T\in\mathbb{N}$, such that
+$\mathbb{P}\left(\|\overline{\bm{\Delta}}_{T}\|_{2}\geq\mathsf{poly}\left(T,\log\frac{1}{\delta}\right)\right)\geq\delta.$
+Proof: Consider the following MDP, where $|\mathcal{S}|=3$, $\gamma=0.9$, while the transition matrix $\bm{P}$ and the reward vector $\bm{r}$ are
+$\bm{P}=\begin{pmatrix}0.1&0.1&0.8\\ 0.1&0.1&0.8\\ 0.1&0.1&0.8\end{pmatrix},\quad\text{and}\quad\bm{r}=\begin{pmatrix}0.1\\ 0.1\\ 1\end{pmatrix}.$
+[…]
+$\bm{A}=0.54475,\quad\bm{b}=0.785,\quad\text{and}\quad\bm{\theta}^{\star}=1.441.$
+For any $T\in\mathbb{N}$, use $\mathcal{E}_{T}$ to denote the event that $s_{t}=1$ and $s_{t}^{\prime}=2$ for all $t\in[T]$.
+In the case of $i.i.d.$ samples, clearly $\mathbb{P}(\mathcal{E}_{T})=0.01^{T}$.
+For any bi-variate polynomial function $\mathsf{poly}(\cdot,\cdot)$, let $\delta=0.01^{T}$, then $\log\frac{1}{\delta}=T\log 100$.
+Therefore, $\mathsf{poly}(T,\log\frac{1}{\delta})=\mathsf{poly}(T,T\log 100)$ reduces to a univariate polynomial function of $T$.
+It suffices to show that under the event $\mathcal{E}_{T}$, the norm of $\overline{\bm{\Delta}}_{T}$ grows faster than any polynomial function of $T$.
+Under the event $\mathcal{E}_{T}$, direct calculation yields $\bm{A}_{t}=-0.2$ and $\bm{b}_{t}=-0.05$. Hence, the estimation error $\bm{\Delta}_{t}$ is featured by the iterative relation
+$\bm{\Delta}_{t}$ $=\bm{\Delta}_{t-1}-\eta_{t}(\bm{A}_{t}\bm{\Delta}_{t-1}+\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})$
+$=(1+0.2\eta_{t})\bm{\Delta}_{t-1}+(0.2\bm{\theta}^{\star}-0.05)\eta_{t}.$
+It is easy to show by induction that, when $\bm{\theta}_{0}=0$ and hence $\bm{\Delta}_{0}=-\bm{\theta}^{\star}$,
+$\bm{\Delta}_{t}$ $=(0.2\bm{\theta}^{\star}-0.05)\sum_{i=1}^{t}\eta_{i}\prod_{k=i+1}^{t}(1+0.2\eta_{k})+\prod_{k=1}^{t}(1+0.2\eta_{k})\bm{\Delta}_{0}$
+$=5(0.2\bm{\theta}^{\star}-0.05)\sum_{i=1}^{t}\left(\prod_{k=i}^{t}-\prod_{k=i+1}^{t}\right)(1+0.2\eta_{k})+\prod_{k=1}^{t}(1+0.2\eta_{k})(-\bm{\theta}^{\star})$
+$=5(0.2\bm{\theta}^{\star}-0.05)\left(\prod_{k=1}^{t}(1+0.2\eta_{k})-1\right)-\prod_{k=1}^{t}(1+0.2\eta_{k})\bm{\theta}^{\star}$
+$=-0.25\prod_{k=1}^{t}(1+0.2\eta_{k})+(0.25-\bm{\theta}^{\star})$
+$=-0.25\prod_{k=1}^{t}(1+0.2\eta_{k})-1.191.$
+Consequently, the averaged TD estimation error is featured by
+$\overline{\bm{\Delta}}_{T}=-1.191-0.25\frac{1}{T}\sum_{t=1}^{T}\prod_{k=1}^{t}(1+0.2\eta_{k}),$
+and therefore
+$\left\|\overline{\bm{\Delta}}_{T}\right\|_{2}=1.191+0.25\frac{1}{T}\sum_{t=1}^{T}\prod_{k=1}^{t}(1+0.2\eta_{k})$
+We now consider the growth of $\left\|\overline{\bm{\Delta}}_{T}\right\|_{2}$ with polynomial-decaying stepsizes $\eta_{t}=\eta_{0}t^{-\alpha}$, where $\alpha\in(\frac{1}{2},1)$.
+In this scenario, the norm of $\overline{\bm{\Delta}}_{T}$ is bounded by
+$\left\|\overline{\bm{\Delta}}_{T}\right\|_{2}$ $=1.191+0.25\frac{1}{T}\sum_{t=1}^{T}\prod_{k=1}^{t}(1+0.2\eta_{k})$
+$=1.191+\frac{0.25}{T}\cdot\sum_{t=1}^{T}\frac{1}{0.2\eta_{t}}\left(\prod_{k=1}^{t}-\prod_{k=1}^{t-1}\right)(1+0.2\eta_{k})$
+$>1.191+\frac{0.25}{0.2\eta_{T}T}\sum_{t=1}^{T}\left(\prod_{k=1}^{t}-\prod_{k=1}^{t-1}\right)(1+0.2\eta_{k})$
+$=1.191+\frac{5}{4\eta_{0}}T^{-\alpha-1}\left(\prod_{k=1}^{T}(1+0.2\eta_{0}k^{-\alpha})-1\right).$
+Here, since $k^{-\alpha}\leq 1$, it can be guaranteed that
+$\prod_{k=1}^{T}(1+0.2\eta_{0}k^{-\alpha})$ $>\prod_{k=1}^{T}(1+0.2\eta_{0})^{k^{-\alpha}}$
+$=\exp\left(\log(1+0.2\eta_{0})\sum_{k=1}^{T}k^{-\alpha}\right)$
+$>\exp\left(\log(1+0.2\eta_{0})\cdot\frac{T^{1-\alpha}}{1-\alpha}\right).$
+Take $u=T^{1-\alpha}$, then the norm of $\overline{\bm{\Delta}}_{T}$ is bounded below by
+$\left\|\overline{\bm{\Delta}}_{T}\right\|_{2}>1.191+\frac{5}{4\eta_{0}}u^{-\frac{\alpha+1}{1-\alpha}}\left[\exp\left(\frac{\log(1+0.2\eta_{0})}{1-\alpha}u\right)-1\right],$
+which grows exponentially with $u$ and hence faster than any polynomial function of $T$. This completes the proof of the proposition. ∎
+B.2 Proof of Theorem 3.2
+By triangle inequality, we firstly decompose the convex distance between $\sqrt{T}\overline{\bm{\Delta}}_{T}$ and its asymptotic distribution $\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})$ by
+$d_{\mathsf{C}}(\sqrt{T}\overline{\bm{\Delta}}_{T},\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\leq d_{\mathsf{C}}(\sqrt{T}\overline{\bm{\Delta}}_{T},\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T}))+d_{\mathsf{C}}(\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T}),\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})).$ (79)
+Here, recall that $\overline{\bm{\Lambda}}_{T}$ is defined as in (49).
+The following two theorems bound the two terms on the right-hand-side of (79) respectively; Theorem 3.2 follows directly by combining them with (79).
+Theorem B.3.
+Consider the TD with Polyak-Ruppert averaging with independent samples and decaying stepsizes $\eta_{t}=\eta_{0}t^{-\alpha}$ for $\alpha\in(\frac{1}{2},1)$ and $\eta_{0}<1/2$. Then,
+$d_{\mathsf{C}}(\sqrt{T}\overline{\bm{\Delta}}_{T},\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T}))\lesssim\sqrt{\frac{\eta_{0}}{(1-\alpha)(1-\gamma)\lambda_{0}}}\mathsf{Tr}(\bm{\Gamma})\|\bm{\Gamma}^{-1}\|T^{-\frac{\alpha}{2}}+\tilde{C}\;T^{-\frac{1}{2}},$
+where $\tilde{C}$ is a problem related quantity independent of $T$.
+Theorem B.4.
+Let $\bm{\Lambda}^{\star}$ and $\overline{\bm{\Lambda}}_{T}$ be defined as in (12) and (49), respectively, with the stepsizes $\eta_{t}=\eta_{0}t^{-\alpha}$, for all $t>0$, any $\alpha\in(\frac{1}{2},1)$ and $\eta_{0}<1/2$.
+Then,
+$d_{\mathsf{C}}(\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T}),\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\leq d_{\mathsf{TV}}(\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T}),\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\lesssim\frac{\sqrt{d\mathsf{cond}(\bm{\Gamma})}}{(1-\gamma)\lambda_{0}\eta_{0}}T^{\alpha-1}+\tilde{C}^{\prime}\;T^{2\alpha-2},$
+where $\tilde{C}^{\prime}$ is a problem related quantity independent of $T$.
+B.2.1 Proof of Theorem B.3
 This proof is based on Theorem A.13, the Berry-Esseen bound for multivariate non-linear statistics.
@@
 $=\frac{1}{T}\mathbb{E}\sum_{j=i}^{T}\left\|\bm{AQ}_{j}(\bm{\omega}_{j}-\bm{\omega}_{j}^{(i)})\right\|_{2}^{2}+\frac{2}{T}\mathbb{E}\left\{\sum_{i\leq j<j^{\prime}<T}\left[\bm{AQ}_{j}\left(\bm{\omega}_{j}-\bm{\omega}_{j}^{(i)})\right)\right]^{\top}\left[\bm{AQ}_{j^{\prime}}\left(\bm{\omega}_{j^{\prime}}-\bm{\omega}_{j^{\prime}}^{(i)}\right)\right]\right\}.$ (88)
-For the last term on the right-hand-side of (B.3), recall from the definitions of $\bm{\omega}_{j}$ (78) and $\bm{\omega}_{j}^{(i)}$ (83), for any $j>i$,
+For the last term on the right-hand-side of (B.2.1), recall from the definitions of $\bm{\omega}_{j}$ (82) and $\bm{\omega}_{j}^{(i)}$ (87), for any $j>i$,
 $\bm{\omega}_{j}-\bm{\omega}_{j}^{(i)}$ $=\left[(\bm{A}_{j}-\bm{A})(\bm{\theta}_{j-1}-\bm{\theta}^{\star})\right]-\left[(\bm{A}_{j}-\bm{A})\left(\bm{\theta}_{j-1}^{(i)}-\bm{\theta}^{\star}\right)\right]$
@@
 since $\mathbb{E}_{j^{\prime}-1}[\bm{A}_{j^{\prime}}-\bm{A}]=\bm{0}$.
-This effectively means that the last term on the right-hand-side of (B.3) is 00.
+This effectively means that the last term on the right-hand-side of (B.2.1) is 00.
 Therefore, the difference bewteen $\Delta$ and $\Delta^{(i)}$ is bounded by
@@
 $\lesssim\mathbb{E}\left\|\bm{\theta}_{j-1}-\bm{\theta}_{j-1}^{(i)}\right\|_{2}^{2}.$ (92)
-By combining (B.3), (B.3),(B.3) with (86b), we obtain
+By combining (B.2.1), (B.2.1),(B.2.1) with (90b), we obtain
 $\mathbb{E}|\Delta-\Delta^{(i)}|^{2}$ $\leq\frac{1}{T}\left((2+O(i^{\alpha-1}))\mathbb{E}\|\bm{\omega}_{i}-\bm{\omega}_{i}^{(i)}\|_{2}^{2}+\sum_{j=i+1}^{T}(2+O(j^{\alpha-1}))\mathbb{E}\|\bm{\omega}_{j}-\bm{\omega}_{j}^{(i)}\|_{2}^{2}\right)$
@@
 $+\frac{1}{T}\underset{I_{2}}{\underbrace{\sum_{j=i+1}^{T}(2+O(j^{\alpha-1}))\prod_{k=i+1}^{j}\left(1-\frac{1-\gamma}{2}\lambda_{0}\eta_{k}\right)\mathbb{E}\|\bm{\theta}_{i}-\bm{\theta}_{i}^{(i)}\|_{2}^{2}}}.$
-With $I_{1}$ bounded by (B.3), the proof now boils down to bounding $I_{2}$. Notice that (86a) implies
+With $I_{1}$ bounded by (B.2.1), the proof now boils down to bounding $I_{2}$. Notice that (90a) implies
 $\sum_{j=i+1}^{T}\prod_{k=i+1}^{j}\left(1-\frac{1-\gamma}{2}\lambda_{0}\eta_{k}\right)\mathbb{E}\|\bm{\theta}_{i}-\bm{\theta}_{i}^{(i)}\|_{2}^{2}$
@@
 $=\frac{\eta_{0}\mathsf{Tr}(\bm{\Gamma})}{(1-\alpha)(1-\gamma)\lambda_{0}}T^{-\alpha}+O\left(\frac{\log T}{T}\right).$ (95)
-A direct combination of (89), (90) and (B.3) yields the following bound for the third term of (79):
+A direct combination of (93), (94) and (B.2.1) yields the following bound for the third term of (83):
 $\mathbb{E}\sum_{i=1}^{T}\|\bm{\xi}_{i}\||\Delta-\Delta^{(i)}|\lesssim\sqrt{\frac{\eta_{0}}{(1-\alpha)(1-\gamma)\lambda_{0}}}\mathsf{Tr}(\bm{\Gamma})T^{-\frac{\alpha}{2}}+O\left(\sqrt{\frac{\log T}{T}}\right).$ (96)
@@
 Taking supremum over $\mathcal{A}$ completes the proof. ∎
-B.4 Proof of Theorem 3.3
+B.2.2 Proof of Theorem B.4
 For any Borel set $\mathcal{A}\subset\mathbb{R}^{d}$, define $\mathcal{A}^{\prime}=\{\bm{Ax}:\bm{x}\in\mathbb{R}^{d}\}$.
@@
 $d_{\mathsf{C}}(\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T}),\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))$ $\leq\|\bm{Y}\|_{\mathsf{F}}T^{\alpha-1}+\|\bm{O}(T^{2\alpha-2})\|_{\mathsf{F}}=\frac{\sqrt{d\mathsf{cond}(\bm{\Gamma})}}{2(1-\gamma)\lambda_{0}\eta_{0}}T^{\alpha-1}+O(T^{2\alpha-2}).$
-B.4.1 Tightness of Theorem 3.3.
+B.2.3 Tightness of Theorem B.4
 As is indicated by Lemma A.11, the difference between $\overline{\bm{\Lambda}}_{T}$ and $\bm{\Lambda}^{\star}$ is featured by
@@
 Meanwhile, Theorem A.14 indicates the TV distance between $\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T})$ and $\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))$ converges at the same rate, i.e.
-$\left\|\bm{\Lambda}^{\star-\frac{1}{2}}\overline{\bm{\Lambda}}_{T}\bm{\Lambda}^{\star-\frac{1}{2}}-\bm{I}\right\|_{\mathsf{F}}=\left\|\bm{\Lambda}^{\star-\frac{1}{2}}(\overline{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star})\bm{\Lambda}^{\star-\frac{1}{2}}\right\|_{\mathsf{F}}=O(T^{\alpha-1}).$
+$d_{\mathsf{TV}}(\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T}),\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))$ $\approx\left\|\bm{\Lambda}^{\star-\frac{1}{2}}\overline{\bm{\Lambda}}_{T}\bm{\Lambda}^{\star-\frac{1}{2}}-\bm{I}\right\|_{\mathsf{F}}$
+$=\left\|\bm{\Lambda}^{\star-\frac{1}{2}}(\overline{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star})\bm{\Lambda}^{\star-\frac{1}{2}}\right\|_{\mathsf{F}}=O(T^{\alpha-1}).$
 Consequently, the $O(T^{\alpha-1})$ rate by which $\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T})$ converges to $\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})$ cannot be further improved.
-B.5 Proof of Theorem 3.4
-Applying the same reasoning as in the proof of Theorem 3.3,
+As a direct consequence, when $\alpha\in(\frac{2}{3},1)$, since $\alpha-1>-\frac{1}{3}>-\frac{\alpha}{2}$, it can be guaranteed through triangle inequality that
+$d_{\mathsf{C}}(\sqrt{T}\overline{\bm{\Delta}}_{T},\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))$ $\geq d_{\mathsf{C}}(\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T}),\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))-d_{\mathsf{C}}(\sqrt{T}\overline{\bm{\Delta}}_{T},\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T}))$
+$\geq\Theta(T^{\alpha-1})-O(T^{-\frac{\alpha}{2}})\geq\Theta(T^{-\frac{1}{3}}).$
+B.3 Proof of Theorem 3.3
+Applying the same reasoning as in the proof of Theorem B.4,
 $d_{\mathsf{TV}}(\mathcal{N}(\bm{0},\widehat{\bm{\Lambda}}_{T}),\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))$ $=d_{\mathsf{TV}}(\mathcal{N}(\bm{0},\bm{A}\widehat{\bm{\Lambda}}_{T}\bm{A}^{\top}),\mathcal{N}(\bm{0},\bm{\Gamma}))$
@@
 In order to bound the right-hand-side of the inequality above, we firstly present the difference between $\widehat{\bm{\Lambda}}$ and $\bm{\Lambda}^{\star}$ as
-$\widehat{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star}$ $=\overline{\bm{A}}^{-1}\widehat{\bm{\Gamma}}_{T}\overline{\bm{A}}^{-\top}-\bm{A}\bm{\Gamma}\bm{A}$
-$=\overline{\bm{A}}^{-1}(\widehat{\bm{\Gamma}}_{T}-\bm{\Gamma})\overline{\bm{A}}^{-\top}+\left[\overline{\bm{A}}^{-1}\bm{\Gamma}\overline{\bm{A}}^{-\top}-\bm{A}^{-1}\bm{\Gamma}\bm{A}^{-\top}\right].$ (93)
+$\widehat{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star}$ $=\overline{\bm{A}}_{T}^{-1}\widehat{\bm{\Gamma}}_{T}\overline{\bm{A}}_{T}^{-\top}-\bm{A}\bm{\Gamma}\bm{A}$
+$=\overline{\bm{A}}_{T}^{-1}(\widehat{\bm{\Gamma}}_{T}-\bm{\Gamma})\overline{\bm{A}}_{T}^{-\top}+\left[\overline{\bm{A}}_{T}^{-1}\bm{\Gamma}\overline{\bm{A}}_{T}^{-\top}-\bm{A}^{-1}\bm{\Gamma}\bm{A}^{-\top}\right].$ (97)
 The proof now boils down to bounding the estimation error of $\overline{\bm{A}}^{-1}$ and $\widehat{\bm{\Gamma}}$, for which the following lemmas come in handy.
 Lemma B.7.
-Let $\overline{\bm{A}}$ be defined as in (20) with $\{(s_{t},s_{t}^{\prime})\}_{0\leq t\leq T}$ being $i.i.d.$ samples. It can then be guaranteed with probability at least $1-\delta$ that
-$\|(\overline{\bm{A}}^{-1}-\bm{A}^{-1})\bm{A}\|_{\mathsf{F}}\leq\frac{4\sqrt{2}}{\lambda_{0}(1-\gamma)\sqrt{T}}\sqrt{\log\frac{3}{\delta}},$
+Let $\overline{\bm{A}}_{T}$ be defined as in (14) with $\{(s_{t},s_{t}^{\prime})\}_{0\leq t\leq T}$ being $i.i.d.$ samples. It can then be guaranteed with probability at least $1-\delta$ that
+$\|(\overline{\bm{A}}_{T}^{-1}-\bm{A}^{-1})\bm{A}\|_{\mathsf{F}}\leq\frac{4\sqrt{2}}{\lambda_{0}(1-\gamma)\sqrt{T}}\sqrt{\log\frac{3}{\delta}},$
 with the proviso that
@@
 Proof: See Appendix C.4. ∎
-Based on Lemmas B.4 and B.5, we now consider the estimation error of $\widehat{\bm{\Lambda}}_{T}$. By triangle inequality, (B.5) implies that
-$\left\|\bm{A}(\widehat{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star})\bm{A}^{\top}\right\|_{\mathsf{F}}\leq\underset{I_{1}}{\underbrace{\|\overline{\bm{A}}^{-1}\bm{A}\|^{2}\|\widehat{\bm{\Gamma}}_{T}-\bm{\Gamma}\|_{\mathsf{F}}}}+\underset{I_{2}}{\underbrace{\left\|\bm{A}\overline{\bm{A}}^{-1}\bm{\Gamma}\overline{\bm{A}}^{-\top}\bm{A}^{\top}-\bm{\Gamma}\right\|_{\mathsf{F}}}}.$
+Based on Lemmas B.7 and B.8, we now consider the estimation error of $\widehat{\bm{\Lambda}}_{T}$. By triangle inequality, (B.3) implies that
+$\left\|\bm{A}(\widehat{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star})\bm{A}^{\top}\right\|_{\mathsf{F}}\leq\underset{I_{1}}{\underbrace{\|\overline{\bm{A}}_{T}^{-1}\bm{A}\|^{2}\|\widehat{\bm{\Gamma}}_{T}-\bm{\Gamma}\|_{\mathsf{F}}}}+\underset{I_{2}}{\underbrace{\left\|\bm{A}\overline{\bm{A}}_{T}^{-1}\bm{\Gamma}\overline{\bm{A}}_{T}^{-\top}\bm{A}^{\top}-\bm{\Gamma}\right\|_{\mathsf{F}}}}.$
 In what follows, we bound $I_{1}$ and $I_{2}$ in order.
@@
 The norm of $\widehat{\bm{\Gamma}}_{T}-\bm{\Gamma}$ is bounded by Lemma B.8, while the norm of $\bm{A}\overline{\bm{A}}^{-1}$ is bounded by the fact that
-$\|\bm{A}\overline{\bm{A}}^{-1}\|\leq\|\bm{I}\|+\|\bm{A}(\overline{\bm{A}}^{-1}-\bm{A}^{-1})\|\leq 3$
+$\|\bm{A}\overline{\bm{A}}_{T}^{-1}\|\leq\|\bm{I}\|+\|\bm{A}(\overline{\bm{A}}_{T}^{-1}-\bm{A}^{-1})\|\leq 3$
 with probability at least $1-\delta$, according to Lemma B.7 under condition (98). In combination, $I_{1}$ can be bounded by
@@
 The norm of $I_{2}$ is bounded by estimation error of $\overline{\bm{A}}^{-1}$. Namely, we further decompose $I_{2}$ by triangle inequality as
-$I_{2}$ $\leq\|(\bm{A}\overline{\bm{A}}^{-1}-\bm{I})\bm{\Gamma}\|_{\mathsf{F}}+\|\bm{\Gamma}(\overline{\bm{A}}^{-\top}\bm{A}^{\top}-\bm{I})\|_{\mathsf{F}}+\|(\overline{\bm{A}}^{-1}-\bm{A}^{-1})\bm{\Gamma}(\overline{\bm{A}}^{-\top}-\bm{A}^{-\top})\|_{\mathsf{F}}$
-$\leq 2\|\bm{\Gamma}\|\|\bm{A}(\overline{\bm{A}}^{-1}-\bm{A}^{-1})\|_{\mathsf{F}}+\|\bm{\Gamma}\|\|\bm{A}(\overline{\bm{A}}^{-1}-\bm{A}^{-1})\|_{\mathsf{F}}^{2},$
+$I_{2}$ $\leq\|(\bm{A}\overline{\bm{A}}_{T}^{-1}-\bm{I})\bm{\Gamma}\|_{\mathsf{F}}+\|\bm{\Gamma}(\overline{\bm{A}}_{T}^{-\top}\bm{A}^{\top}-\bm{I})\|_{\mathsf{F}}+\|(\overline{\bm{A}}_{T}^{-1}-\bm{A}^{-1})\bm{\Gamma}(\overline{\bm{A}}_{T}^{-\top}-\bm{A}^{-\top})\|_{\mathsf{F}}$
+$\leq 2\|\bm{\Gamma}\|\|\bm{A}(\overline{\bm{A}}_{T}^{-1}-\bm{A}^{-1})\|_{\mathsf{F}}+\|\bm{\Gamma}\|\|\bm{A}(\overline{\bm{A}}_{T}^{-1}-\bm{A}^{-1})\|_{\mathsf{F}}^{2},$
 where we can apply Lemma B.7 to conclude that
@@
 with probability at least $1-T^{-\frac{1}{2}}$. Theorem 3.3 follows by combining (99) and (100).
-B.6 Proof of Theorem 3.5
+B.4 Proof of Theorem 3.4
 We approach the theorem in two steps: first measuring the difference between the distribution of $\sqrt{T}\overline{\bm{\Delta}}_{T}$ and $\mathcal{N}(\bm{0},\widehat{\bm{\Lambda}}_{T})$, and then bounding the probability that $\bm{\theta}$ is within the region $\mathcal{C}_{\delta}$.
@@
 +|($\quad+\underset{I_{2}}{\underbrace{\left|\mathbb{P}(\overline{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}\in\mathcal{A})-\mathbb{P}({\bm{\Lambda}}^{\star\frac{1}{2}}\bm{z}\in\mathcal{A})\right|}}+\underset{I_{3}}{\underbrace{\left|\mathbb{P}({\bm{\Lambda}}^{\star\frac{1}{2}}\bm{z}\in\mathcal{A})-\mathbb{P}(\widehat{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}\in\mathcal{A})\right|}}.$
-Here, $I_{1}$ and $I_{2}$ measures the difference between the non-asymptotic and asymptotic distributions of $\overline{\bm{\Delta}}_{T}$, and are bounded by Theorems 3.2 and 3.3 respectively.
+Here, $I_{1}$ and $I_{2}$ measures the difference between the non-asymptotic and asymptotic distributions of $\overline{\bm{\Delta}}_{T}$, and are bounded by Theorems B.3 and B.4 respectively.
 Specifically,
@@
 $\leq\mathsf{Tr}\left([\bm{\phi}(s_{t})-\gamma\bm{\phi}(s_{t}^{\prime})][\bm{\phi}(s_{t})-\gamma\bm{\phi}(s_{t}^{\prime})]^{\top}\right)<4$
-holds uniformly for all $0\leq t\leq T$. Hence the vector Azuma’s inequality A.16 directly implies that
-$\|\overline{\bm{A}}-\bm{A}\|_{\mathsf{F}}$ $=\left\|\mathbf{vec}(\overline{\bm{A}}-\bm{A})\right\|\leq\frac{4\sqrt{2}}{\sqrt{T}}\sqrt{\log\frac{3}{\delta}}$
+holds uniformly for all $0\leq t\leq T$. Hence the vector Azuma’s inequality A.18 directly implies that
+$\|\overline{\bm{A}}_{T}-\bm{A}\|_{\mathsf{F}}$ $=\left\|\mathbf{vec}(\overline{\bm{A}}_{T}-\bm{A})\right\|\leq\frac{4\sqrt{2}}{\sqrt{T}}\sqrt{\log\frac{3}{\delta}}$
 with probability at least $1-\delta$.
-In what follows, we aim to transform this error bound to the error bound for $\overline{\bm{A}}^{-1}$.
-Towards this end, we observe that condition 94 guarantees with probability at least $1-\delta$,
-$\left\|\bm{A}^{-1}(\overline{\bm{A}}-\bm{A})\right\|$ $\leq\|\bm{A}^{-1}\|\|\overline{\bm{A}}-\bm{A}\|$
+In what follows, we aim to transform this error bound to the error bound for $\overline{\bm{A}}_{T}^{-1}$.
+Towards this end, we observe that condition (98) guarantees with probability at least $1-\delta$,
+$\left\|\bm{A}^{-1}(\overline{\bm{A}}_{T}-\bm{A})\right\|$ $\leq\|\bm{A}^{-1}\|\|\overline{\bm{A}}_{T}-\bm{A}\|$
 $\leq\lambda_{0}^{-1}(1-\gamma)^{-1}\frac{4\sqrt{2}}{\sqrt{T}}\sqrt{\log\frac{3}{\delta}}<\frac{1}{2}$
@@
 Consequently, the difference between $\overline{\bm{A}}^{-1}\bm{A}$ and $\bm{I}$ is bounded by
-$\|(\overline{\bm{A}}^{-1}-\bm{A}^{-1})\bm{A}\|_{\mathsf{F}}\leq 2\|(\overline{\bm{A}}-\bm{A})\bm{A}^{-1}\|_{\mathsf{F}}$ $\leq\frac{4\sqrt{2}}{\sqrt{T}}\sqrt{\log\frac{3}{\delta}}\|\bm{A}^{-1}\|$
+$\|(\overline{\bm{A}}_{T}^{-1}-\bm{A}^{-1})\bm{A}\|_{\mathsf{F}}\leq 2\|(\overline{\bm{A}}_{T}-\bm{A})\bm{A}^{-1}\|_{\mathsf{F}}$ $\leq\frac{4\sqrt{2}}{\sqrt{T}}\sqrt{\log\frac{3}{\delta}}\|\bm{A}^{-1}\|$
 $\leq\frac{4\sqrt{2}}{\lambda_{0}(1-\gamma)\sqrt{T}}\sqrt{\log\frac{3}{\delta}}.$
@@
 Since
-$\mathbb{E}\|\overline{\bm{\Delta}}_{T}\|_{2}^{2}$ $=\frac{1}{T}\mathsf{Tr}(\overline{\bm{\Lambda}}_{T})+\widetilde{C}T^{-\alpha-1}=\frac{1}{T}\mathsf{Tr}(\bm{\Lambda}^{\star})+o(T^{-1})\leq\frac{\mathsf{Tr}(\bm{\Gamma})}{\lambda_{0}^{2}(1-\gamma)^{2}T}+o(T^{-1}),$
+$\mathbb{E}\|\overline{\bm{\Delta}}_{T}\|_{2}^{2}$ $=\frac{1}{T}\mathsf{Tr}(\overline{\bm{\Lambda}}_{T})+\tilde{C}T^{-\alpha-1}=\frac{1}{T}\mathsf{Tr}(\bm{\Lambda}^{\star})+o(T^{-1})\leq\frac{\mathsf{Tr}(\bm{\Gamma})}{\lambda_{0}^{2}(1-\gamma)^{2}T}+o(T^{-1}),$
 the Markov’s inequality indicates
@@
 Meanwhile, the norm of $I_{12}$ is bounded by
-$\left\|\frac{1}{T}\sum_{t=1}^{T}\left[(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}+(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}\right]\right\|_{\mathsf{F}}$
-$\quad\leq\frac{1}{T}\sum_{t=1}^{T}\left\|(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}+(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}\right\|_{\mathsf{F}}$
-$\quad\leq\frac{2}{T}\sum_{t=1}^{T}\left\|(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}\right\|_{\mathsf{F}}$
-$\quad\leq\frac{2}{T}\sum_{t=1}^{T}\sqrt{\mathsf{Tr}\left((\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}\right)}$
-$\quad\leq\frac{4}{T}\sum_{t=1}^{T}4(2\|\bm{\theta}^{\star}\|_{2}+1)\|\overline{\bm{\Delta}}_{T}\|_{2}$
-$\quad=4(2\|\bm{\theta}^{\star}\|_{2}+1)\|\overline{\bm{\Delta}}_{T}\|_{2}.$ (105)
+$\left\|\frac{1}{T}\sum_{t=1}^{T}\left[(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}+(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}\right]\right\|_{\mathsf{F}}$ (109)
+$\leq\frac{1}{T}\sum_{t=1}^{T}\left\|(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}+(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}\right\|_{\mathsf{F}}$
+$\leq\frac{2}{T}\sum_{t=1}^{T}\left\|(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}\right\|_{\mathsf{F}}$
+$\leq\frac{2}{T}\sum_{t=1}^{T}\sqrt{\mathsf{Tr}\left((\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}\right)}$
+$\leq\frac{4}{T}\sum_{t=1}^{T}4(2\|\bm{\theta}^{\star}\|_{2}+1)\|\overline{\bm{\Delta}}_{T}\|_{2}$
+$=4(2\|\bm{\theta}^{\star}\|_{2}+1)\|\overline{\bm{\Delta}}_{T}\|_{2}.$ (110)
 Therefore, the Chebyshev’s inequality indicates
@@
 Appendix D Experimental details
-D.1 The Markov decision process and the feature map
+D.1 Online algorithm to construct the variance estimator
+Algorithm 1 Online algorithm for computing $\overline{\bm{\theta}}_{T}$ and $\widehat{\bm{\Lambda}}_{T}$
+1:MDP transition kernel $\mathcal{P}$, stationary distribution $\mu$, discount factor $\gamma$, feature map $\bm{\phi}$, initial stepsize $\eta_{0}$, stepsize decay parameter $\alpha\in(0.5,1)$, sample size $T$
+2:Initialize $\bm{\theta}=\bm{0}$, $\overline{\bm{\theta}}=\bm{0}$, $\overline{\bm{A}}=\bm{0}$, $\overline{\bm{AA}}=\bm{0}$, $\overline{\bm{Ab}}=\bm{0}$, $\overline{\bm{bb}}=\bm{0}$
+3:for $t=0:(T-1)$ do
+4:     Sample $s\sim\mu$, $s^{\prime}\sim\mathcal{P}(\cdot\mid s)$
+5:     Compute:
+$\bm{A}\leftarrow\bm{\phi}(s)[\bm{\phi}(s)-\gamma\bm{\phi}(s^{\prime})],\quad\bm{b}\leftarrow r(s)\bm{\phi}(s)$
+6:     Perform TD update:
+$\bm{\theta}\leftarrow\bm{\theta}-\eta_{0}t^{-\alpha}(\bm{A}\bm{\theta}-\bm{b})$
+7:     Perform online updates:
+$\overline{\bm{\theta}}\leftarrow\overline{\bm{\theta}}+\frac{1}{t}(\bm{\theta}-\overline{\bm{\theta}})$
+$\overline{\bm{A}}\leftarrow\overline{\bm{A}}+\frac{1}{t}(\bm{A}-\overline{\bm{A}})$
+$\overline{\bm{AA}}\leftarrow\overline{\bm{AA}}+\frac{1}{t}(\bm{A}\otimes\bm{A}-\overline{\bm{AA}})$
+$\overline{\bm{Ab}}\leftarrow\overline{\bm{Ab}}+\frac{1}{t}(\bm{A}\otimes\bm{b}+\bm{b}\otimes\bm{A}-\overline{\bm{Ab}})$
+$\overline{\bm{bb}}\leftarrow\overline{\bm{bb}}+\frac{1}{t}(\bm{b}\otimes\bm{b}-\overline{\bm{bb}})$
+8:end for
+9:Construct $\widehat{\bm{\Gamma}}$:
+$\widehat{\bm{\Gamma}}\leftarrow\mathsf{reshape}(\overline{\bm{AA}}(\overline{\bm{\theta}}\otimes\overline{\bm{\theta}})-2\overline{\bm{Ab}}(\overline{\bm{\theta}})+\overline{\bm{bb}},(d\times d))$
+10:Construct the variance estimator:
+$\widehat{\bm{\Lambda}}\leftarrow\overline{\bm{A}}^{-1}\widehat{\bm{\Gamma}}\overline{\bm{A}}^{-\top}$
+11:Averaged TD estimator $\overline{\bm{\theta}}$, covariance matrix estimator $\widehat{\bm{\Lambda}}$
+The basic properties of Kronecker products indicate
+$\textbf{vec}\left(\bm{A}_{t}\overline{\bm{\theta}}_{T}\overline{\bm{\theta}}_{T}^{\top}\bm{A}_{t}\right)=(\bm{A}_{t}\otimes\bm{A}_{t})(\overline{\bm{\theta}}_{T}\otimes\overline{\bm{\theta}}_{T});$
+$\textbf{vec}\left(\bm{A}_{t}\overline{\bm{\theta}}_{T}\overline{\bm{b}}_{t}^{\top}\right)=(\bm{A}_{t}\otimes\bm{b}_{t})\overline{\bm{\theta}}_{T};$
+$\textbf{vec}\left(\bm{b}_{t}\bm{\overline{\theta}}_{T}^{\top}\bm{A}_{t}^{\top}\right)=(\bm{b}_{t}\otimes\bm{A}_{t})\overline{\bm{\theta}}_{T};$
+$\textbf{vec}\left(\bm{b}_{t}\bm{b}_{t}^{\top}\right)=\bm{b}_{t}\otimes\bm{b}_{t}.$
+Therefore, the vectorization of $\widehat{\bm{\Gamma}}_{T}$ can be represented as
+$\textbf{vec}(\widehat{\bm{\Gamma}}_{T})$ $=\frac{1}{T}\sum_{t=1}^{T}\textbf{vec}\left(\bm{A}_{t}\overline{\bm{\theta}}_{T}\overline{\bm{\theta}}_{T}^{\top}\bm{A}_{t}-\bm{A}_{t}\overline{\bm{\theta}}_{T}\overline{\bm{b}}_{t}^{\top}-\bm{b}_{t}\bm{\overline{\theta}}_{T}^{\top}\bm{A}_{t}^{\top}+\bm{b}_{t}\bm{b}_{t}^{\top}\right)$
+$=\frac{1}{T}\sum_{t=1}^{T}\left((\bm{A}_{t}\otimes\bm{A}_{t})(\overline{\bm{\theta}}_{T}\otimes\overline{\bm{\theta}}_{T})-(\bm{A}_{t}\otimes\bm{b}_{t})\overline{\bm{\theta}}_{T}-(\bm{b}_{t}\otimes\bm{A}_{t})\overline{\bm{\theta}}_{T}+\bm{b}_{t}\otimes\bm{b}_{t}\right)$
+=($=\underset{I_{1}}{\underbrace{\left(\frac{1}{T}\sum_{t=1}^{T}\bm{A}_{t}\otimes\bm{A}_{t}\right)}}(\overline{\bm{\theta}}_{T}\otimes\overline{\bm{\theta}}_{T})-\underset{I_{2}}{\underbrace{\left(\frac{1}{T}\sum_{t=1}^{T}[\bm{A}_{t}\otimes\bm{b}_{t}+\bm{b}_{t}\otimes\bm{A}_{t}]\right)}}\overline{\bm{\theta}}_{T}$
+$+\underset{I_{3}}{\underbrace{\frac{1}{T}\sum_{t=1}^{T}\bm{b}_{t}\otimes\bm{b}_{t}}},$
+where $I_{1}$, $I_{2}$ and $I_{3}$ are sample means and can hence be calculated using online updates.
+Based on this, we propose Algorithm 1 for the calculation $\widehat{\bm{\Lambda}}_{T}$ in an online manner.
+Since Algorithm 1 iterates from $t=0$ to $t=T-1$ in a single loop and does not store the samples for each iteration, it has a space complexity is $O(Td^{4})$ and a space complexity of $O(d^{4})$.
+Here, the $d^{4}$ term stems from calculating the Kronecker product $\bm{A}\otimes\bm{A}$.
+D.2 The Markov Decision Process and the feature map
 Given a finite state space $\mathcal{S}$, define a sequence of MDP $\{\mathcal{M}_{\bm{q}}\}$ indexed by $\bm{q}\in\mathcal{Q}\subset\{q_{+},q_{-}\}^{d-1}$ where for each $\bm{q}$, the transition kernel equals to
@@
 From the expressions above, we remark that, the values of $q$ and $V^{\star}(s)$ with $s\geq d$ are fixed for all $\bm{q}\in\mathcal{Q}$ which is ensured by the construction (117).
-It is shown in Theorem 2 of Li et al., 2024b that the sequence $\{\mathcal{M}_{\bm{q}}\}_{q\in\mathcal{Q}}$ is hard to distinguish for a considerably large set $\mathcal{Q}$.
-D.2 Parameter specification
+It is shown in Theorem 2 of Li et al.
+(2024) that the sequence $\{\mathcal{M}_{\bm{q}}\}_{q\in\mathcal{Q}}$ is hard to distinguish for a considerably large set $\mathcal{Q}$.
+D.3 Parameter specification
 In our experiments, we fix the size of the state space to $\|\mathcal{S}\|=10$, the discount factor to be $\gamma=0.2$, and let $\varepsilon=0.01$.
@@
   2. 2.
-    In order to illustrate the Berry-Esseen bound as indicated by Theorems 3.2 and 3.3, we measure the difference between the empirical and asymptotic distributions of $\overline{\bm{\Delta}}_{T}$ by
+    In order to illustrate the Berry-Esseen bound as indicated by Theorems B.3 and B.4, we measure the difference between the empirical and asymptotic distributions of $\overline{\bm{\Delta}}_{T}$ by
     $d(\sqrt{T}\overline{\bm{\Delta}}_{T},\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})):=\max_{1\leq j\leq d}\sup_{x\in\mathbb{R}}\left|\mathbb{P}(\sqrt{T}(\overline{\bm{\Delta}}_{T})_{j}\leq x)-\mathbb{P}((\bm{\Lambda}^{\star\frac{1}{2}}\bm{z})_{j}\leq x)\right|,$
@@
   3. 3.
-    In order to illustrate the efficiency of our variance esitmator $\widehat{\bm{\Lambda}}_{T}$, we measure the averaged Frobenius norm of the difference between $\widehat{\bm{\Lambda}}_{T}$ and $\bm{\Lambda}^{\star}$ among $100$ independent trials.
-As is indicated by A.12, this is a direct proxy for the TV distance between the zero-mean Gaussian distributions with the two corresponding variance matrices.
+    In order to illustrate the efficiency of our variance estimator $\widehat{\bm{\Lambda}}_{T}$, we measure the averaged Frobenius norm of the difference between $\widehat{\bm{\Lambda}}_{T}$ and $\bm{\Lambda}^{\star}$ among $100$ independent trials.
+As is indicated by A.14, this is a direct proxy for the TV distance between the zero-mean Gaussian distributions with the two corresponding variance matrices.
   4. 4.
     In order to validate our statistical inference procedure, we run Algorithm 1 on the MDP with $d=3$ using $\alpha=\frac{2}{3}$, construct both individual and simultaneous confidence intervals for $\bm{\theta}^{\star}$, and measure the coverage rates of these confidence intervals in $1000$ independent trials.
-Results are shown in section 4.2.
-D.3 Full experimental results
-D.3.1 High-probability convergence guarantee for $\overline{\bm{\Delta}}_{T}$
+Details to follow in section D.4.4.
+D.4 Full experimental results
+D.4.1 High-probability convergence guarantee for the TD estimator
 Figure 3 shows the decay of the $95\%$ quantile of the $L_{2}$ norm of $\overline{\bm{\Delta}}_{T}$; the slope of this log-log plot verifies our theoretical claim in Theorem 3.1 that $\|\overline{\bm{\Delta}}_{T}\|_{2}$ converges by the rate of $O(T^{-\frac{1}{2}})$ with high probability.
@@
 Both are in accordance with our conclusion in Theorem (1) that $\|\overline{\bm{\Delta}}_{T}\|_{2}$ converges by the rate of
-$O\left(\sqrt{\frac{\mathsf{Tr}(\bm{\Lambda}^{\star})}{T}\log\frac{d}{\delta}}\right)$
+$O\left(\sqrt{\frac{\mathsf{Tr}(\bm{\Lambda}^{\star})}{T}\log\frac{1}{\delta}}\right)$
 with probability at least $1-\delta$.
@@
 Figure 3: The decay of the $95\%$ quantile of $\|\overline{\bm{\Delta}}_{T}\|_{2}^{2}$ when (a) $d=3$, for different choices of $\alpha$; and (b) $\alpha=\frac{2}{3}$, for different choices of $d$.
-D.3.2 Berry-Esseen bound
+D.4.2 Berry-Esseen bound
 Figure 4 illustrates how the finite-sample distribution of $\sqrt{T}\overline{\bm{\Delta}}_{T}$ converges to its asymptotic Gaussian distribution $\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})$.
-Notably, plot (a), which is a replication of Figure 2(b) in Section 4, indicates that the convergence is slower when $\alpha=\frac{1}{2}$.
-This directly contradicts the suggestion by Samsonov et al., (2024) to choose $\alpha=\frac{1}{2}$ in order to yield a faster in-distribution convergence of the TD estimation error.
+Notably, plot (a), which is a replication of Figure 2 in Section 4, indicates that the convergence is slower when $\alpha=\frac{1}{2}$.
+This directly contradicts the suggestion by Samsonov et al.
+(2024) to choose $\alpha=\frac{1}{2}$ in order to yield a faster in-distribution convergence of the TD estimation error.
 In fact, with $d=3$ fixed and other choices of $\alpha$, the convergence rate are similar when $T$ is large.
@@
 Figure 4: The decay of $d(\sqrt{T}\overline{\bm{\Delta}}_{T},\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))$, when (a) $d=3$, for different choices of $\alpha$; and (b) $\alpha=\frac{2}{3}$, for different choices of $d$.
-D.3.3 Estimator of the covariance matrix
+D.4.3 Estimator of the covariance matrix
 Figure 5 illustrates the rate by which our variance estimator $\widehat{\bm{\Lambda}}_{T}$ converges to the asymptotic variance $\bm{\Lambda}^{\star}$.
@@
 Refer to caption Refer to caption
-Figure 7: The decay of the difference bewteen estimated and asymptotic covariance matrices, measured in squared Frobenius norm, when (a) $d=3$, for different choices of $\alpha$; and (b) $\alpha=\frac{2}{3}$, for different choices of $d$.
+Figure 5: The decay of the difference between estimated and asymptotic covariance matrices, measured in squared Frobenius norm, when (a) $d=3$, for different choices of $\alpha$; and (b) $\alpha=\frac{2}{3}$, for different choices of $d$.
+D.4.4 Confidence intervals
+Finally, we validate our statistical inference framework on the MDP with $d=3$ and TD learning with $\alpha=\frac{2}{3}$.
+We run 1000 independent trials of Algorithm 1, and at each iteration $T$, we construct both the individual and the simultaneous confidence intervals.
+Specifically, the individual confidence intervals for the three entries of $\bm{\theta}^{\star}$ are constructed by
+$\mathcal{C}_{j}=\left[\overline{\bm{\theta}}_{T}-\Phi(1-\delta/2)\cdot\sqrt{\frac{(\widehat{\bm{\Lambda}}_{T})_{jj}}{T}},\overline{\bm{\theta}}_{T}+\Phi(1-\delta/2)\cdot\sqrt{\frac{(\widehat{\bm{\Lambda}}_{T})_{jj}}{T}}\right],\quad\text{for}\quad j=1,2,3.$
+Here, $\Phi(1-\delta/2)$ is the $1-\frac{\delta}{2}$ quantile of the standard Gaussian distribution.
+In our experiment, we take $\delta=0.05$ and therefore $\Phi(1-\delta/2)=0.96$.
+Figure 6 illustrates the emergence of $\overline{\bm{\theta}}_{T}$ and $\{\mathcal{C}_{j}\}_{j=1,2,3}$ in one single trial.
+As $T$ grows larger, the three entries of the point estimate $\overline{\bm{\theta}}_{T}$ (blue curves) converges to the corresponding entries of the optimal coefficients $\bm{\theta}^{\star}$ (black dotted lines), while the individual confidence intervals $\{\mathcal{C}_{j}\}_{j=1,2,3}$ (blue shades) also concentrate around $\bm{\theta}^{\star}$.
+Figure 6: The emergence of the point estimate $\overline{\bm{\theta}}_{T}$ (blue curves) and individual confidence intervals (blue shades) on the three entries of the optimal coefficients $\bm{\theta}^{\star}$ (black dotted lines) for the MDP with $d=3$, using TD with Polyak-Ruppert averaging, $\alpha=\frac{2}{3}$
+Meanwhile, the simultaneous confidence intervals are constructed as
+$\mathcal{C}=\{\bm{\theta}:\|\bm{\theta}-\overline{\bm{\theta}}_{T}\|_{\infty}\leq x^{\star}\},\quad\text{where}$
+$x^{\star}:=\sup\left\{x:\mathbb{P}_{z\sim\mathcal{N}(\bm{0},\bm{I}_{3})}\left(\left\|\widehat{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}\right\|_{\infty}\leq x\right)\leq 1-\delta\right\}.$
+Essentially, $x^{\star}$ is the $1-\delta$ quantile of the $\ell_{\infty}$ norm of the $\mathcal{N}(\bm{0},\widehat{\bm{\Lambda}}_{T})$ random variable.
+Calculating this quantity is a non-trivial task, so we use $1000$ independent simulations from $\mathcal{N}(\bm{0},\widehat{\bm{\Lambda}}_{T})$ and take the $1-\delta$ quantile of the $\ell_{\infty}$ norm of the samples as an estimate.
+Figure 7 shows the coverage rates of the independent and simultaneous confidence intervals, defined as the percentage of times when the optimal coefficient $\bm{\theta}^{\star}$ (or its entries) are covered by the confidence intervals in the $1000$ independent trials.
+As is indicated by the plot, the coverage rates all approach the target confidence level of $1-\delta=95\%$ when $T$ grows large.
+Figure 7: The emergence coverage rates of independent and simultaneous confidence intervals on the three entries of the optimal coefficients $\bm{\theta}^{\star}$ of the MDP with $d=3$, using TD with Polyak-Ruppert averaging, $\alpha=\frac{2}{3}$.
+Black dotted line represents the target confidence level of $95\%$; blue, orange and green curves represent the coverage rates of the individual confidence intervals $\mathcal{C}_{1}$, $\mathcal{C}_{2}$ and $\mathcal{C}_{3}$ correspondingly; the red curve represents the coverage rate of the confidence region (simultaneous confidence intervals) $\mathcal{C}$.

### v4

@@
 We investigate the statistical properties of Temporal Difference (TD) learning with Polyak-Ruppert averaging, arguably one of the most widely used algorithms in reinforcement learning, for the task of estimating the parameters of the optimal linear approximation to the value function.
-We make three significant contributions that improve the current state-of-the-art results: (i) we derive sharper high probability convergence guarantee that depend explicitly on the asymptotic variance and hold under weaker conditions than those normally assumed; (ii) we establish refined high-dimensional Berry-Esseen bounds over the class of convex sets, achieving faster rates than those previously established in the literature, and (iii) we propose and analyze a novel, computationally efficient online plug-in estimator of the asymptotic covariance matrix.
+Assuming independent samples, we make three theoretical contributions that improve upon the current state-of-the-art results: (i) we derive sharper high probability convergence guarantees that depend explicitly on the asymptotic variance and hold under weaker conditions than those adopted in the literature; (ii) we establish refined high-dimensional Berry-Esseen bounds over the class of convex sets, achieving faster rates than the best known results, and (iii) we propose and analyze a novel, computationally efficient online plug-in estimator of the asymptotic covariance matrix.
 These results enable the construction of confidence regions and simultaneous confidence intervals for the linear parameters of the value function approximation, with guaranteed finite-sample coverage.
@@
 As RL models are based on Markov Decision Processes (MDPs), statistical inference must accommodate the online nature of the sampling mechanism.
-TD learning (see, e.g., Sutton, 1988), an iterative procedure that can be framed as an instance of stochastic approximation (SA), is arguably the most widely used and effective algorithm for policy evaluation in RL.
+TD learning (see, e.g., Sutton,, 1988), an iterative procedure that can be framed as an instance of stochastic approximation (SA), is arguably the most widely used and effective algorithm for policy evaluation in RL.
 The last several years have witnessed a flurry of activities within the statistical machine learning community to derive finite sample consistency rates of TD iterates, leading to deep results and a much improved understanding of the performance and limitations TD learning.
-See, e.g., Khamaru et al.
-(2020); Li et al.
-(2024); Bhandari et al.
-(2021); Dalal et al.
-(2018); Srikant and Ying (2019); Lakshminarayanan and Szepesvari (2018); Mou et al.
-(2020); Durmus et al.
-(2022); Li et al.
-(2024); Samsonov et al.
-(2024); Prashanth et al.
-(2013); Borkar et al.
-(2025) for a partial and growing list of relevant references.
-Though for the tabular case (finite state and action spaces) the TD algorithm is known to be minimax rate optimal Li et al.
-(2024), its properties – and, in particular, its sample complexity – in more complex and high-dimensional settings have not been fully understood.
+See, e.g., Khamaru et al., (2020); Li et al., 2024a ; Bhandari et al., (2021); Dalal et al., (2018); Srikant and Ying, (2019); Lakshminarayanan and Szepesvari, (2018); Mou et al., (2020); Durmus et al., (2022); Li et al., 2024b ; Samsonov et al., 2024b ; Prashanth et al., (2013); Borkar et al., (2025) for a partial and growing list of relevant references.
+Though for the tabular case (finite state and action spaces) the TD algorithm is known to be minimax rate optimal Li et al., 2024a , its properties – and, in particular, its sample complexity – in more complex and high-dimensional settings have not been fully understood.
 paper stepsize convergence guarantee constraints on $t,T,\delta$
 […]
-Li et al. (2024) $\eta_{t}=\eta$ $\|\overline{\bm{\Delta}}_{T}\|_{2}\leq O\left(T^{-\frac{1}{2}}\log^{\frac{1}{2}}\frac{T}{\delta}\right)$ $\eta\sqrt{\log\frac{T}{\delta}}\leq\tilde{C}$
-Samsonov et al. (2024) $\eta_{t}=\eta$ $\|\overline{\bm{\Delta}}_{T}\|_{2}\leq O\left(T^{-\frac{1}{2}}\log\frac{T}{\delta}\right)$ $\eta\log\frac{T}{\delta}\leq\tilde{C}$
-Theorem 3.1 $\eta_{t}=\eta_{0}t^{-\alpha},\alpha\in\left(\frac{1}{2},1\right)$ $\|\overline{\bm{\Delta}}_{T}\|_{2}\leq O\left(T^{-\frac{1}{2}}\sqrt{\log\frac{d}{\delta}}\right)$ $\eta_{0}\sqrt{\log\frac{1}{\delta\eta_{0}}}\leq\tilde{C}$
+Li et al., 2024b $\eta_{t}=\eta$ $\|\overline{\bm{\Delta}}_{T}\|_{2}\leq O\left(T^{-\frac{1}{2}}\log^{\frac{1}{2}}\frac{T}{\delta}\right)$ $\eta\sqrt{\log\frac{T}{\delta}}\leq\widetilde{C}$
+Samsonov et al., 2024b $\eta_{t}=\eta$ $\|\overline{\bm{\Delta}}_{T}\|_{2}\leq O\left(T^{-\frac{1}{2}}\log\frac{T}{\delta}\right)$ $\eta\log\frac{T}{\delta}\leq\widetilde{C}$
+Theorem 3.1 $\eta_{t}=\eta_{0}t^{-\alpha},\alpha\in\left(\frac{1}{2},1\right)$ $\|\overline{\bm{\Delta}}_{T}\|_{2}\leq O\left(T^{-\frac{1}{2}}\sqrt{\log\frac{1}{\delta}}\right)$ $\eta_{0}\sqrt{\log\frac{1}{\delta\eta_{0}}}\leq\widetilde{C}$
 Table 1: Comparisons between our results on TD convergence rate with prior work.
-The quantity $\tilde{C}$, whose value may change at each occurrence, may depend on problem-related parameters other than $T$ and $\delta$.
-In a more specialized line of work, concerned with establishing distributional approximations of the TD iterates, it has been shown that, using Polyak-Ruppert averaging (Polyak and Juditsky, 1992; Ruppert, 1998), the TD estimation error converges to a Gaussian distribution in the large sample limit; see, e.g., Fort (2015); Mou et al.
-(2020); Li et al.
-(2023).
-While these results provide a formal justification for the validity of statistical inference, the guarantees provided in those contributions are non-quantitative, in the sense that they are of asymptotic nature and hold only in fixed dimensions.
+The quantity $\widetilde{C}$, whose value may change at each occurrence, may depend on problem-related parameters other than $T$ and $\delta$.
+The conditions required in Theorem 3.1 are weaker than those in Samsonov et al., 2024b and Li et al., 2024b while delivering the same rates; see discussion in the paragraph below Theorem 3.1.
+In a more specialized line of work, concerned with establishing distributional approximations of the TD iterates, it has been shown that, using Polyak-Ruppert averaging (Polyak and Juditsky,, 1992; Ruppert,, 1998), the TD estimation error converges to a Gaussian distribution in the large sample limit; see, e.g., Fort, (2015); Mou et al., (2020); Li et al., (2023).
+While these results provide a formal justification for the validity of statistical inference, the guarantees from those contributions are non-quantitative, in the sense that they are of asymptotic nature and hold only in fixed dimensions.
 As a result, they do not adequately capture the dependence on the sample size and other problem-related quantities, most notably the dimensionality, all of which affect the accuracy of statistical inference.
@@
 We derive a new high-dimensional Berry-Esseen bound for the estimation error over the class of convex sets and using an estimated variance, achieving an $O(T^{-\frac{1}{3}})$ rate as a function of the number of iterations, which appears to be the fastest in the literature.
-While completing this work, we became aware of the recent related paper by Samsonov et al.
-(2024), which, under the same settings, derived a high-dimensional Berry-Esseen bound for linear SA with Polyak-Ruppert averaging, obtaining a slower rate, of order $O(T^{-\frac{1}{4}})$.
-See Section 3.2 for detailed comparisons.
+While completing this work, we became aware of the recent related paper by Samsonov et al., 2024a , which, under the same settings, derived a high-dimensional Berry-Esseen bound for linear SA with Polyak-Ruppert averaging, obtaining the slower rate $O(T^{-\frac{1}{4}})$.
+See Section 3.2 for a discussion and detailed comparisons.
   3. 3.
@@
 We develop a computationally efficient online algorithm to construct an estimator $\widehat{\bm{\Lambda}}_{T}$ of the asymptotic variance $\bm{\Lambda}^{\star}$.
-We also show that the total variation distance between the (computable) distribution $\mathcal{N}(\bm{0},\widehat{\bm{\Lambda}}_{T})$ and the asymptotic distribution $\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})$ vanishes with high probability at a rate $O(T^{-\frac{1}{3}})$.
+We also show that the total variation distance between the (computable) distribution $\mathcal{N}(\bm{0},\widehat{\bm{\Lambda}}_{T})$ and the asymptotic distribution $\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})$ vanishes with high probability at a rate $O(T^{-\frac{1}{3}})$, matching the order of our Berry-Esseen rate.
+As a result, estimating the variance, a crucial step for constructing confidence sets, does not led to worse rates.
 Figure 1: Emergence of the first entry of TD estimation error on an MDP with a 3-dimensional feature map.
-We verify our theoretical findings with some numerical simulations, discussed in the appendix.
+We verify our theoretical findings with numerical simulations, discussed in the appendix.
 As an illustration, Figure 1 shows how the empirical distribution of TD estimation error in $1,000$ independent trials (as shown in blue histograms), scaled by the square-root of the number of iterations $T$, converges to its asymptotic Gaussian distribution (shown by the red curves).
@@
 condition.
-Our results are derived under the simplifying but standard assumption of i.i.d.
+Our results are derived under the standard assumption of i.i.d.
 data, which has been widely used in the recent literature on statistical analyses of TD learning, SA and stochastic gradient descent (see, e.g.
-Sheshukova et al., 2025).
-Analogous results, obtained using different techniques and yielding slower rates, have been recently obtained for the more challenging case of Markovian data: see Srikant (2024) and Wu et al.
-(2025).
+Sheshukova et al.,, 2025).
+While unrealistic, the theoretical investigation of this simplified scenario has proved challenging; the present work addresses some of the gap still existing in the literature and provide state-of-the art rates.
+We mentioon that analogous results, obtained using different techniques and yielding slower rates, have been recently obtained for the more challenging case of Markovian data: see Srikant, (2024) and Wu et al., (2025).
 Notation.
@@
 For a vector $\bm{x}\in\mathbb{R}^{d}$ and a symmetric matrix $\bm{M}\in\mathbb{R}^{d\times d}$, we let $\|\bm{x}\|_{\bm{M}}=\sqrt{\bm{x}^{\top}\bm{Mx}}$.
-We will use the notation $\tilde{C}$ to express a quantity independent of $T$, the number of iterations/sample size, but possibly dependent on other problem-related parameters.
-Similarly, for positive sequences $\{f_{t}\}_{t\in\mathbb{N}}$ and $\{g_{t}\}_{t\in\mathbb{N}}$, we write $f_{t}\lesssim g_{t}$, or $f_{t}=O(g_{t})$, to signify that $f_{t}\leq\tilde{C}g_{t}$, for some $C>0$ independent of $t$.
+We will use the notation $\widetilde{C}$ to express a quantity independent of $T$, the number of iterations/sample size, but possibly dependent on other problem-related parameters.
+Similarly, for positive sequences $\{f_{t}\}_{t\in\mathbb{N}}$ and $\{g_{t}\}_{t\in\mathbb{N}}$, we write $f_{t}\lesssim g_{t}$, or $f_{t}=O(g_{t})$, to signify that $f_{t}\leq\widetilde{C}g_{t}$, for some $C>0$ independent of $t$.
 The convex distance between two probability distributions on $\mathbb{R}^{d}$, say $P$ and $Q$, is defined as
@@
 Markov decision process.
-An infinite-horizon Markov Decision Process (MDP) (Bertsekas, 2017; Sutton and Barto, 2018) is the tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},P,r,\gamma)$ where $\mathcal{S}$ and $\mathcal{A}$ represent the state space of the environment and the action space available to an agent, respectively, both assumed for simplicity to be finite (though our results hold for countable spaces).
+An infinite-horizon discounted Markov Decision Process (MDP) (Bertsekas,, 2017; Sutton and Barto,, 2018) is the tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},P,r,\gamma)$ where $\mathcal{S}$ and $\mathcal{A}$ represent the state space of the environment and the action space available to an agent, respectively, both assumed for simplicity to be finite (though our results hold for countable spaces).
 The function $P:\mathcal{S}\times\mathcal{A}\mapsto\Delta(\mathcal{S})$ is a transition kernel that associates to any state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$ the probability distribution $P(\cdot\mid s,a)$ of the next state of the environment when the agent executes action $a$ while the environment is in state $s$.
@@
 For any policy $\pi$, its associated reward function $r:\mathcal{S}\mapsto[0,1]$ is the expected reward of the next action selected according to the policy $\pi$, i.e.
-$r(s):=\mathbb{E}_{a\sim\pi(\cdot\mid s)}[r(s,a)],$ for $s\in\mathcal{S}.$ Any policy $\pi$ induces a transition kernel $P$, representing the probability distribution of the next state given the current state $s$ and the policy $\pi$, given by $P(s^{\prime}\mid s)=\mathbb{E}_{a\sim\pi(\cdot\mid s)}P(s^{\prime}\mid s,a)$ for $s,s^{\prime}\in\mathcal{S}.$ We assume throughout that there exists a unique stationary distribution $\mu$ for the transition matrix ${P}$; that is, $\mu(i)=\mathbb{P}_{s\sim\mu,s^{\prime}\sim P(\cdot\mid s)}(s^{\prime}=i)$, $i\in\mathcal{S}$.
+$r(s):=\mathbb{E}_{a\sim\pi(\cdot\mid s)}[r(s,a)],$ for $s\in\mathcal{S}.$ Any policy $\pi$ induces a transition kernel $P$, representing the probability distribution of the next state given the current state $s$ and the policy $\pi$, given by $P(s^{\prime}\mid s)=\mathbb{E}_{a\sim\pi(\cdot\mid s)}P(s^{\prime}\mid s,a)$ for $s,s^{\prime}\in\mathcal{S}.$ We assume throughout that there exists a unique stationary distribution $\mu$ for the transition matrix P{P}; that is, $\mu(i)=\mathbb{P}_{s\sim\mu,s^{\prime}\sim P(\cdot\mid s)}(s^{\prime}=i)$, $i\in\mathcal{S}$.
 The value function and the Bellman equation.
@@
 A fundamental, defining feature of the value function is that it arises as the unique solution to the the Bellman equation, a functional fixed-point equation of the form
-$V(s)=r(s)+\mathbb{E}_{s^{\prime}\sim P(\cdot\mid s)}V(s^{\prime})=:\mathcal{T}V(s).$ (1)
+$V(s)=r(s)+\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot\mid s)}V(s^{\prime})=:\mathcal{T}V(s).$ (1)
 where $\mathcal{T}$ is known as Bellman’s operator.
 Linear policy approximation and the projected Bellman equation.
-In most practical scenarios, the value function is very complex and impossible to compute due to the prohibitively large size of the state space $\mathcal{S}$ A common way to overcome this issue is to consider a parametric approximation of $V$ through a linear function, which is both computationally tractable and amenable to statistical estimation.
-In detail, for a given feature map $\bm{\phi}:\mathcal{S}\mapsto\mathbb{R}^{d}$, mapping each state in $\mathcal{S}$ to a centered feature vector in $\mathbb{R}^{d}$, where $d$ is typically small(er than $|\mathcal{S}|$), a linear value function approximation is given by
+In most practical scenarios, the value function is very complex and impossible to compute due to the prohibitively large size of the state space $\mathcal{S}$.
+A common way to overcome this issue is to consider a parametric approximation of $V$ through a linear function, which is both computationally tractable and amenable to statistical estimation.
+In detail, for a given feature map $\bm{\phi}:\mathcal{S}\mapsto\mathbb{R}^{d}$, mapping each state in $\mathcal{S}$ to a centered feature vector in $\mathbb{R}^{d}$, where $d$ is typically smaller than $|\mathcal{S}|$, a linear value function approximation is given by
 $s\in\mathcal{S}\mapsto V_{\bm{\theta}}(s):={\bm{\phi}}(s)^{\top}{\bm{\theta}},$ (2)
 where ${\bm{\theta}}\in\mathbb{R}^{d}$ is a vector of linear coefficients.
-Throughout this paper, we assume that the Eculidean norm of the features are uniformly bounded, i.e., $\|\bm{\phi}(s)\|_{2}\leq 1$ for all $s\in\mathcal{S}$.
-The best approximating linear function is the one parametrized by the coefficient $\bm{\theta}^{\star}$ solving the projected Bellman equation (Tsitsiklis and Van Roy, 1997) $V_{\bm{\theta}^{\star}}=\Pi\mathcal{T}V_{\bm{\theta}^{\star}},$ with $\Pi$ the orthogonal projection onto the subspace of functions that can be represented by linear combinations of the features.
-See, e.g., Duan et al.
-(2021).
+Throughout this paper, we assume that the Euclidean norm of the features are uniformly bounded, i.e., $\|\bm{\phi}(s)\|_{2}\leq 1$ for all $s\in\mathcal{S}$.
+The best approximating linear function is the one parametrized by the coefficient $\bm{\theta}^{\star}$ solving the projected Bellman equation (Tsitsiklis and Van Roy,, 1997) $V_{\bm{\theta}^{\star}}=\Pi\mathcal{T}V_{\bm{\theta}^{\star}},$ with $\Pi$ the orthogonal projection onto the subspace of functions that can be represented by linear combinations of the features.
+See, e.g., Duan et al., (2021).
 In turn, this (unique!) solution is given by
@@
 where
-$\bm{A}:=\mathop{\mathbb{E}}\limits_{s\sim\mu,s^{\prime}\sim P(\cdot\mid s)}\left[\bm{\phi}(s)\left(\bm{\phi}(s)-\gamma\bm{\phi}(s^{\prime})\right)^{\top}\right]\in\mathbb{R}^{d\times d}\quad\text{and}\quad\bm{b}:=\mathop{\mathbb{E}}\limits_{s\sim\mu}\left[\bm{\phi}(s)r(s)\right]\in\mathbb{R}^{d}.$ (4)
+$\bm{A}:=\mathop{\mathbb{E}}\limits_{s\sim\mu,s^{\prime}\sim P(\cdot\mid s)}\left[\bm{\phi}(s)\left(\bm{\phi}(s)-\gamma\bm{\phi}(s^{\prime})\right)^{\top}\right]\in\mathbb{R}^{d\times d}$ (4)
+$\text{and}\quad\bm{b}:=\mathop{\mathbb{E}}\limits_{s\sim\mu}\left[\bm{\phi}(s)r(s)\right]\in\mathbb{R}^{d}.$ (5)
 Throughout, we let
-$\bm{\Sigma}:=\mathbb{E}_{s\sim\mu}[\bm{\phi}(s)\bm{\phi}^{\top}(s)],\quad\lambda_{0}:=\lambda_{\min}(\bm{\Sigma}),\quad\text{and}\quad\lambda_{\Sigma}:=\lambda_{\max}(\bm{\Sigma})$ (5)
-be the feature covariance matrix and its smallest and largest eigenvalues, respectively.
+$\bm{\Sigma}:=\mathbb{E}_{s\sim\mu}[\bm{\phi}(s)\bm{\phi}^{\top}(s)]$ (6)
+be the feature covariance matrix and use $\lambda_{0}$, $\lambda_{\Sigma}$ to denote its smallest and largest eigenvalues, respectively.
 The Temporal Difference (TD) learning with Polyak-Ruppert averaging.
-While in most real-world RL scenarios the state-action-reward tuples follow a Markov decision process, many of the most recent theoretical contributions – including those delivering the sharpest guarantees listed in Table 1 and Samsonov et al.
-(2024) – have made the simplifying but technically very convenient assumption of $i.i.d.$ data.
+While in most real-world RL scenarios the state-action-reward tuples follow a Markov decision process, many of the most recent theoretical contributions – including those delivering the sharpest guarantees listed in Table 1 and Samsonov et al., 2024a – have made the simplifying but technically very convenient assumption of $i.i.d.$ data.
 Here we follow the same convention and analyze the properties of the TD learning algorithm applied to a sequence of $T$ $i.i.d.$ state-action-reward tuples of the form
 $\{(s_{t},a_{t},s_{t}^{\prime},r_{t})\}_{t=1}^{T},$ (7)
-where $(s_{1},\ldots,s_{T})\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mu$, the stationary distribution, and, for eact $t$, $a_{t}\sim\pi(s_{t}),r_{t}=r(s_{t},a_{t})$ and $s_{t}^{\prime}\sim P(\cdot\mid s_{t},a_{t})$.
+where $(s_{1},\ldots,s_{T})\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mu$, the stationary distribution, and, for each $t$, $a_{t}\sim\pi(s_{t}),r_{t}=r(s_{t},a_{t})$ and $s_{t}^{\prime}\sim P(\cdot\mid s_{t},a_{t})$.
 The TD learning algorithm is a stochastic approximation method to find in an iterative manner the solution (3).
@@
 $\overline{\bm{\theta}}_{T}$ $=\frac{1}{T}\sum_{t=1}^{T}\bm{\theta}_{t}.$ (10)
-This estimator is an instance of the Polyak-Ruppert averaging scheme (Polyak and Juditsky, 1992; Ruppert, 1998), a stochastic approximation methodology that provably reduces the variability and accelerates convergence.
+This estimator is an instance of the Polyak-Ruppert averaging scheme (Polyak and Juditsky,, 1992; Ruppert,, 1998), a stochastic approximation methodology that provably reduces the variability and accelerates convergence.
 Furthermore, with a proper choice of the stepsizes, the averaged iterates converge to a Gaussian distribution, thus enabling statistical inference.
-To date, the best Berry-Esseen bounds rates for stochastic approximation methods and TD learning with Polyak-Ruppert averaging are due to Samsonov et al.
-(2024), which in turn are based on general non-linear Berry-Esseeen bounds by Shao and Zhang (2022).
+To date, the best Berry-Esseen bounds rates for stochastic approximation methods and TD learning with Polyak-Ruppert averaging are due to Samsonov et al., 2024a , which in turn are based on general non-linear Berry-Esseen bounds by Shao and Zhang, (2022).
 In this paper we improve on these results by deriving novel and sharper high-probability and Berry-Esseen bounds for the estimation error of the TD algorithm
@@
 3 RESULTS
-It is known (Polyak and Juditsky, 1992; Fort, 2015) that the TD estimation error of (10) satisfies
-[…]
-where the asymtoptic covariance matrix $\bm{\Lambda}^{\star}$ is given by
-$\bm{\Lambda}^{\star}=\bm{A}^{-1}\bm{\Gamma}\bm{A}^{-\top},\quad\text{with}\quad\bm{\Gamma}=\mathbb{E}[(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}].$ (12)
-Thus, in the fixed-dimensional setting, the above result implies the rate
+It is known (Polyak and Juditsky,, 1992; Fort,, 2015) that, in fixed-dimensional settings, the TD estimation error of (11) satisfies
+$\sqrt{T}\,\overline{\bm{\Delta}}_{T}\xrightarrow{d}\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}),$ (12)
+where the asymptotic covariance matrix $\bm{\Lambda}^{\star}$ is given by
+$\bm{\Lambda}^{\star}=\bm{A}^{-1}\bm{\Gamma}\bm{A}^{-\top},\quad\text{with}$ (13)
+$\bm{\Gamma}=\mathbb{E}[(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}].$ (14)
+Thus, the above result implies the familiar paranetric rate
 $\|\overline{\bm{\Delta}}_{T}\|_{2}=O_{P}\left(\sqrt{\frac{\mathsf{Tr}(\bm{\Lambda}^{\star})}{T}}\right),$ (15)
 which may serve as a benchmark.
-Our main goal is to demonstrate how the TD algorithm can be used to construct non-asymptotic confidence regions and simultaneous confidence intervals for $\bm{\theta}^{*}$ in high dimensions.
-Towards this goal, in the subsequent sections we will establish a high-probability convergence guarantee for $\overline{\bm{\Delta}}_{T}$, develop high-dimensional Berry-Esseen bounds over the class of convex sets, and finally build confidence regions/intervals for $\bm{\theta}^{\star}$ using the estimator $\overline{\bm{\theta}}_{T}$.
+Below, we will extend these guarantees to a high-dimensional scenario by demonstrating an explicit dependence on all model-related parameters including, in particular, the dimension of the feature space.
+Our main goal is to show how the TD algorithm can be used to construct non-asymptotic confidence regions and simultaneous confidence intervals for $\bm{\theta}^{*}$ in high dimensions.
+To that effect, we will establish a high-probability convergence guarantee for $\overline{\bm{\Delta}}_{T}$, develop high-dimensional Berry-Esseen bounds over the class of convex sets, and finally build confidence regionsfor $\bm{\theta}^{\star}$ using $\overline{\bm{\theta}}_{T}$ and a novel, online estimator of the asymptotic variance.
 3.1 High-probability bounds
@@
 For any $\delta\in(0,1)$, there exists $\eta_{0}=\eta_{0}(\delta)>0$, such that
-$\|\overline{\bm{\Delta}}_{T}\|_{2}$ $\lesssim\sqrt{\frac{\mathsf{Tr}(\bm{\Lambda}^{\star})}{T}\log\frac{1}{\delta}}+\tilde{C}T^{-\frac{\alpha+1}{2}}\log\frac{T}{\delta}$
-with probability at least $1-\delta$, where $\tilde{C}$ is a problem-related quantity independent of $T$ and $\delta$.
-The proof can be found in Appendix B.1, where the value of $\tilde{C}$ is given in equation (78).
+$\|\overline{\bm{\Delta}}_{T}\|_{2}$ $\lesssim\sqrt{\frac{\mathsf{Tr}(\bm{\Lambda}^{\star})}{T}\log\frac{1}{\delta}}+\widetilde{C}(T^{-\frac{\alpha+1}{2}}+T^{\frac{\alpha}{2}-1})\log\frac{T}{\delta}$
+with probability at least $1-\delta$, where $\widetilde{C}$ is a problem-related quantity independent of $T$ and $\delta$.
+The proof of this result can be found in Appendix B.1, where the value of $\widetilde{C}$ is given in equation (81).
 Throughout the paper, for the sake of readability, we have opted to express unwieldy problem-dependent quantities that do not directly depend on $T$ in implicit form; their values can be tracked in the proofs.
 Accordingly, when writing our finite-sample bounds, we will first present the leading terms in $T$.
-In a nutshell, Theorem 3.1 states that $\overline{\bm{\theta}}_{T}$ converges to $\bm{\theta}^{*}$ (in Euclidean distance) at the rate of $O(\sqrt{{\mathsf{Tr}(\bm{\Lambda}^{\star})}/{T}})$ up to a log factor.
+In a nutshell, Theorem 3.1 states that $\overline{\bm{\theta}}_{T}$ converges to $\bm{\theta}^{*}$ (in Euclidean distance) at the rate of order $O(\sqrt{{\mathsf{Tr}(\bm{\Lambda}^{\star})}/{T}})$, up to a log factor.
 It is worth comparing the high-probability sample complexity bound of Theorem 3.1 with those in the literature.
-To the best of our knowledge, the tightest results were attained by Dalal et al.
-(2018) for inverse-decaying stepsizes and Li et al.
-(2024); Samsonov et al.
-(2024) for constant stepsize as summarized in Table 1.
+To the best of our knowledge, the tightest results were attained by Dalal et al., (2018) for inverse-decaying stepsizes and Li et al., 2024b ; Samsonov et al., 2024b for constant stepsize as summarized in Table 1.
 Theorem 3.1 is the first result to achieve tight high-probability convergence guarantees up to log factors using polynomial-decaying stepsizes.
 Another subtle but consequential difference concerns the restriction on the step size.
-For our bound to hold, the choice of the initial stepsize $\eta_{0}$ depends on $\delta$; specifically, $\eta_{0}$ is required to be sufficiently small such that $\eta_{0}\sqrt{\log{1}/(\delta\eta_{0}})\leq\tilde{C}$, where $\tilde{C}$ is independent of $T$ and $\delta$ but depends on the other parameters of interest; its specific form is shown in (70) and (71) in the Appendix.
+For our bound to hold, the choice of the initial stepsize $\eta_{0}$ depends on $\delta$ – specifically, $\eta_{0}$ is required to be sufficiently small such that $\eta_{0}\sqrt{\log{1}/(\delta\eta_{0}})\leq\widetilde{C}$, where $\widetilde{C}$ is independent of $T$ and $\delta$ but depends on the other parameters of interest (its specific form is shown in (73) and (74) in the Appendix).
 Though not ideal, this condition is necessary for a Chernoff-type convergence guarantee; we refer to Appendix B.1.3 for a detailed discussion.
-Nonetheless, our stepsize requirements are considerably weaker than those imposed by Samsonov et al.
-(2024) and Li et al.
-(2024), which depend on both $\delta$ and $T$; see Table 1.
+Nonetheless, our stepsize requirements are considerably weaker than those imposed by Samsonov et al., 2024b and Li et al., 2024b , which depend on both $\delta$ and $T$; see Table 1.
 3.2 High-dimensional Berry-Esseen bounds
@@
 Consider the TD algorithm with Polyak-Ruppert averaging (8), independent samples and decaying stepsizes $\eta_{t}=\eta_{0}t^{-\alpha}$ for $\alpha\in(\frac{1}{2},1)$ and $\eta_{0}<1/2$. Then,
-$d_{\mathsf{C}}(\sqrt{T}\overline{\bm{\Delta}}_{T},\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))$ $\lesssim\sqrt{\frac{\eta_{0}}{(1-\alpha)(1-\gamma)\lambda_{0}}}\mathsf{Tr}(\bm{\Gamma})\|\bm{\Gamma}^{-1}\|T^{-\frac{\alpha}{2}}+\frac{\sqrt{d\mathsf{cond}(\bm{\Gamma})}}{(1-\gamma)\lambda_{0}\eta_{0}}T^{\alpha-1}$
-$+\tilde{C}\;T^{-\frac{1}{2}}+\tilde{C}^{\prime}\;T^{2\alpha-2}$
-where $\tilde{C},\tilde{C}^{\prime}$ are problem-related quantities independent of $T$.
+$d_{\mathsf{C}}(\sqrt{T}\overline{\bm{\Delta}}_{T},\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\lesssim\sqrt{\frac{\eta_{0}}{(1-\alpha)(1-\gamma)\lambda_{0}}}\mathsf{Tr}(\bm{\Gamma})\|\bm{\Gamma}^{-1}\|T^{-\frac{\alpha}{2}}+\frac{\sqrt{d\mathsf{cond}(\bm{\Gamma})}}{(1-\gamma)\lambda_{0}\eta_{0}}T^{\alpha-1}+\widetilde{C}\;T^{-\frac{1}{2}}+\widetilde{C}^{\prime}\;T^{2\alpha-2}$
+where $\widetilde{C},\widetilde{C}^{\prime}$ are problem-related quantities independent of $T$.
 On the final rate.
 By choosing $\alpha=2/3$, we obtain a rate in $T$ of order $O(T^{-1/3})$.
-In fact, in Theorem B.3 in the Appendix we establish the more refined result that the Berry-Esseen bound (in the convex distance) between the distribution of $\sqrt{T}\overline{\bm{\Delta}}_{T}$ and the Gaussian distribution with (approximately) matching first two moments is of order $\sqrt{\frac{\eta_{0}}{(1-\alpha)(1-\gamma)\lambda_{0}}}\mathsf{Tr}(\bm{\Gamma})\|\bm{\Gamma}^{-1}\|T^{-\frac{\alpha}{2}}$.
+In fact, in Theorem B.3 in the Appendix we establish the more refined result that the Berry-Esseen bound (in the convex distance) between the distribution of $\sqrt{T}\overline{\bm{\Delta}}_{T}$ and a Gaussian distribution with (approximately) matching first two moments is of order $\sqrt{\frac{\eta_{0}}{(1-\alpha)(1-\gamma)\lambda_{0}}}\mathsf{Tr}(\bm{\Gamma})\|\bm{\Gamma}^{-1}\|T^{-\frac{\alpha}{2}}$.
 As $\alpha\to 1$, this rate approaches the optimal rate of $1/\sqrt{T}$.
 The additional term $\frac{\sqrt{d\mathsf{cond}(\bm{\Gamma})}}{(1-\gamma)\lambda_{0}\eta_{0}}T^{\alpha-1}$ in the above bound, which is ultimately the reason for the final rate $O(T^{-1/3})$, originates from replacing an approximation to the finite sample variance of $\sqrt{T}\overline{\bm{\Delta}}_{T}$ with its asymptotic variance $\bm{\Lambda}^{\star}$.
-Comparison with Samsonov et al.
-(2024).
-Prior to this work, the best Berry-Esseen bounds for stochastic approximation with polynomial-decaying stepsizes and Polyak-Ruppert averaging are those obtained by Samsonov et al.
-(2024), thus covering the settings considered here.
+Comparison with Samsonov et al., 2024a .
+Prior to this work, the best Berry-Esseen bounds for stochastic approximation with polynomial-decaying stepsizes and Polyak-Ruppert averaging are those obtained by Samsonov et al., 2024a , thus covering the settings considered here.
 Specifically, the authors show in their Theorem 3 that when $\eta_{0}\leq\min\{\lambda_{0}(1-\gamma),\frac{1-\gamma}{(1+\gamma)^{2}}\}$ and $\alpha\in[\frac{1}{2},1)$,
 $d_{\mathsf{C}}(\sqrt{T}\overline{\bm{\Delta}}_{T},\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\lesssim O(T^{-\frac{\alpha}{2}}+T^{\frac{\alpha-1}{2}}).$
-Choosing $\alpha=\frac{1}{2}$ leads to a convergence rate of order $O(T^{-\frac{1}{4}})$.
+Choosing $\alpha=\frac{1}{2}$ leads to a convergence rate in $T$ of order $O(T^{-\frac{1}{4}})$.
 In contrast, Theorem 3.2 shows that, by selecting $\alpha=\frac{2}{3}$, we obtain the faster rate $O(T^{-\frac{1}{3}})$.
-As shown in Appendix B.2.3, this rate is tight for $\alpha\in(\frac{2}{3},1)$.
-See also Figure 2 below, showing the output of a numerical experiment demonstrating that the choice of $\alpha=\frac{2}{3}$ leads to a smaller approximation error.
-Though worked out independently, the proofs of our Theorem 3.2 and of Theorem 3 of Samsonov et al.
-(2024) bear similarities, as they both rely on the general Berry-Esseen bounds for non-linear statistics developed by Shao and Zhang (2022).
-We obtain sharper rates by a more targeted analysis of the TD settings, while, in comparison, the results of Samsonov et al.
-(2024) apply more broadly to SA problems.
-Specifically we (i) adopt a different decomposition of the TD estimation error, (ii) handle in a more refined way certain terms appearing in the bound of Theorem 3.1 of Shao and Zhang (2022) [^2] Specifically, the term $I_{3}$ in Appendix B.2.1.
-and (iii) carry out a more specialized analysis of the properties of the matrices $\bm{A}_{t}$ and $\bm{A}$ (see, e.g., Lemma A.4 in the Appendix) originating from TD learning, which has allowed us to tighten problem-related coefficient, as well as to work with a substantially larger intial stepsize $\eta_{0}$.
-This non-trivial sharpening may be potentially generalized to analyze stochastic gradient algorithms applied to non-linear functions, as, e.g., in Theorem 3.4 of Shao and Zhang (2022).
-3.3 Online estimation of the covariance matrix
+As demonstrated in Appendix B.2.3, this rate is tight for $\alpha\in(\frac{2}{3},1)$, since the $O(T^{\alpha-1})$ term is inevitable.
+See also Figure 2 below, showing the output of a numerical experiment indicating that, in fact, the choice of $\alpha=\frac{2}{3}$ leads to a smaller approximation error.
+Though worked out independently, the proofs of our Theorem 3.2 and of Theorem 3 of Samsonov et al., 2024a bear similarities, as they both rely on the general Berry-Esseen bounds for non-linear statistics developed by Shao and Zhang, (2022).
+We obtain sharper rates by a more targeted analysis of the TD settings, while, in comparison, the results of Samsonov et al., 2024a apply more broadly to SA problems.
+Specifically we (i) adopt a different decomposition of the TD estimation error, (ii) handle in a more refined way certain terms appearing in the bound of Theorem 3.1 of Shao and Zhang, (2022) [^2] Specifically, the term $I_{3}$ in Appendix B.2.1.
+and (iii) carry out a more specialized analysis of the properties of the matrices $\bm{A}_{t}$ and $\bm{A}$ (see, e.g., Lemma A.4 in the Appendix) originating from TD learning, which has allowed us to tighten problem-related coefficient, as well as to work with a substantially larger initial stepsize $\eta_{0}$.
+This non-trivial sharpening may be potentially generalized to analyze stochastic gradient algorithms applied to non-linear functions, as, e.g., in Theorem 3.4 of Shao and Zhang, (2022).
+3.3 Online estimation of the covariance
 In the last section, we obtained finite sample bounds on the convex distance between $\sqrt{T}\overline{\bm{\Delta}}_{T}$ and its Gaussian approximation, which features an unknown variance.
 To construct confidence regions for $\bm{\theta}^{\star}$, it is necessary to estimate the covariance matrix $\bm{\Lambda}^{\star}$.
-Two types of approaches have been recently proposed: plug-in estimation (Chen et al., 2020) and bootstraping (White and White, 2010; Hanna et al., 2017; Hao et al., 2021; Ramprasad et al., 2022; Zhu et al., 2023; Samsonov et al., 2024).
+Two types of approaches have been recently proposed: plug-in estimation (Chen et al.,, 2020) and bootstrapping (White and White,, 2010; Hanna et al.,, 2017; Hao et al.,, 2021; Ramprasad et al.,, 2022; Zhu et al.,, 2023; Samsonov et al., 2024a, ).
 Here we propose to use a computationally feasible online plug-in estimator for $\bm{\Lambda}^{\star}$ with superior computational efficiency than existing methodologies and establish its non-asymptotic rate of convergence.
-We will use
-$\overline{\bm{A}}_{T}=\frac{1}{T}\sum_{t=1}^{T}\bm{A}_{t},\quad\text{and}\quad\widehat{\bm{\Gamma}}_{T}=\frac{1}{T}\sum_{t=1}^{T}(\bm{A}_{t}\overline{\bm{\theta}}_{T}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\theta}}_{T}-\bm{b}_{t})^{\top},$ (14)
-as estimators for $\bm{A}$ and $\bm{\Gamma}$ respectively, and set
+Specifically, we consider the following
+$\overline{\bm{A}}_{T}=\frac{1}{T}\sum_{t=1}^{T}\bm{A}_{t},\quad\text{and}$ (16)
+$\widehat{\bm{\Gamma}}_{T}=\sum_{t=1}^{T}w_{t,T}(\bm{A}_{t}\overline{\bm{\theta}}_{t}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\theta}}_{t}-\bm{b}_{t})^{\top}$ (17)
+as estimators for $\bm{A}$ and $\bm{\Gamma}$ respectively, where the weights $\{w_{t,T}\}_{1\leq t\leq T}$ in $\widehat{\bm{\Gamma}}_{T}$ are chosen as
+$w_{t,T}:=\frac{t}{\sum_{t=1}^{T}t}=\frac{2t}{T(T+1)},$
+following the intuition that as $t$ grows larger, $\overline{\bm{\theta}}_{t}$ is a more accurate estimator of $\bm{\theta}^{\star}$ and hence the corresponding term should be assigned higher weights.
+We then consider
 $\widehat{\bm{\Lambda}}_{T}=\overline{\bm{A}}_{T}^{-1}\widehat{\bm{\Gamma}}_{T}\overline{\bm{A}}_{T}^{-\top}$ (18)
 as the final estimator of $\bm{\Lambda}^{\star}$.
-Crucially, because of the linear nature of the problem, the estimator $\widehat{\bm{\Lambda}}_{T}$ can be evaluated in an online manner.
-See Algorithm 1 in Appendix D.1 for a detailed description of the procedure, which has time and space complexities of order $O(Td^{4})$ and $O(d^{4})$, respectively.
-The estimator $\widehat{\bm{\Lambda}}_{T}$ also enjoys good statistical properties.
-This is made precise by the following Gaussian comparison result, whose proof can be found in Appendix B.3, showing that the unknown asymptotic variance $\bm{\Lambda}^{\star}$ in the Berry-Esseen bound of Theorem B.4 can be replaced by $\widehat{\bm{\Lambda}}_{T}$ without compromising the accuracy of the approximation (and without affecting the rate in $T$).
+Crucially, the estimator $\widehat{\bm{\Lambda}}_{T}$ can be evaluated in an online fashion.
+See Algorithm 1 in Appendix D.1 for a detailed description of the procedure, which does not require a priori knowledge of the total round of iterations $T$, and has time and space complexities of order $O(Td^{2}+d^{3})$ and $O(d^{2})$, respectively.
+The estimator $\widehat{\bm{\Lambda}}_{T}$ also enjoys nice statistical properties.
+This is made precise by the following Gaussian comparison result, showing that the unknown asymptotic variance $\bm{\Lambda}^{\star}$ in the Berry-Esseen bound of Theorem B.4 can be replaced by $\widehat{\bm{\Lambda}}_{T}$ without compromising the accuracy of the approximation (i.e., without affecting the rate in $T$).
+The proof of this result can be found in Appendix B.3.
 Theorem 3.3.
@@
 with probability at least $1-3T^{-\frac{1}{3}}-\frac{\sqrt{\mathsf{Tr}(\bm{\Gamma})}}{\lambda_{0}(1-\gamma)\sqrt{T}}$.
-Above, the term $o(T^{-\frac{1}{3}})$ depends on all the parameters of the problem (and, in particular, to $\alpha$) but, as a function of $T$ alone, is of smaller order than $T^{-\frac{1}{3}}$.
+Above, the term $o(T^{-\frac{1}{3}})$ depends on all the parameters of the problem (and, in particular, on $\alpha$) but, as a function of $T$ alone, it is of order $O(T^{-\frac{1}{3}})$.
 See Appendix B.3 for the proof and details.
-Comparison with Chen et al.
-(2020).
-In different settings, Chen et al.
-(2020) considered the problem of carrying out statistical inference for the true model parameters using stochastic gradient descent and assuming a strongly convex and smooth population loss function.
+Comparison with Chen et al., (2020).
+In different settings, Chen et al., (2020) considered the problem of carrying out statistical inference for the true model parameters using stochastic gradient descent and assuming a strongly convex and smooth population loss function.
 In their work, they consider the slightly different plug-in estimator $\widehat{\bm{\Lambda}}_{T}^{\prime}=\overline{\bm{A}}_{T}^{-1}\widehat{\bm{\Gamma}}_{T}^{\prime}\overline{\bm{A}}_{T}^{\top}$, where $\widehat{\bm{\Gamma}}_{T}^{\prime}:=\frac{1}{T}\sum_{t=1}^{T}(\bm{A}_{t}{\bm{\theta}}_{t}-\bm{b}_{t})(\bm{A}_{t}{\bm{\theta}}_{t}-\bm{b}_{t})^{\top}$.
@@
 Comparing (19) with (20), we have sharpened the rate from $O\left(T^{-\frac{\alpha}{2}}\right)$ to $O\left(T^{-\frac{1}{2}}\right)$, independent of the choice of $\alpha$.
-It is noteworthy that the online Algorithm 1 takes advantage of the linearity of the problem; in general, potentially non-linear settings, replacing $\bm{\theta}_{t}$ in $\widehat{\bm{\Gamma}}_{T}^{\prime}$ by $\overline{\bm{\theta}}_{T}$ in $\widehat{\bm{\Gamma}}_{T}$ would require storing all the samples throughout the iterative trajectory, resulting in a larger space complexity.
-Comparison with Samsonov et al. (2024).
-Leveraging the multiplier-bootstrap technique proposed by Zhu et al.
-(2023), Samsonov et al.
-(2024) constructs a randomized estimator $\overline{\bm{\theta}}_{T}^{(b)}$ based on the samples $\mathcal{S}_{T}:=\{s_{t},s_{t}^{\prime}\}_{t=1}^{T}$.
-Their Theorems 4 and 5 imply that, in our notation,
+Comparison with Samsonov et al., 2024a .
+Leveraging the multiplier-bootstrap technique proposed by Zhu et al., (2023), Samsonov et al., 2024a construct a randomized estimator $\overline{\bm{\theta}}_{T}^{(b)}$ based on the samples $\mathcal{S}_{T}:=\{s_{t},s_{t}^{\prime}\}_{t=1}^{T}$.
+Their Theorems 4 and 5 yield that
 $d_{\mathsf{C}}(\sqrt{T}(\overline{\bm{\theta}}_{T}^{(b)}-\bm{\theta}^{\star})|\mathcal{S}_{T},\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\lesssim\frac{\|\bm{\Gamma}^{-1}\|}{[\lambda_{0}(1-\gamma)]^{\frac{3}{2}}}(2\|\bm{\theta}^{\star}\|_{2}+1)^{2}\frac{\log T}{T^{\frac{1}{4}}}+o(T^{-\frac{1}{4}})$ (21)
@@
 (ii) The leading term in the upper bound in Theorem 3.3 entails a smaller problem-related coefficient than that of (21).
-(iii) Perhaps most importantly, in order to estimate the bootstrap distribution $\sqrt{T}(\overline{\bm{\theta}}_{T}^{(b)}-\bm{\theta}^{\star})|\mathcal{S}_{T}$ with an $O(T^{-\frac{1}{4}})$ error rate, it is necessary to store the sample trajectory $\mathcal{S}_{T}$ and generate $O(T^{\frac{1}{2}})$ perturbed trajectories.
+(iii) Perhaps most importantly, in order to estimate the bootstrap distribution with an $O(T^{-\frac{1}{4}})$ error rate, it is necessary to store the sample trajectory $\mathcal{S}_{T}$ and generate $O(T^{\frac{1}{2}})$ perturbed trajectories.
 Consequently, the bootstrap procedure has a time complexity of order $O(T^{\frac{3}{2}}d^{2})$ and a space complexity of order $O(T)$.
-When $d\ll T$, this amounts to a much higher computational burden than our online Algorithm 1.
+This amounts to a much higher computational burden than our online Algorithm 1.
 3.4 Statistical inference
-Combining the results from Sections 3.2 and 3.3, we can now propose a general methodology for constructing convex confidence sets for $\bm{\theta}^{\star}$ with guaranteed finite sample coverage.
-Based on the TD estimator $\overline{\bm{\theta}}_{T}$ (9) and the variance estimator $\widehat{\bm{\Lambda}}_{T}$ (15), and given any tolerance level $\delta\in(0,1)$, it is enough to pick any convex set $\mathcal{C}_{\delta}$ that has probability at least $1-\delta$ under the (random!) Gaussian distribution $\mathcal{N}(\overline{\bm{\theta}}_{T},\frac{\widehat{\bm{\Lambda}}_{T}}{T})$.
-That is, ssattisfying, the condition
+Combining the results from Sections 3.2 and 3.3, we describe a general methodology for constructing convex confidence sets for $\bm{\theta}^{\star}$ with guaranteed finite sample coverage.
+Based on the TD estimator $\overline{\bm{\theta}}_{T}$ in (10) and the variance estimator $\widehat{\bm{\Lambda}}_{T}$ (18), and given any tolerance level $\delta\in(0,1)$, it is enough to pick any convex set $\mathcal{C}_{\delta}$ that has probability at least $1-\delta$ under the Gaussian distribution $\mathcal{N}(\overline{\bm{\theta}}_{T},\frac{\widehat{\bm{\Lambda}}_{T}}{T})$ with random mean and covariance; that is, satisfying the condition
 $\mathbb{P}_{\bm{\theta}\sim\mathcal{N}(\overline{\bm{\theta}}_{T},\frac{\widehat{\bm{\Lambda}}_{T}}{T})}\left(\bm{\theta}\in\mathcal{C}_{\delta}\bigg|\overline{\bm{\theta}}_{T},\widehat{\bm{\Lambda}}_{T}\right)\geq 1-\delta.$ (22)
-Natural choices of $\mathcal{C}_{\delta}$ are (i) hyperctangles, leading to simultaneous confidence intervals of the form
-$\mathcal{C}_{\delta}=\prod_{j=1}^{d}[(\overline{\bm{\theta}}_{T})_{j}-R/\sqrt{T},(\overline{\bm{\theta}}_{T})_{j}+R/\sqrt{T}],$ (20)
-where $R>0$ is the quantile of the $\ell_{\infty}$ norm of a centered Gaussian random vectors with covariance $(\widehat{\bm{\Lambda}}_{T}/{T})$, which can be calculated via simulations; and (ii) ellipsoids or balls, for example
+Natural choices of $\mathcal{C}_{\delta}$ are (i) hyper-rectangles, leading to simultaneous confidence intervals of the form
+$\mathcal{C}_{\delta}=\prod_{j=1}^{d}[(\overline{\bm{\theta}}_{T})_{j}-\widehat{R}/\sqrt{T},(\overline{\bm{\theta}}_{T})_{j}+\widehat{R}/\sqrt{T}],$ (23)
+where $\widehat{R}>0$ is the (random) quantile of the $\ell_{\infty}$ norm of a centered Gaussian random vectors with covariance $(\widehat{\bm{\Lambda}}_{T}/{T})$, which can be calculated via simulations; and (ii) ellipsoids or balls, for example
 $\mathcal{C}_{\delta}=\left\{\bm{\theta}:\sqrt{T}\|\bm{\theta}-\overline{\bm{\theta}}_{T}\|_{\widehat{\bm{\Lambda}}_{T}^{-1}}^{2}\leq R\right\},$ (24)
 where $R$ is the $1-\delta$ quantile of the $\chi^{2}(d)$ distribution.
-The next result, whose proof is in Appendix B.4, derives a coverage rate for $\mathcal{C}_{\delta}$, which converges to the nominal level of $1-\delta$ at a rate of order $O(T^{-1/3})$, as a function of $T$ alone.
+The next result, whose proof is in Appendix B.4, shows that the coverage of $\mathcal{C}_{\delta}$ converges to the nominal level of $1-\delta$ at a rate $O(T^{-1/3})$, as a function of $T$ alone.
 Theorem 3.4.
@@
 We highlight two important features of the above result.
-First, the coverage guarantee holds uniformly over any choice of the convex set $\mathcal{C}_{\delta}$ meeting condition (19), including the common choices like (20) and (21).
-Secondly, the size (i.e., volume) of the confidence sets vanishes at the rate of $O(T^{-1/2})$, keeping all the other variable constant, suggesting that, in this regard, there is little or no room for improvements, at least in terms of rates.
+First, the coverage guarantee holds uniformly over all choices of the convex set $\mathcal{C}_{\delta}$ meeting condition (22), such as those in (23) and (24).
+Secondly, the size (i.e., volume) of the confidence sets vanishes at the standard (and optimal) parametric rate of $O(T^{-1/2})$, keeping all the other variables constant.
 4 SIMULATIONS
-We illustrate our theoretical results with some numerical experiments.
-[^3] The code for our experiments can be found at https://www.dropbox.com/scl/fo/tbw256dlx4gzxzgb4t3hq/AOUaiLFj79iSnc29dFTPpvQ?rlkey=nsgwlcbfx0b73uebwv8xlovag&st=1axbxaa1&dl=0 In order to test our proposed framework for statistical inference, we consider a member of a family of MDPs that are designed to be difficult to distinguish among each other.
-The specifics of the MDP are the same as described in Li et al.
-(2024), and we refer readers to Appendix D for the construction details.
-With $10,000$ independent trials, we examine the high-probability convergence of the $\ell_{2}$ norm of TD estimation errors, the rate by which they converge to their asymptotic Gaussian distributions, and the precision of our estimator for the covariance matrix.
-Figure 2 reveals that the empirical distribution of TD estimation error converges to the asymptotic Gaussian distribution slowly when $\alpha$ is set to ${1}/{2}$ as suggested by Samsonov et al.
-(2024).
-Notice that, for computational simplicity, the difference bewteen the empirical distribution of $\sqrt{T}\overline{\bm{\Delta}}_{T}$ and the asymptotic distribution $\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})$ is measured by
-$d(\sqrt{T}\overline{\bm{\Delta}}_{T},\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})):=\max_{1\leq j\leq d}\sup_{x\in\mathbb{R}}\left|\mathbb{P}(\sqrt{T}(\overline{\bm{\Delta}}_{T})_{j}\leq x)-\mathbb{P}((\bm{\Lambda}^{\star\frac{1}{2}}\bm{z})_{j}\leq x)\right|$
-instead of the convex distance we invoked in the theoretical results.
-Figure 2: Distance between empirical and asymptotic distributions of the TD estimation error.
+We corroborate our theoretical findings with numerical experiments.
+Towards that goal, we consider a family of MDPs that are constructed to be difficult to distinguish from each other.
+The specification of this family of MDPs, along with detailed descriptions of our experiments, is provided in Appendix D.
+Here we summarize some of our results.
+We ran $10,000$ independent trials, each for $T=2^{20}$ iterations, and recorded data at iterations $2^{1},2^{2},2^{3},...,2^{19},2^{20}$.
+We report the rate of convergence of $\|\overline{\bm{\Delta}}_{T}\|_{2}^{2}$, the rate of converge of $\overline{\bm{\Delta}}_{T}$ to its asymptotic Gaussian approximation, and the precision of our estimator for the covariance matrix, under different choices of the parameter $\alpha=\frac{1}{2},\frac{2}{3},\frac{3}{4},1$.
+The results of these experiments confirm that the convergence speed is in accordance with our theoretical findings.
+In particular, for the choices of $\alpha=\frac{2}{3}$, the statistical inference procedure described in Section 3.4 achieves the target coverage rate when $T$ is sufficiently large.
+Figure 2(a) illustrates the behavior of $\|\overline{\bm{\Delta}}_{T}\|_{2}^{2}$ under different choices of $\alpha$.
+As a reference, the black dotted line represents the $95\%$ quantile of $\{\bm{Z}_{T}\}$ where $\bm{Z}_{T}\sim\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}/T)$.
+Our theory predicts that when $T$ grows large, this should be the asymptotic line for the empirical $95\%$ quantiles when $\alpha\in(\frac{1}{2},1)$.
+This is indeed what is shown for $\alpha=\frac{2}{3}$ and $\frac{3}{4}$.
+When $\alpha=1$, however, the empirical quantiles do not appear to converge to the asymptotic quantiles.
+This fact is also consistent with our theory, as reflected in the $O(T^{\frac{\alpha}{2}-1})$ term in the bound of Theorem 3.1, which when $\alpha=1$, converges at the $O(T^{-\frac{1}{2}})$ rate, just like the leading term.
+As for the case of $\alpha=\frac{1}{2}$, we observe that when $T$ is small, $\|\overline{\bm{\Delta}}_{T}\|_{2}$ is relatively large and the empirical and asymptotic lines begin to overlap only when $T$ is very large (i.e., $T>10^{5}$).
+This phenomenon is also predicted by the $O(T^{-\frac{\alpha+1}{2}})$ term in the bound of Theorem 3.1, which vanishes at a slower rate when $\alpha=\frac{1}{2}$.
+Notice that for ease of computation, we use a different measure of the distributional difference than the convex distance, as specified in Eq.
+(124) in Appendix D.
+Figure 2(b) depicts the speed of convergence of the empirical distribution of the rescaled TD estimation error $\sqrt{T}\overline{\bm{\Delta}}_{T}$ to its asymptotic distribution $\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})$.
+When $\alpha=1$, this distance does not converge to 0 as $T$ grows large, attesting to our claim that the $O(T^{\alpha-1})$ term in the distributional difference is inevitable.
+On the other hand, choosing $\alpha=\frac{1}{2}$, as done in Samsonov et al., 2024a , leads to a slower convergence than with $\alpha=\frac{2}{3}$ (our choice) and $\alpha=\frac{3}{4}$.
+Figure 2: (a) Empirical $95\%$ quantiles of $\|\overline{\bm{\Delta}}_{T}\|_{2}$, and (b) distance between the empirical and asymptotic distributions of $\overline{\bm{\Delta}}_{T}$.
 5 DISCUSSION
-We develop high-probability convergence guarantees and Berry-Esseen bounds for TD learning with linear function approximation and $i.i.d.$ samples in a non-asymptotic manner.
+We develop high-probability convergence guarantees and Berry-Esseen bounds for TD learning with linear function approximation and $i.i.d.$ samples.
 We also propose an efficient way to construct a covariance estimator and derive its consistency rates.
 These results lead to a practicable methodology for constructing convex confidence sets for the target parameter, with its size vanishing at a parametric rate $O(T^{-1/2})$.
-An important direction for future research is to investigate whether the order $O(T^{-1/3})$ for the convergence speed of coverage rate in Theorem 3.4 are tight.
+An important direction for future research is to investigate whether the order $O(T^{-1/3})$ for the convergence speed of coverage rate in Theorem 3.4 is tight.
 Another interesting extension would be to derive specialized Berry-Esseen bounds for the sub-class of hyper-rectangles, where one would expect to find a better dependence on the dimension $d$ of the feature space, possibly of poly-logarithmic, as opposed to polynomial, order.
-We believe our analysis and results can be adapted or extended to cover other iterative algorithms, such as stochastic gradient descent and SA methods.
-Recent analyses (see, e.g., Sheshukova et al., 2025; Samsonov et al., 2024) have demonstrated Berry-Esseen bounds for the convex distance (using an estimated variance and assuming i.i.d.
-data), as well as valid bootstrap schemes.
-Acknowledgments and Disclosure of Funding
+Our analysis and results can be generalized to cover other iterative algorithms, such as stochastic gradient descent and stochastic approximation methods.
+In Appendix E, we describe conditions under which our bounds and rates apply to the output of a general linear stochastic approximation algorithm for i.i.d.
+data, a setting that has received attention in the recent literature (see, e.g., Sheshukova et al.,, 2025; Samsonov et al., 2024a, ).
+Acknowledgements
 W.
@@
 For a vector $\bm{x}\in\mathbb{R}^{d}$, we use $\{x_{i}\}_{1\leq i\leq d}$ to denote its entries; for a matrix $\bm{X}\in\mathbb{R}^{m\times n}$, we use $\{X_{ij}\}_{1\leq i\leq m,1\leq j\leq n}$ to denote its entries.
-The vectorization of matrix $\bm{X}\in\mathbb{R}^{m\times n}$, denoted as $\textbf{vec}(\bm{X})$, is defined as a vector in $\mathbb{R}^{mn}$ with entries $[\textbf{vec}(\bm{X})]_{(i-1)m+j}=\bm{X}_{ij}$ for every $1\leq i\leq m$ and $1\leq j\leq n$.
-For any matrices $\bm{X}\in\mathbb{R}^{m\times n}$, and $\bm{Y}\in\mathbb{R}^{p\times q}$, their Kronecker product, denoted as $\bm{X}\otimes\bm{Y}$, is a matrix of the size of $mp\times nq$, and can be written in blocked form as
-$\bm{X}\otimes\bm{Y}=\begin{bmatrix}X_{11}\bm{Y}&X_{12}\bm{Y}&\ldots&X_{1n}\bm{Y}\\ X_{21}\bm{Y}&X_{22}\bm{Y}&\ldots&X_{2n}\bm{Y}\\ \ldots&\ldots&\ldots&\ldots\\ X_{m1}\bm{Y}&X_{m2}\bm{Y}&\ldots&X_{mn}\bm{Y}\end{bmatrix}.$
-With subscripts, it can be represented as $(\bm{X}\otimes\bm{Y})_{(i-1)m+j,(k-1)p+\ell}=X_{ik}Y_{j\ell}$.
+The vectorization of matrix $\bm{X}\in\mathbb{R}^{m\times n}$, denoted as $\textbf{vec}(\bm{X})$, is defined as a vector in $\mathbb{R}^{mn}$ with entries $\textbf{vec}(\bm{X})]_{(i-1)m+j}=\bm{X}_{ij}$ for every $1\leq i\leq m$ and $1\leq j\leq n$.
 For a matrix sequence $\{\bm{F}_{t}\}_{t\in\mathbb{N}}$, a number sequence $\{f_{t}\}_{t\in\mathbb{N}}$ and a fixed matrix $\bm{X}$, if there exists a constant $\widetilde{C}$ independent of $t$ and $\bm{X}$ such that
@@
 $\bm{F}_{t}\bm{X}=O(\|\bm{F}_{t}\|)\bm{X},\quad\text{and}\quad\bm{X}\bm{F}_{t}=O(\|\bm{F}_{t}\|)\bm{X}.$ (28)
-The following lemma, due to Chen et al. (2020)[Lemma 4.1], bounds the difference between the inverse matrices and would be useful for our analysis of the plug-in variance estimator.
+The following lemma, due to Chen et al., (2020)[Lemma 4.1], bounds the difference between the inverse matrices and would be useful for our analysis of the plug-in variance estimator.
 Lemma A.3.
@@
 A direct expansion yields
-00 0 $\leq\mathbb{E}[f^{2}(s)]-2\mathbb{E}[f(s)\bm{\phi}^{\top}(s)\bm{y}]+\bm{y}^{\top}\mathbb{E}[\bm{\phi}(s)\bm{\phi}(s)^{\top}]\bm{y}$
+$0$ $\leq\mathbb{E}[f^{2}(s)]-2\mathbb{E}[f(s)\bm{\phi}^{\top}(s)\bm{y}]+\bm{y}^{\top}\mathbb{E}[\bm{\phi}(s)\bm{\phi}(s)^{\top}]\bm{y}$
 $=\mathbb{E}[f^{2}(s)]-2\mathbb{E}[f(s)\bm{\phi}^{\top}(s)\bm{y}]+\bm{y}^{\top}\bm{\Sigma}\bm{y}$
@@
 $-\beta C(k+1)^{-\alpha}\left(\frac{k+1}{k}\right)^{\nu-\alpha}+(k+1)^{-\alpha}$
-<$\overset{(i)}{<}C+C\left(\frac{k+1}{k}\right)^{\nu-\alpha}\cdot\frac{\nu-\alpha}{k+1}-\beta C(k+1)^{-\alpha}\left(\frac{k+1}{k}\right)^{\nu-\alpha}+(k+1)^{-\alpha}$
-<$\overset{(ii)}{<}C-\frac{\beta}{2}C(k+1)^{-\alpha}\left(\frac{k+1}{k}\right)^{\nu-\alpha}+(k+1)^{-\alpha}$
-<$\overset{(iii)}{<}C-(k+1)^{-\alpha}\left(\frac{k+1}{k}\right)^{\nu-\alpha}+(k+1)^{-\alpha}<C.$
+$\overset{(i)}{<}C+C\left(\frac{k+1}{k}\right)^{\nu-\alpha}\cdot\frac{\nu-\alpha}{k+1}-\beta C(k+1)^{-\alpha}\left(\frac{k+1}{k}\right)^{\nu-\alpha}+(k+1)^{-\alpha}$
+$\overset{(ii)}{<}C-\frac{\beta}{2}C(k+1)^{-\alpha}\left(\frac{k+1}{k}\right)^{\nu-\alpha}+(k+1)^{-\alpha}$
+$\overset{(iii)}{<}C-(k+1)^{-\alpha}\left(\frac{k+1}{k}\right)^{\nu-\alpha}+(k+1)^{-\alpha}<C.$
 Throughout the process, we applied the following facts:
@@
 $Z_{t}<\underset{I_{1}}{\underbrace{t^{\alpha}\prod_{k=1}^{t}\left(1-\beta k^{-\alpha}\right)Z_{1}}}+\underset{I_{2}}{\underbrace{2t^{\alpha}\sum_{i=1}^{t}\left(\frac{1}{\beta}i^{-\alpha-1}+i^{-2\alpha-1}\right)\prod_{k=i+1}^{t}(1-\beta k^{-\alpha})}},$
-where $Z_{1}=1^{\alpha}X_{1}-\frac{1}{\beta}=1-\frac{1}{\beta}$. The proof now boils down to showing that both $I_{1}$ and $I_{2}$ converges to 00 as $t\to\infty$. On one hand, $I_{1}$ is bounded by
+where $Z_{1}=1^{\alpha}X_{1}-\frac{1}{\beta}=1-\frac{1}{\beta}$. The proof now boils down to showing that both $I_{1}$ and $I_{2}$ converges to 0 as $t\to\infty$. On one hand, $I_{1}$ is bounded by
 $t^{\alpha}\exp\left(-\beta\sum_{k=1}^{t}k^{-\alpha}\right)<\left(\frac{1}{\beta}\right)^{\frac{1}{1-\alpha}}t^{\alpha-1}.$
@@
 $=\frac{1}{\beta}\left[\left(\prod_{k=t+1}^{t}(1-\beta k^{-\alpha})+\sum_{j=t+1}^{\infty}\prod_{k=t+1}^{j}(1-\beta k^{-\alpha})\right)-\sum_{j=t}^{\infty}\prod_{k=t+1}^{j+1}(1-\beta k^{-\alpha})\right]$
-=$\overset{(i)}{=}\frac{1}{\beta}\left(1+\sum_{i=t}^{\infty}\prod_{k=t+1}^{i+1}(1-\beta k^{-\alpha})-\sum_{j=t}^{\infty}\prod_{k=t+1}^{j+1}(1-\beta k^{-\alpha})\right)=\frac{1}{\beta}.$ (45)
+$\overset{(i)}{=}\frac{1}{\beta}\left(1+\sum_{i=t}^{\infty}\prod_{k=t+1}^{i+1}(1-\beta k^{-\alpha})-\sum_{j=t}^{\infty}\prod_{k=t+1}^{j+1}(1-\beta k^{-\alpha})\right)=\frac{1}{\beta}.$ (48)
 Notice that in (i) we invoked the convention
 $\prod_{k=t+1}^{t}(1-\beta k^{-\alpha})=1,$
-and a change of variable $i=j-1$.
+and a change of variable i=j1i=j-1.
 Therefore, the proof boils down to showing that the second term is bounded by
@@
 $I_{2}$ $=\underset{I_{21}}{\underbrace{\frac{1}{T}\sum_{t=1}^{T}\bm{S}_{t}\bm{\Gamma}\bm{S}_{t}^{\top}}}+\underset{I_{22}}{\underbrace{\frac{1}{T}\sum_{t=1}^{T}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A})\bm{\Lambda}^{\star}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})}}$
-($-\underset{I_{23}}{\underbrace{\left(\frac{1}{T}\sum_{t=1}^{T}\bm{S}_{t}\bm{\Gamma}\bm{A}^{-\top}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})+\frac{1}{T}\sum_{t=1}^{T}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A})\bm{A}^{-1}\bm{\Gamma}\bm{S}_{t}^{\top}\right)}}.$
+$-\underset{I_{23}}{\underbrace{\left(\frac{1}{T}\sum_{t=1}^{T}\bm{S}_{t}\bm{\Gamma}\bm{A}^{-\top}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A}^{\top})+\frac{1}{T}\sum_{t=1}^{T}\prod_{k=t}^{T}(\bm{I}-\eta_{k}\bm{A})\bm{A}^{-1}\bm{\Gamma}\bm{S}_{t}^{\top}\right)}}.$
 Since $\|\bm{S}_{t}\|=O(T^{\alpha-1})$, it is easy to verify that $I_{21}=O(T^{2\alpha-2})\bm{\Gamma}$; meanwhile, theorem A.9 indicates that $I_{22}=T^{\alpha-1}\bm{X}(\bm{\Lambda}^{\star})+O(T^{2\alpha-2})\bm{\Lambda}^{\star}$.
@@
 A.2.1 Berry-Esseen bounds
-The following theorem, due to Shao and Zhang (2022), gives a Berry-Esseen bound for non-linear statistics which would be crucial for our analysis of the non-asymptotic distribution of the TD estimation error.
+The following theorem, due to Shao and Zhang, (2022), gives a Berry-Esseen bound for non-linear statistics which would be crucial for our analysis of the non-asymptotic distribution of the TD estimation error.
 Theorem A.13.
-(Theorem 2.1 of Shao and Zhang (2022)) Let $\bm{X}_{1},...,\bm{X}_{n}$ be independent random variables, and let $\bm{T}(\bm{X}_{1},...,\bm{X}_{n})$ be a $d$-dimensional statistic that can be represented as
+(Theorem 2.1 of Shao and Zhang, (2022)) Let $\bm{X}_{1},...,\bm{X}_{n}$ be independent random variables, and let $\bm{T}(\bm{X}_{1},...,\bm{X}_{n})$ be a $d$-dimensional statistic that can be represented as
 $\bm{T}=\bm{W}+\bm{D},$
@@
 A.2.2 Properties of multivariate Gaussian distributions
-The following theorem by Devroye et al. (2018) gives an upper bound for the total variation (TV) distance between two Gaussian random variables with same means and different covariance matrices.
-Theorem A.14 (Devroye et al. (2018)).
+The following theorem by Devroye et al., (2018) gives an upper bound for the total variation (TV) distance between two Gaussian random variables with same means and different covariance matrices.
+Theorem A.14 (Devroye et al., (2018)).
 Let $\bm{\Lambda}_{1},\bm{\Lambda}_{2}\in\mathbb{S}^{d\times d}$ be two positive definite matrices, and $\bm{\mu}$ be any vector in $\mathbb{R}^{d}$.
@@
 A.3 Concentration inequalities
-The following Azuma’s inequality for matrix martingales is shown in Tropp (2011).
+The following Azuma’s inequality for matrix martingales is shown in Tropp, (2011).
 Theorem A.16.
@@
 $\mathbb{P}\left(\left\|\sum_{k}\bm{X}_{k}\right\|\geq\varepsilon\right)\leq d\exp\left(-\frac{\varepsilon^{2}}{8\sigma^{2}}\right).$
-Specifically, for vector-valued martingales, the following proposition is shown by Kallenberg and Sztencel (1991).
+Specifically, for vector-valued martingales, the following proposition is shown by Kallenberg and Sztencel, (1991).
 Proposition A.17.
@@
 Let $\{\bm{x}_{i}\}_{i\geq 1}$ be a martingale in $\mathbb{R}^{d}$, and let $W_{\max}$ be a positive constant that satisfies
-$W_{\max}\geq\sum_{i=1}^{t}\|\bm{x}_{i}-\bm{x}_{i-1}\|_{2}^{2}\quad\text{almost surely.}$.
+$W_{\max}\geq\sum_{i=1}^{t}\|\bm{x}_{i}-\bm{x}_{i-1}\|_{2}^{2}\quad\text{almost surely.}$
 Then it can be guaranteed with probability at least $1-\delta$ that
 $\|\bm{x}_{t}\|_{2}\leq 2\sqrt{2W_{\max}\log\frac{3}{\delta}}.$
-The following Freedman’s inequality on Hilbert spaces was shown by Peng et al. (2024) in their Theorem 2.1.
+The following Freedman’s inequality on Hilbert spaces was shown by Peng et al., (2024) in their Theorem 2.1.
 Theorem A.19.
@@
 and assume that
-$16\eta_{0}\sqrt{\frac{2}{2\alpha-1}}\sqrt{\log\frac{9t^{\star}}{\delta}}\leq 1$ (71)
+$16\eta_{0}\sqrt{\frac{2}{2\alpha-1}}\sqrt{\log\frac{9t^{\star}}{\delta}}\leq 1.$ (74)
 Furthermore, we denote, for simplicity,
-$\mathcal{H}_{i}=\{\max_{1\leq j\leq i}\|\bm{\Delta}_{j}\|\leq R\},\quad\text{and}\quad\tilde{\bm{\zeta}}_{i}=\bm{\zeta}_{i}\mathds{1}(\mathcal{H}_{i-1}).$
+$\mathcal{H}_{i}=\{\max_{1\leq j\leq i}\|\bm{\Delta}_{j}\|\leq R\},\quad\text{and}\quad\widetilde{\bm{\zeta}}_{i}=\bm{\zeta}_{i}\mathds{1}(\mathcal{H}_{i-1}).$
 It can then be easily verified that
-$\|\tilde{\bm{\zeta}}_{i}\|_{2}\leq 4R+2\|\bm{\theta}^{\star}\|_{2}+1.$
+$\|\widetilde{\bm{\zeta}}_{i}\|_{2}\leq 4R+2\|\bm{\theta}^{\star}\|_{2}+1.$
 In what follows, we prove the desired result in the following 4 steps:
@@
 Since $\bm{\theta}_{0}=0$, the norm of $\bm{\Delta}_{0}$ is equal to that of $\bm{\theta}^{\star}$, and therefore $\mathbb{P}(\mathcal{H}_{0})=1$ is trivial.
-When (72) holds for $t-1$, we consider the probability of $\mathcal{H}_{t}$.
+When (75) holds for t1t-1, we consider the probability of $\mathcal{H}_{t}$.
 We firstly observe that
-$\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\tilde{\bm{\zeta}}_{i}\right\|_{2}^{2}$ $\leq\sum_{i=1}^{t}\|\bm{R}_{i}^{t}\|^{2}\|\tilde{\bm{\zeta}}_{i}\|_{2}^{2}$
+$\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\widetilde{\bm{\zeta}}_{i}\right\|_{2}^{2}$ $\leq\sum_{i=1}^{t}\|\bm{R}_{i}^{t}\|^{2}\|\widetilde{\bm{\zeta}}_{i}\|_{2}^{2}$
 $<\eta_{0}^{2}(4R+2\|\bm{\theta}^{\star}\|_{2}+1)^{2}\underset{W_{t}}{\underbrace{\sum_{i=1}^{t}i^{-2\alpha}\prod_{k=1}^{t}(1-\beta k^{-\alpha})}}.$
@@
 Therefore, the vectors Azuma’s inequality (Corollary A.18) shows that, with probability at least $1-\frac{\delta}{3t^{\star}}$,
-$\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\tilde{\bm{\zeta}}_{i}\right\|_{2}$ $<2\eta_{0}(4R+2\|\bm{\theta}^{\star}\|_{2}+1)\sqrt{\frac{2}{2\alpha-1}\log\frac{9t^{\star}}{\delta}}$
+$\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\widetilde{\bm{\zeta}}_{i}\right\|_{2}$ $<2\eta_{0}(4R+2\|\bm{\theta}^{\star}\|_{2}+1)\sqrt{\frac{2}{2\alpha-1}\log\frac{9t^{\star}}{\delta}}$
 $\leq\frac{1}{8}(4R+2\|\bm{\theta}^{\star}\|_{2}+1),$
@@
 $\|\bm{\Delta}_{t}\|$ $\leq\left\|\prod_{k=1}^{t}(\bm{I}-\eta_{k}\bm{A})\bm{\Delta}_{0}\right\|_{2}+\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\bm{\zeta}_{i}\right\|_{2}$
-$=\left\|\prod_{k=1}^{t}(\bm{I}-\eta_{k}\bm{A})\bm{\Delta}_{0}\right\|_{2}+\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\tilde{\bm{\zeta}}_{i}\right\|_{2}$
+$=\left\|\prod_{k=1}^{t}(\bm{I}-\eta_{k}\bm{A})\bm{\Delta}_{0}\right\|_{2}+\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\widetilde{\bm{\zeta}}_{i}\right\|_{2}$
 $<\|\bm{\theta}^{\star}\|_{2}+\frac{1}{8}(4R+2\|\bm{\theta}^{\star}\|_{2}+1)=R,$
@@
 it can then be seen that with probability at least $1-\delta_{t}$,
-$\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\tilde{\bm{\zeta}}_{i}\right\|_{2}$ $\leq 2\eta_{0}(4R+2\|\bm{\theta}^{\star}\|_{2}+1)\sqrt{2W_{t}\log\frac{3}{\delta_{t}}}$
+$\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\widetilde{\bm{\zeta}}_{i}\right\|_{2}$ $\leq 2\eta_{0}(4R+2\|\bm{\theta}^{\star}\|_{2}+1)\sqrt{2W_{t}\log\frac{3}{\delta_{t}}}$
 $\leq 2\eta_{0}(4R+2\|\bm{\theta}^{\star}\|_{2}+1)\cdot\frac{1}{16}=\frac{1}{2}R+\frac{1}{4}(2\|\bm{\theta}^{\star}\|_{2}+1).$
@@
 Meanwhile, the vector Azuma’s inequality (Corollary A.18) implies with probability at least $1-\frac{\delta}{3T}$,
-$\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\tilde{\bm{\zeta}}_{i}\right\|_{2}\leq 2\eta_{0}\left(\frac{2\alpha}{\beta}\right)^{\frac{\alpha}{2(1-\alpha)}}\sqrt{\frac{2}{2\alpha-1}\log\frac{9T}{\delta}}(4R+2\|\bm{\theta}^{\star}\|_{2}+1)t^{-\frac{\alpha}{2}}.$
+$\left\|\sum_{i=1}^{t}\bm{R}_{i}^{t}\widetilde{\bm{\zeta}}_{i}\right\|_{2}\leq 2\eta_{0}\left(\frac{2\alpha}{\beta}\right)^{\frac{\alpha}{2(1-\alpha)}}\sqrt{\frac{2}{2\alpha-1}\log\frac{9T}{\delta}}(4R+2\|\bm{\theta}^{\star}\|_{2}+1)t^{-\frac{\alpha}{2}}.$
 The bound (69) then follows by invoking the triangle inequality and a union bound, and plugging in the definition of $R$ and $\beta$. ∎
@@
 $\mathcal{H}_{t}=\left\{\max_{1\leq j\leq t}j^{\frac{\alpha}{2}}\|\bm{\Delta}_{j}\|_{2}\leq R\right\},\quad\text{and}$
-$\tilde{\bm{\Delta}}_{t}=\bm{\Delta}_{t}\mathds{1}(\mathcal{H}_{t}).$
+$\widetilde{\bm{\Delta}}_{t}=\bm{\Delta}_{t}\mathds{1}(\mathcal{H}_{t}).$
 Then by definition, the probability of $\mathcal{H}_{t}$ is bounded by
@@
 Meanwhile, the vector Azuma’s inequality can be invoked to obtain that with probability at least $1-\frac{\delta}{3}$,
-$\left\|\frac{1}{T}\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}-\bm{A})\tilde{\bm{\Delta}}_{i-1}\right\|_{2}\leq 2\sqrt{2W_{\max}\log\frac{9}{\delta}},$
+$\left\|\frac{1}{T}\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}-\bm{A})\widetilde{\bm{\Delta}}_{i-1}\right\|_{2}\leq 2\sqrt{2W_{\max}\log\frac{9}{\delta}},$
 where
-$W_{\max}=\sup\frac{1}{T^{2}}\sum_{i=1}^{T}\|\bm{Q}_{i}(\bm{A}_{i}-\bm{A})\tilde{\bm{\Delta}}_{i-1}\|_{2}^{2}$ $\lesssim\frac{1}{T^{2}}3\eta_{0}^{2}\left(\frac{2}{\beta}\right)^{\frac{1}{1-\alpha}}\cdot 4R^{2}\sum_{i=1}^{T}i^{-\alpha}$
+$W_{\max}=\sup\frac{1}{T^{2}}\sum_{i=1}^{T}\|\bm{Q}_{i}(\bm{A}_{i}-\bm{A})\widetilde{\bm{\Delta}}_{i-1}\|_{2}^{2}$ $\lesssim\frac{1}{T^{2}}3\eta_{0}^{2}\left(\frac{2}{\beta}\right)^{\frac{1}{1-\alpha}}\cdot 4R^{2}\sum_{i=1}^{T}i^{-\alpha}$
 $\asymp\frac{12\eta_{0}^{2}}{1-\alpha}\left(\frac{2}{\beta}\right)^{\frac{1}{1-\alpha}}R^{2}T^{-1-\alpha}.$
 In other words, it can be guaranteed with probability at least $1-\frac{\delta}{3}$ that
-$\left\|\frac{1}{T}\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}-\bm{A})\tilde{\bm{\Delta}}_{i-1}\right\|_{2}$ $\lesssim\eta_{0}\left(\frac{2}{\beta}\right)^{\frac{1}{2(1-\alpha)}}R\sqrt{\frac{1}{1-\alpha}\log\frac{9}{\delta}}$
+$\left\|\frac{1}{T}\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}-\bm{A})\widetilde{\bm{\Delta}}_{i-1}\right\|_{2}$ $\lesssim\eta_{0}\left(\frac{2}{\beta}\right)^{\frac{1}{2(1-\alpha)}}R\sqrt{\frac{1}{1-\alpha}\log\frac{9}{\delta}}$
 $\lesssim\eta_{0}^{2}\left(\frac{2}{\beta}\right)^{\frac{2+\alpha}{2(1-\alpha)}}(6\|\bm{\theta}^{\star}\|_{2}+1)\sqrt{\frac{1}{(1-\alpha)(2\alpha-1)}}T^{-\frac{\alpha+1}{2}}\log\frac{27T}{\delta}.$ (80)
@@
 $+\eta_{0}^{2}\left(\frac{2}{\beta}\right)^{\frac{2+\alpha}{2(1-\alpha)}}(6\|\bm{\theta}^{\star}\|_{2}+1)\sqrt{\frac{1}{(1-\alpha)(2\alpha-1)}}T^{-\frac{\alpha+1}{2}}\log\frac{27T}{\delta}.$
+In the mean time, since the difference between $\overline{\bm{\Lambda}}_{T}$ and $\bm{\Lambda}^{\star}$ is bounded by
+$\mathsf{Tr}(\overline{\bm{\Lambda}}_{T})-\mathsf{Tr}(\bm{\Lambda}^{\star})\lesssim T^{\alpha-1}\mathsf{Tr}(\bm{X}(\bm{\Lambda}^{\star}))\lesssim\frac{\mathsf{Tr}(\bm{\Lambda}^{\star})}{\beta}T^{\alpha-1},$
 Theorem 3.1 follows by taking
-$\tilde{C}$ $=\eta_{0}^{2}\left(\frac{2}{\beta}\right)^{\frac{2+\alpha}{2(1-\alpha)}}(6\|\bm{\theta}^{\star}\|_{2}+1)\sqrt{\frac{1}{(1-\alpha)(2\alpha-1)}}.$ (78)
+$\widetilde{C}$ $=\eta_{0}^{2}\left(\frac{2}{\beta}\right)^{\frac{2+\alpha}{2(1-\alpha)}}(6\|\bm{\theta}^{\star}\|_{2}+1)\sqrt{\frac{1}{(1-\alpha)(2\alpha-1)}}.$ (81)
 B.1.3 The necessity of the condition (74)
@@
 The relationship is specified in Equation (74).
-This constraint, albeit only related to $\delta$ in logarithm terms, prevents $\delta$ to converge to 00 for any fixed $\eta_{0}$.
+This constraint, albeit only related to $\delta$ in logarithm terms, prevents $\delta$ to converge to 0 for any fixed $\eta_{0}$.
 One might naturally wonder if such a constraint could be relaxed and the convergence of $\overline{\bm{\Delta}}_{T}$ can be guaranteed with high probability for any fixed $\eta_{0}$ and arbitrarily small $\delta$.
@@
 $\bm{P}=\begin{pmatrix}0.1&0.1&0.8\\ 0.1&0.1&0.8\\ 0.1&0.1&0.8\end{pmatrix},\quad\text{and}\quad\bm{r}=\begin{pmatrix}0.1\\ 0.1\\ 1\end{pmatrix}$.
-[…]
+It is easy to verify that $\bm{\mu}=(0.1,0.1,0.8)^{\top}$.
+Furthermore, let d=1d=1333In this case, many vectors and matrices reduce to scalars; however, we keep the boldface notation for simplicity., $\bm{\phi}(1)=-0.5$$\bm{\phi}(2)=-1$ and $\bm{\phi}(3)=1$.
+Direct calculation then yields
 $\bm{A}=0.54475,\quad\bm{b}=0.785,\quad\text{and}\quad\bm{\theta}^{\star}=1.441.$
@@
 Consider the TD with Polyak-Ruppert averaging with independent samples and decaying stepsizes $\eta_{t}=\eta_{0}t^{-\alpha}$ for $\alpha\in(\frac{1}{2},1)$ and $\eta_{0}<1/2$. Then,
-$d_{\mathsf{C}}(\sqrt{T}\overline{\bm{\Delta}}_{T},\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T}))\lesssim\sqrt{\frac{\eta_{0}}{(1-\alpha)(1-\gamma)\lambda_{0}}}\mathsf{Tr}(\bm{\Gamma})\|\bm{\Gamma}^{-1}\|T^{-\frac{\alpha}{2}}+\tilde{C}\;T^{-\frac{1}{2}},$
-where $\tilde{C}$ is a problem related quantity independent of $T$.
+$d_{\mathsf{C}}(\sqrt{T}\overline{\bm{\Delta}}_{T},\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T}))\lesssim\sqrt{\frac{\eta_{0}}{(1-\alpha)(1-\gamma)\lambda_{0}}}\mathsf{Tr}(\bm{\Gamma})\|\bm{\Gamma}^{-1}\|T^{-\frac{\alpha}{2}}+\widetilde{C}\;T^{-\frac{1}{2}},$
+where $\widetilde{C}$ is a problem related quantity independent of $T$.
 Theorem B.4.
@@
 Then,
-$d_{\mathsf{C}}(\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T}),\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\leq d_{\mathsf{TV}}(\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T}),\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\lesssim\frac{\sqrt{d\mathsf{cond}(\bm{\Gamma})}}{(1-\gamma)\lambda_{0}\eta_{0}}T^{\alpha-1}+\tilde{C}^{\prime}\;T^{2\alpha-2},$
-where $\tilde{C}^{\prime}$ is a problem related quantity independent of $T$.
+$d_{\mathsf{C}}(\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T}),\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\leq d_{\mathsf{TV}}(\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T}),\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\lesssim\frac{\sqrt{d\mathsf{cond}(\bm{\Gamma})}}{(1-\gamma)\lambda_{0}\eta_{0}}T^{\alpha-1}+\widetilde{C}^{\prime}\;T^{2\alpha-2},$
+where $\widetilde{C}^{\prime}$ is a problem related quantity independent of $T$.
 B.2.1 Proof of Theorem B.3
@@
 Throughout the proof, we will firstly define corresponding statistics, and then bound the terms on the right-hand-side of (68).
-Defining $\bm{W}$,$\bm{D}$ and other statistics.
+Defining $\bm{W}$$\bm{D}$ and other statistics.
 Recall from (76) that the averaged TD error can be represented as
@@
 since $\mathbb{E}_{j^{\prime}-1}[\bm{A}_{j^{\prime}}-\bm{A}]=\bm{0}$.
-This effectively means that the last term on the right-hand-side of (B.2.1) is 00.
-Therefore, the difference bewteen $\Delta$ and $\Delta^{(i)}$ is bounded by
+This effectively means that the last term on the right-hand-side of (B.2.1) is 0.
+Therefore, the difference between $\Delta$ and $\Delta^{(i)}$ is bounded by
 $\mathbb{E}|\Delta-\Delta^{(i)}|^{2}$ $\leq\frac{1}{T}\mathbb{E}\sum_{j=i}^{T}\left\|\bm{AQ}_{j}(\bm{\omega}_{j}-\bm{\omega}_{j}^{(i)})\right\|_{2}^{2}$
@@
 The proof now boils down to bounding the expected difference between $\bm{\omega}_{j}$ and $\bm{\omega}_{j}^{(i)}$, for which we recall from definition that
-$\bm{\omega}_{j}-\bm{\omega}_{j}^{(i)}$ =$=\begin{cases}(\bm{A}_{i}-\check{\bm{A}}_{i})(\bm{\theta}_{i-1}-\bm{\theta}^{\star}),\quad\text{for}\quad j=i;\\ (\bm{A}_{j}-\bm{A})\left(\bm{\theta}_{j-1}-\bm{\theta}_{j-1}^{(i)}\right),\quad\text{for}\quad j>i.\end{cases}$
+$\bm{\omega}_{j}-\bm{\omega}_{j}^{(i)}$ $=\begin{cases}(\bm{A}_{i}-\check{\bm{A}}_{i})(\bm{\theta}_{i-1}-\bm{\theta}^{\star}),\quad\text{for}\quad j=i;\\ (\bm{A}_{j}-\bm{A})\left(\bm{\theta}_{j-1}-\bm{\theta}_{j-1}^{(i)}\right),\quad\text{for}\quad j>i.\end{cases}$
 The norm of $(\bm{\theta}_{i-1}-\bm{\theta}^{\star})$ can be bounded by Lemma B.5; for the norm of $\left(\bm{\theta}_{j-1}-\bm{\theta}_{j-1}^{(i)}\right)$, the following proposition comes in handy.
@@
 Let $\lambda$ be any eigenvalue of $\bm{Y}$ and $\bm{x}$ be its corresponding eigenvector with $\|\bm{x}\|_{2}=1$. Then by definition,
-11 11 $=\eta_{0}\bm{x}^{\top}(\bm{\Gamma}^{-\frac{1}{2}}\bm{A}\bm{\Gamma}^{\frac{1}{2}}\bm{Y}+\bm{Y}\bm{\Gamma}^{\frac{1}{2}}\bm{A}^{\top}\bm{\Gamma}^{-\frac{1}{2}})\bm{x}$
+$1$ $=\eta_{0}\bm{x}^{\top}(\bm{\Gamma}^{-\frac{1}{2}}\bm{A}\bm{\Gamma}^{\frac{1}{2}}\bm{Y}+\bm{Y}\bm{\Gamma}^{\frac{1}{2}}\bm{A}^{\top}\bm{\Gamma}^{-\frac{1}{2}})\bm{x}$
 $=\eta_{0}\lambda\bm{x}^{\top}(\bm{\Gamma}^{-\frac{1}{2}}\bm{A}\bm{\Gamma}^{\frac{1}{2}}+\bm{\Gamma}^{\frac{1}{2}}\bm{A}^{\top}\bm{\Gamma}^{-\frac{1}{2}})\bm{x}$
@@
 We firstly observe that by triangle inequality, for any convex set $\mathcal{A}\subset\mathbb{R}^{d}$ and a $d$-dimensional standard Gaussian random variable $\bm{z}$,
-$\left|\mathbb{P}(\sqrt{T}\overline{\bm{\Delta}}_{T}\in\mathcal{A})-\mathbb{P}(\widehat{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}\in\mathcal{A})\right|$ |($\leq\underset{I_{1}}{\underbrace{\left|\mathbb{P}(\sqrt{T}\overline{\bm{\Delta}}_{T}\in\mathcal{A})-\mathbb{P}(\overline{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}\in\mathcal{A})\right|}}$
-+|($\quad+\underset{I_{2}}{\underbrace{\left|\mathbb{P}(\overline{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}\in\mathcal{A})-\mathbb{P}({\bm{\Lambda}}^{\star\frac{1}{2}}\bm{z}\in\mathcal{A})\right|}}+\underset{I_{3}}{\underbrace{\left|\mathbb{P}({\bm{\Lambda}}^{\star\frac{1}{2}}\bm{z}\in\mathcal{A})-\mathbb{P}(\widehat{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}\in\mathcal{A})\right|}}.$
+$\left|\mathbb{P}(\sqrt{T}\overline{\bm{\Delta}}_{T}\in\mathcal{A})-\mathbb{P}(\widehat{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}\in\mathcal{A})\right|$ $\leq\underset{I_{1}}{\underbrace{\left|\mathbb{P}(\sqrt{T}\overline{\bm{\Delta}}_{T}\in\mathcal{A})-\mathbb{P}(\overline{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}\in\mathcal{A})\right|}}$
+$\quad+\underset{I_{2}}{\underbrace{\left|\mathbb{P}(\overline{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}\in\mathcal{A})-\mathbb{P}({\bm{\Lambda}}^{\star\frac{1}{2}}\bm{z}\in\mathcal{A})\right|}}+\underset{I_{3}}{\underbrace{\left|\mathbb{P}({\bm{\Lambda}}^{\star\frac{1}{2}}\bm{z}\in\mathcal{A})-\mathbb{P}(\widehat{\bm{\Lambda}}_{T}^{\frac{1}{2}}\bm{z}\in\mathcal{A})\right|}}.$
 Here, $I_{1}$ and $I_{2}$ measures the difference between the non-asymptotic and asymptotic distributions of $\overline{\bm{\Delta}}_{T}$, and are bounded by Theorems B.3 and B.4 respectively.
@@
 We firstly represent $\widehat{\bm{\Gamma}}_{T}-\bm{\Gamma}$ as
-$\widehat{\bm{\Gamma}}_{T}-\bm{\Gamma}$ $=\underset{I_{1}}{\underbrace{\frac{1}{T}\sum_{t=1}^{T}\left[(\bm{A}_{t}\overline{\bm{\theta}}_{T}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\theta}}_{T}-\bm{b}_{t})^{\top}-(\bm{A}_{t}{\bm{\theta}}^{\star}-\bm{b}_{t})(\bm{A}_{t}{\bm{\theta}^{\star}}-\bm{b}_{t})^{\top}\right]}}$
-$+\underset{I_{2}}{\underbrace{\frac{1}{T}\sum_{t=1}^{T}\left[(\bm{A}_{t}{\bm{\theta}}^{\star}-\bm{b}_{t})(\bm{A}_{t}{\bm{\theta}^{\star}}-\bm{b}_{t})^{\top}]-\mathbb{E}[(\bm{A}_{t}{\bm{\theta}}^{\star}-\bm{b}_{t})(\bm{A}_{t}{\bm{\theta}^{\star}}-\bm{b}_{t})^{\top}\right]}},$
+$\widehat{\bm{\Gamma}}_{T}-\bm{\Gamma}$ $=\underset{I_{1}}{\underbrace{\sum_{t=1}^{T}w_{t,T}\left[(\bm{A}_{t}\overline{\bm{\theta}}_{t}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\theta}}_{t}-\bm{b}_{t})^{\top}-(\bm{A}_{t}{\bm{\theta}}^{\star}-\bm{b}_{t})(\bm{A}_{t}{\bm{\theta}^{\star}}-\bm{b}_{t})^{\top}\right]}}$
+$+\underset{I_{2}}{\underbrace{\sum_{t=1}^{T}w_{t,T}\left[(\bm{A}_{t}{\bm{\theta}}^{\star}-\bm{b}_{t})(\bm{A}_{t}{\bm{\theta}^{\star}}-\bm{b}_{t})^{\top}]-\mathbb{E}[(\bm{A}_{t}{\bm{\theta}}^{\star}-\bm{b}_{t})(\bm{A}_{t}{\bm{\theta}^{\star}}-\bm{b}_{t})^{\top}\right]}},$
 where $I_{1}$ is mainly controlled by the estimation error of $\overline{\bm{\theta}}_{t}$, and $I_{2}$ only involves the randomness of $\{(\bm{A}_{t},\bm{b}_{t})\}$.
@@
 By definition, $I_{1}$ can be further decomposed as
-$I_{1}$ $=\frac{1}{T}\sum_{t=1}^{T}\left[(\bm{A}_{t}\overline{\bm{\theta}}_{T}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\theta}}_{T}-\bm{b}_{t})^{\top}-(\bm{A}_{t}{\bm{\theta}}^{\star}-\bm{b}_{t})(\bm{A}_{t}{\bm{\theta}^{\star}}-\bm{b}_{t})^{\top}\right]$
-$=\underset{I_{11}}{\underbrace{\frac{1}{T}\sum_{t=1}^{T}\left[(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}\right]}}+\underset{I_{12}}{\underbrace{\frac{1}{T}\sum_{t=1}^{T}\left[(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}+(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}\right]}}.$
+$I_{1}$ $=\sum_{t=1}^{T}w_{t,T}\left[(\bm{A}_{t}\overline{\bm{\theta}}_{t}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\theta}}_{t}-\bm{b}_{t})^{\top}-(\bm{A}_{t}{\bm{\theta}}^{\star}-\bm{b}_{t})(\bm{A}_{t}{\bm{\theta}^{\star}}-\bm{b}_{t})^{\top}\right]$
+$=\underset{I_{11}}{\underbrace{\sum_{t=1}^{T}w_{t,T}\left[(\bm{A}_{t}\overline{\bm{\Delta}}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{t})^{\top}\right]}}+\underset{I_{12}}{\underbrace{\sum_{t=1}^{T}w_{t,T}\left[(\bm{A}_{t}\overline{\bm{\Delta}}_{t})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}+(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{t})^{\top}\right]}}.$
 By triangle inequality, the Frobenius norm of $I_{11}$ is bounded by
-$\left\|\frac{1}{T}\sum_{t=1}^{T}\left[(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}\right]\right\|_{\mathsf{F}}$ $\leq\frac{1}{T}\sum_{t=1}^{T}\left\|(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}\right\|_{\mathsf{F}}$
-$\leq\frac{1}{T}\sum_{t=1}^{T}\sqrt{\mathsf{Tr}\left((\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}\right)}$
-$\leq\frac{1}{T}\sum_{t=1}^{T}4\|\overline{\bm{\Delta}}_{T}\|_{2}^{2}=4\|\overline{\bm{\Delta}}_{T}\|_{2}^{2}.$ (107)
+$\left\|\sum_{t=1}^{T}w_{t,T}\left[(\bm{A}_{t}\overline{\bm{\Delta}}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{t})^{\top}\right]\right\|_{\mathsf{F}}$ $\leq\sum_{t=1}^{T}w_{t,T}\left\|(\bm{A}_{t}\overline{\bm{\Delta}}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{t})^{\top}\right\|_{\mathsf{F}}$
+$\leq\sum_{t=1}^{T}w_{t,T}\sqrt{\mathsf{Tr}\left((\bm{A}_{t}\overline{\bm{\Delta}}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{t})^{\top}(\bm{A}_{t}\overline{\bm{\Delta}}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{t})^{\top}\right)}$
+$\leq\sum_{t=1}^{T}w_{t,T}\cdot 4\|\overline{\bm{\Delta}}_{t}\|_{2}^{2}.$ (110)
 Since
-$\mathbb{E}\|\overline{\bm{\Delta}}_{T}\|_{2}^{2}$ $=\frac{1}{T}\mathsf{Tr}(\overline{\bm{\Lambda}}_{T})+\tilde{C}T^{-\alpha-1}=\frac{1}{T}\mathsf{Tr}(\bm{\Lambda}^{\star})+o(T^{-1})\leq\frac{\mathsf{Tr}(\bm{\Gamma})}{\lambda_{0}^{2}(1-\gamma)^{2}T}+o(T^{-1}),$
-the Markov’s inequality indicates
-$\left\|\frac{1}{T}\sum_{t=1}^{T}\left[(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}\right]\right\|_{\mathsf{F}}\leq\frac{\sqrt{\mathsf{Tr}(\bm{\Gamma})}}{\lambda_{0}(1-\gamma)\sqrt{T}}+o(T^{-\frac{1}{2}})$ (108)
+$\mathbb{E}\|\overline{\bm{\Delta}}_{t}\|_{2}^{2}$ $=\frac{1}{t}\mathsf{Tr}(\overline{\bm{\Lambda}}_{t})+\widetilde{C}t^{-\alpha-1}=\frac{1}{t}\mathsf{Tr}(\bm{\Lambda}^{\star})+o(t^{-1})\leq\frac{\mathsf{Tr}(\bm{\Gamma})}{\lambda_{0}^{2}(1-\gamma)^{2}t}+o(t^{-1}),$
+the expectation of this bound can be bounded by
+$\mathbb{E}\left\|\sum_{t=1}^{T}w_{t,T}\left[(\bm{A}_{t}\overline{\bm{\Delta}}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{t})^{\top}\right]\right\|_{\mathsf{F}}$ $\leq 4\sum_{t=1}^{T}w_{t,T}\cdot\mathbb{E}\|\overline{\bm{\Delta}}_{t}\|_{2}^{2}$
+$\leq 4\sum_{t=1}^{T}\frac{2t}{T(T+1)}\cdot\left(\frac{\mathsf{Tr}(\bm{\Gamma})}{\lambda_{0}^{2}(1-\gamma)^{2}t}+o(t^{-1})\right)$
+$\asymp\frac{\mathsf{Tr}(\bm{\Gamma})}{\lambda_{0}^{2}(1-\gamma)^{2}T}+o(T^{-1});$
+Consequently, the Markov’s inequality indicates
+$\left\|I_{11}\right\|_{\mathsf{F}}\leq\frac{\sqrt{\mathsf{Tr}(\bm{\Gamma})}}{\lambda_{0}(1-\gamma)\sqrt{T}}+o(T^{-\frac{1}{2}})$ (111)
 with probability at least $1-\frac{\sqrt{\mathsf{Tr}(\bm{\Gamma})}}{\lambda_{0}(1-\gamma)\sqrt{T}}$.
 Meanwhile, the norm of $I_{12}$ is bounded by
-$\left\|\frac{1}{T}\sum_{t=1}^{T}\left[(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}+(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}\right]\right\|_{\mathsf{F}}$ (109)
-$\leq\frac{1}{T}\sum_{t=1}^{T}\left\|(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}+(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}\right\|_{\mathsf{F}}$
-$\leq\frac{2}{T}\sum_{t=1}^{T}\left\|(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}\right\|_{\mathsf{F}}$
-$\leq\frac{2}{T}\sum_{t=1}^{T}\sqrt{\mathsf{Tr}\left((\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}\right)}$
-$\leq\frac{4}{T}\sum_{t=1}^{T}4(2\|\bm{\theta}^{\star}\|_{2}+1)\|\overline{\bm{\Delta}}_{T}\|_{2}$
-$=4(2\|\bm{\theta}^{\star}\|_{2}+1)\|\overline{\bm{\Delta}}_{T}\|_{2}.$ (110)
+$\left\|\sum_{t=1}^{T}w_{t,T}\left[(\bm{A}_{t}\overline{\bm{\Delta}}_{t})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}+(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{t})^{\top}\right]\right\|_{\mathsf{F}}$
+$\leq\sum_{t=1}^{T}w_{t,T}\left\|(\bm{A}_{t}\overline{\bm{\Delta}}_{t})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}+(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{t})^{\top}\right\|_{\mathsf{F}}$
+$\leq 2\sum_{t=1}^{T}w_{t,T}\left\|(\bm{A}_{t}\overline{\bm{\Delta}}_{t})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}\right\|_{\mathsf{F}}$
+$\leq 2\sum_{t=1}^{T}w_{t,T}\sqrt{\mathsf{Tr}\left((\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{t})^{\top}(\bm{A}_{t}\overline{\bm{\Delta}}_{t})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}\right)}$
+$\leq 4\sum_{t=1}^{T}w_{t,T}(2\|\bm{\theta}^{\star}\|_{2}+1)\|\overline{\bm{\Delta}}_{t}\|_{2}$
+$=4(2\|\bm{\theta}^{\star}\|_{2}+1)\sum_{t=1}^{T}w_{t,T}\|\overline{\bm{\Delta}}_{t}\|_{2}.$ (112)
+Furthermore, since $\sum_{t=1}^{T}w_{t,T}=1$, the Jensen’s inequality guarantees
+$\mathbb{E}\left(\sum_{t=1}^{T}w_{t,T}\|\overline{\bm{\Delta}}_{t}\|_{2}\right)^{2}$ $\leq\sum_{t=1}^{T}w_{t,T}\mathbb{E}\|\overline{\bm{\Delta}}_{t}\|_{2}^{2}$
+$\leq\sum_{t=1}^{T}\frac{2t}{T(T+1)}\left(\frac{\mathsf{Tr}(\bm{\Gamma})}{\lambda_{0}^{2}(1-\gamma)^{2}t}+o(t^{-1})\right)$
+$\asymp\frac{\mathsf{Tr}(\bm{\Gamma})}{\lambda_{0}^{2}(1-\gamma)^{2}T}+o(T^{-1})$
 Therefore, the Chebyshev’s inequality indicates
-$\left\|\frac{1}{T}\sum_{t=1}^{T}\left[(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}+(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}\right]\right\|_{\mathsf{F}}$
-$\leq 4(2\|\bm{\theta}^{\star}\|_{2}+1)\frac{\sqrt{\mathsf{Tr}(\bm{\Gamma})}}{\lambda_{0}(1-\gamma)}T^{-\frac{1}{3}}+o(T^{-\frac{1}{3}})$ (111)
+$\left\|I_{22}\right\|_{\mathsf{F}}\lesssim 4(2\|\bm{\theta}^{\star}\|_{2}+1)\frac{\sqrt{\mathsf{Tr}(\bm{\Gamma})}}{\lambda_{0}(1-\gamma)}T^{-\frac{1}{3}}+o(T^{-\frac{1}{3}})$ (113)
 with probability at least $1-T^{-\frac{1}{3}}$.
-Hence by combining (108) and (C.4), we obtain
+Hence by combining (111) and (113), we obtain
 $\|I_{1}\|_{\mathsf{F}}\lesssim 4(2\|\bm{\theta}^{\star}\|_{2}+1)\frac{\sqrt{\mathsf{Tr}(\bm{\Gamma})}}{\lambda_{0}(1-\gamma)}T^{-\frac{1}{3}}+o(T^{-\frac{1}{3}})$ (114)
@@
 holds uniformly for all $t\in[T]$. Consequently, the Frobenius norm of $I_{2}$ can be bounded by directly applying the vector Azuma’s inequality: for every $\delta\in(0,1)$,
-$\left\|I_{2}\right\|_{\mathsf{F}}\leq\frac{2\sqrt{2}}{\sqrt{T}}(2\|\bm{\theta}^{\star}\|_{2}+1)^{2}\sqrt{\log\frac{3}{\delta}}$ (113)
-with probability at least $1-\delta$. Lemma B.8 follows by combining (112) and (113) with $\delta=T^{-\frac{1}{3}}$.∎
+$\left\|I_{2}\right\|_{\mathsf{F}}\leq 2\sqrt{2}\sqrt{\sum_{t=1}^{T}w_{t,T}^{2}}(2\|\bm{\theta}^{\star}\|_{2}+1)^{2}\sqrt{\log\frac{3}{\delta}};$ (115)
+with probability at least $1-\delta$. Since
+$\sum_{t=1}^{T}w_{t,T}^{2}$ $=\frac{4}{T^{2}(T+1)^{2}}\sum_{t=1}^{T}t^{2}$
+$=\frac{4}{T^{2}(T+1)^{2}}\cdot\frac{T(T+1)(2T+1)}{6}\asymp\frac{1}{T},$
+Lemma B.8 follows by combining (114) and (115) with $\delta=T^{-\frac{1}{3}}$.∎
 Appendix D EXPERIMENTAL DETAILS
 D.1 Online algorithm to construct the variance estimator
+The code for our experiments can be found at https://www.dropbox.com/scl/fo/u28832upzxu7l24ewb0ar/AKWC7X68R7u3UEmgZ8PY5k0?rlkey=znzky05vh7q8gwc6bdmzhq5ms&st=4zsj1f02&dl=0.
+Experiments were performed on a personal laptop (MacOS).
+No external GPU clusters or cloud resources were used.
+1:MDP transition kernel $\mathcal{P}$, stationary distribution $\mu$, discount factor $\gamma$, feature map $\bm{\phi}$, initial stepsize $\eta_{0}$, stepsize decay parameter $\alpha\in(0.5,1)$, sample size $T$
+2:Initialize $\bm{\theta}=\bm{0}$, $\overline{\bm{\theta}}=\bm{0}$, $\overline{\bm{A}}=\bm{0}$, $\widehat{\bm{\Gamma}}=\bm{0}$.
+3:for $t=1,\dots,T$ do
+4: Sample $s\sim\mu$, $s^{\prime}\sim\mathcal{P}(\cdot\mid s)$.
+5: Compute: $\bm{A}\leftarrow\bm{\phi}(s)[\bm{\phi}(s)-\gamma\bm{\phi}(s^{\prime})],\ \bm{b}\leftarrow r(s)\,\bm{\phi}(s)$.
+6: Perform TD update: $\bm{\theta}\leftarrow\bm{\theta}-\eta_{0}t^{-\alpha}(\bm{A}\bm{\theta}-\bm{b})$.
+7: Perform online updates:
+[…]
+[…]
+[…]
+11:end for
+12:Construct the variance estimator:
+[…]
+14:Averaged TD estimator $\overline{\bm{\theta}}$, covariance matrix estimator $\widehat{\bm{\Lambda}}$.
 Algorithm 1 Online algorithm for computing $\overline{\bm{\theta}}_{T}$ and $\widehat{\bm{\Lambda}}_{T}$
-1:MDP transition kernel $\mathcal{P}$, stationary distribution $\mu$, discount factor $\gamma$, feature map $\bm{\phi}$, initial stepsize $\eta_{0}$, stepsize decay parameter $\alpha\in(0.5,1)$, sample size $T$
-2:Initialize $\bm{\theta}=\bm{0}$, $\overline{\bm{\theta}}=\bm{0}$, $\overline{\bm{A}}=\bm{0}$, $\overline{\bm{AA}}=\bm{0}$, $\overline{\bm{Ab}}=\bm{0}$, $\overline{\bm{bb}}=\bm{0}$
-3:for $t=0:(T-1)$ do
-4:     Sample $s\sim\mu$, $s^{\prime}\sim\mathcal{P}(\cdot\mid s)$
-5:     Compute:
-$\bm{A}\leftarrow\bm{\phi}(s)[\bm{\phi}(s)-\gamma\bm{\phi}(s^{\prime})],\quad\bm{b}\leftarrow r(s)\bm{\phi}(s)$
-6:     Perform TD update:
-$\bm{\theta}\leftarrow\bm{\theta}-\eta_{0}t^{-\alpha}(\bm{A}\bm{\theta}-\bm{b})$
-7:     Perform online updates:
-$\overline{\bm{\theta}}\leftarrow\overline{\bm{\theta}}+\frac{1}{t}(\bm{\theta}-\overline{\bm{\theta}})$
-$\overline{\bm{A}}\leftarrow\overline{\bm{A}}+\frac{1}{t}(\bm{A}-\overline{\bm{A}})$
-$\overline{\bm{AA}}\leftarrow\overline{\bm{AA}}+\frac{1}{t}(\bm{A}\otimes\bm{A}-\overline{\bm{AA}})$
-$\overline{\bm{Ab}}\leftarrow\overline{\bm{Ab}}+\frac{1}{t}(\bm{A}\otimes\bm{b}+\bm{b}\otimes\bm{A}-\overline{\bm{Ab}})$
-$\overline{\bm{bb}}\leftarrow\overline{\bm{bb}}+\frac{1}{t}(\bm{b}\otimes\bm{b}-\overline{\bm{bb}})$
-8:end for
-9:Construct $\widehat{\bm{\Gamma}}$:
-$\widehat{\bm{\Gamma}}\leftarrow\mathsf{reshape}(\overline{\bm{AA}}(\overline{\bm{\theta}}\otimes\overline{\bm{\theta}})-2\overline{\bm{Ab}}(\overline{\bm{\theta}})+\overline{\bm{bb}},(d\times d))$
-10:Construct the variance estimator:
-$\widehat{\bm{\Lambda}}\leftarrow\overline{\bm{A}}^{-1}\widehat{\bm{\Gamma}}\overline{\bm{A}}^{-\top}$
-11:Averaged TD estimator $\overline{\bm{\theta}}$, covariance matrix estimator $\widehat{\bm{\Lambda}}$
-The basic properties of Kronecker products indicate
-$\textbf{vec}\left(\bm{A}_{t}\overline{\bm{\theta}}_{T}\overline{\bm{\theta}}_{T}^{\top}\bm{A}_{t}\right)=(\bm{A}_{t}\otimes\bm{A}_{t})(\overline{\bm{\theta}}_{T}\otimes\overline{\bm{\theta}}_{T});$
-$\textbf{vec}\left(\bm{A}_{t}\overline{\bm{\theta}}_{T}\overline{\bm{b}}_{t}^{\top}\right)=(\bm{A}_{t}\otimes\bm{b}_{t})\overline{\bm{\theta}}_{T};$
-$\textbf{vec}\left(\bm{b}_{t}\bm{\overline{\theta}}_{T}^{\top}\bm{A}_{t}^{\top}\right)=(\bm{b}_{t}\otimes\bm{A}_{t})\overline{\bm{\theta}}_{T};$
-$\textbf{vec}\left(\bm{b}_{t}\bm{b}_{t}^{\top}\right)=\bm{b}_{t}\otimes\bm{b}_{t}.$
-Therefore, the vectorization of $\widehat{\bm{\Gamma}}_{T}$ can be represented as
-$\textbf{vec}(\widehat{\bm{\Gamma}}_{T})$ $=\frac{1}{T}\sum_{t=1}^{T}\textbf{vec}\left(\bm{A}_{t}\overline{\bm{\theta}}_{T}\overline{\bm{\theta}}_{T}^{\top}\bm{A}_{t}-\bm{A}_{t}\overline{\bm{\theta}}_{T}\overline{\bm{b}}_{t}^{\top}-\bm{b}_{t}\bm{\overline{\theta}}_{T}^{\top}\bm{A}_{t}^{\top}+\bm{b}_{t}\bm{b}_{t}^{\top}\right)$
-$=\frac{1}{T}\sum_{t=1}^{T}\left((\bm{A}_{t}\otimes\bm{A}_{t})(\overline{\bm{\theta}}_{T}\otimes\overline{\bm{\theta}}_{T})-(\bm{A}_{t}\otimes\bm{b}_{t})\overline{\bm{\theta}}_{T}-(\bm{b}_{t}\otimes\bm{A}_{t})\overline{\bm{\theta}}_{T}+\bm{b}_{t}\otimes\bm{b}_{t}\right)$
-=($=\underset{I_{1}}{\underbrace{\left(\frac{1}{T}\sum_{t=1}^{T}\bm{A}_{t}\otimes\bm{A}_{t}\right)}}(\overline{\bm{\theta}}_{T}\otimes\overline{\bm{\theta}}_{T})-\underset{I_{2}}{\underbrace{\left(\frac{1}{T}\sum_{t=1}^{T}[\bm{A}_{t}\otimes\bm{b}_{t}+\bm{b}_{t}\otimes\bm{A}_{t}]\right)}}\overline{\bm{\theta}}_{T}$
-$+\underset{I_{3}}{\underbrace{\frac{1}{T}\sum_{t=1}^{T}\bm{b}_{t}\otimes\bm{b}_{t}}},$
-where $I_{1}$, $I_{2}$ and $I_{3}$ are sample means and can hence be calculated using online updates.
-Based on this, we propose Algorithm 1 for the calculation $\widehat{\bm{\Lambda}}_{T}$ in an online manner.
-Since Algorithm 1 iterates from $t=0$ to $t=T-1$ in a single loop and does not store the samples for each iteration, it has a space complexity is $O(Td^{4})$ and a space complexity of $O(d^{4})$.
-Here, the $d^{4}$ term stems from calculating the Kronecker product $\bm{A}\otimes\bm{A}$.
+It is easy to verify that by definition (17), the estimators $\overline{\bm{A}}_{T}$ and $\widehat{\bm{\Gamma}}_{T}$ can be updated by
+$\overline{\bm{A}}_{T}=\overline{\bm{A}}_{T-1}+\frac{1}{T}(\bm{A}_{T}-\overline{\bm{A}}_{T-1}),\quad\text{and}$
+$\widehat{\bm{\Gamma}}_{T}=\widehat{\bm{\Gamma}}_{T-1}+\frac{2}{T+1}\left((\bm{A}_{T}\overline{\bm{\theta}}_{T}-\bm{b}_{T})(\bm{A}_{T}\overline{\bm{\theta}}_{T}-\bm{b}_{T})^{\top}-\widehat{\bm{\Gamma}}_{T-1}\right).$
+Algorithm 1 reveals the complete process of generating both the averaged TD estimator $\overline{\bm{\theta}}_{T}$ and the variance estimator $\widehat{\bm{\Lambda}}_{T}$.
+Since Algorithm 1 iterates from $t=0$ to t=T1t=T-1 in a single loop and does not store the samples for each iteration, it has a space complexity is $O(Td^{2}+d^{3})$ and a space complexity of $O(d^{2})$.
+Here, the $d^{3}$ term stems from calculating matrix inverse.
 D.2 The Markov Decision Process and the feature map
@@
 $P_{\bm{q}}(s^{\prime}\mid s)$
-=$=\left\{\begin{array}[]{lcc}q_{s}\mathds{1}(s^{\prime}=s)+\frac{1-q_{s}}{|\mathcal{S}|-d+1}\mathds{1}(s^{\prime}\geq d)&\text{for}&s<d;\\[5.69046pt] \frac{\gamma}{|\mathcal{S}|-d+1}\mathds{1}(s^{\prime}\geq d)+\frac{1-q_{s^{\prime}}}{d-1}\mathds{1}(s^{\prime}<d)&\text{for}&s\geq d.\end{array}\right.$ (116)
+$=\left\{\begin{array}[]{lcc}q_{s}\mathds{1}(s^{\prime}=s)+\frac{1-q_{s}}{|\mathcal{S}|-d+1}\mathds{1}(s^{\prime}\geq d)&\text{for}&s<d;\\[5.69046pt] \frac{\gamma}{|\mathcal{S}|-d+1}\mathds{1}(s^{\prime}\geq d)+\frac{1-q_{s^{\prime}}}{d-1}\mathds{1}(s^{\prime}<d)&\text{for}&s\geq d.\end{array}\right$.
+(118)
 and the reward function equals to $r(s)=\mathds{1}(s\geq d)$.
@@
 From the expressions above, we remark that, the values of $q$ and $V^{\star}(s)$ with $s\geq d$ are fixed for all $\bm{q}\in\mathcal{Q}$ which is ensured by the construction (119).
-It is shown in Theorem 2 of Li et al.
-(2024) that the sequence $\{\mathcal{M}_{\bm{q}}\}_{q\in\mathcal{Q}}$ is hard to distinguish for a considerably large set $\mathcal{Q}$.
+It is shown in Theorem 2 of Li et al., 2024b that the sequence $\{\mathcal{M}_{\bm{q}}\}_{q\in\mathcal{Q}}$ is hard to distinguish for a considerably large set $\mathcal{Q}$.
 D.3 Parameter specification
 In our experiments, we fix the size of the state space to $\|\mathcal{S}\|=10$, the discount factor to be $\gamma=0.2$, and let $\varepsilon=0.01$.
-For the TD learning algorithm, we set the initial stepsize to be $\eta_{0}=5$, and total number of iterations to be $T=10^{5}$.
-Data is recorded every $100$ iterations.
-In order to show the role played by the stepsize decaying speed $\alpha$, we fix $d=3$ and $\alpha$ iterate through the group $(\frac{1}{2},\frac{2}{3},\frac{3}{4},\frac{4}{5})$; and in order to illustrate the influence of the dimension of the feature map, we fix $\alpha=\frac{2}{3}$ and let $d$ iterate through the group of $(3,5,7,9)$.
+For the TD learning algorithm, we set the initial stepsize to be $\eta_{0}=5$, and total number of iterations to be $T=2^{20}$.
+Data is recorded at every power of 2, so that the log plot has constant intervals on the x-axis.
+In order to show the role played by the stepsize decaying speed $\alpha$, we fix $d=3$ and $\alpha$ iterate through the group $(\frac{1}{2},\frac{2}{3},\frac{3}{4},1$; and in order to illustrate the influence of the dimension of the feature map, we fix $\alpha=\frac{2}{3}$ and let $d$ iterate through the group of $(3,5,7,9)$.
 For every given $d$, we set $\bm{q}$ to be
@@
   3. 3.
-    In order to illustrate the efficiency of our variance estimator $\widehat{\bm{\Lambda}}_{T}$, we measure the averaged Frobenius norm of the difference between $\widehat{\bm{\Lambda}}_{T}$ and $\bm{\Lambda}^{\star}$ among $100$ independent trials.
+    In order to illustrate the efficiency of our variance estimator $\widehat{\bm{\Lambda}}_{T}$, we measure the averaged Frobenius norm of the difference between $\widehat{\bm{\Lambda}}_{T}$ and $\bm{\Lambda}^{\star}$ among $10,000$ independent trials.
 As is indicated by A.14, this is a direct proxy for the TV distance between the zero-mean Gaussian distributions with the two corresponding variance matrices.
   4. 4.
-    In order to validate our statistical inference procedure, we run Algorithm 1 on the MDP with $d=3$ using $\alpha=\frac{2}{3}$, construct both individual and simultaneous confidence intervals for $\bm{\theta}^{\star}$, and measure the coverage rates of these confidence intervals in $1000$ independent trials.
+    In order to validate our statistical inference procedure, we run Algorithm 1 on the MDP with $d=3$ using $\alpha=\frac{2}{3}$, construct both individual and simultaneous confidence intervals for $\bm{\theta}^{\star}$, and measure the coverage rates of these confidence intervals in $10,000$ independent trials.
 Details to follow in section D.4.4.
@@
 D.4.1 High-probability convergence guarantee for the TD estimator
-Figure 3 shows the decay of the $95\%$ quantile of the $L_{2}$ norm of $\overline{\bm{\Delta}}_{T}$; the slope of this log-log plot verifies our theoretical claim in Theorem 3.1 that $\|\overline{\bm{\Delta}}_{T}\|_{2}$ converges by the rate of $O(T^{-\frac{1}{2}})$ with high probability.
-Plot (a) shows that the convergence rate is similar for different choices of $\alpha>\frac{1}{2}$ when $T$ is large, while plot(b) shows that the convergence is slower when $d$ is larger.
+Figure 3 shows the decay of the $95\%$ quantile of $\|\overline{\bm{\Delta}}_{T}\|_{2}^{2}$; the slope of this log-log plot verifies our theoretical claim in Theorem 3.1 that $\|\overline{\bm{\Delta}}_{T}\|_{2}$ converges by the rate of $O(T^{-\frac{1}{2}})$ with high probability.
+Plot (a) illustrates the behavior of the estimation error under different choices of $\alpha$.
+As a reference, the black dotted line represents the $95\%$ quantile of $\{\bm{Z}_{T}\}$ where $\bm{Z}_{T}\sim\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}/T)$.
+Our theory predicts that when $T$ grows large, this should be the asymptotic line for the empirical $95\%$ quantiles when $\alpha\in(\frac{1}{2},1)$.
+This is exactly what we observe that $\alpha=\frac{2}{3}$ and $\alpha=\frac{3}{4}$.
+On the contrary, when $\alpha=1$, the empirical quantiles does not converge to the asymptotic quantiles.
+This is also predicted by our theory that since $\overline{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star}=\bm{O}(T^{\alpha-1})$ (which does not converge when $\alpha=1$), the variance of $\overline{\bm{\Delta}}_{T}$ will not converge to $\bm{\Lambda}^{\star}$ when $\alpha=1$.
+This fact is also reflected in the $O(T^{\frac{\alpha}{2}-1})$ term in the bound of Theorem 3.1, which when $\alpha=1$, converges by $O(T^{-\frac{1}{2}})$, the same rate as the leading term.
+As for the case of $\alpha=\frac{1}{2}$, we observe that when $T$ is small, $\|\overline{\bm{\Delta}}_{T}\|_{2}$ is relatively large; it is not until a very large number of iterations (i.e., $T>10^{5}$), that the empirical and asymptotic lines begin to overlap.
+This phenomenon is also predicted by the $O(T^{-\frac{\alpha+1}{2}})$ term in the bound of Theorem 3.1, which converges slower when $\alpha=\frac{1}{2}$.
+A more detailed explanation follows in Appendix D.5.
+Meanwhile, plot(b) shows that, fixing $\alpha=\frac{2}{3}$, the convergence is slower when $d$ is larger.
 Both are in accordance with our conclusion in Theorem (1) that $\|\overline{\bm{\Delta}}_{T}\|_{2}$ converges by the rate of
@@
 Figure 4 illustrates how the finite-sample distribution of $\sqrt{T}\overline{\bm{\Delta}}_{T}$ converges to its asymptotic Gaussian distribution $\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})$.
-Notably, plot (a), which is a replication of Figure 2 in Section 4, indicates that the convergence is slower when $\alpha=\frac{1}{2}$.
-This directly contradicts the suggestion by Samsonov et al.
-(2024) to choose $\alpha=\frac{1}{2}$ in order to yield a faster in-distribution convergence of the TD estimation error.
-In fact, with $d=3$ fixed and other choices of $\alpha$, the convergence rate are similar when $T$ is large.
+For ease of computation, instead of the convex distance we evaluated the distance
+$\max_{1\leq j\leq d}\sup_{x\in\mathbb{R}}\left|\widehat{\mathbb{P}}(\sqrt{T}(\overline{\bm{\Delta}}_{T})_{j}\leq x)-\mathbb{P}_{\bm{z}}((\bm{\Lambda}^{\star\frac{1}{2}}\bm{z})_{j}\leq x)\right|,$ (124)
+where $\widehat{\mathbb{P}}$ refers to the empirical distribution in our $10,000$ trials, and $\bm{z}$ is the $d$-dimensional standard Gaussian.
+Notably, plot (a) indicates that the convergence is slower when $\alpha=\frac{1}{2}$, compared to the choices of $\alpha=\frac{2}{3}$ and $\alpha=\frac{3}{4}$.
+This directly contradicts the suggestion by Samsonov et al., 2024a to choose $\alpha=\frac{1}{2}$ in order to yield a faster in-distribution convergence of the TD estimation error.
+On the other hand, when $\alpha=1$, the empirical distribution does not converge to $\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})$; that verifies our claim in Appendix B.2.3 that the $O(T^{\alpha-1})$ term in the Berry-Esseen bound is unavoidable.
 Meanwhile, plot (b) shows the convergence generally is slower when $d$ is larger.
@@
 Figure 5 illustrates the rate by which our variance estimator $\widehat{\bm{\Lambda}}_{T}$ converges to the asymptotic variance $\bm{\Lambda}^{\star}$.
-Plot (a) shows that the convergence rate is similar for different choices of $\alpha$ when the number of interations $T$ is sufficiently large; plot(b) indicates that the convergence is slower when $d$ is larger, by the nearly parallel curves in the log-log plot demonstrate that the convergence rate with respect to $T$ is similar among different choices of $d$.
+Plot (a) shows that the convergence rate is similar for different choices of $\alpha$ when the number of iterations $T$ is sufficiently large; plot(b) indicates that the convergence is slower when $d$ is larger, by the nearly parallel curves in the log-log plot demonstrate that the convergence rate with respect to $T$ is similar among different choices of $d$.
 Refer to caption Refer to caption
@@
 Calculating this quantity is a non-trivial task, so we use $1000$ independent simulations from $\mathcal{N}(\bm{0},\widehat{\bm{\Lambda}}_{T})$ and take the $1-\delta$ quantile of the $\ell_{\infty}$ norm of the samples as an estimate.
-Figure 7 shows the coverage rates of the independent and simultaneous confidence intervals, defined as the percentage of times when the optimal coefficient $\bm{\theta}^{\star}$ (or its entries) are covered by the confidence intervals in the $1000$ independent trials.
+Figure 7 shows the coverage rates of the individual and simultaneous confidence intervals, defined as the percentage of times when the optimal coefficient $\bm{\theta}^{\star}$ (or its entries) are covered by the confidence intervals in the $10,000$ independent trials.
 As is indicated by the plot, the coverage rates all approach the target confidence level of $1-\delta=95\%$ when $T$ grows large.
@@
 Black dotted line represents the target confidence level of $95\%$; blue, orange and green curves represent the coverage rates of the individual confidence intervals $\mathcal{C}_{1}$, $\mathcal{C}_{2}$ and $\mathcal{C}_{3}$ correspondingly; the red curve represents the coverage rate of the confidence region (simultaneous confidence intervals) $\mathcal{C}$.
+D.5 Decomposition of TD estimation error
+In this section, we aim to further explore the emergence of the TD estimation error $\overline{\bm{\Delta}}_{T}$ in numerical simulations, in order to verify our key claims in the proofs of our theoretical results.
+Specifically, we aim to explore the role played by the choice of $\alpha$ through decomposing the TD estimation error.
+Recall that throughout our proofs, we repetitively used the decomposition
+$\overline{\bm{\Delta}}_{T}=\underset{I_{1}}{\underbrace{\frac{1}{T\eta_{0}}\bm{Q}_{0}\bm{\Delta}_{0}}}-\underset{I_{2}}{\underbrace{\frac{1}{T}\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i})}}-\underset{I_{3}}{\underbrace{\frac{1}{T}\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}-\bm{A})\bm{\Delta}_{i-1}}},$
+where the matrices $\{\bm{Q}_{i}\}_{0\leq i\leq T}$ are defined by (53), and therefore decided by the matrix $\bm{A}$ as well as the choice of stepsizes $\{\eta_{t}\}_{1\leq t\leq T}$.
+Of the three terms on the right-hand-side:
+  1. 1.
+    The term $I_{1}$ is related only to the initial estimator $\bm{\theta}_{0}$ through $\bm{\Delta}_{0}$ and not to the samples $\{(\bm{A}_{t},\bm{b}_{t})\}_{1\leq t\leq T}$.
+Therefore, we refer to this term as the initialization error.
+This term is deterministic, and according to our analysis, converges by the rate of
+    $\left\|\frac{1}{T\eta_{0}}\bm{Q}_{0}\bm{\Delta}_{0}\right\|_{2}\lesssim O(T^{-1}).$
+  2. 2.
+    The term $I_{2}$ is related only to the samples $\{(\bm{A}_{t},\bm{b}_{t})\}_{1\leq t\leq T}$ and not to any of the iterations $\{\bm{\theta}_{t}\}_{0\leq t\leq T}$.
+Therefore, we refer to this term as the sampling error.
+According to our analysis, this term corresponds to the asymptotic error:
+    $\frac{1}{\sqrt{T}}\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i})\xrightarrow{d}\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})$
+    when $\alpha\in(\frac{1}{2},1)$; non-asymptotically speaking, the variance of this term is $\overline{\bm{\Lambda}}_{T}/T$ by definition, and $\overline{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star}=\bm{O}(T^{\alpha-1})$;
+  3. 3.
+    The term $I_{3}$ is related to both the samples $\{(\bm{A}_{t},\bm{b}_{t})\}_{1\leq t\leq T}$ and the intermediate iterations $\{\bm{\theta}_{t}\}_{0\leq t\leq T}$.
+Hence, we refer to this term as the iterative error.
+According to our analysis, this term converges by the rate of
+    $\left\|\frac{1}{T}\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}-\bm{A})\bm{\Delta}_{i-1}\right\|_{2}\lesssim O(T^{-\frac{\alpha}{2}}).$
+Comparing the magnitudes of the components.
+Figure 8 illustrates the convergence of these terms in our $10,000$ independent trials when $d=3$ and $\alpha=\frac{2}{3}$.
+In the plot, the blue, orange and green lines represent the sample averages of $\|I_{1}\|_{2}^{2},\|I_{2}\|_{2}^{2}$ and $\|I_{3}\|_{2}^{2}$ respectively; the red line represents the evolvement of $\|\overline{\bm{\Delta}}_{T}\|_{2}^{2}$, and the black dotted line represents the asymptotic variance $\mathsf{Tr}(\bm{\Lambda}^{\star})/T$.
+As is shown in the graph, the initialization error $I_{1}$ decreases the fastest; when $T<10^{3}$, the iterative error $I_{2}$ is the dominant component in the estimation error, as the orange and red lines almost overlap.
+However, the slope of the orange line is steeper than that of the green line representing the sampling error $I_{3}$, an illustration that the iterative error converges faster as a function of $T$.
+In the mean time, the green line converges quickly to the black dotted line, verifying our claim that the sampling error corresponds to the asymptotic distribution.
+Finally, as $T$ becomes sufficiently large ($10^{5}$ in this example), the green, red, and black dotted lines overlap with each other, an illustration that the sampling error dominates the estimation error, which converges to its asymptotic distribution.
+Figure 8: Decomposition of the TD estimation error through $10,000$ independent trials, using MDP with $d=3$ and stepsizes $\eta_{t}=\eta_{0}t^{-\frac{2}{3}}$
+Difference between $\overline{\bm{\Lambda}}_{T}$ and $\bm{\Lambda}^{\star}$.
+Theorem A.11 features the difference between the "non-asymptotic variance" $\overline{\bm{\Lambda}}_{T}$ and the asymptotic variance $\bm{\Lambda}^{\star}$ as
+$\overline{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star}\asymp\bm{O}(T^{\alpha-1}),$
+which results in our conclusion that the difference between the non-asymptotic and asymptotic distributions of $\overline{\bm{\Delta}}_{T}$ is lower bounded by $O(T^{\alpha-1})$.
+In order to illustrate this difference in variance through numerical simulations, we use $\check{\bm{\Lambda}}_{T}$ to denote the sample mean (averaged over our $10,000$ independent trials) of
+$\frac{1}{T}\left(\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i})\right)\left(\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i})\right)^{\top}.$
+By definition, $\check{\bm{\Lambda}}_{T}$ is an unbiased estimator of $\overline{\bm{\Lambda}}_{T}$.
+Figure 9 demonstrates the difference between $\check{\bm{\Lambda}}_{T}$ and $\bm{\Lambda}^{\star}$, measured by the Frobenius norm of their difference, as $T$ increases.
+It is clear from the plot that the difference converges slower when $\alpha$ is larger, and especially that it does not converge to 0 when $\alpha=1$.
+Notice that the plot is truncated before $T=10^{4}$ as when $T$ becomes larger than the number of trials, the randomness within $\check{\bm{\Lambda}}_{T}$ dominates the difference, making the plot uninformative.
+Figure 9: The difference between $\check{\bm{\Lambda}}_{T}$ and $\bm{\Lambda}^{\star}$, for $d=3$ and different choices of $\alpha$
+Appendix E EXTENSION TO GENERAL LINEAR STOCHASTIC APPROXIMATION
+In this section, we explore on the conditions under which our theoretical analysis and statistical inference procedure can be extended to general Linear Stochastic Approximation (LSA) for the linear system
+$\bm{A\theta}=\bm{b},$
+where a sample matrix $\bm{A}_{t}$ and a sample vector $\bm{b}_{t}$ are observed $i.i.d.$ at each time $t$, with expectations being $\bm{A}$ and $\bm{b}$ respectively.
+Similar to TD learning, we also run
+$\bm{\theta}_{t}=\bm{\theta}_{t-1}-\eta_{t}(\bm{A}_{t}\bm{\theta}_{t-1}-\bm{b}_{t}),$
+with the same choice of stepsizes $\eta_{t}=\eta_{0}t^{-\alpha}$, $\alpha\in(\frac{1}{2},1)$.
+We first make the following assumptions on the matrix $\bm{A}$:
+Assumption E.1 (Stability).
+There exists a $d\times d$ symmetric matrix $\bm{U}\succeq\bm{0}$ and a constant $\beta_{0}>0$, such that $\bm{AU}+\bm{UA}^{\top}\succeq\beta_{0}\bm{I}$;
+Notice that in our setting of TD learning, this assumption is satisfied with $\bm{U}=\bm{I}$, and $\beta_{0}=2(1-\gamma)\lambda_{0}$; for general LSA, it is guaranteed as long as $\bm{A}$ is Hurwitz, i.e., all the eigenvalues of $\bm{A}$ has strictly negative real-parts.
+Under that condition, $\beta_{0}$ can be set to $1$, and $\bm{U}$ is the unique solution to the Lyapunov equation $\bm{AU}+\bm{UA}^{\top}=\bm{I}$.
+Meanwhile, we need the following constraints on the samples $\{(\bm{A}_{t},\bm{b}_{t})\}$:
+Assumption E.2 (Sampling constraints).
+With $\bm{U}$ specified as in Assumption E.1,
+  1. 1.
+    There exists a constant $c_{U}>0$, such that $\mathbb{E}[\bm{A}_{t}\bm{U}\bm{A}_{t}^{\top}]\preceq c_{U}(\bm{AU}+\bm{UA}^{\top})$;
+  2. 2.
+    There exists constants $c_{A}>0,c_{b}>0$, such that $\|\bm{A}_{t}\|_{\bm{U}}\leq c_{A}$ and $\|\bm{b}_{t}\|_{\bm{U}}\leq c_{b}$ almost surely.
+Under our setting of TD learning, Assumption E.2 is satisfied with $c_{U}=1,c_{A}=2$ and $c_{b}=1$ by Lemma A.4.
+For general LSA with $\bm{A}$ being Hurwitz, as long as (2) is satisfied, (1) can also be guaranteed by setting $c_{U}=c_{A}\|\bm{U}\|$.
+As long as both Assumption E.2 and Assumption E.2 are satisfied, it is immediate that $\|\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t}\|_{\bm{U}}\leq c_{A}\|\bm{\theta}^{\star}\|_{\bm{U}}+c_{b}$ almost surely, and that for every $\bm{x}\in\mathbb{R}^{d}$,
+$\|(\bm{I}-\eta\bm{A})\bm{x}\|_{\bm{U}}^{2}$ $\leq\mathbb{E}\|(\bm{I}-\eta\bm{A}_{t})\bm{x}\|_{\bm{U}}^{2}$
+$=\|\bm{x}\|_{\bm{U}}^{2}-\eta\bm{x}^{\top}(\bm{AU}+\bm{UA}^{\top})\bm{x}+\eta^{2}\bm{x}^{\top}\mathbb{E}[\bm{A}_{t}\bm{U}\bm{A}_{t}^{\top}]\bm{x}$
+$\leq\|\bm{x}\|_{\bm{U}}^{2}-(\eta-c_{U}\eta^{2})\bm{x}^{\top}(\bm{AU}+\bm{UA}^{\top})\bm{x};$
+as long as
+$\eta\leq\frac{1}{2c_{U}}=:\eta_{\max},$
+this bound can be further elaborated as
+$\|(\bm{I}-\eta\bm{A})\bm{x}\|_{\bm{U}}^{2}$ $\leq\mathbb{E}\|(\bm{I}-\eta\bm{A}_{t})\bm{x}\|_{\bm{U}}^{2}$
+$\leq\|\bm{x}\|_{\bm{U}}^{2}-\frac{\eta}{2}\bm{x}^{\top}(\bm{AU}+\bm{UA}^{\top})\bm{x}$
+$\leq\|\bm{x}\|_{\bm{U}}^{2}-\frac{\beta_{0}\eta}{2}\|\bm{x}\|_{2}^{2}\leq\left(1-\frac{\beta_{0}\eta}{2\|\bm{U}\|}\right)\|\bm{x}\|_{\bm{U}}^{2};$
+Consequently, as long as $\eta\leq\eta_{\max}$,
+$\|\bm{I}-\eta\bm{A}\|_{\bm{U}}\leq\sqrt{1-\frac{\beta_{0}\eta}{2\|\bm{U}\|}}\leq 1-\frac{\beta_{0}\eta}{4\|\bm{U}\|}.$
+This is a direct generalization of Lemma A.4.
+While the detailed derivations are well beyond the scope of this paper, our theoretical analysis in Theorems 3.1, 3.2, 3.3 can all be extended to general LSA with this generalized version of Lemma A.4.
+Specifically:
+  1. 1.
+    For every $\delta\in(0,1)$, there exists a proper choice of $\eta_{0}$, such that with probability at least $1-\delta$,
+    $\|\overline{\bm{\Delta}}_{T}\|_{\bm{U}}\lesssim\sqrt{\frac{\mathsf{Tr}(\bm{U\Lambda}^{\star})}{T}\log\frac{1}{\delta}};$
+  2. 2.
+    As long as $\eta_{0}\leq\eta_{\max}$, it can be guaranteed that
+    $d_{\mathsf{C}}(\sqrt{T}\overline{\bm{\Delta}}_{T},\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\lesssim C_{1}T^{-\frac{\alpha}{2}}+C_{2}T^{\alpha-1},$
+    with coefficients $C_{1}$ and $C_{2}$ related to $\bm{U},\beta_{0},c_{U},c_{A},c_{b}$. Therefore, choosing $\alpha=\frac{2}{3}$ would still yield an $O(T^{-1/3})$ convergence rate;
+  3. 3.
+    With variance estimator $\widehat{\bm{\Lambda}}_{T}$ defined in the same way as (18),
+    $d_{\mathsf{TV}}(\mathcal{N}(\bm{0},\widehat{\bm{\Lambda}}_{T}),\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\lesssim O(T^{-\frac{1}{3}})$
+    with probability at least $1-O(T^{-\frac{1}{3}})$;
+  4. 4.
+    Under the same statistical inference procedure as described in Section 3.4, the confidence region $\mathcal{C}_{\delta}$ still satisfies
+    $\mathbb{P}(\bm{\theta}^{\star}\in\mathcal{C}_{\delta})$ $\geq 1-\delta-O(T^{-\frac{1}{3}}).$